A Complete Bibliography of ACM Queue: Tomorrow’s Computing Today

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 June 2024 Version 2.117

Title word cross-reference

#1 [O’H06j].

! = [BNB^{+}05]. + [Ano03f]. $10 [Lim20c]. $109 [Nag04]. $125 [And04]. $14.95 [Bal03]. $15.00 [Imr03]. $165.00 [Ber09]. $174.00 [Pet04]. $23.95 [Kau03]. $25 [Kil04]. $25.00 [Goo07]. $29.99 [Hil08, Men04, Put08]. 3 [Fra20e, Sta05b]. $34.95 [Che03]. $34.99 [Gib07, Kar05]. $39.95 [Eng05, Sch04b, Sta06n, Ste03, Whi08]. $39.99 [Bal07a, Git06, Hor03, Kuc06, Mer04, Ram08, Spi07]. 42 [Wad05]. $42.99 [ZM09]. $44.95 [Joy04, Kha05, Mar04b, Wol06]. $44.99 [Bal09, Ber05, Edw05, Hin08, Lec07, Mor06, Sta07v, Wal07, ZM09]. $45.00 [Rao04b]. $49.95 [Jen05b, Mar05b, Sta06m]. $49.99 [Bal07b, Bel05b, Bel08, Cha05a, Fox08, Gum04b, Hid04, Jov06, Llo06, Mar05a, Nau08a, Pos04, Sta05k, Sta06l, Sta07u, Sta06k, Tho05, Tom05, Tom06]. $54.95 [Ber07, Zen04]. $54.99 [Hac06, Nau08b]. $55.00 [Ber03b, Sta05j]. $59.95 [Bha07b, Bha07a, Por04b]. $59.99 [Cor07, Jaf06, Jen05a, Koh06, Par07, Put07, Sta06a]. $64.95 [Ber03c, Paq06]. $69.95 [Ekt05]. $69.96 [Lla07]. $74.99 [Mil08]. $75.00 [Cha05b, Wol08]. $79.00 [Gum04a]. $79.95 [Par06]. $89.00 [Men05]. $89.95 [Las06]. $99.00 [Hen08]. < [SM05b]. = [Ano03f, WHT^{+}05]. > [SM05b]. n [Har03b].

-Spectre [BCJ^{+}18].

.DOC [Sta05c]. .NET [Cha05b, Jen05a, Nau08b, Sta06a, Tit05, Mac08, Rob04, Zhu04]. .ODF [Sta05c].

0 [And04, Bal03, Bel05b, Ber03b, Ber05, Cha05a, Cha05b, Che03, Edw05, Ekt05, Fox08, Gro04o, Gum04b, Hac06, Hen08, Hid04, Hil08, Hin08, Hor03, Imr03, Jen05a, Jov06, Joy04, Kar05, Kau03, Kil04, Koh04, Koh06, Las06, Lla07, Llo06, Mar04b, Mar05a, Mar05b, Men04, Mor06, Par06, Par07, Pos04, Puz04, Ram08, Rao04b, Sch04b, Spi07, Sta05k, Sta06l, Sta07u, Sta06k, Sta06n, Sta07v, Ste03, Tho05, Tom06, Wal07, ZM09]. 0-07-222742-7 [Gum04b]. 0-07-225709-1 [Sta05k]. 0-07-226299-0 [Sta07u]. 0-13-100152-3 [Koh04]. 0-13-101415-3 [Hid04]. 0-13-117705-2 [Ber05]. 0-13-144853-6 [Tho05]. 0-13-146716-6 [Edw05]. 0-13-147941-5 [Mor06]. 0-13-167984-8 [Sta07v]. 0-13-224206-0 [Ram08]. 0-13-234971-X [Fox08]. 0-13-235088-2 [ZM09]. 0-13-597444-5 [Ber03b]. 0-13-713012-0 [Hil08]. 0-201-74868-1 [Pos04]. 0-201-78695-8 [Mar05a]. 0-262-19487-2 [Rao04b]. 0-262-69276-7 [Bal03]. 0-306-48189-8 [Gro04o]. 0-321-15078-3 [Hor03]. 0-321-18612-5 [Men04]. 0-321-20217-1 [Puz04]. 0-321-22409-4 [Cha05a]. 0-321-24362-5 [Par07]. 0-321-26838-5 [Kar05]. 0-321-28608-1 [Spi07]. 0-321-29353-3 [Jov06]. 0-321-33643-7 [Tom06]. 0-321-34995-4 [Llo06]. 0-321-34996-2 [Sta06l]. 0-321-35670-5 [Sta06k]. 0-321-37577-7 [Hac06]. 0-321-38555-1 [Wal07]. 0-321-48079-1 [Hin08]. 0-385-49532-3 [Imr03]. 0-385-50860-3 [Kau03]. 0-387-68766-1 [Hen08]. 0-470-84402-7 [And04]. 0-471-64446-3 [Las06]. 0-521-52583-7 [Cha05b]. 0-521-52749-X [Kil04]. 0-596-00290-4 [Che03]. 0-596-00462-1 [Ste03]. 0-596-00482-6 [Sch04b].

0-596-00545-8 [Mar04b]. 0-596-00641-1 [Joy04]. 0-596-00868-6 [Sta06n]. 0-7645-5746-7 [Bel05b]. 0-7821-4428-4 [Koh06]. 0-8493-2800-4 [Par06].

1 [Bal07a, Bal07b, Ber09, Ber03c, Ber07, Bha07a, Cor07, Ekt05, Eng05, Gib07, Git06, Goo07, Gum04a, Hen08, Hin08, Jaf06, Jen05a, Jen05b, Joy04, Kha05, Kuc06, Lec07, Mar05b, Men05, Mer04, Nau08b, Pet04, Por04b, Pos04, Put08, Puz04, Spi07, Sta06a, Sta05k, Sta05j, Sta06m, Ste03, Tom05, Wal07, Whi08, Wol06, ZM09]. 1-4020-3704-X [Bha07a]. 1-4020-7528-6 [Pet04]. 1-55558-318-0 [Mar05b]. 1-55860-755-2 [Ber03c]. 1-55860-866-4 [Ber07]. 1-580-53369-8 [Gum04a]. 1-58053-749-9 [Men05]. 1-58705-115-X [Sta05j]. 1-590-59096-1 [Mer04]. 1-59059-222-0 [Jen05a]. 1-59059-438-X [Jaf06]. 1-59059-452-5 [Sta06a]. 1-59059-503-3 [Kuc06]. 1-59059-580-7 [Bal07b]. 1-59059-582-3 [Git06]. 1-59059-596-3 [Bal07a]. 1-59059-631-5 [Cor07]. 1-59059-753-2 [Gib07]. 1-59059-755-9 [Lec07]. 1-59059-779-6 [Put08]. 1-59059-909-8 [ZM09]. 1-59059-955-1 [Nau08b]. 1-59102-415-3 [Goo07]. 1-59327-036-4 [Jen05b]. 1-59749-030-X [Sta06m]. 1-84628-344-2 [Whi08]. 1-84821-003-5 [Ber09]. 1-85233-672-2 [Por04b]. 1-85233-785-0 [Ekt05]. 1-932266-47-X [Tom05]. 1-932394-20-6 [Eng05]. 1-932394-24-9 [Kha05]. 1-932394-35-4 [Wol06]. 1.1 [Tit05]. 101 [GK22]. 1st [Ber09, Hil08, Mil08].

2 [Bel08, Ber03c, Ber05, Bha07b, Gib07, Mau05f, Por04b, Put07, Rao04b, Sta06l, Whi08, Wol08, ZM09]. 2.0 [Dwy08, JSK16, Kam15a, Kam15b, PRM09, Sta05i]. 2005 [BW07, Lho06]. 2008 [Mac08, Nau08a]. 2020s [GS23]. 21st [Wil04]. 24-year-old [Cra20]. 2nd [Nau08b, Sta07u, Zen04].

3 [Bal07a, Bal09, Bha07b, Git06, Goo07, Gro04p, Hid04, Hor03, Imr03, Jov06, Kau03, Koh04, Kuc06, Las06, Lla07, Nag04, Paq06, Put07, Zen04]. 3-540-00896-9 [Gro04p]. 3-540-25995-3 [Paq06]. 3-540-28713-2 [Bha07b]. 3-540-32180-2 [Put07]. 3-540-40704-9 [Zen04]. 3-540-43330-9 [Nag04]. 3-540-44492-0 [Lla07]. 3.5 [Mac08, Nau08b]. 3rd [Wol08].

4 [Ber07, Cha05a, Che03, Jen05b, Koh06, Llo06, Par06, Sta04a, Wol06]. 4th [Gum04b, Sch04b].

5 [Ber09, Ber03b, Cor07, Kar05, Men04, Mor06, Nau08a, Par07, Sta06a, Sta06k, Zan04]. 5.2 [MNN04].

6 [Edw05, Eng05, Pet04, Put08, Sch04b, Sta06n, Tho05]. 64-bit [Sta06f].

7 [And04, Bal07b, Bal03, Bel05b, Cha05b, Gum04b, Hac06, Mil08, Tom06]. 7th [Par07].

8 [Gro04o, Gum04a, Mar04b, Mar05a, Sta07v, ZM09].

9 [Gro04p, Kha05, Lec07, Men05, Nag04, Zen04]. 978 [Bal09, Bel08, Mil08, Nau08a, Wol08]. 978-0-321-44561-2 [Bel08]. 978-0-521-86697-2 [Wol08]. 978-1-4302-0973-7 [Mil08]. 978-1-59059-965-5 [Nau08a]. 978-1-59059-985-3 [Bal09].

Être [KB09a].

Aaron [Sch04b]. ability [Lev23]. Able [O’H06g]. Above [Coo19]. absolute [Kam23]. Abstracting [AT17]. abstraction [Gro05e]. Absurdity [Ano21]. Academic [RB17, NC17, Pet04]. Academies [NN19a]. accelerated [DGK^{+}23]. acceptable [O’H05d]. Access [EMC19, FKB^{+}03, Sam14, Vac23a, Wan24, Wat13]. Access-control [Wat13]. Accessibility [Sta06f]. Accessorizing [O’H06h]. accident [Kam21c]. Accountability [Dia15]. Accounts [Hil03]. Achieving [BB18]. ACID [Hel21a, Pri08]. ACM [SMM^{+}04, BS04, Cer14, Cre10c, Cre10e, Cre11a]. acmqueue [Mau09]. Acoustical [Sta05c]. across [DG15, Hel22a]. Action [CLNM03, FKN^{+}23, Wol06, WB05]. actionable [ESTT22]. Active [Tom05]. activism [Mau05f]. Acts [NN11e]. Actually [NSFG23]. Ad [Ber09, Swe13, Lab08]. Ada [Wol08, BW07, Wol04b]. Adam [AHC22, McK03]. Adamatzky [Rao04b]. Adaptation [EK12]. Adddison [Sta06l]. Adddison-Wesley [Sta06l]. Addison [Bel08, Cha05a, Edw05, Hac06, Hin08, Hor03, Jov06, Kar05, Llo06, Mar05a, Men04, Par07, Pos04, Spi07, Sta06k, Tom06]. Addison-Wesley [Bel08, Cha05a, Edw05, Hac06, Hin08, Hor03, Jov06, Kar05, Llo06, Mar05a, Men04, Par07, Pos04, Spi07, Sta06k, Tom06]. Additive [Fra20e]. Address [Kam22b]. Addressing [LNV14]. Administration [HKM10, Lea11]. Administrators [Edw05, HN05]. Adopt [Sta05h]. Adopting [VD24, GTK24]. Advances [GT03, Hen08, Nag04, SMM^{+}04, Wol04c, Tel23]. advantages [Coo23]. Advice [NN09b, NN07c]. Affair [Gro04h]. Affine [KB08a]. Afterthought [SL04]. Again [Cra20, O’D09, NN04c]. Against [Mer04, SR03]. Age [Fer03, Fra18, Wan24]. Agenda [Kel23a]. Agile [Ber03b, CG10, Coh06, Hor03, JSS16, Mar03, Mor06, PP03b, Spi07, ZM09, ATM22, Kru07, M^{+}09, Tat06, Var21]. Agility [Men04, BT03]. Agnew [AHC22]. ahead [Gor07]. Ahmed [Pet04]. AI [BC06, BFZ^{+}22, BS04, GHM15, GPPV24, Lev23, Nar04, SC21]. AI-powered [BFZ^{+}22]. Aikido [BNNC03]. Ailing [NN22c]. Ajax [Bal07b, Git06, Ram08, JWC08, Nor09, Sch09, AS06c, JF06]. Akeley [Sta08g]. al [Sta06a]. Alan [Wol08, Fel04a]. Alexander [BS04, SMM^{+}04]. Alexei [Ram08]. Alfred [SNW^{+}23]. algorithm [WWG22]. Algorithmic [Dia15]. Algorithms [Ano04n, Kel20d, LHML21, NN18c, Pet18, Sha04a, SNW^{+}23, Gro04o, VT03]. Alike [KB11]. Alison [Gum04a]. All-Optical [BP09]. All-Seeing [Tox14]. Allen [Sta05k]. Allman [Sta08h, Sta08i]. Allocation [Kel22b]. Alloneword [KB07b]. Already [CLNM03, GTNNM21]. Also [NN13e, DLdFV^{+}23, Kam21a]. Alternative [Pri08, SD23]. alternatives [SD23]. Always [Ano20a]. Always-on [Ano20a]. am [Gro05e]. Amazon [Kil20]. Ambler [Jov06]. AMD [Kap23]. amidst [Bel06]. Amine [Pet04]. among [Vid21]. Amplification [Hel19c]. Analog [MB04]. Analysis [CT21, DB18, Jov20, Nad21, NG22, Tho21, TDA21, WRS^{+}09, ESTT22]. Analytics [KNR13, Hel20a]. Anatomy [Cor12]. Ancient [Imr03]. Anderson [Gum04a, Mau05c]. Andre [Ram08]. Andreas [Ber07]. Andrei [Puz04]. Andrew [Puz04, Rao04b]. Andy [Las06, Wol08]. Angela [Tom05]. Anirban [Lla07]. Anja [Paq06]. announces [Ano03d]. Anomaly [Gra19]. Anonymity [DRW^{+}14]. ANSI [Mar04a]. answer [Wad05]. Anti [Mau05f, Sta06c, Sma06, Sta07s]. Anti-P2P [Sta07s]. anti-patterns [Sma06]. Anti-Spam [Sta06c, Mau05f]. antipatterns [Lap04a]. Anton [Mar04b]. Any [Coa04a, Hel18b, See04, Kam22b]. Any Body [Ano04m]. Anybody [Mat17d]. Anymore [Rie03, Lap06a]. Anyone [Mac03a]. Anything [KB05a]. Apache [Ano03d]. API [Mil08, Hen07, Lim19a, SW14, Tul08]. APIs [NN07d]. APL [WHT^{+}05]. appetite [NN07d]. Apple [Bla14]. Application [Bal09, CL11, CB23, Dwy08, NH20, Ram08, SGN04, Sta07u]. Applications [Ala17, And04, Bro15, GF08, Hac06, Kle14, Sha19, vD15, Nag04, Ber06, Fin03, GT03, JWC08, NN07f, SSS06, Sta05a]. Apply [KB09b]. Approach [Nad21, Sta07v, Hal07, NSFG23]. approaches [Tel23]. Apps [Mei14c, Wol04g]. APress [Git06, Mer04, Bal07a, Bal07b, Bal09, Cor07, Gib07, Jaf06, Jen05a, Kuc06, Lec07, Mil08, Nau08a, Nau08b, Put08, Sta06a, ZM09]. April [Par06, Lim17b]. Arbitrary [CP14]. Architect [Ano11, Mil08, Tul08, Bel07b]. Architecting [Hac06, Ber06]. Architectural [Mon10]. Architecture [Bha07b, Cha05b, KR03, Mat16a, Put07, GDD06, Gor06, Zhu04]. Architectures [Bha07a, Gro04o, Par07, Sha04a, Sta05j, Bir05, Con04, Mar08, MSJL06, VT03, VB07]. Archives [BG11]. aren’t [Hel21b]. Arm [GTK24]. arms [GC05]. arrays [MAB^{+}20]. Arrival [LVDM22]. Art [BNNC03, Dec04, GL11, Mat16b]. Artech [Gum04a, Men05]. Arthorne [Kar05]. Arthur [Can09]. Ashish [Bha07a, Nag04]. ASICs [Mat21]. Asleson [Git06]. aspiring [NN06h]. ASPs [Tak06]. assertions [NN21c]. assistant [Bor06]. Assisted [MB04]. Associates [Joy04, Sta06n, Ste03]. Assurance [Hac06, Ber06, Fel05b, Ker24]. Assured [Lev20]. Astronomy [BG11]. Asymmetric [FSSP09]. Asynchronous [Don03]. Atefeh [SBS^{+}23]. ATM [Ano03f]. Attached [Mat19c]. Attack [NN17f, Sch05, Sta06g]. attacker [SP21]. Attackers [SP21, Sta06b]. Attacks [Ano04k, GAM18]. Attestation [GTK24]. Auctions [Kel20d]. Audio [Han09b, BLA21]. Auerbach [Par06].

Aurelius [Hin08]. Aussie [Ano03d]. Australia [Ano03g]. authenticate [DLdFV^{+}23]. Authentication [Jov20, Sch04a, YD16, All06b]. Authoritative [CP14]. authority [Vac23a]. Automated [CDH14, Don05, NN15b, NN23d]. Automatically [AR23]. Automating [Mur04]. Automation [Lim15a]. Autonomous [Hel22a]. Autonomy [Mat15b]. Autotools [NN11c]. Availability [SDRB17]. available [Kop05]. Aversion [NN21a]. Avoid [Lim18c, Lim23c]. Avoiding [NN10b]. Away [AT17, Fra20b, Sta04a, Sta07q]. Azure [JVG^{+}24].

B [Eng05]. B.Y.O.C. [Kam11a]. Babbin [Tom05]. Babble [Hel16b]. Back [Ano04k, Col19a, Hay04, Leh03]. Backend [WWG22]. Bad [Coa04b, HD22, Mat16a, Ele07, Sim04]. bag [Lap04a]. Bail [Bel03]. balances [Mat23b]. Balancing [BT03, Lim16b, Men04]. Baleen [Hel20a]. Bang [Boe03]. Bank [AhCD22]. barn [KB07c]. Baron [SBS^{+}23]. barrier [Rix08]. Barriers [Har15]. Barry [Men04]. Base [RR18, Pri08]. Baseboard [Fra19b]. Based [CCG^{+}16, BLK15, Gro04p, Ish04, Jov20, Kle14, VTA23, VD24]. Bases [NN03a]. basics [Spa07]. Batches [Lim16c]. Battery [Sta04b, Fra20a]. Battery-powered [Sta04b]. Battle [NN12d, Smi09]. Battlefield [Ham14]. battles [NN05a]. Bazel [Lam18]. BBR [CCG^{+}16]. BCS [Cre10e]. Be [Ano12b, Dur10, Fra20d, Hel22b, KW04, Lim15a, Mat17a, NN22f, Ros10, CPG^{+}04, KB07d, Lim23a, Lim23b, MBNN^{+}23, NN09c, NN21e, NN23d, Vac23b, VD24]. Beam [Ste03]. Bear [NN23f]. beats [KB08c]. Beautiful [NN08b]. Becky [Tom05]. Becomes [LHML21, Wol03a]. Beginning [SB06, Bal07a]. behavior [KGM21, O’H05d]. Behind [Pet18]. Being [Mat16d, KB08b, Mat23b, MR21]. Bejtlich [Sta06l]. Bell [Lim20c]. Below [Coo19]. Ben [SNW^{+}23]. Benchmarking [Sit18]. benefits [Fra21a, Mat21]. Beowulf [PBK07]. Berg [Hac06]. Berkeley [SOM21]. Berlin [Gib07]. Best [CFL23, Mie06, NH20, Par07, Ano21, VB07]. Betrayed [Phi10a]. Better [Ano04h, Car07, Cre06b, KA23, Mat17a, WKGD08, DSBS23, KB05a]. Between [Sha03, LHML21, Pop24]. Beyond [CFL23, Hel16a, Lev09, PBK07, Rao04a, Ree19, Sel05, TB03, Vog08a, vD15, Spa07]. Bias [Kel24, LPS24]. Biases [SC21]. Bible} [Bel05b, CPM04]. Big [Ano04h, Boe03, CBB07, Coa03d, DRW^{+}14, Dat16, HWH17, Jac09, Kam21b, KNR13, NN14e, Sta04b, Gro04i, KB07c, Lim23a, NN07e, Sto05, GAI^{+}21]. Big-data [KNR13]. Bigot [Coa04a, CPG^{+}04]. Bigots [GBG^{+}04]. Bikeshed [Kam14a]. Bill [Goo07, Sta07l]. Biological [DFZ^{+}09]. Biology [KM09]. Biometrics [Sta05b, Sta06d]. Biomolecular [PS09]. Biomorphic [Lod04]. bisection [NN21b]. bit [Sta06f]. Bitcoin [NC17, SZ17]. Bits [Ros10, Mas06]. Black [NN22f, Pos04, Sta04b, WT03, KB07e]. Blackburn [Par07]. Blakeley [Coa08]. Blast [Ano03e]. Blast-Off [Ano03e]. Blaster [Mor04]. Blessing [Kot10]. Bliss [WNN03]. Blockchain [Kuz19, Wal18]. Blockchains [NMHB16]. Blow [Sta05c]. blues [KB06b]. Bluetooth [Ano03d]. Blurring [Sha03]. Board [Kam22a, Kam21c]. Bob [Che03]. Boba [Ano04i]. Bodden [Bal07a]. bodies [O’H06g]. Boehm [Men04]. Bones [Ano04n]. Bonwick [Sta07l]. Book [And04, Ano04h, Bal09, Bal03, Ber03b, Ber03c, Che03, Gum04a, Gum04b, Hid04, Hor03, Imr03, Kau03, Kil04, Mar04b, Men04, Mer04, Mil08, Nag04, Pet04, Pos04, Rao04b, Sch04b, Ste03, Wol08, ZM09, Zen04, Mar04a, Sin99]. bookmarking [MFK05]. Books [Goo07, KB11, Sch04b]. Boot [Fra19c]. Bootle [SMM^{+}04]. bootloader [Vac21]. Borg [BGO^{+}16]. Bosworth [McK03]. Botnet [Hen08, LWD07]. bots [KB07a]. Bound [NN10f]. boundaries [Hel22a, Ses04]. Bourne [Sta08h, Sta08i]. Boutell [Koh06]. bowl [MGL22]. Box [Bos06, NN22f, WT03]. BPM [Mie06, Ver04]. Bragging [NN13b]. brain [BC06]. Brandt [SBS^{+}23]. Bray [Gra05]. Breadth [Mat17b]. Break [Ano12b, Mar05a, HM04]. Breaking [Poo06, Rix08]. Breidenbach [Wol06]. Brewster [Fel04b]. Brian [Bal07a]. Bridging [Rus08]. bring [BLA21]. Bringing [CP14]. brings [DGK^{+}23]. brink [Lim23a]. Broadcast [Jan03]. broader [NN21d, SCQT20]. Broken [Kro17, NN10c, NN20a]. Brother [Kam21b]. Brown [O’H06p]. brownfield [HJ08, Hil08]. Browser [RBP09, WG09]. Bruce [Che03, Bou04]. Bryan [Sta08h, Sta08i]. buckets [NN22d]. Buffer [Bat04b, CKG^{+}04, CPG^{+}04]. Bufferbloat [Get12, CJWG11]. Buffers [EK12]. Bug [Coa03b, Lim18b, Ano03g, Ull03, Kau03]. Bugs [NN12c, NN13b, NN06d]. Build [KNR13, Kop05]. Building [Bal07b, BKMR21, CC04, CdSH^{+}03, Hin08, KW04, Kil08a, Luc05, Mat15a, NN07f, Put08, Ram08, Sta06k, Tho05, HP08, JF06, JWC08, McG06, SP21, SS07]. Builds [NN10c]. Built [Coa06, Sha09, Fra21a, VCR15]. bulk extractor [GS23]. bullets [Bel06]. Bunch [Rie03]. burning [Lap04a]. Burns [Wol08]. Business [Cor07, Cre10e, Cre10f, Goo07, SMM^{+}04, Lho06, PZ06, Vac23b]. Bust [Sta04a]. Butler [Sta07k]. Buy [Coa03d]. Byron [Gum04a]. Byte [Fra20d].

C [Ber03b, Cha05a, Chi18, God19, Hel21a, KP23]. C# [Nau08a, Nau08b, Mac08]. Cache [KS08, Lam18, Lov17, NN17a]. Caching [TR03]. CAD [Fra21b]. Calculus [SDRB17, KB06c]. Caleb [Sta07u]. Calendar [Ano03a, Ano03b, Ano03c, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano09, Gro04a, Gro04b, Gro04c, Gro05a, Gro05b, Gro05c, Gro05d, Mau05a, Mau05b, O’H05a, O’H05b, O’H07a, O’H06a, O’H06b, O’H06c, O’H06d, O’H06e, O’H06f, Sta07b, Sta07c, Sta07d, Sta07e, Sta07f, Sta07g, Sta07h, Sta08a, Sta08b, Sta08c, Sta08d, Sta08e, Sta09a]. Calipers [NN16c]. Call [KB05c, GC05]. called [Cre22]. Calling [Gro04e]. Cambridge [Cha05b, Kil04, Wol08]. Campaign [Boe09]. Can [Ano03h, Lov17, NN12c, Ros10, SMM^{+}04, Wea03, ESTT22, KB05a, KB06d, Lim23b, NN23d, SP21, VD24]. Canary [DB18]. Cannon [Cha05a]. Cantrill [Sta08h, Sta08i]. capability [Var21]. capture [JC23]. Cardboard [Ano04h, Hel17b]. Cards [And04, SJK20, Fin03]. Care [Can10, EMC19, JNSM14, NN22d, Cre06b]. Careers [Gro04d]. Cargo [Lev23]. Carpal [Sta05i]. Carrier [NFG04]. Carrying [Sta06c]. Cars [Gro04e]. Case [Asc04, EK12, cF03, MQ09, Mer04, Ano21, FoDF22, SR03, JSK16]. Cases [JC23, BLA21]. Cast [Lev20]. Catch [Ano20a, KP23, NN22c]. Catch-23 [KP23]. Catmull [Cre10a]. Cautionary [EGBA03]. Cavernous [NN03a]. CCA [GTK24]. celebrate [Lim20c]. Cellphone [Ano04m]. cellular [Jov20]. Center [AF12, Hel17a, Kam13a, Put08, SS07, Woo10]. Centered [Nad21]. Centers [Alo18, Sch17b]. Central [AhCD22]. centric [NSFG23]. cents [FT07]. Century [Wil04]. certain [Mat21]. Certificate [Lau14]. chain [KMNNM22]. Chakrabarti [Lla07]. Challenge [Can10, May03, Sta06j, Ver04, XMGM21, Tak06, Tho21]. Challenges [Bha07a, Bor03, GLF^{+}15, HD22, NGJ^{+}19, VD24, Jov20, Mat21, MSJL06]. Chance [NN18a, Vac21]. Chances [SD23]. Chang [O’H06t]. Change [Hel21a, Med16]. changed [Kri08]. Changes [Gra08a, Hel15, Hel21c, Lim23c]. chaos [Noy05]. characteristics [Hel20b]. Charland [Ram08]. ChatGPT [LPS24]. Cheap [Kam21b]. checks [Mat23b]. Cherry [NN13c]. Cherry-picking [NN13c]. Chess [NN16a]. Chilling [NN16b]. China [Ano03e]. Chip [Fra21a, Sta06g]. Chipping [Fra20b]. Chips [Ano03h, Bor03, Sta05e, Fra21a]. Chisnall [Fox08]. Choosing [Mat18e]. chosen [SD23]. Chris [Kar05, Sta05k, All03c, Ano20a, SNW^{+}23]. Christopher [Sta07v]. Chrome [RBP09]. Chuck [O’H06o]. churn [WWG22]. Chuvakin [Mar04b]. Circuits [MB04, RCN03]. Cisco [Sta05j]. Clark [Bel05b, Tom05]. Classes [Bat04c, GBG^{+}04]. Classroom [Sta05c]. Clean [M^{+}09, ZM09]. Clearing [Kel20d]. Clearly [Mil10]. ClickShift [O’H06t]. Client [Mat21]. Cliff [Hen08]. Clifford [Hac06]. Clip [Han09b]. Clocks [BP16]. Closed [Leh03, For07]. closer [FH08, Fra20a]. Cloud [Dat16, Dur10, FAHB18, Kap23, LSB17, NN16c, NN18b, Owe10, RCF^{+}21, Rus23, VOT^{+}09, Hel22b, Cre09a, PRM09]. Clouds [SCQT20]. Cluster [CP18, Jen05b, Sch17b, Sin16, Wol04d, Kop05, Kop05]. Cluster-level [Sin16]. Clustered [Tho05, Luc05]. clusters [GSI07, PBK07]. CNNs [FoDF22]. Coatta [GDWC20, Mau05e]. Cocoa [Ste03, BD03]. Code [Ano03d, Ber05, CGB^{+}18, DVR10, DLdFV^{+}23, EKG^{+}05, Gil14, Imr03, KGM21, Lim19b, NN03a, NN08h, NN08b, NN11a, NN12c, NN12a, NN15a, NN15c, NN18d, PP03a, Spi03, SSN14, ZM09, Mar05a, DGK^{+}23, Fea05, Gro04j, HM04, KB07f, KB07d, M^{+}09, NN06h, NN21a, Ras05, Sin99, SK05, Var21, VTA23, Vid21, WG06]. code-based [VTA23]. code-dependency [NN21a]. Code-generating [SSN14]. CodeFlow [CGB^{+}18]. Coding [EGBA03, GL11, See03, SK05]. Coffee [NN20a]. Cognitive [Gra19]. Cohen [O’H06s]. cohesive [Mat24]. Cohn [Mor06]. Cold [NN17a]. Cole [AHC22]. Collaboration [CdSH^{+}03, HKM10, Paq06, Ebe05]. Collaborative [BB03]. Collapse [AAVV14]. Collecting [NN10d]. Collection [DEE^{+}16, DLP19, Bac07]. Collie [Gum04a]. COM [Ols06]. combination [DSBS23, Mat23a]. Combinatorial [Kel20d]. Combining [NG22, Tel23]. Come [Mat18b, Sta05i, Kam11c]. Comes [Ano04l, Fra20c, NA22]. Coming [KM09, Sta05c, Wol04d, VTA23]. Comma [Ras04a]. Commentary [Chr09]. Comments [Kam22a, Ras05, WHT^{+}05]. Commerce [Mei14c, Sta07a]. commercial [Cre22]. Commercializing [Kar03]. Commitment [NN10e]. commodity [Kop05, O’H07e]. Common [BNB^{+}05, Lim23c, Mar04a, Ols06]. Communicate [Lim20a]. Communication [BLA21, KW17, Lim23a]. Communications [Ano04j, DL09, Ekt05, Ste07a]. Communities [DMBB17]. Community [XMGM21]. Compact [Wol03c]. Company [NN22e]. Compare [Sta04b]. Comparison [Ull21]. competing [GPPV24]. Competitiveness [Cre10e]. Compilers [God19]. complete [Vac23a]. Complex [MH21]. Complexity [Mon10, Tho21, Vix07]. Compliance [CB06, All06a, Nol06, Rag06]. Complying [All06a]. Component [Bal07b, Las06, WQ05]. Component-Oriented [Las06, WQ05]. Componentizing [Sav15]. components [JF06, Ses04]. Composition [Ano04m].

Compromises [AHMT03]. Computation [SBS^{+}23]. computations [MBNN^{+}23]. Compute [CP14, Mat21, Vid17, Hel21c, Hel22a]. Computer [Ano04h, RB17, BS04, Ekt05, Gum04a, MAC^{+}03b, Nar04, Vid21, Can06a, Coa06, NN23d]. Computers [Can10, Kel23b, NN19f]. Computing [BP09, BR10, Cre09a, Dur10, Gra08b, Hel22a, Kel20a, Koh04, Lla07, Mor18, Nag04, Pop24, RCF^{+}21, Rus23, VOT^{+}09, YRZB16, Zah16, Zen04, Rao04b, Zen04, Boy08, Cha07, DLdFV^{+}23, DGK^{+}23, Fel05a, GTK24, GT03, Gof04, GPPV24, Hir04, JVG^{+}24, KA23, Kap23, SARC03, Str05]. Con [Hel21b]. Concepts [Bha07a, MSJL06]. concise [JC23]. Concurrency [CB08, ATKS07, Gat04, SL05]. Concurrent [BW07, Lar08, Wol08]. Confessions [Mil08, Tul08]. Confidential [DLdFV^{+}23, DGK^{+}23, Pop24, RCF^{+}21, Rus23, GTK24, GPPV24, JVG^{+}24, Kap23, GPPV24, JVG^{+}24]. confidentiality [DGK^{+}23]. Conflict [Mat16e]. confluence [Hel21b]. Congestion [CCG^{+}16]. Congestion-Based [CCG^{+}16]. Connect [Coo23]. Connected [Vas17]. connection [WWG22]. Connectivity [Ber03a]. Conquer [NN21b]. Conrad [Rao04b]. Consensus [BFAP16]. considerations [SCQT20]. Considered [All03b, Ros11, BT06, Gil04]. Consistency [AA22, Hel21b, LFKA14, PCA^{+}15]. Consistent [GRA^{+}14, Vog08b]. Consistently [Hel18a]. consolidation [Vog08a]. Consortium [Sta06g]. Constrained [Sta06c]. Construction [RR18]. consumer [Cre07b]. Contactless [And04, Fin03]. Container [JVG^{+}24]. Containers [Kro17, LSB17, Sin16]. Contest [Fra21a]. Context [Ekt05, Vid21, Lon04]. Continuing [Bor03]. Continuous [Hum17, Ker24]. Contract [KF19, SW14]. Contracts [NMHB16]. contrasted [Coo23]. contrivances [Kel21b]. Control [Ano04j, CCG^{+}16, PSBG11, Bel07b, Wat13]. Controller [Fra19b]. Controlling [ERR18]. Controls [EMC19, Sam14]. Conventional [Kel19]. Convergence [KA22, Hel21b]. Conversation [All03c, All03d, All04, Ano10, Bro04a, Bro03, Bro10, Can09, Cre10a, DiB04, Dit03, Fel03, Fel04b, Han09a, Han09b, Har03a, Kil20, Kir04, Pat03, Rus04, San04, SOM21, Sta08i, Sta09b, Was03, Bou04, Coa08, Fel04a, Gra05, Ham05, Mau05c, Mau05d, Mau05e, McK05, NN04a, O’H05f, O’H05g, O’H06o, O’H06p, O’H06q, O’H07c, O’H06r, O’H06s, O’H06t, O’H06u, O’H06v, Sta07j, Sta07i, Sta07k, Sta07m, Sta07l, Sta07o, Sta07n, Sta08f, Sta08g, Sta08h]. Conversational [Tel23, McT04]. Conversations [Mat17c, Lim20b]. Converse [Bel05b]. Convery [Sta05j]. Cooked [Sta05c]. Cool [Ano03h, CPG^{+}04]. Cooling [Woo10]. Coordination [Mag19]. Copilot [BFZ^{+}22]. Coping [Bro04b]. CORBA [Hen06]. Core [Hub04, Wol04g]. Corp [FAHB18]. Corporate [Fer03, Kil04, Jen02]. correctly [Ull21]. Cory [Sta07i]. Cost [SD23, Sta05f, Dre08a, Suv05]. Costs [Kam14a, Mag19]. could [BLA21, Hel21a, Lim23a, WWG22]. Couldn’t [NN16a]. Count [FPWM09]. Countering [Hen08, LWD07]. Counters [NN10d]. Counting [Hil03, Kam11a]. Coupling [NN18d]. course [Wad05]. Covid [BY22]. CPT [ESTT22]. CPU [DKM^{+}12]. CPUs [Cre05, Mac03a]. Crackberries [Sta05f]. Craft [JSS16]. craftsmanship [M^{+}09, ZM09]. Craig [Wol06]. Crash [AA22, PCA^{+}15]. crashes [AA22]. Crashproofing [Kel21a]. Create [Mat19b]. Creating [DGK^{+}23, Var21, Mat24]. Creation [LHML21]. creators [LHML21]. Crime [Ano04m, Sta05g, Ano03f]. Criminal [WG06, O’H06l]. Critical [Bla04, ESTT22, Pos04, WA19]. Cron [DG15]. Cross [Ano12a]. Crossing [SMM^{+}04]. Crowdworkers [KLL23]. Crushing [Sta05c]. Cryptocurrencies [NMHB16]. Cryptographic [Pop24, RCF^{+}21]. Cryptography} [Imr03, Cre22, Pop24, Sin99]. CS&IT [Lim17a]. CSO [NN19c]. CSRB [Kam22a]. CTO [Bou08, Cre08, Cre09b, Cre10b, Cre10c, Cre10d, Cre11a, Cre11b, Kil20, VOT^{+}09]. Cubes [Ano03h]. CUDA [NBGS08]. Cullen [Sta07j]. Cult [Lev23]. Culture [Kro17, Mat19b, OO03]. Curation [LHML21]. Cure [Sta05d]. Curmudgeon [All03a, All03b, Bat04a, Bat05, Bel06, Bel07a, Bro04c, Coa04a, Coa04b, Coa05, Cre05, Cre06a, KB04a, KB04c, KB04b, KB04e, KB04d, KB05a, KB05c, KB05b, KB06c, KB06a, KB06b, KB07a, KB06d, KB07b, KB07c, KB07f, KB07d, KB07e, KB07g, KB08a, KB08b, KB08d, KB08c, KB08e, KB09c, KB09a, KB09b, KB10, Lap05, Lap06a, Lap06b, Nie03, Nol06, Pat05, Ras05, Sta07p, Suv05]. Curran [Sta06a]. Currencies [AhCD22]. Curricula [Lim17a]. Curried [Ano04j]. Curse [Kot10, Mei14a]. Customers [Lim19a]. CXL [SMM23]. Cyber [Ano03f, Kam22a]. Cyber-Crime [Ano03f]. Cybercrime [Cym06, PRM09]. cybercriminal [WG06]. Cybersecurity [SD23]. Cycles [NN05c]. Cyrus [Mar04b].

D [Tom05, Fra20e, Sta05b]. DAFS [Kle03]. Dagon [Hen08]. Dale [Edw05]. Damien [GDWC20]. DAML [KF19]. Damnéd [KB04a]. Dan [Dit03]. Dance [Wol04a]. Daniel [Gib07]. Danilo [Kha05]. Dargan [Men05]. Dark [NMCK20, PRM09]. darned [Sto05]. Dart [MMB15]. Data [AF12, AHC22, Alo18, ABW15, BG11, Boy08, CP14, Coa03c, Cor17, DRW^{+}14, Dat16, EMC19, FKB^{+}03, Fra20d, GAM18, GBM15, Hel20b, HSN13, Hyd09, Jac09, Lim19b, Mat15c, NN22d, Pet09, Sch17b, SKC^{+}04, Tom06, Woo10, AA22, BKMR21, DSBS23, DGK^{+}23, Flo05, Hal05, Hel20a, JR05, Kel21b, KNR13, McC05, MBNN^{+}23, RCF^{+}21, Suv05, Wan24, SM05b]. Data-Center [AF12]. Data-parallel [Boy08]. Database [AR23, Ano20a, BFAP16, GAM18, Jov06, Par06, AS06a, Kel21a, Wel05]. Databases [Jov06, Ost05, Ric14, AS06a, Sel05]. Datacenter [Col19b]. Datapath [Vac14]. David [Fox08, Ram08, Sta06a, Han09a, Han09b, Lim24, Mau05c, O’H06p, O’H07c]. Davidson [Ste03]. Day [Ano03h, Coa03b, Fra20a]. DB [DKM^{+}12, SOM21]. DBMS [AR23]. De-identify [ABW15]. Dead [BP09, NN20e]. deadline [Vac21]. deadly [Tox07]. Dear [NN23a]. Death [Bel04a, Gro04k]. Deathmatch [Sta04a]. Debates [RB17]. Debt [All12]. Debuggers [NN21c]. Debuggery [NN11e]. Debugging [AGB17, Ber07, BWB^{+}16, CC20, Don03, NN08a, O’D17, Pac11, Phi10b, Sch09, WT03, Gat04, NA22, Zel06]. Decade [RBSH14, Wat13]. decades [Fra21b]. decades-old [Fra21b]. Decentralized [Kel20a, BKMR21]. Decentralizing [BL07]. Deception [SP21]. Decision [Dia15, Var21]. Decision-making [Dia15]. Declarative [MR21]. deconstructed [CB06]. Deduplicating [Ros11]. Deep [DSBS23, NMHB16]. Deep-learning [DSBS23]. Default [GDWC20]. Defects [HPE03]. Defense [Cre10b]. deficiencies [Jov20]. Defined [Sta06i]. defining [Lev23]. Definitive [Chi08, Fox08]. Delacorte [Imr03]. Delegation [Mat16b]. deleted [Vac21]. Deliberate [Bur03]. Deliman [KTS^{+}04, NN04a]. Deliver [Lap03, SMM^{+}04]. Delivered [Ano04j]. delivering [Ger05]. Delivers [Sta05c]. Delivery [Hum17, Swe13]. delusion [KB07f]. Demise [LN04]. Demo [Lim19b]. Demonstration [Get12]. Demystifying [Ano20b, BM22]. Depend [LCW18, MR21]. Dependencies [Cox19, ERR18]. dependency [NN21a]. depends [Nad21]. Deploying [BS16, Tom05]. Depth [Mat17b]. Derivation [Mub18]. Descends [Sta05g]. Describing [FT05]. desert [Coa05]. Desfontaines [GDWC20]. Design [Ber03c, Cha05b, Chi23, CG10, Hac06, Kil04, Mat18a, Mil08, Par06, SSN14, Wie03, Jov06, AS06a, Ber06, Fra21b, Hen07, Ish04, Jen02, RCN03, TDA21, Tul08, Wel05, Zhu04]. Designers [Ras03b]. Designing [BB03, CP18, Ekt05, Lon04, Tel23, TDA21, TCR03, WWG22, Zel07, Chi23, Kil20, Wal07]. Desktop [Dec04, Sta04a]. Desktops [FAHB18]. Despite [Mat21]. details [NN21e]. Detecting [Sta06d]. Detection [Col19e, Hen08, Sta06l, Bej06, LWD07]. Deterministic [Hel22b, WWG22]. Devalue [Lim15b]. Devedzic [Put07]. Developer [Ano03d, BNNC03, FSM^{+}21, Ker24, Nad21, NN09c, NSFG23]. developer-centric [NSFG23]. Developers [Cha05a, Lap04b, Can05]. Development [Bal09, Ber03b, Blo04, CL11, Coa03a, DVR10, Dwy08, EGBA03, Fox08, Gro04f, Hil08, Hor03, Lec07, LNV14, Mat15a, Men05, NN16e, OO03, PP04, PP03b, Put07, Spi07, Tur03, Var21, VTA23, VD24, ATM22, Bel06, GDD06, HJ08, Mar03, Tat06, VW07]. DevEx [FKN^{+}23, NSFG23]. Device [Coo23]. Devices [Cre11a, Cre11b, FKB^{+}03, NN08a, Ros11, VBH03, Wol03c]. devising [Kol04]. DevOps [FK17, Lim18c, Lim20c, Sch17a, WFW^{+}19]. DHTML [BM13]. diagnosis [Bel05a]. diagnostics [CLK^{+}21]. Dialogue [Por04b, McT04]. Diary [NN23a]. DiBona [All03c]. Dickory [NN15b]. Didn’t [Mat16d]. Diederik [Pet04]. different [FT05, Hel20b]. Differential [GDWC20, DSBS23]. Difficult [NN13e, VD24]. Digging [GAI^{+}21]. Digital [Ano14c, AhCD22, BB18, Lon15, Mar05b, RCN03, Mau08b]. Digitally [MB04]. Digits [KB04a]. Dignity [KB11]. direct [NN23c]. dirty [NN05b]. Disabled [Sta06f]. Disambiguating [Ric14]. Disassembler [NN21e]. Discipline [Men04, BT03]. Discontinuum [Hel19b]. discovery [Ost05]. Discrimination [Swe13]. discuss [Kil20]. discussed [Hel22a]. discussion [AHC22, Ano20a, GDWC20, SBS^{+}23, SNW^{+}23]. Disk [Ano04j, CLNM03, GF08, Ele07]. Disks [And03, McK12, Rie03]. Dismantling [Har15]. Disruptive [Var21]. Dissection [Col19b]. Distance [Ano04n]. distilling [McC05]. distinction [LHML21]. Distributed [AGB17, BFAP16, BWB^{+}16, Bro15, CC20, Coa03a, ESTT22, Fel05a, Gra08b, Hel16a, Hel16d, KF19, Koh04, Mad15, McC15, Tur03, ZAB16, Gof04, Hel21c]. distributed-systems [Hel21c]. Diverse [FKB^{+}03]. Divide [NN21b, Rus08]. Divided [NN13d]. Dividends [Kil16]. Division [God03, NN13d]. Djurić [Put07]. DNS [Vix07, Vix09]. Do [FPWM09, Lap04b, Lim10, SBS^{+}23, KB05a, Lim23c, NN21d]. DOA [Bel07a]. Dobberpuhl [Dit03]. Doc [NN15b]. Doctor [NN05g]. Doctorow [Sta07i]. Document [Gar07a]. documentation [NN15b]. Documents [NR18]. Does [BLK15, LVDM22, Mat17d, Wol04a]. Doing [Kam10, Cob22]. Doku [KB05a]. Domain [Gil14, Vid21]. Domain-specific [Gil14]. domino [O’H06m]. Donald [San04]. Done [Mat18c]. Don’t [ATM22, And03, CB23, Goo07, GBG^{+}04, Hel21b, Kam23, LFKA14, Mat18c, NN09c, NN21d, NN22c, SF04, SCB09, Sta05g, FM05, NN24b, PZ06]. Don’ts [Lim10]. doomed [CB23]. Door [Wea03]. Dorothea [Hid04]. Doubleday [Kau03]. Doug [Mer04, Sta07j]. Douglas [O’H06q]. Down [JS18, Sta07a, Ano03i, Gar07b]. Dragan [Put07]. Dreams [Ano03e]. Drift [NN24a].

Drink [Sta06g]. drinks [Ano03i]. Drive [Ano04o, Ano04p, Cor12, Gro04q, Gro04r, Gro04s, Gro05g, Gro05h, Gro05i, Gro05j, Mau05g, Mau05h, Mau05i, O’H05j, O’H05h, O’H05i, O’H06y, O’H06w, O’H06x, O’H06z, O’H06-27, O’H06-28, O’H06-29, O’H06-30, O’H06-31, O’H07d, Sta07t]. Driven [CDH14, Nad21, Put07, GDD06]. Drives [Ano04j, NSFG23, Ele07]. DRM [Soh07]. Drupal [Lec07, VW07]. DSL [Gho11]. DSLs [SSN14]. DSP [FS04, Lyo04, SMM^{+}04, Sha04a]. DSPs [Hay04]. DST [Sta07q]. Duncan [Ste03]. dung [Lap04a]. During [Gra19]. Dust [Ano03e]. Duty [Wol04e]. Dwivedi [Llo06]. Dwyer [Bal09]. Dynamic [Gro04o, NG22, Pac11, Ric08, VT03, YD16]. Dynamics [MW03, Med16].

E-Commerce [Mei14c]. E-mail [All06b]. EA [CDH14]. Each [Goo07, Coo23, PZ06]. Eagle [Sta05k]. earlier [Tel23]. Early [BFZ^{+}22, KB04d, NN09a]. Easier [KNR13]. easily [JC23]. Easter [BTM22]. Easy [BP16, Coo23, Sta05i]. eat [NN23b]. Eating [Hil08, HJ08]. Ebersbach [Paq06]. Echoes [Vid23]. Eclipse [Kar05, Joy04, Rub06, AL04, Hol04, Wol03a]. Economics [All03a, Gra08b, RW06]. Ecosystems [Ker24, CFL23]. Ed [Ekt05, Ber09, Gum04b, Hil08, Nau08b, Sch04b, Zen04, Cre10a]. Edge [Cre10c, Cre10d, Fre03, Mor18, Sta05f, FoDF22]. edited [Gum04a, Pet04, Rao04b]. Edition [Gib07, GWE^{+}05, Cor07, Mil08, Par07, Sta07u, Wal07, Wol08, ZM09]. editors [Bro05, Fel05a, Har05a, Mau08a, O’H05c, O’H05d, O’H06h, O’H06g, O’H06l, O’H06m, O’H06i, O’H06j, O’H06k, AM03, AM04, Bou03, BNNC03, Coa03c, Gro03a, Gro04i, Gro04e, Gro04f, Gro04j, Gro04k, Gro04g, Gro04h, Gro04l, Gro05e, Gro05f, Kel03, Lyo04, Mac03a, Mau07, Mau08b, NN03b, O’H05e, O’H06n, O’H07b, Sta05a].

Education [Puz04]. EDVAC [Wan15]. Effect [NN17d, O’H06m]. Effective [NG22]. Effectively [Ber05, Fea05]. effectiveness [Lim23b]. Effects [Hel17a, MMB15, VW19]. Efficiency [DKK^{+}04, FSSP09, Lim20b]. Efficient [BR10, cF03, Kel20b, NG22, Sax10]. effort [Mat23a]. Eggs [BTM22]. Egypt [Imr03]. Eilert [Hid04]. Eisenhaendler [Hid04]. either [Hel21b]. Elasticity [Owe10]. Electronic [Sim04]. electronics [Fra20a]. Elephant [Hil08, LPS24, NN22b, FT05, HJ08]. Elevating [GTK24, Rus23]. Elision [Kle14]. Ellen [Kau03, Sch04b]. Embedded [God03, Hal07, JYWV03, NN03b, Pet04, Sta07v]. Embedding [Vac14]. Embrace [Sta06f]. Embracing [Ano03i, Cre07a, Hel22b, MK14]. Emergence [Gol03]. Emerging [Sta06f]. Emulating [Lim20b]. enabler [Vac23b]. enables [DLdFV^{+}23]. Enabling [Kap23]. Enclaves [SCQT20, Pop24]. encourages [Lim20c]. Encryption [Cre22, Kam13b, Kam16, Lim23a, NN15b]. End [KB04e, Lim23a, Kam24, O’D09]. End-to-End [Lim23a]. Energy [BR10, Fra20c, VBH03]. Energy-Efficient [BR10]. enforced [RCF^{+}21]. Engine [Pat04, Sta06h]. Engineer [Mat16d]. Engineering [Cha05b, JS14, JSS16, Kil19, LNV14, Mat23b, SP21, Tho21, Zhu04]. Engineers [Har16]. Enhanced [Phi10b]. Enlightenment [Fer03]. Enough [LV03]. ensure [Hel22a]. Enter [Vac23a]. Enterprise [Ano14a, BS16, Cre11a, Cre11b, Hac06, Jac05, Jen05b, Kop05, MSJL06, MK14, MW08, MM04, Sha19, Zen05, Ber06, Bra05, BM22, Fel05a, MFK05, O’H07e, TCR03, JWC08, Ric06, Str05, Bha07a, Ram08]. Enterprise-grade [Zen05]. Enterprising [Gro04h]. entire [Kam21b]. Entry [Har15]. Environment [Koh06, MB05]. environmental [Lap04a]. Environments [Pac11, Hel21c, Hel22b, SP21]. epidemic [Cym06]. equal [KB08b]. Era [Fra21b, Mau08a, RR18]. Eric [Sta08h, Sta08i]. Erik [Coa08, Mat17c]. Erlang [Lar08]. Errata [BS04, SMM^{+}04, SKC^{+}04]. Error [MK04, Bro04b]. Errors [Che10]. Essence [SJK20, Lev23]. Essential [Gor06, JC23, Ste07b, Bha07b, Whi08]. Essentials [SJK20]. Estimating [Mor06, Coh06]. Ethical [Sta05k, Har05b, KLL23]. Evaluating [TDA21]. Event [Kam22a, KBS19]. Events [CDH14]. Eventual [Hel18a, Hel21b, LFKA14]. Eventually [Ano04l, GRA^{+}14, Vog08b]. ever [BLA21, NN23e]. Every [Har03b, NN18b]. Everyone [CKG^{+}04, Vac20, NH20]. Everything [Cra20, Hel15]. evidence [DLdFV^{+}23]. Evolution [Bro15, Chi14a, Cre06a, GBM15, Mat19a, MQ09, Mei14b, Gee07, Kil20, Sin99]. Evolutionary [Jov06, Nag04, AS06a, GT03]. Examples [AGB17, Par07, VB07]. Excellence [Lim17b, Mat23b]. Exception [NN17b, NN22c]. Exception-Handling [NN17b]. Exchange [AHC22]. Excluded [Mei14a]. Excuse [Lim18c]. Executables [Kel22a]. Executive [JNSM14]. Executives [Lim19c]. Existing [Kle14]. Exists [NN08b]. expanded [BLA21]. Expecting [GRA^{+}14]. expensive [Kam21b]. experience [CT21]. Experiences [SJK20]. Experimenting [Wan24]. Expert [Jen05a, Rob04, Lho06, Cor07]. explain [JC23]. exploit [SP21]. exploitation [Gar07a]. Exploiting [HM04, Mar05a]. exploits [Jov20]. Exploration [Gra19, SSN14]. Explorer [Eng05, Pas04]. Exploring [NN03a]. Exposed [Gum04b, Sta07u, Sta06m, Jam05, MSK03, SSS06]. Exposing [KS08]. Extended [GWE^{+}05]. Extending [RCF^{+}21]. Extensibility [Wat13]. Extensible [GWE^{+}05, Wil04, Bir05]. extra [Mat23a]. Extract [Hel19a]. extraction [McC05]. Extreme [McD05, Mer04, SR03, Var21]. Extrusion [Bej06, Sta06l]. Eye [IdFC11, Tox14].

F [Por04b]. Fabio [Sta06a]. Fabrication [DMBB17, KB08d]. Facebook [HOSC16]. faces [FT05]. Facing [KB10, NN06a]. Factor [Jov20]. Fail [Ber07, HWH17, Hel21c, Mau15, NN14e, CB23, Zel06]. Fail-fast [Hel21c]. Failing [Hel21c]. Failure [AT17, Mur04]. Failures [Kam24]. fall [Hen06]. Fallacies} [Koh04, Gof04]. Fallows [Bal07b]. Family [MW03]. FAQs [Kar05, AL04]. Far [Mat17e, MBNN^{+}23]. Farm [Sta07a]. Farmer [Ano04n]. Fashion [ERF^{+}03, Ras03b]. Fast [MQ09, Hel21c, Var21]. Fast-forward [MQ09]. Faster [Ano13, Mat13, Sch16, Car07]. Father [Cof04, Lap06a]. Faucet [BS16]. Faye [Ano04k]. FDO [Coo23]. FDO’s [Coo23]. Feathers [Ber05]. Features [Sta06f, NN20e]. Federated [BKMR21, GPPV24]. Feel [Mat18c]. feeling [Har05a]. Fernanda [Ano10]. Fernandez [Bel08]. Ferracchiati [Sta06a]. Fett [Ano04i]. Fever [Bel04a, Boo04, NN20c, SMM^{+}04, WG08, Bel05a]. Fewer [NN12c]. FHE [Cre22]. FHIR [AHC22]. FI [Ano03f, Ano04i, Coo23]. fiction [Gro04g]. Fictional [Asc04]. Field [Ano04j, Lap05, NN13f, MAB^{+}20]. field-programmable [MAB^{+}20]. Figgins [Sch04b]. Fighting [PO05, Smi09]. Fights [Leh03]. File [Ano04m, KB05b, Kle03, McK12]. files [O’H06y, O’H06w, O’H06x]. filtering [Hel20a]. Finance [Bro15]. Find [Wea03]. Finding [Bla14, Bou03, GL11, NNB15, Woo04]. Finish [Mat18d]. Finkenzeller [And04].

Fire [KP23]. Firmware [Fra19a]. First [Ano04l, DGK^{+}23, Lap04b, Sta07a]. First-Life [Sta07a]. Fit [CG10]. Fitscapes [Koh04, Gof04]. Five [Gra08a, Lim20b, Lim22]. Five-minute [Gra08a]. Fix [Kro17, Ree19]. Fix-it [Ree19]. Flaky [All03b]. Flame [Gre16]. Flash [Can08, Gra08a, GF08, Lev08]. Flaws [KB09c]. Fleck [Che03]. Flexible [SMM23]. Flight [BFZ^{+}22, Hyd09]. floor [Rog05]. flow [De 06]. flows [AhCD22]. Focus [Coa06, De 06]. Foo [NN13f, Puz04, VGM04]. food [NN23b]. Fools [Lim17b]. Foot [Wea03]. forbidden [ARB07]. Forced [NN17b]. Ford [All03d]. Forensics [Lau04, Gum04a, MAC^{+}03b]. Forget [GBG^{+}04]. Forked [NN14a]. forking [NN24a]. formally [KB06c]. format [JC23]. Formats [DFZ^{+}09]. forms [Fra20e]. Formula [Mat18e]. Fortran [GWE^{+}05, See04]. Forward [O’H07b, Fra21b, Mau07, MQ09]. found [Cra20]. Foundation [Nau08b, Mac08]. Foundations [AS06c, Git06]. Four [Lim17a, NN22c, BY22]. Fox [Sta07a]. FPGAs [Alo18, Mat21, MAB^{+}20]. Framework [Mil08, Tel23, KLL23, Tul08, Wol03c]. Frameworks [NH20, SGN04]. Franca [WHT^{+}05]. Francae [KB04c]. Franciscus [Kha05]. Franglais [Bat04a]. freaking [NN07i]. Free [Ano03f, CPG^{+}04, DM15, Dur10, JS18, Kam14a, Kam24, Mic04, Kop05, CPG^{+}04]. FreeBSD [MNN04]. French [Ano04h]. Frequency [Kel20d, Mub18]. frequently [Hel22a]. Fresh [Mat16c]. Friction [AHC22]. Fried [SNW^{+}23]. Front [Hel17a, O’D09]. Front-end [O’D09]. FTC [All03e, ERF^{+}03]. Fully [Cre22]. Fun [Che10, NA22, PP04]. Functional [Eri16, HOSC16]. functions [NN20e]. Fund [Gro04f]. fundamental [AA22, Tho21]. Fundamentals [And04, Fin03]. Furber [Bro10]. Future [Hay04, Mar08, MAB^{+}20, NMCK20, Rit03, Sta05i, Sup04, AhCD22, Can06a, GK22, MR21, Nad21, OH05k, Rus23]. Fuzzer [Guo17]. Fuzzing [NA22, AR23]. Fuzzy [Ses04].

G [O’H06q, Rao04b]. gain [SD23]. Gains [Sta06d]. Game [Blo04, Hel21b, SKC^{+}04, Kol04, Rag06]. Games [BS04, Coa04b, Gro04f, Nar04, PP04, Wal08, WKGD08, CBB07]. Gaming [Por04a, Sta06e]. GAN [Col19b]. Gap [Goo07, PZ06]. Garbage [DEE^{+}16, DLP19, Bac07]. Garbage-Collection [DEE^{+}16]. Gardener [NN18a]. Garrett [Gib07]. Gary [Mar05a, Sta06k]. Gasević [Put07]. gate [MAB^{+}20]. Gates [Ano03f]. gdbm [Kel21a]. Geek [Goo07, Cre06b, Cre06c, Cre07a, PZ06]. General [Chi14b]. General-purpose [Chi14b]. generating [SSN14]. Generation [SP21]. Generators [HD22]. Geniuses [AT17]. George [Gum04a, Gum04b, Kha05]. Geronimo [Ano03d]. Get [ATM22, Ano03f, Hel21b, Mat18c, NN08c, SMM^{+}04, Sta05e, Sta06g, Wol04c, NN22b, Sta05a]. Gets [Sta06d, Sta06g, BC06, NN05b, Rou07]. Gettin [NN06b, NN07g]. Getting [Bor03, NN21c]. GFS [MQ09]. Ghosh [Nag04]. GHz [Sta04a]. Giants [Hel16d]. gibberish [KB05c]. Gigascale [Bor03]. GIGO [Ras04b]. Gilani [Sta06a]. GitHub [CT21, NN22e]. GitOps [Lim18a]. Give [Mat24]. Glaser [Paq06]. Global [Cre10e, Fra20c, VCR15]. GNL [NN16d]. GNU [Wol03b]. Go [Vix15a, Vix15b, NN20d]. goal [Kam24]. Goes [Ano04i, Sta04a, Sta05b, Mat24]. Goff [Koh04]. Going [Mat17e, Mau07, MGL22, Sta04a, De 06, Mau08b]. goldfish [MGL22]. Good [AE04, Coa04b, NN19c, Ele07, Nad21, NN22d, NN23c, Sim04]. Google [CC20, FAHB18, LPS24, RBP09, Sta06j, VCR15, WWG22]. Googlebombs [Sta07q]. Googling [Sta06j]. Gorton [Bha07b]. Gosling [All04]. govern [KLL23]. Government [Ano03d, Sta06d, Wea03]. GPUs [DGK^{+}23, FH08]. grade [Zen05]. Graham [Tom05]. Grail [Cre22]. Graph [Gre16, Kel20b, Ove17]. Graphics [PS09, Por04a, SMM^{+}04, Wol04c, Mar08]. Graphs [NGJ^{+}19]. Gravrilenko [Puz04]. Gray [Har05b, Sta05k, All08, Ano11, Pat03]. Great [Hum17, Mat18b, Mat18d, Mat19b, Mat23b]. Greenfield [Hil08, HJ08]. Greg [Mar05a]. Grid [Par06, Wel05, Cha07, Str05, Wol04d, Lla07]. Gridding [Sta05d]. Ground [Bel07b]. Groupon [Ano14b]. Groups [JVG^{+}24]. Grown [CPG^{+}04]. Guarded [Wan24]. Guevara [GDWC20]. Guide [Ber07, Chi08, Edw05, Ekt05, Eng05, Fox08, Lap05, Llo06, Lod04, Mat23b, Men04, NN20d, Par07, Wal18, BT03, Dwi06, HN05, Lon04, Pas04, VB07, Zel06]. Guided [AF12]. Guidelines [GL11, NN20c, TDA21]. Gurovich [Kha05]. Guru [EKG^{+}05, Gro04j, Ano11]. GWT} [Bal09, Dwy08].

habit [Poo06]. Hack [Mir19]. Hacker [Ano04n, MK14, Sta05k, Har05b]. Hackers [Sta06e]. Hacking [Gum04b, MSK03, SSS06, Sta05k, Puz04, Har05b, VGM04, Sta07u]. Hadoop [GPT15]. Hagan [Was03]. Hagino [Mar05b]. Hal [Sta07i]. Halfway [NN23b]. Hall [Ber03b, Ber05, Fox08, Hid04, Koh04, Mor06, Ram08, Sta07v, Tho05, ZM09]. Hallinan [Sta07v]. Hand [Ano04j]. Hand-Delivered [Ano04j]. Handbook [And04, Sta05k, ZM09, Fin03, Har05b, M^{+}09]. Handheld [VBH03]. handle [Ull21]. Handling [NN17b]. Hands [Koh06, MR21, MB05]. Hands-on [Koh06, MB05]. Hanrahan [Sta08g].

Happen [Coa04b]. Hard [Ano04o, Ano04p, Ele07, Kro17, NN17a, Pat04, Ros10, Gro04q, Gro04r, Gro04s, Gro05g, Gro05h, Gro05i, Gro05j, Mau05g, Mau05h, Mau05i, O’H05j, O’H05h, O’H05i, O’H06y, O’H06w, O’H06x, O’H06z, O’H06-27, O’H06-28, O’H06-29, O’H06-30, O’H06-31, O’H07d, Sta07t]. Harder [Blo04, VM04]. Hardware [Kap23, Kel19, Lev20, Mat21, NMHB16, NN22a, Sha03, WNN03, Cre07b, Kop05, NN21e, Pop24, RCF^{+}21]. hardware-enforced [RCF^{+}21]. Harm [Lap04b]. Harmful [Ros11, Gil04]. Harmony [ATM22]. Harper [Sta05k]. Harris [Sta05k]. Hartl [Hin08]. Harvey [DiB04]. Haskell [Gil14]. Haskins [Edw05]. Hate [Lap03, Lim19a]. having [KB06a]. Hazy [KNR13]. HCI [GHM15, Nad21]. HDF5 [DFZ^{+}09]. head [NN07g]. Headache [SM04]. Headlines [Ano03d]. Healing [GWE^{+}05, PBA03, Sha04b]. Health [EMC19, Cre06b]. Healthcare [AHC22]. heart [Rub06]. Heartbleed [Kam14a]. Hearts [NN20a]. Heavy [Wol04e]. Heavy-Duty [Wol04e]. Heer [Ano10]. Heigl [Paq06]. Helicopters [Ros03]. Helland [AHC22]. Hello [Sit18]. Hennessy [O’H07c]. Here [CLNM03, Hum17, Zah16, NN21d]. Heretics [ERF^{+}03]. heroin [Bat05]. Heterogeneous [Zah16]. Hickory [NN15b]. Hidden [Can06b, Kel23a, Kil16, Mag19, Coa06]. High [Vac14, Ber06, Hac06, Kle03, Ram08, Sou08, Bee06, JWC08]. High-Assurance [Hac06, Ber06]. High-Performance [Kle03, Vac14, Ram08, Bee06]. Highlights [KW17]. highly [Kop05]. Hill [Gum04b, Sta05k, Sta07u]. Himanshu [Llo06]. hinderer [Vac23b]. Hippocratic [Lap04b]. Hire [Mir19]. Hirvensalo [Zen04]. History [DKM^{+}12, MAB^{+}20, AR23, SOM21]. Hitchhiker [Lod04, VB07, Wal18, Par07].

Hitting [MAB^{+}20]. Hoarding [NN15a]. Hoc [Ber09, Lab08]. Hoffman [Sta08f]. Hofmeyr [Mau05d]. Hoglund [Mar05a]. Hole [Sta04b]. Holy [Cre22]. Holzner [Joy04]. Home [CLNM03, Get12, Lim23c, Sta07r, Vix15a, Vix15b, Cre06c, NN20c, Cre06b, Cre07a, Cre06c]. Homomorphic [Cre22]. Hones [Sta06h]. honest [NN23c]. Hootsuite [SBB^{+}17]. Hop [SB12]. Horsemen [NN22c]. Horses [Lim23c]. Host [Tom05]. Hostility [TF04]. Hosts [NN13e]. Hot [Ano03h]. Houda [Ber09]. House [Gum04a, Men05]. HPC [Loh10]. HTTP [Kam15b, Kam15a]. HTTP/ [Kam15a]. HTTP/2.0 [Kam15b]. HTTPS [AAVV14]. hubris [NN12d]. Human [KLL23, Nad21, NN23c, Sta06c, VW19, WA19, Bro04b, Can06a, Kam23, SP21, NN06c]. Human-Centered [Nad21]. human-computer [Can06a]. Human-KV [NN06c]. Human-subjects [KLL23]. Humans [NN19f]. Hunch [Sta06h]. Hundred [BM22]. Hurts [Bro04c]. Hybrid [Mat23a]. Hypervisor [Chi08, Fox08]. Hypothesis [Gra19].

Ian [Bha07b, Whi08]. IBM [Hil08, Sta05h]. ichiro [Mar05b]. IDAR [Ove17]. Idea [NN19d]. Ideal [Loh10]. Ideas [Mat18b]. Identification} [And04, Fin03]. identify [ABW15]. Identity [Hel18b, Vac20]. IDEs [Boe03, CdSH^{+}03, Ras03c]. Idle [DEE^{+}16]. Idle-Time [DEE^{+}16]. IETF [Kam15a, Kam15b]. If [Lim19a, Lov17, NN08b]. ignorance [Ano21]. II [Lon15]. Illusion [NN15f]. illustrate [Coo23]. I’m [Hel22b, Hil03, Nie03]. Image [DFZ^{+}09]. imagine [BLA21]. Immersive [Lim17a]. Imminent [LN04]. Immunity [Gro04p, Ish04]. Immunity-based [Gro04p, Ish04]. Immutability [Hel15]. impacts [FKN^{+}23]. Implementation [Ano03d]. Implementing [Ein11, JVG^{+}24].

Implications [BFAP16, BM22, Hel22a, SCQT20]. imply [Vac23a]. Importance [Ano12a, Mat18d]. Important [WHT^{+}05, CB23, Fra20c, Ras05]. Improve [NN11b]. improved [Lim23b]. Improvement [Lim23a]. Improving [CGB^{+}18, DSBS23, Lei08, SJK20, NSFG23]. in-memory [FoDF22]. in-person [Lim20b]. Incentives [SZ17]. Incidents [CC20]. including [Jov20]. incog [MGL22]. independence [Hel22a]. India [Ano03i, KH05]. Industrial [JSS16]. Industry [Kam11c, Kam21a, NFG04, NGJ^{+}19]. Industry-scale [NGJ^{+}19]. Inevitability [TS03]. inferencing [FoDF22]. Inferno [Bel13]. Information [Hen08, McC05, MGL22, Nie03, CSL^{+}06]. Ingol [Hid04]. Injecting [Che10]. Innovation [Cre10e, Cre10f]. Innovative [Kil04, Jen02, Rus23]. ins [Bir05]. Insider [TF04]. Insight [Can08]. insights [BFZ^{+}22]. Inspiring [SM05a]. Installation [Lim23c]. Installer [Coo23]. instances [JVG^{+}24]. Instant [SM04, TB03]. instruction [Chi23]. integral [BT06, GTK24]. Integrated [MB04, RCN03]. Integrating [Sar04]. Integration [Bur03, Cha05b, Koh06, Zhu04, MB05, Tak06]. Integration-Ready [Cha05b, Zhu04]. integrity [DGK^{+}23]. Intel [Wol04e]. Intelligence [Sta06c, Vid23, Lev23]. interaction [Can06a, KLL23, NN06c]. Interactive [Ala17]. Interface [Liu14, Ras03b, GAI^{+}21, McT04]. Interfaces [Tel23]. Internal [Sam14, Sta06l, Bej06]. Internet [Bal07b, Kam23, CJWG11, CP18, Gol14, JF06, JSN17, Lei08, Lim23c, NN17c, RV10]. Internet-Scale [CP18]. Interoperability [Ano04h]. Interpretability [Lip18]. Interpretable [CLK^{+}21]. interpretation [Vid23]. Interscience [Las06]. Intersection [GHM15]. Interview [All03c, All03d, All04, Bou04, Bro04a, Bro03, Bro10, Can09, Coa08, DiB04, Dit03, Fel03, Fel04a, Fel04b, Gra05, Ham05, Har03a, Kir04, Mau05c, Mau05d, Mau05e, McK03, McK05, NN04a, O’H05f, O’H05g, O’H06o, O’H06p, O’H06q, O’H07c, O’H06r, O’H06s, O’H06t, O’H06u, O’H06v, Pat03, Rus04, San04, Sta07j, Sta07i, Sta07k, Sta07m, Sta07l, Sta07o, Sta07n, Sta08f, Sta08g, Sta08h, Sta08i, Sta09b, Was03]. Interviewer [Mat17a]. Interviews [Kel23a]. Intrinsics [PSBG11]. Introducing [BSP16, LSB17, Mau09]. Introduction [Tho21]. Intrusion [Gum04a, Tom05, MAC^{+}03b]. Intrusions [Sta06l, Bej06]. Investigating [LPS24]. investigation [Kam21c]. Invisible [NN20b, Bor06]. IoT [HD22, NN17c]. IP [Nie03, Sta05b]. iPod [Ano04l]. IPS [Tom05]. IPv6 [Mar05b, Sta05h, Sta06d, iiH05]. IR [Rao04a]. Irish [Ano04k]. Iron [Lim15a]. IRS [Sta07a]. ISA [Chi23]. ISBN [And04, Bal07a, Bal07b, Bal09, Bal03, Bel05b, Bel08, Ber09, Ber03b, Ber03c, Ber05, Ber07, Bha07b, Bha07a, Cha05a, Cha05b, Che03, Cor07, Edw05, Ekt05, Eng05, Fox08, Gib07, Git06, Goo07, Gro04o, Gro04p, Gum04a, Gum04b, Hac06, Hen08, Hid04, Hil08, Hin08, Hor03, Imr03, Jaf06, Jen05a, Jen05b, Jov06, Joy04, Kar05, Kau03, Kha05, Kil04, Koh04, Koh06, Kuc06, Las06, Lec07, Lla07, Llo06, Mar04b, Mar05a, Mar05b, Men04, Men05, Mer04, Mil08, Mor06, Nag04, Nau08a, Nau08b, Paq06, Par06, Par07, Pet04, Por04b, Pos04, Put07, Put08, Puz04, Ram08, Rao04b, Sch04b, Spi07, Sta06a, Sta05k, Sta05j, Sta06l, Sta07u, Sta06m, Sta06k, Sta06n, Sta07v, Ste03, Tho05, Tom05, Tom06, Wal07, Whi08, Wol06, Wol08, ZM09, Zen04]. iSCSI [Gol03]. Ishida [Gro04p]. Isn’t [Ano03f, Sta04a, Gro04g, Lap06a]. Isolation [Kap23]. issue [Gro04i]. Issues [NN10e, NN21a]. Items [YD16]. itojun [Mar05b].

J [Ber03c, Cha05a]. J2EE [Ano03d]. J2ME [Vir05]. Jack [And03, CB23, SF04, SCB09, FM05]. Jacobi [Bal07b]. Jacobson [Sta09b]. Jain [Bha07a]. Jake [Tom05]. James [Bha07a, Sta06m, Ste03, AHC22, All04]. Jamie [Sta07k]. Jarod [O’H06r]. Jaroslav [Mil08]. Jason [Sta08f]. Java [Bal03, Cha05b, Kuc06, Mil08, Joh06, Ric06, Ses02, Tul08, Vir05, Wol04f, Zhu04]. JavaScript [Guo17, HOSC16, Liu14, Mei14b]. Jawbreaker [Vac23a]. Jeannette [SNW^{+}23]. Jeff [Bal09, Cha05b, Ano10, Sta07l]. Jeffrey [Wal07]. Jennings [Sta07j]. Jenson [Kil04, O’H06r]. Jepsen [Sta07n]. Jermy [Koh06]. Jerraya [Pet04]. Jerry [Gro04o]. Jesper [Tom06]. Jewels [NN22e]. Jim [All08, Ano11, GDWC20, Har03a, Pat03]. Job [Kel23a, O’H06j]. Jobs [Sta04a]. jockeys [NN06h]. Joe [Sta06m]. Joel [Gum04b, Sta07u, Sta07m]. Johansson [Tom06]. John [And04, Bal07b, Ber09, Hid04, Kar05, Lec07, O’H07c]. Johnson [Ram08]. Joint [DLP19]. Jonas [Bal07b]. Jonathan [Sta05k]. Jones [O’H06q]. Jordan [O’H06s, SBS^{+}23]. Jose [Coa08]. Joshua [SBS^{+}23]. joy [Lap06b]. Joyce [Bel05b]. JSF [Bal07b, JF06]. JSON [Hel17b]. Ju [Las06]. Judgment [VW19]. Jumpin [Wol04f]. Jun [Mar05b]. Jun-ichiro [Mar05b]. Just [ATM22, Rie03, SNW^{+}23, Sta06f, NN21d]. Justin [Ano20a].

Kabin [NN20c]. Kahle [Fel04b]. Kai [Las06]. Karl [Jen05b]. Kate [Coa13]. Kaufmann [Ber03c, Ber07]. Kay [BNB^{+}05, Fel04a]. Kay-OK [BNB^{+}05]. Keep [Ano04n]. Keeping [AA22, Ano20a, CW10, Rag06, Ros10, Sta06j, Vac20, NN23a]. keeps [Gee06]. Kelly [SMM^{+}04]. Kelly-Bootle [SMM^{+}04]. kept [Hel20b]. Kernel [Sta05e]. Kernels [NN17e]. Kevin [Spi07]. Key [Kel21a, Vac21]. Keys [Vac21]. KFH [NN20c]. Killalea [Kil20]. kind [Fra20d]. Kingdom [Vac21]. Klaus [And04]. Kluwer [Pet04]. Knives [NN23f]. Know [And03, Bel03, Lim19c, Lim23b, Mat16d, NN08b, NN18c, SF04, SCB09, Cha05a, Can05, CB23, FM05, Fra20c, GK22, NN20d]. Knowing [CB23, Lim23b]. Knowledge [Cha05b, NGJ^{+}19, RR18, Vid21, Zhu04]. Known [NN18d, KB06c]. Kode [NN05b, NN06a, NN07a, NN06h, NN06i, NN07d, NN07c, NN07g, NN07h, NN07b, NN07e, NN07i, NN08b, NN08c, NN08d, NN08e, NN08f, NN08g, NN09c, NN09d, NN09a, NN09b, NN10e, NN10b, NN10c, NN10d, NN10a, NN13b, NN13c, NN13d, NN13f, NN13a, NN13e, NN14a, NN14c, NN14b, NN15b, NN15c, NN15d, NN15e, NN20e, NN06b, KTS^{+}04, NN04b, NN04d, NN04c, NN05a, NN05g, NN05c, NN05e, NN05d, NN05f, NN05h, NN05j, NN05i, NN06b, NN06c, NN06d, NN06e, NN06f, NN06g]. Koding [NN19a, NN20c]. Konqueror [NN05e]. Konstantin [Puz04]. Kopper [Jen05b]. Kraken [Ano14a]. Kubernetes [BGO^{+}16]. Kudos [Sta07p]. Kurt [Sta08g]. Kurtz [Gum04b]. KV [NN05e, NN06c, NN06e, NN07h, NN20c].

L [Gro04o]. Lab [Sta05c, Don05]. Labiod [Ber09]. Labor [God03]. laboratory [NN23a]. Labs [Lim20c]. Lack [Gil04]. Laffra [Kar05]. Lamboozling [SP21]. Lance [Sta06m]. Landscape [AhCD22]. Language [Chi18, GHM15, IdFC11, KF19, Loh10, PP04, See04, LPS24, NN23b, Vid23]. Languages [Gil14, Mas04, Ric08, Sha09, Gro04i]. Large [Mei14c, Hel20a, Vid23]. Large-scale [Hel20a]. Largest [Hen08, LWD07]. Larry [Lim24]. last [Vac21]. Latency [ESTT22, Gre10, NN08d, Bro15]. Later [Gra08a]. Latest [Ano03d, Tel23]. Law [Ano03f, Mub18, Bor03, Fra20b]. laws [Kam11c]. layer [Gro05e]. Lazarus [NN15c]. Lead [Lim23a]. leader [Mat23b]. Leaders [Mat17c, Mat23a]. league [JNSM14]. Leaking [Mit13]. Lean [Hor03, Mat15a, PP03b]. Learn [Mat16d, NN23b, NN20d, NN24b]. Learned [Tur03, NN21e]. Learning [BKMR21, Bos05, CLK^{+}21, Col19c, KLL23, Kel21b, LS19, MR21, NMHB16, VW19, DSBS23, GPPV24, NN23e, RR18, Tel23, GPPV24]. Leave [Ano04k]. led [Chi23]. Ledgers [KF19]. Lee [Hen08]. Leffler [Rus04]. left [Wol04b]. Legacy [Ber05, Fea05]. Legal [SCQT20]. Legit [Sta06j]. Legitimacy [Col19e]. lens [BY22]. Leo [O’H06t]. Less [Hel22b, Kam16, Ano03g]. Lesson [GWE^{+}05, NN13a]. Lessons [NN10g, NGJ^{+}19, RBP09, Rog05, Tur03, NN21d]. Lester [Sta05k]. Let [Fra20c, NN20d, NN22c, FPWM09]. Letter [NN10g]. Letters [BS04, BNB^{+}05, CKG^{+}04, CPG^{+}04, CLNM03, ERF^{+}03, EGBA03, EKG^{+}05, Gro04m, GBG^{+}04, GWE^{+}05, KTS^{+}04, MSNL04, NFG04, RW06, SM05a, SMM^{+}04, SKC^{+}04, Sta07p, WHT^{+}05]. Level [Sta06c, Chi18, NN21e, Sin16]. levels [Mas04]. Leverage [Kel22b]. Leveraging [NG22, SGN04]. Lhotka [Cor07]. liability [Kam11c, Kam21a]. libraries [Mas04]. license [NN07b]. Licenses [Sta05d]. Life [Fra20d, Hel16a, Sta07a, Sta06d, Kri08, Wol04b, SKC^{+}04]. Light [Ano04j, NN10f]. Light-Speed [Ano04j]. Lighter [Ano13]. Like [Coa03c, Hel17b, Lim15a, Mat18c, Sta04b, KB06b]. limiting [Vix14]. limits [NN21b]. Lindsay [Bou04]. Line [Coo19]. Linear [Kam22b, Lim16a]. Lines [Sha03]. Lingua [WHT^{+}05]. Linguae [KB04c]. Lining [NN18b]. LinkedIn [Ano13]. LINQ [Nau08a, Ein11]. Linux [Hid04, Jen05b, Koh06, Sch04b, Sta07v, Tho05, Dec04, EEMS03, Hal07, Kop05, Luc05, MB05, NN16d, SFW03, Sta04b, Sta05f, Sta05h, Sta06j, Sta07q, Tox07, Wol04a]. Lisp [Mar04a]. Listen [Mat17d]. Literate [Kel22a]. Little [Bha07a, Lim23b]. Live [BTM22, BCJ^{+}18, Sta07p]. livelocks [NN08d]. Lives [GWE^{+}05]. Living [Bar03]. Load [Hel19a, Lim16b]. Location [SKC^{+}04]. Lock [Kle14]. Lock-based [Kle14]. Log4j [Kam22a]. Logging [NN06e, NN14c, Sin16]. Logic [NN14c]. Logical [BP16]. Long [BTM22, Gol14, Sta07p, Mas06, Mat24]. longevity [Mas04]. Longhorn [KTS^{+}04, Wol04g]. Longoria [Ekt05]. Look [CKG^{+}04, NN08b, FH08, Fra20a, KB06d]. Looking [Sta05i]. Looks [Ano03d]. looming [Vac21]. loose [Sta05h]. lose [KB07g]. Losing [Sta05f]. loss [SD23]. Lost [Ras04a, Sta05d]. Lot [KB08e]. LoTech [KB06b]. LOTR [Ano04l]. Lou [Sta07n]. loudmouth [NN07h]. Low [Bro15, Chi18, Kel20d, VTA23, Var21]. Low-code [VTA23, Var21]. Low-Frequency [Kel20d]. Low-latency [Bro15]. Low-level [Chi18]. lower [NN21e]. lower-level [NN21e]. Lucke [Tho05]. Lunch [Mic04]. lunchbox [Ano03g]. lution [Bar06].

Mac [Sta06g]. Machine [CLK^{+}21, Col19c, KLL23, Kel21b, LS19, MR21, RR18, VW19, BKMR21, Tel23]. Machine-learning [RR18]. Machines [PSBG11, PS09, Ras04b, Ros04, Sup04]. Macintosh [Sta05b]. Macs [Sta06b]. Mad [NN21c]. Madness [Bat04b, CPG^{+}04, Sta04a]. magic [Wan04]. mail [All06b]. Mainframe} [Hid04, EEMS03]. Maintain [KNR13]. Maintenance [SCB09]. Major [JNSM14, Poo06]. Major-league [JNSM14]. Make [Ano04n, Coa03d, Lim17a, Lim24, FT07]. makes [Fra20a]. Making [cF03, FT07, KNR13, Mat13, Mei17, Nor09, Dia15, Kit05, WG06]. Mal [Lap05]. Malware [Cre10b, Sta07r]. Man [KTS^{+}04, Lim15a]. Manage [DM15]. Management [BFAP16, Fra19b, GLF^{+}15, Gro03b, HSN13, Mat19a, NN13a, SMM^{+}04, VBH03, JW06, Kum05, Var21]. Manager [Chi14a, Mat16d, Mat23b]. Managerium [Lap05]. Managers [ATM22]. Managing [All12, Flo05, Mag19, Mat15c, Mat18a, Mat23a, NN20d]. Manning [Eng05, Kha05, Wol06]. Manual [Lim18b]. Manufacturing [Fra20e]. many [Chi23, Fra20e]. Map [Sta04b]. Mapping [AhCD22, Nau08a, Sha04a]. Maps [DVR10]. Margo [SOM21, Sta07o]. Maria [Hid04]. Mario [Fel03]. Mark [Bha07a]. Market [AAVV14, Kam24]. Markets [Sta06f]. Markings [NN20b]. Markus [Paq06]. Marsland [McK05]. Martin [Ber03b, Ano10]. Mary [Hor03, Sin99, Sta07n]. Mashatan [SBS^{+}23]. Masses [Ano04j, Jan03, Sta05g, Cre05]. Massively [Hen04]. Master [NN19e]. Math [Mei17, SBS^{+}23]. Matsudaira [Coa13]. Matt [Lec07, Mer04, Kir04]. Matter [NN13e, NR18, SNR18]. Matters [Mat19b, Med16, Hen07]. Matthaeu [Hid04]. Max [Koh04]. Maximizing [FSSP09]. May [KB09b, Lim23a]. maybe [KB06d]. Mazzola [Fel03]. McClure [Gum04b]. McCubbin [Ano20a]. McGovern [Bha07a]. McGraw [Gum04b, Mar05a, Sta05k, Sta07u, Sta06k]. McGraw-Hill [Gum04b, Sta05k, Sta07u]. McKemmish [Gum04a]. McManis [O’H06o]. McTear [Por04b]. MDD [RW06]. Me [FPWM09, LCW18, Lov17]. Mean [LVDM22, NN12c, KB06b]. Meaning [Ras04a, Vid21]. Means [Kam16, Kam24]. measurement [NN07i]. Measuring [Fra21a, Sta06e, NSFG23]. Mechanical [Fra21b]. mechanism [Coo23]. Media [Ano14c, Put08, Sta05k, Sta07u, Gar07a, Kit05, SS07]. mediation [Vac23a]. Medical [Sta06f]. medium [Vid21]. Meet [Ano04l, Kil08b, NN23b]. Meetings [Lap03, SKC^{+}04]. Meets [Sta07a, Whi05]. Mehta [Nau08a]. Meijer [Coa08, Mat17c]. Meltdown [BCJ^{+}18]. memeplex [Kru07]. Memory [CBM^{+}08, Dre08b, GLF^{+}15, Gra08a, Kel19, Kel20c, Kel22b, MBNN^{+}23, SMM23, FoDF22]. Mend [Sta06e]. Meng [Bro04a]. Merge [NN09a]. Mersenne [Ano04i]. messages [MK04]. Messaging [Jan03, SM04, TB03]. Messenger [NN16b]. META [Lon15]. Metamorphosis [KM09]. Metaphors [Vid17]. Meter [Kam21b]. Method [JSN17, JS18, JC23, NN13c]. Methods [Ber03c, Wie03]. Metrics [FK17, SNR18]. Michael [Ber05, Hin08, Por04b, Put08, Rao04b, Sta05k, Ste03, Tom05, SNW^{+}23, Sta07o]. microcontent [LHML21]. Microprocessor [DKM^{+}12]. microprocessors [OH05k]. Microservice [ERR18, VD24]. Microservice-based [VD24]. Microservices [Kil16]. Microsoft [Par07, Ano03d, Ano04i, CGB^{+}18, Wol03c]. mid [Lim23c]. mid-process [Lim23c]. Middle [Mei14a]. Middleware [Hen04, O’H07e, GK22]. Miguel [GDWC20]. Mika [Zen04]. Mike [Mor06, Sta07u, NN04a, SOM21]. Mikhailovsky [Puz04]. Mile [BY22]. Milk [Ano04n]. Mind [AM03, Ano04j, Hel18c, O’H06l]. Minda [Goo07]. Mindset [O’D17]. minute [Gra08a]. MIPS [Mac03a]. Mirror [LPS24]. Misery [Kam14b]. Mismatch [Vac23b]. miss [Hel21a]. Missed [SD23]. mix [Hal05]. Mixed [Koh06, MB05]. Mixing [VW19]. ML [KLL23]. Mobile [Mei14c, CL11, Cre11a, Cre11b, Ekt05, Kit05, Mat13, Sta05a, Sta06j, Sta06e, YRZB16, Ger05, Lon04]. Mobility [TR03].

Model [BLK15, Coo23, FPWM09, GDD06, LN04, Lip18, Cra20, Put07]. Model-based [BLK15]. Modeling [Mon10]. models [Vid23]. Modern [Fra18, GLF^{+}15, Gro03b, Pet18, Pur06, Fel05a, HD22, KA23, Kel21b, Sha04b]. Mohay [Gum04a]. Molecular [Rao04b, SARC03]. Molecules [Ano04i]. Moment [KB04b]. Money [Kam14a, Mei17]. MongoDB [Guo17]. Monitoring [Hof05, Pur06, Sch17a, Sta06l, Bej06]. Mood [Sta06e]. Moore [Ano03f, Bor03, Fra20b, Sta07l]. morale [Mat24]. Morgan [Bel05b, Ber03c, Ber07]. Moskowitz [Koh06]. Mothra [NN05j]. Move [Nor09, VCR15, Fra21b, Kel22b]. Moving [CLK^{+}21, Cre10c, Cre10d, Hil08, HJ08]. Much [LV03, CSL^{+}06, Fel05b, Kam24, NN21e]. Mugs [NN20a]. Multi [PP04]. Multi-Language [PP04]. Multicore [Cre05, FSSP09, Mor16]. multifactor [Jov20]. Multimedia [EK12]. Multiparty [SBS^{+}23]. Multipath [PB14]. Multiplayer [Hen04]. Multitier [SB12]. Musican [Ano04m]. Must [Wad14, Vac23b, NN19b]. mutation [DSBS23]. My [Ano03f, Hel21a, Sta06g, Ano03g, Kri08]. MySQL [Bel05b, CPM04]. Mythos [Lip18, CLK^{+}21]. MythTV [Put08, SS07]. Myware [Sta06i].

N [Ano14b]. Naivety [TF04]. Name

[Hel18b, Mat24, Coa04a]. Naming [NN13e]. Nan [Kau03]. Nareyek [BS04]. NAS [Llo06, Dwi06]. Nathaniel [Git06]. Native [CL11]. Natural [CG10, GHM15, Zen04]. Nature [KB09c, SP21]. NBOR [Ano04l]. Near [Ano04j, Sta05c, Wol04d, Cre22]. necessary [Hel22b]. necks [NN22b]. Need [Goo07, Sta05g, DLdFV^{+}23, Kam21c, NN24b, PZ06]. Needle [IdFC11]. Needs [LNV14]. Neeed [Lim23b]. nerve [MAB^{+}20]. Ness [Sta05k]. ’Net [Sta06f, Kuz19, NN03c]. NetAccel [Col19d]. Netflix [LSB17, Liu14]. Network [Ala17, Ano04k, Vac14, BK14, Con04, Cre10c, Gof04, Gum04b, Koh04, Lau04, Mar05b, NN12d, O’D09, Rix08, Sta05j, Tom05, Tom06, VCR15, Vas17, AS06b, FM05, JR05, MSK03, iiH05, Ste15]. Networked [Kle03]. Networking [AF12, All03b, Hin08, NN15e, HP08]. Networks [BB03, BP09, Ber09, Ekt05, PBA03, ZAB16, Cre07a, Lab08, TCR03]. Neumann [O’H06q]. Neutrality [Kuz19]. Never [Coa03a, Dur10, KB04e]. newbie [NN07c]. News [Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano04h, Ano04i, Ano04j, Ano04n, Ano04l, Ano04m, Ano04k, Gro04n, Mau05f, Sta04a, Sta04b, Sta05c, Sta05d, Sta05e, Sta05f, Sta05g, Sta05b, Sta05h, Sta05i, Sta06b, Sta06c, Sta07a, Sta06d, Sta06j, Sta06f, Sta06e, Sta06g, Sta06h, Sta06i, Sta07q, Sta07r, Sta07s]. Next [Mat18e, Sta06c, Lim23a, NN07e]. Nice [NN12a]. Nicholas [Rao04b]. Nielsen [Edw05]. Nigel [SBS^{+}23]. Nightmare [WNN03]. Nine [Hil03, Mat16d]. Nineties [Sta05i]. nirvana [Nol06]. NJ [Bha07b]. No [Ano20a, Chi14b, Jen05b, Kam23, Lap04b, Lim18c, Mic04, PP03a, She15, Sta05f, Sta05g, Sta06g, NN22d, NN23c]. Node [Ano12a, Ano13]. Node.js [Vac14, Ham14, Ano14c, Ano14a, Mei14c]. Non [NSWW15, NN20d]. Non-psychopath [NN20d]. Non-volatile [NSWW15]. Nonobvious [Lim20b]. Norbert [Pet04]. Norvig [SNW^{+}23]. NoSQL [Kel21a]. notebook [NN23a]. Notes [Ham14]. Nothing [KTS^{+}04, Sta04b]. Notion [SKC^{+}04]. novel [Ull03]. Novice [Bal07a, SB06]. NSA [Tox14]. null [NN15b, KB04e]. NUMA [GLF^{+}15]. Number [HD22]. Numbers [Lim20a, NN19f, HD22]. Nutch [CC04].

Nutshell} [Ste03, BD03, SFW03, Sch04b]. NVIDIA [DGK^{+}23]. NVM [BFAP16].

Oath [Lap04b]. Obama [Boe09]. Obie [Bel08]. Obje [Ano04h]. Object [Cor07, Nau08a, Rus08]. Object-Relational [Rus08]. Objects [ZM09, vD15, Lho06, Ses04, Zan04, Zan08]. Obscene [NN18d]. Observer [NN17d]. Obsolescence [NN10b]. OCCAM [NG22, KB08c]. OCCAM-v2 [NG22]. Ode [Car08, Ano14b, KB07d]. Off [Ano03e, NN21c, Bos06, NN22b]. Office [Ano04l, CLNM03]. Official [AL04, Kar05]. Offline [Kel20d]. Offload [Cur04]. Often [Cre22, KB04d, NN09a]. Oh [Sta06g]. OK [BNB^{+}05]. ol [KB06b]. old [BTM22, Cra20, Fra21b]. Oliver [Bha07a]. Olivier [Gum04a]. Olson [SOM21]. Omega [BGO^{+}16]. on-site [Mat23a]. Onboarding [Coo23]. once [BLA21]. One [Bla14, BM22, Cre09b, Gor07, Kam11b, KB09a, Lim22, SD23, Tho21, Kam22a]. One-second [Kam11b]. ones [VD24]. Online [KBS19, KW17, Sta06j, Swe13]. only [CBM^{+}08]. Ontology [Put07, GDD06]. Oops [Bro04b]. Open [ATM22, AM03, AM04, Ber03a, CC04, CPG^{+}04, DM15, Fer03, For07, Fox08, Fra19a, Hub04, Kam24, Kar03, KR03, Kuc06, Men05, NN20d, NN24a, Sta06g, Wan24, Wea03, Wol03a, XMGM21, Ano03i, CFL23, Vir05, BLA21, Mau05f]. Open-Source [Wol03a, ATM22, Fra19a, NN20d, XMGM21]. OpenGL [SMM^{+}04, Wol04c]. Opening [Fra19b]. OpenOffice [CLNM03]. OpenSSL [Kam14b]. Operating [Col19e, Wol04g, Sha04b]. Operational [Lim17b, Mat23b]. Opinion [Bat04b, Bat04c, BP09, Bel07b, Coa03d, Fre03, Lap03, LN04, Lap04b, Ras03a, Ras03b, Ras04b, Ras04a, Ros03]. Opportunities [Bor03, BFZ^{+}22].

Opportunity [GF08, SD23]. Optical [BP09]. Optimizations [God19, Lim16a]. Optimizing [SD23]. Option [Sta04b]. Opus [Kam22a]. oracles [AR23]. Orchestrating [Don05]. Order [Gro03a, Noy05]. Orebaugh [Tom05]. O’Reilly [Che03, Joy04, Mar04b, Sch04b, Sta06n, Ste03]. organization [MH21]. Organizational [VD24, Lim23b]. organized [MBNN^{+}23]. Oriented [Bha07a, Las06, MSJL06, WQ05]. Original [Kel21a]. ORM [KS08, Ric08]. Osborne [Sta05k, Sta07u]. OSGi [Kri08]. OSS [KMNNM22]. OSx86 [Sta05g]. Other [Coa04a, Goo07, Hel18b, Kam24, Kro17, Pet09, Lap04a, PZ06, SD23]. Our [RC04, NN22b, Sta05f]. Out-of-this-World [Fra20e]. outages [Lim20c]. Outfoxing [Ano03g]. Outsourcing [Kol04, NN14b, RW06, Ano03g, Pat05]. overdue [Kam21a]. overhaul [Kam21b]. Overkill [NN22a]. Overload [EKG^{+}05]. Overly [Mat19c]. Overrun [Bat04b, CPG^{+}04]. Overruns [CKG^{+}04]. overstrike [Gro04e]. Overview [Cre09a, Cre10d, Cre10f, Cre11b]. Overwhelmed [Lim22]. Own [Pat04, NA22, NN23e]. Owns [EMC19]. Ozzie [O’H05g].

P [Men05]. P2P [NN07a, Sta07s]. Page [vD15]. Pain [Ein11, Gus05]. pair [BFZ^{+}22]. pair-programming [BFZ^{+}22]. pale [NN22c]. pandemic [BY22]. Paradox [Mat15b, Sta06b]. Parallel [Dre08b, Rat15, RBSH14, Boy08, NBGS08]. Paranoia [NN08g]. Parchment [NN23e]. Parity [Lev09]. Park [Bel05b]. Part [Cre09a, Cre09b, ATM22, Cre08]. Partnership [Ano04i]. Pass [NN16a]. passage [KH05]. Passin [Eng05]. Passing [IdFC11]. Passports [WRS^{+}09]. passwords [Ano03i]. Past [KB10, NMCK20, Sta05i, Sup04]. Pat [AHC22, Ham05, Sta08g]. patch [Vac21]. Patches [NN15d]. Patching [Bra05, Dad05]. Patent [Ano21, Sta05b]. Path [Ber03a, ESTT22, Lim18a, MH21, NN21c, NN23e]. Pathologies [Jac09]. Patient [Can10]. Patterns [Ber03b, Mat18a, NMCK20, ZM09, Hel22a, Mar03, Sma06, Zan04, Zan08]. Patterson [O’H07c]. Paucity [NN10a]. payment [AhCD22]. PayPal [Ano14a]. Payroll [Ano04i]. Pays [BS04]. PBX [Cof04]. PC [Put08, SS07]. Peachpit [Wal07]. Pearson [Puz04]. Pedigree [NC17]. Peerless [NN07a]. Peikari [Mar04b]. People [Cha06, Coa04b, Fra20f, Mat16a, Pet09, RC04, See03, Arn05, BM22, NN23b, NN24b]. perception [Cro06, O’H06k]. Perfect [TF04, Bla04, Pos04]. Perform [Pos04, Bla04]. Performance [CB23, Kle03, Lei08, Mil10, Pur06, Sma06, Sou08, SW14, WA19, Vac14, Bar05, Bee06, FM05, JWC08, Mat21, Ram08, Rix08]. perils [NN13e]. Perimeter [Tom06, JR05]. Perl [Jaf06, Wai05]. Permanence [BB18]. Perplexed} [Men04, BT03]. Persistence [Cob22]. Persistent [Kel19, Kel20c, Kel22b]. person [Lim20b]. Personal [Ano03h, DMBB17, Hel21a, Lim23b, MGL22]. Perspectival [LPS24]. Perspective [Boe09, McK12, Spi07, Ish04, RCN03, Tat06]. Perspiration [Hel19c]. Pervasive [YD16]. Peter [Bal03, Jaf06, Par07, All03d, Mau05d, O’H06q, SNW^{+}23]. Peterson [San04]. Peut [KB09a]. Peut-Être [KB09a]. Pfleging [Goo07]. Phenomenon [WFW^{+}19]. Phil [O’H05f]. Phishing [ARB07, Jam05, NN06f, Sta06m]. Phone [Ano03f]. Phoning [Kam15a, Kam15b]. Photons [Ano04i]. Photoshop [CW10]. PHP [ZM09, Zan04, Zan08]. PHP5 [Bel05b, CPM04]. Physical [Sta05e, YD16, ZAB16]. Physics [Smi09]. pi [KB06c]. picking [NN13c]. Pickled [NN15d]. Piece [NN12a]. PiM [FoDF22].

Pinkard [Tom05]. Pioneer [Can08]. PKI [NN06g]. Place [Bel04b]. placing [Hel21c]. plain [Can06b]. Plan [NN18f, Pos04, Bla04, Kol04]. Planet [DG15]. Planning [Mor06, NN22d, Coh06]. Platform [Ano12a, Wol03a, Wol04g, BLA21]. platforms [LHML21]. Play [Fra20c, Sta06h]. Player [NN16a]. Playing [Gee06, O’H06i]. Plea [Lim10]. Please [Kam14b]. Plenum [Gro04o]. Plug [SM05a, Bir05]. Plug-In [SM05a]. plug-ins [Bir05]. Pocket [Vac20]. podcast [KB06b]. pointer [NG22]. Pointers [MBNN^{+}23]. Pointless [NN06g]. Poisonous [NN08e]. POJOs [Bal07a, SB06]. policies [KA23]. policy [Ano21]. Polish [Vir05]. Pollution [Nie03]. Poppendieck [Hor03]. Popping [NN17e]. popularity [Chi23]. Port [NN14d]. Portable [BB03, Fra20a]. Porting [NN11c]. Portrait [Coa13, NN12b]. Ports [NN10a]. POSIX [NN22b]. Possible [WHT^{+}05, Fra20a]. Post [BCJ^{+}18]. Post-Meltdown [BCJ^{+}18]. Postmortem [Pac11]. potential [SD23]. Potter [Che03]. Power [FSSP09, Fra20f, Gro03b, Hel16b, Sax10, Sta07s, Fra20c]. Power-Efficient [Sax10]. powered [BFZ^{+}22, Sta04b]. powerful [NH20]. Powering [Gar07b]. Practical [BR06, Bue14, Llo06, SS07, Sta07v, Tul08, Dwi06, Hal07, Mil08, Gib07, Put08]. Practice [AGB17, BSP16, BFAP16, Bro15, DMBB17, KA22, NMHB16, NH20, Par07, YRZB16, ZAB16, ZM09, KGM21, Mie06, VB07, Zan04, Zan08]. Practices [JS18, Lim19a, Ber03b, CFL23, Mar03]. Practitioner [Ekt05, Lon04]. practitioners [SC21]. Praise [NN21e]. Pramodkumar [Jov06]. Pranks [Lim17b]. precise [ESTT22]. precisely [Ses02, Bal03]. Predecessors [SM05a]. Prediction [CG18]. Prediction-Serving [CG18]. Prejudice [NN09d]. premature [NN13e]. Prentice [Ber03b, Ber05, Fox08, Hid04, Koh04, Mor06, Ram08, Sta07v, Tho05, ZM09]. preparation [Lim23b]. Prepare [Pos04, Bla04]. prepared [MH21]. presence [AA22]. Present [NMCK20]. Presentation [Mac08, Nau08b]. preserving [BKMR21]. Press [Bal03, Cha05b, Hil08, Imr03, Jen05b, Kil04, Mar05b, Rao04b, Sta05j, Wal07, Wol08]. pressure [Hel21c]. Prevention [Tom05]. price [Bar05]. Pride [NN09d]. Prime [Ano04i]. Primer [CAB^{+}14, Sta07v, BG06, Hal07]. Primitives [Bue14]. Principle [Lim16c, Ras04b, Vac23a]. Principles [Ber03b, RV10, Mar03]. Print [Ano03e]. printing [Fra20e]. Priority [EKG^{+}05, Gil04]. Prisons [JS18]. Privacy [AhCD22, BKMR21, Can05, Cha05a, DRW^{+}14, DL09, GDWC20, Kam16, MGL22, Rus23, SBS^{+}23, Lim23a, Wan24]. privacy-preserving [BKMR21]. Private [KW17, O’H06n, Vac21]. Pro [Bal09, Jaf06, Nau08a, Nau08b, Sta06a, Wai05, Dwy08, JF06, Mac08, Tit05, VW07, Vir05, Bal07b, Lec07]. Probably [Hel22b, McM15]. Probing [PS09]. Problem [Kam11c, Kam21a, KTS^{+}04, Mac04, Mat16a, PP03a, Ran04, AA22, MK04, NN06i, NN24b, Vie05]. Problems [CKG^{+}04]. Process [CGB^{+}18, Fra19c, Pos04, SMM^{+}04, BM22, Cha06, Lim23c]. processes [Bla04]. Processing [FS04, KBS19, Rat15]. Processor [Chi14b]. Processors [DKK^{+}04, FS04, O’D09, PS09]. Prochazka [Hin08]. Product [Chi14a, Men05]. Production [Col19c, Sch09]. Productivity [FSM^{+}21, NSFG23, RBSH14, VTA23]. Profession [Sta06f]. Professional [Bal07a, Bel08, Cer14, Cha05a, Edw05, Hac06, Hid04, Hin08, Jov06, Kar05, Koh04, Llo06, LNV14, Men04, Par07, Pos04, Spi07, Sta06l, Sta06k, Tom06, Kam21a, SB06].

Professionals [Cha05a, Goo07, Can05, PZ06]. Profiling [ESTT22]. Profit [Che10]. Program [ATM22, NG22]. Programmability [DKK^{+}04]. programmable [MAB^{+}20]. Programmer [Boe09, Cer14, Kel23a, PSBG11]. Programmers [Arn05, Sta06h, KB07d, NN08e, Vid21]. Programming [AGB17, Bat04a, Cob22, Dre08b, GWE^{+}05, Kel19, Lar08, Las06, Loh10, Mer04, NN03c, RBSH14, Rob04, SB12, Sta06c, Wol08, Jen05a, Mar05b, BFZ^{+}22, BW07, Gro04i, NBGS08, SR03, WQ05, Wil04, iiH05]. Programs [Ber07, Mor16, Vid21, Zel06]. Progress [RBSH14]. Project [Mat24, NN20d, NN22c, CB23, Var21]. proliferating [HD22]. Prometheus [Goo07]. Promise [Can10, Ver04]. Promises [AHMT03, Rat15]. Protect [JR05, Tom06]. Protecting [Kel23b]. Protection [Col19e, SMM23, RCF^{+}21]. Protections [GDWC20]. Protocol [NN12d, Ste15]. Prototypes [Vin03]. proven [JC23]. Provenance [CAB^{+}14, HSN13, GAI^{+}21, GAI^{+}21]. provide [ESTT22, JC23, Mat21]. Providers [Ein11]. Provides [Sta04b, Hel20a]. PS3 [Sta07s]. psychopath [NN20d]. PTR [Ber05, Fox08, Mor06, Ram08, Sta07v, Tho05, ZM09]. Public [RB17, GAM18]. Publications [Eng05, Kha05, Par06, Wol06]. Publishers [Pet04]. Publishing [Gro04o, Sta06m, Tom05]. Pull [NN16f]. Punishment [Ano04m]. Purpose [Coa06, Sha09, VCR15, Chi14b, Fra21a]. Purpose-Built [Coa06, Sha09, VCR15, Fra21a]. Pursuit [Ano13, GDWC20, SBB^{+}17]. pushes [Ano03g]. Put [Kam14b]. Puttin [Gro05f]. Putting [Col19c, Ern03]. PVR [Put08, SS07].

Q [Coa06, De 06]. QA

[BNB^{+}05, Coa05, CDH14, Gro05f, Whi05]. Qian [Las06]. Quality [DM15, Fel05b, Kam14a, O’H06j, Sta07r]. Quantity [Sta07r]. quantized [FoDF22]. Quantum [GTNNM21, Hir04, Imr03, MH21, Zen04, Sin99]. Quantum-safe [GTNNM21]. Queen [Sin99]. querying [GAI^{+}21]. Questioning [RW06]. Questions [Bou03, NNB15]. Queue [BS04, Coa13, SMM^{+}04, Gro05f, NN12b]. Queueing [EKG^{+}05, EKG^{+}05]. queuing [Gil04].

R [Ber03c]. race [GTNNM21]. RAID [Lev09]. Rails [Bel08, HP08, Hin08]. RailsSpace [Hin08, HP08]. Ramachandran [Gro04o]. Rambidi [Rao04b]. Random [HD22]. randomness [Hel22b]. Ransom [Sta05h]. Ransom-ware [Sta05h]. rarely [Hel22a]. Rash [Tom05]. Rate [Vix14]. Rate-limiting [Vix14]. Rationalizing [Cre06c]. Raw [NN15e]. Ray [O’H05g]. razor [KB08c]. RC1 [Ano03d]. React [HOSC16]. Reactive [Ber03c, SBB^{+}17, Wie03]. Reactivity [Med16]. Read [Hel19c]. Readability [EGBA03]. Reader [O’H06y, O’H06w, O’H06x, WHT^{+}05]. Readers [Gro04e]. Reading [NFG04, Spi03]. Reads [Sea16]. Ready [Cha05b, Zhu04, Har03a]. Real [Boo04, CB08, HD22, NN08c, Sta05f, Sta05g, Sta06h, Sta07v, Ull21, Wol08, BW07, Hal07, KB07a, Sta05a]. Real-time [Wol08, BW07]. Real-world [CB08, Sta07v, Ull21, Hal07]. Reality [KB08d, O’H06k, CB07, Kit05]. Really [Ano03h, CPG^{+}04, SKC^{+}04, NN24b, O’H05e, O’H06j]. Realtime [Bac07, BLA21, NN08c, Lap06a]. Rearranging [NN22d]. reason [NN22d]. reboot [KA22]. rebooting [NN13e]. Reception [Sta05e]. Recipes [Kha05, FG05]. Recognition [Mat15b].

Recommendation [Bha07a]. recommendations [MSJL06]. Reconfigurable [TS03]. Reconfiguration [Gro04o, VT03]. Reconsidered [Sta07s]. Reconstruction [GAM18]. Recording [DKM^{+}12]. Records [EMC19]. recovery [Bel05a]. reduced [WWG22]. Reducing [AHC22, NN17f]. Reduction [Sta05b]. Redux [KTS^{+}04, MSNL04, NN08h]. Refactored [Mer04, SR03]. Refactoring [AS06a, Jov06]. Reference [Hid04, Koh04, Mar04a]. Reflection [Sta06a, Tit05]. Reincarnation [Ros04]. Reinventing [WWG22]. Reject [Sta05e]. Relational [Nau08a, Rus08, Sel05]. Release [Ano14a, Poo06]. Relevant [Wol03b]. Reliability [Sha19]. Reliable [BK14, DG15, Hac06, Ber06]. reloaded [NN05f]. Remote [Lam18, Lim20b, OO03, Mat23a]. Remoting [Sta06a, Tit05]. Removing [NN20e]. RenderMan [Whi08, Ste07b]. replace [WWG22]. replacement [NN22f]. Replacements [Ano03d]. Replacing [SMM23]. report [CT21, Kam22a]. reporting [Mur04]. Repository [CFL23]. Repurposing [Cre07b]. reputation [PO05]. require [Kam21b]. Required [NFG04]. requirements [BG06, JC23]. Rescue [Cur04]. Research [AGB17, BSP16, BFAP16, DMBB17, KA22, NMHB16, YRZB16, ZAB16, CBM^{+}08]. Researchers [Lim23a, CFL23, KLL23]. Resistance [MH21]. resolution [Lim24]. Resolving [KLL23, Mat16e, NN21a]. Resource [NN13a]. Respect [NN20d]. Respond [May03]. Response [Gra19, Tom05]. Responsibility [NN14b]. responsibly [Wan24]. Responsive [MK14]. rest [MR21]. Restrictions [KB09b]. results [DSBS23]. rethink [MBNN^{+}23]. return [NN04d]. Returning [PSBG11]. reveal [AhCD22]. Revealing [WA19]. Reverse [Ano03f]. Review [And04, Bal07a, Bal07b, Bal03, Bel05b, Bel08, Ber09, Ber03b, Ber03c, Ber05, Ber07, Bha07b, Bha07a, Cha05a, Cha05b, Che03, Cor07, CGB^{+}18, Edw05, Ekt05, Eng05, Fox08, Gib07, Git06, Goo07, Gro04o, Gro04p, Gum04a, Gum04b, Hac06, Hen08, Hid04, Hil08, Hin08, Hor03, Imr03, Jaf06, Jen05a, Jen05b, Jov06, Joy04, Kam22a, Kar05, Kau03, Kha05, Kil04, Koh04, Koh06, Kuc06, Las06, Lec07, Lla07, Llo06, Mar04b, Mar05a, Mar05b, Men04, Men05, Mer04, Mil08, Mor06, Nag04, Nau08a, Nau08b, Paq06, Par06, Par07, Pet04, Por04b, Pos04, Put07, Put08, Puz04, Ram08, Rao04b, Sch04b, Spi07, Sta06a, Sta05k, Sta05j, Sta06l, Sta07u, Sta06m, Sta06k, Sta06n, Sta07v, Ste03, Tho05, Tom05, Tom06, Wal07, Whi08, Wol06, Wol08, Zen04, Bal09]. Reviews [ZM09]. Revisited [AM04, EKG^{+}05, Mei14b, Mor04]. Reviving [Sta05i]. Revolution [Bur03, Lim23a, Por04a, Sta06c, Sta07r, Cre06a, SL05]. rewind [Fra20d]. Rex [Pos04]. RFID [And04, EKG^{+}05, Fin03, Gro04g, Sar04, Sta06h, WRS^{+}09, Wan04]. RFIDs [Ano03f]. Rich [Bal07b, JF06]. Richard [Men04, Paq06, Sta06l]. Ride [Mau05f]. rider [NN22c]. Riding [Ano14b]. Riely [Tom06]. Right [Asc04, Bou03, Mat18e, NNB15, NN22a, SMM^{+}04, Cob22, Kam23, Mat23b]. Rights [NN13b]. RISC [Chi23]. RISC-V [Chi23]. Rise [Cre22, Hen06]. Risk [DL09]. Risks [Sta06h]. Road [FRZ13, Por04a, Mas06]. Roadmap [Chr09]. Robert [Ber03b, Edw05, Kuc06, Tho05, NN12b]. Robinson [Jen05a]. Roboexotica [Sta07r]. robot [Mau05f]. Robotics [Sta07r]. Robots [Sta05d]. Robust [RV10]. Rock [Ano03h]. rockets [Fra20e]. Rockford [Cor07]. Rod [Gum04a]. Roger [Mau05e]. Role [Mat18e, WA19, AS06b]. Roles [Ano03f]. Roman [Ekt05]. Romance [KB08a]. Room [Bar03, NN22b]. Rooney [Gib07]. root [Ran04, Vac21, KTS^{+}04]. Rosenberg [Mer04]. Ross [O’H06u]. Ross-Talbot [O’H06u]. Rossing [Bro03]. Rototilling [NN11a]. Roundtable [Bou08, Cre10b, Cre10c, Cre10d, Cre10e, Cre11a, Cre11b, Cre08, Cre09b, VOT^{+}09]. Routing [Gol14]. RPCs [Col19b]. Ruby [Hin08, HP08]. Rule [Gra08a]. Rules [Gra08a, KB09b, O’H06i]. Run [Ano03h]. runbook [NN22e]. running [DLdFV^{+}23]. Ryan [Git06, Wol06].

Sadalage [Jov06]. Saddle [NN06h]. Safe [Ano12b, Ros10, AA22, GTNNM21]. Safety [Kam22a, Ker24]. SageDB [Col19d]. said [KB06a]. Sailor [Car08]. sake [NN23e]. Salm [Hid04]. Salt [NN16a]. Sam [Bal07a, Rus04]. Sam-Bodden [Bal07a]. Samba [Ano03d, Wol04a]. Same [Lim22, Hel21b]. SAN [Llo06, Dwi06]. sandbox [Whi05]. Sanity [NN20b]. Sanjay [SBS^{+}23]. Saravanan [SBS^{+}23]. Save [Ano03h]. Say [Ano03f, Sta05f]. Scalability [Bue14, CW10, GPT15, Mac04]. Scalable [Kil08a, NBGS08, Rys11, NG22]. Scale [Mei14c, CP18, Eri16, JSS16, Mau15, Hel20a, Ker24, Kil20, NGJ^{+}19]. Scaling [Kle14, Kuz19, Mor16, Mub18, Wal08, McD05]. scalpel [SMM23]. Scam [Ano04k]. Scambray [Gum04b, Sta07u]. Schedulers [CP18]. Scheduling [DEE^{+}16, KA23, Sch17b, MNN04]. Schema.org [GBM15]. Schizoid [Bat04c, GBG^{+}04]. Schlossnagle [Ano20a]. Scholarship [LS19]. Schrödinger [KGM21]. Schutta [Git06]. Science [RB17, GL11, Imr03, Mat15c, BKMR21, Gro04g, Kel21b, NN23d]. Sciences [DRW^{+}14]. Scientific [NN13c]. scientifically [Lev23]. SCM [Whi05]. SCO [Leh03]. score [Rag06]. Scots [Sin99]. Scott [Jov06, Kil04]. Screen [Gro04k]. screens [CBB07]. Scripting [Vin03]. Scriptorium [Lon15]. Scripts [WKGD08]. Scrum [SJK20]. SDN [BS16, FRZ13, VCR15]. sea [KB06b]. Sean [Sta05j]. Search [BT06, Bro04c, CC04, Gro04h, Kel20b, Lim16a, MM04, Pat04, Rao04a, Sta04a, Sta06d]. Search-as-a-Service [Sta06d]. Searching [Woo04]. Secaucus [Bha07b]. Second [Gib07, Jov20, Kil20, Sta07a, Cor07, Cra20, Kam11b, Wal07, ZM09]. Second-Life [Sta07a]. Secrecy [Imr03, Sin99]. Secret [Mat18e, Mat23b]. Secrets [Gum04b, Kel23b, Puz04, VGM04, MSK03]. Secure [Ano12b, Gol14, Hac06, SBS^{+}23, Ber06, Car07, For07, Har05a, NN07f, Rus23, TCR03, Vac20]. Securely [KW04]. Securing [Dwi06, Fra19c, Fre03, Ker14, Llo06, NN22e, Owe10, Ste15, Tox14]. Security [Ano04l, AAVV14, CKG^{+}04, DL09, Fra18, GTK24, Gum04b, Hen08, Jov20, KMNNM22, Lev20, Lla07, Llo06, Mar04b, NN11b, PC04, PF03, Ran04, RBP09, Rus23, SL04, Sta05e, Sta06b, Sta05j, Sta06l, Sta06k, Vac23a, Vac23b, VM04, Vie05, WG09, YRZB16, Bej06, Cha07, Con04, Cra20, Dwi06, Gee07, LWD07, Mat21, MSK03, McG06, NN22e, Tox07, Che03]. See [Ano03d]. Seeing [Tox14]. Seeking [Nol06]. Seetoft [Bal03]. Self [GWE^{+}05, Lim18a, PBA03, Sha04b]. Self-Healing [GWE^{+}05, PBA03, Sha04b]. Self-service [Lim18a]. Selinger [Ham05, SM05a]. Seltzer [SOM21, Sta07o]. Semantic [Eng05, Pas04]. Semantics [Sta06i]. SEMAT [JNSM14]. semi [Flo05, SM05b]. semi-structured [Flo05, SM05b]. Sender [BS04, EK12]. Sender-side [EK12]. Sensible [Sch04a]. Sensor [Ber09, HSN13, Lab08]. Sensors [ZAB16]. Sentient [FKB^{+}03, SKC^{+}04]. sequences [Ull21]. serendipitous [Hel20a]. Series [Fox08, Hin08, Par07, Zen04, Ano20a]. Serious [Sta06d]. Server [Ano03h, Bar03, GF08, Par07, Cre06c, Vog08a, VB07]. Service [Bha07a, DB18, Lam18, SDRB17, Sta06d, Vin03, FT05, Hof05, Jac05, Lim18a, MSJL06]. Service-Oriented [Bha07a]. Services [AHMT03, Bur03, CP18, Kil08a, KR03, TR03, Kum05, Ses04]. Serving [CG18]. Sessions [Mau05e]. Set [Sta06b, JS18]. SETI [Ano04i]. Sets [KP23, Chi23]. Setting [Mat23b]. Settle [LFKA14]. SEV [Kap23]. SEV-SNP [Kap23]. seven [Tox07]. Shall [Sta04b]. Shalt [Ano03e, LCW18]. shape [AhCD22]. Shared [KW04]. Sharing [Ano04m, Vid21]. Sharpening [GS23]. Shaw [Han09a, Han09b]. Sheehy [Ano20a]. Shema [Sta07u]. Shift [Kil04, Jen02]. Shigeyoshi [Nag04]. Ship [Dat16]. Shipping [Mat15a]. Shoehorn [Hel19a]. Shoes [BY22]. Shon [Sta05k]. Should [Cha05a, Dat16, DLdFV^{+}23, JNSM14, Lim15a, Lim19c, NN22f, Can05, KLL23, MBNN^{+}23, Vac23a]. Shoulders [Hel16d]. Side [Hel17a, MMB15, EK12]. Sidebar [LV03, May03, Wea03]. Sides [Har03b]. Sienko [Rao04b]. Siever [Sch04b]. Sifting [Whi05]. sight [Can06b]. Sights [Sta06b]. Signals [Ano04m]. Silent [Lim23a]. Silicon [Ras03a]. Silver [NN18b, Bel06]. SIM [Jov20]. Sima [Sta07u]. SIMD [PSBG11]. Simon [Imr03, Jen05a]. Simple [CW10, Hel21a, NN21d]. Simplicity [Kil04, Phi10a, Jen02]. Sims [Bha07a]. Simulators [GBG^{+}04, Sup04]. Singh [Imr03]. Single [JSN17]. Singular [Hel16c]. Sink [Bel03]. sins [Tox07]. SIP [BL07, FT07, Spa07, Ste07a]. site [Mat23a]. Sites [Sou08]. Sits [Coa03a, Kil20]. Sizing [NN08f]. Sketching [Cor17]. Skills [Lea11]. Skins [NN23f]. Slamming [Edw05, HN05]. Slaves [ERF^{+}03, Ras03b]. sledgehammer [SMM23]. Sliders [NN22a]. Slow [Sta06f]. Slump [Sta06e]. Small [Lim16c, CBB07]. Smart [And04, NMHB16, See03, Wol03c, Fin03, Rou07, CLNM03, SBS^{+}23]. smartphones [Vac20]. Smith [Put08]. Smoke [Ano04m]. Smoot [O’H05f]. SMS [Jov20]. SneakerNet [Ano04l, Sta06b].

Sniffers [Ano03d]. Snowden [Tox14]. Snowflakes [NN18e]. SNP [Kap23]. SOA [Bel07a]. SoC} [Pet04, JYWV03, WNN03]. Social [Cro06, DRW^{+}14, Hin08, MFK05, HP08]. Socially [O’H05d]. Society [FKB^{+}03]. Sockets [NN09e]. Soft [Lea11]. Software [BTM22, Bel06, Bel13, Ber03b, Bha07b, CBM^{+}08, Cha05b, Cox19, DVR10, DM15, Ekt05, EGBA03, Fox08, Gro04f, HPE03, Hor03, JS14, Kam11c, Kam14a, Kam21a, Kam24, Kar03, Ker24, Kil19, Lap04b, Lim10, Lim19c, Lod04, LNV14, Mar05a, Mat15a, Mat16a, Mat23b, McG06, Men05, Mic04, NN09c, NN16e, NN22c, NN24a, OO03, Pet04, PP03b, RC04, Sax10, Sha03, SP21, Spi07, SCB09, Sta06e, Sta07s, Sta06k, SL05, Var21, VD24, WNN03, Wol04e, WA19, ZM09, ATM22, Bro05, Dad05, Gor06, HM04, Jac05, JYWV03, Kel22b, Kop05, Lon04, Mar03, M^{+}09, McD05, Mur04, NN24b, Tat06, Tho21, Whi05, Zhu04]. Solid [Cor12, Sta05e]. Solid-State [Cor12]. Solomon [KB08c]. Solution [Kam13b, Kuz19, Var21]. Solutions [Koh06, Hel21c, MSK03, MB05, NN06f, Gum04b]. solved [Vie05]. Some [Coa03c, KB09b, Wol04b, KB07e]. Something [GBG^{+}04]. Sons [And04, Ber09]. Sounds [Hum17]. Source [AM03, AM04, Ano03d, CC04, CPG^{+}04, DM15, Fer03, Fox08, HPE03, Hub04, Kam24, Kar03, Kuc06, Leh03, PP03a, Sta06g, Wea03, Wol03a, Wol04e, ATM22, Ano03i, CFL23, For07, Fra19a, NN20d, NN24a, Vid21, Vir05, XMGM21, Mau05f]. Source/2 [Mau05f]. SOXes [Bos06]. Space [Hel19b, Mit13, FSM^{+}21]. Spaces [Kam22b, NN20b]. SPADE [GAI^{+}21]. Spaghetti [NN18d]. Spam [All03a, All03e, Ano04j, ERF^{+}03, Edw05, Sta06c, Ano03g, HN05, Lap06b, Mau05f, PO05]. Speaker [Col19e]. Specialization [NG22]. specific [Gil14]. Spector [SNW^{+}23]. Spectre [BCJ^{+}18]. Spectrum [Ber03a]. Speech [Cha05b, Zhu04]. Speed [Ano04j, NN10f, Kam22b]. Spelunking [NN03a, NN08h]. Spicing [MMB15]. Split [Lim22]. Splog [Sta05i]. Spoken [McT04, Por04b]. Spolsky [Sta07m]. Spring [Wol06, WB05]. Springer [Bha07b, Bha07a, Ekt05, Gro04p, Hen08, Lla07, Nag04, Paq06, Por04b, Put07, Whi08, Zen04]. Springer-Verlag [Bha07b, Bha07a, Ekt05, Gro04p, Hen08, Nag04, Por04b, Whi08, Zen04]. Spying [Sta05c]. Spyware [Sta04a, Sta06i]. SQL [Ano11, Par07, Col19a, Hel16c, Hel20b, Lim18c, Rys11, VB07]. Squabbles [Sta05b]. Squatting [NN14d]. Squeeze [Sta04b]. SRE [NR18]. SS7 [Jov20]. SSDs [MW08]. Stablecoins [Ano20b]. staff [NN20d]. Stand [BLK15, Lap03, SKC^{+}04, KB07a]. Stand-Up [Lap03, SKC^{+}04]. Standard [KP23, Wal07]. Standardizing [GSI07]. Standards [Bel04b, Men05, NN09b, ATM22, Ger05, Zel07]. Standing [Hel16d]. Starch [Jen05b]. starting [CB23]. Starts [Mat16c]. State [Cor12, FPWM09, Hel18c, NFG04, Vix14, vD15]. Statemate [Wie03]. Static [CT21, Nad21, NG22, Tho21, TDA21, Vix15a, Vix15b, Vac21]. Static-Analysis-Driven [Nad21]. Statistics [Har16]. Status [KLL23, MAB^{+}20, NNB15]. Stay [Zah16]. Stealth [Wol04e]. Steampunk [Kel21b]. step [Gor07]. Stephen [Sch04b]. Stephens [Mer04]. Stephenson [Whi08]. Steps [Sta05h]. Stern [Sta07i]. Steve [Joy04, Tom06, Bro10, O’H06u, Sta08h, Sta08i, Was03]. Steven [Mau05d]. Stewart [Put08, Sta06m]. Still [NNB15, Put08, Wol03b, HD22, Wol04b, Kam21a]. Stinkin [Sta05g]. Stole [Lim20c]. Stone [NN23f]. Stonebraker [Sta07o]. Stop [Pat05]. Stopping [Kam21b]. Storage [Coa03d, Har03b, Lev08, Llo06, LV03, NSWW15, NN11d, Pet18, Rie03, Rou07, Cre08, Dwi06, GSI07, Bou08]. Store [Kel21a]. Storm [TF04]. Story [Ano14a, BM13, Har03b]. straight [NN07g]. strain [NN06a]. Strategies [Ram08, JWC08]. Stream [DKK^{+}04]. Streams [Ger05]. Stricken [SMM^{+}04]. Strife [NN11d]. Strikes [Ano04l, NN04c]. String [KB04e, Ull21]. Strings [NN19f]. Strong [Ano04n, Mat24]. structure [NN24b]. Structured [GBM15, Flo05, McC05, SM05b]. Structuring [NN24b]. Struts [FG05, Kha05]. Stuart [Gum04b]. Stuck [Hel21b]. Studio [Par07, VB07]. Study [Asc04, MQ09, FKN^{+}23, NN23d]. Stuff [MM04]. stupid [Sim04]. Style [Ano04m, Sta05b]. subjects [KLL23]. Submarines [Ros03]. Subsetting [WWG22]. substitute [NN23c]. Subversion [Gib07, BR06]. Success [Hel16c, Kam13a, NN24b]. Such [Chi14b, Mic04, Hel21a]. Suck [SKC^{+}04]. Sugar [Ano03h]. Suite [Kuc06, Vir05]. Suites [Ano04l]. Summit [Ano12a]. Sumo [BNNC03]. Sun [Coa03a]. Sungjoo [Pet04]. Supercomputing [cF03]. Superlinear [GPT15]. Superstitions [Ras03a]. Supply [KMNNM22]. Supply-chain [KMNNM22]. Support [Sta04b]. Surface [NN17f]. Surprises [OO03, Hel20a]. Surveillance [DL09, Kam21b]. survey [SC21]. Survive [BG11, Goo07, PZ06]. Surviving [Cox19]. Sustainable [Tat06, Spi07]. Sustaining [VD24]. SWAN [TDA21]. swans [KB07e]. swapping [Jov20]. Swim [Bel03]. Switching [Sta06n, Wal05]. sword [KB08c]. Sybex [Koh06]. Syed [Sta06a]. Symantec [Sta06i]. Synchronization [Bue14, Mor16]. Syngress [Sta06m, Tom05]. Syntactic [Bat05]. Synthesis [Gil14]. Sysadmins [Lim10, Lim15b]. System [Edw05, GLF^{+}15, Gre10, Gro03b, HKM10, Kle03, Lea11, Mad15, McC15, McK12, NN08f, NN23d, Par07, Wol04g, HN05, JC23, Kam21b, Vac23a]. Systematic [Ber07, Zel06]. Systems [AGB17, AR23, BFAP16, BWB^{+}16, CC20, Coa06, Col19c, Col19e, CG18, DSBS23, FSSP09, God03, KW04, KM09, KW17, Men05, MR21, NN03b, Pet18, Rie03, SC21, SBB^{+}17, TS03, Tho05, BKMR21, Bro04b, De 06, Hel21c, HD22, Ish04, KA23, Luc05, PO05, Sha04b, Sim04, Wie03, Gro04p, Ber03c].

Tabs [NN20b]. Tackling [Mon10]. Tactics [Kil04, Jen02]. Take [KMNNM22, NN07i]. Takes [XMGM21, Fra20e, Mat23a]. Taking [BFZ^{+}22, Gol14, Sta06c]. Talbot [O’H06u]. tale [NN12d]. Tales [EGBA03]. Talese [Kau03]. Talese/Doubleday [Kau03]. Talk [Ano11]. Tammy [Ano04k]. tangible [FKN^{+}23]. Tangled [Ker14]. Tanya [Rao04b]. Targets [Sta07s, Wol03c]. Tate [Spi07]. TCP [Cur04, PB14]. teacup [Joh06]. Team [Lim22, Mat19b, Bee06, DGK^{+}23, Lim23b, Mat23a, Mat24, Ano03d]. Teams [Lim22, Mat23a, OO03, SJK20, SP21]. Teapot [BM13]. Tear [JS18]. Tech [Ano03d, Ano03e, BY22]. Technical [All12, Hid04, Koh04, VD24]. technique [NA22]. Techniques [Bue14, Lim20b, Sta05c]. Technologies} [Cha05b, GPPV24, VTA23, Zhu04]. Technology [DMBB17, Goo07, Mat17c, Por04b, BM22, Cre06b, Fra20a, Kap23, McT04, PZ06, Var21]. Ted [Sta06n]. TEEs [DLdFV^{+}23]. Telemanagers [NFG04]. Ten [Lim19c, Lim22]. terabyte [Cre06c]. Teresa [Bro04a]. Terror [NN17c]. Terry [GDWC20, Mau05e]. Test [NN18f, AR23, DSBS23, Don05, Sto05]. Testing [AT17, AR23, BLK15, BNB^{+}05, CDH14, DSBS23, Mad15, Pos04, vD15, Bla04, Fel05b, Ing07]. text [McC05]. Textual [Vid23]. Thee [Sta04b]. Theft [Cre10e, Cre10f]. Their [BY22, Bos06, Vac20]. Themselves [Lim15b]. Theo [Ano20a]. Theory [Boe03, Nag04, GT03, KGM21]. There [Ano03g, Ano20a, Chi14b, FSM^{+}21, JSN17, KB08e, Mic04, NN21e, NN23c, NN23d, She15, Sta04b, Sta06i, Wol04b, NN23e]. Thermonuclear [Fra20c]. Thing [Chi14b, Mic04, Hel21a, NN07e]. thingamabobs [Fra20e]. Things [Lim19c, Mat16d, Mat18c, Vas17, KB08b, JSN17]. Think [Blo04, BNB^{+}05, Kam23, VM04, FSM^{+}21, Lev23]. Thinking [Leh03, Mil10, O’H07b]. Thinner [Ano13]. Thomas [Eng05, Koh06]. Thou [Ano03e, Dec04, LCW18]. Thread [MNN04]. Threading [Sta06a, Tit05]. Threads [Gus05]. Threat [WRS^{+}09, LWD07, Hen08]. Threats [Cre10e]. Three [ATM22, Ano04l, KB09a]. Three-part [ATM22]. Throttling [Ano04k]. Thumbs [Sta06d]. Tickets [NN19e]. Ties [Wol04g]. Tim [Bel05b, Gra05, McK05]. Time [Ano20a, Bel03, Bel04b, Col19e, DEE^{+}16, Fra21b, Hel19b, Kam11b, Lim20c, NN15f, NN19d, Sch16, SMM^{+}04, Ste15, AhCD22, BW07, Kam11c, Kam21a, NN22b, NN23e, Wol08]. Time-series [Ano20a]. Times [Kam11a, SMM^{+}04, Var21, Wol04c]. Timing [RV10]. Tingley [SNW^{+}23]. Tippett [Mau05d]. Titus [LSB17]. TiVo [Bar06]. TiVo-lution [Bar06]. TLS [HD22]. Today [BNNC03, Lev08]. Together [Ern03]. Tolerance [Kel24]. Tom [Hor03, Bel07b, Kil20, Lim19c]. Too [CSL^{+}06, HWH17, Kam21b, NN14e, Sto05, WHT^{+}05, Arn05, Lim24]. tool [TDA21]. Toolkit [PP03b, SMM^{+}04, Vin03, Wol03a, Wol03b, Wol03c, Wol04d, Wol04c, Wol04e, Wol04f, Wol04g, Wol04a, Hor03]. Tools [DM15, EGBA03, GS23, Kuc06, Nad21, See03, TDA21, Wol03b, Wol04d, Wol04e, BFZ^{+}22, Coa05, Vir05]. Top [Lim19c, Mat17f]. Torch [Sta06c]. Touch [NN23c]. Tough [MM04, Smi09]. Tour [AF12, HBO10a, HBO10b]. toy [CBM^{+}08]. Traces [Phi10b]. Tracing [AGB17, ESTT22]. Tracking [ERR18]. Traction [Sta06d, Wea03]. Tradeoffs [Pop24]. Trading [Kel20d]. Trahan [Gro04o]. Train [Ano14b]. Training [Ano03d, NN23c, Sta05h]. Trains [Lim23c]. Traipsing [Coa05]. transaction [AR23]. Transactional [CBM^{+}08, Dre08b]. Transactions [Hel16a, ZAB16]. Transformation [KM09]. Transforming [Bur03]. Translation [GHM15]. Translational [KM09]. Transparency [Lau14]. Transparent [NN22f]. Treadmill [Ree19]. Trends [LS19, Sch05]. Trials [Gat04]. tribulations [Gat04]. Tribute [All08]. tried [Hel21c]. tried-and-true [Hel21c]. Trinity [NN16e]. Trip [Chr09]. Triple [Lev09]. Triple-Parity [Lev09]. Trips [Lim24]. Trojan [Sta04a]. Trojan-Spyware [Sta04a]. Troubling [LS19]. true [Hel21c]. Trust [DLdFV^{+}23, GTNNM21, LVDM22, Wad14, Wan24, BM22, Vac21]. Trustworthy [GPPV24, DLdFV^{+}23]. Truth [NN03b]. Truths [Kro17]. try [Chi23]. Trying [NN16f]. Tsunami [BG11]. Tsutsui [Nag04]. Tulach [Mil08]. Tunnel [Sta05i]. Turn [HOSC16]. Turner [Men04]. Turns [PRM09]. Two [Cre09a, KB09a, KB11, Lim22, Lim24, Mat15a, Cre08]. Typecast [NN09c].

U.S. [Ano03d]. Ubiquitous [Ber03a]. ugly [Ele07]. UIs [TDA21]. Ullman [Kau03]. Ultron [Lim15a]. Umbrella [Col19a]. UML [Bel04a, Bel05a, SMM^{+}04, Wie03]. Uncertain [KB10]. Uncrackable [Ano03i]. Undefined [KGM21]. Undergraduate [LNV14]. Underlying [SZ17]. Underserved [DMBB17]. Understand [Goo07, PZ06]. Understanding [Dad05, GAM18, NN06i, Soh07]. understood [JC23]. Unexpected [Kuz19].

Unholy [NN16e]. Unicode [Ull21]. Unified [Ste07a]. Unifying [DFZ^{+}09]. Uninitialized [Sea16]. Uninitiated [Gho11]. Universe [Ano14c, Wal18, O’H06n]. Universities [Sta07s]. University [Cha05b, Kil04, Wol08]. unknowable [KB05b]. Unleashed [NN05h]. Unlocking [ATKS07]. Unplugged [Ano04m, NN21d]. Unsafe [Kam22b, Har05a]. unscripted [NN05d]. Unstructured [O’H05e, McC05]. Untangling [Ric06]. until [Ano12b, NN24b]. Untrusted [Coo23]. update [Bro05]. updates [Bro05]. Updating [GS23]. upgrade [Kel21a]. Upload [Dat16]. upon [Sta05g]. Uprooting [HPE03]. Usability [Ing07]. Use [JC23, Mei14c, BLA21, NN21b, RCF^{+}21, Vac23a, JSK16, KB07g]. Use-Case [JSK16]. Used [Hel22b, KB07d]. User [Liu14, Ras03b, DGK^{+}23, GAI^{+}21, McT04]. users [DLdFV^{+}23, Vac20]. Using [DM15, Gil14, Lam18, Lim20a, Mei17, Coo23, GPPV24, Vid21]. UX [CG10].

V [Puz04, Chi23]. v2 [NG22]. Vaidyanathan [Gro04o]. Value [Kel21a, NG22, NN23e]. VanDyk [Lec07]. vanishing [LHML21]. Vasa} [NN09d]. Vaugh [Par07]. VB [Cor07, Lho06]. Vehicles [GTNNM21]. Vel [Gum04a]. Vellum [Lon15]. Velocity [Kil19]. Vendors [Lim10]. Venture [DLP19]. Verification [KW17, McC15, Mub18, GTK24]. Verkest [Pet04]. Verlag [Bha07b, Bha07a, Ekt05, Gro04p, Hen08, Nag04, Por04b, Whi08, Zen04]. Version [NNB15]. Versions [Mat15a, NN21a]. Versus [Hel19c]. via [FKB^{+}03]. Vicious [NN05i, NN07a, NN09c, NN09a, NN09b, NN10e, NN10b, NN10c, NN10d, NN10a, NN13b, NN13c, NN13d, NN13f, NN13a, NN13e, NN14a, NN14c, NN14b, NN15b, NN15c, NN15d, NN15e, NN05b, NN06a, NN06h, NN06i, NN07d, NN07c, NN07g, NN07h, NN07b, NN07e, NN07i, NN08b, NN08c, NN08d, NN08e, NN08f, NN08g, NN09d, KTS^{+}04, NN04b, NN04d, NN04c, NN05a, NN05g, NN05c, NN05e, NN05d, NN05f, NN05h, NN05j, NN05i, NN06b, NN06c, NN06d, NN06e, NN06f, NN06g]. Victorian [Kel21b]. Video [Sta06j, BLA21, Ger05]. Videos [Get12]. Viégas [Ano10]. Viewing [AM03]. Views [Mat17f]. Vigilantism [Sta06c, Mau05f]. Vigorous [RB17]. Vijay [Nau08a]. Virkus [Kuc06]. virts [Kil08b]. Virtual [FAHB18, PSBG11, Ros04, Sta05g, Sup04, Wal08]. Virtualization [Cre09b, Cre10c, Kot10, Sta07s, CB07, Dre08a, Rix08, Rou07]. Virtually [Gro04l]. Virtue [NN08g]. Visual [Par07, VB07]. Visualization [HBO10a, HBO10b]. Visualizing [Gre10]. Vladan [Put07]. Vladimirov [Puz04]. VM [Kap23]. Vogels [Kil20, O’H06v]. VoIP [Sta06n, AE04, Ano03f, SF04, SL04, Wal05]. volatile [NSWW15]. Vote [GBG^{+}04, KB04d]. Voting [Ras04b, Sim04]. Voyage [Kru07]. VPN [Cra20]. vs [CL11, For07, Hel20b, Leh03, NN05j, NN18e, NN20b, O’H06k, See03, Woo04]. Vulnerable [Sta05e].

W [O’H06q, SBS^{+}23, Tho05]. Wadkins [Sta07j]. Wainwright [Jaf06]. Waldo [GDWC20]. Walk [BY22]. Wallingford [Sta06n]. Walls [Wol06]. Walmart [Ham14]. Wang [Hen08, Las06]. Want [KTS^{+}04, Ras04a]. Wanted [Sta05d]. Wanton [NN11e]. Wants [KTS^{+}04]. WAPI [Ano04i]. War [Kam11b]. ware [Sta05h]. Warm [Sta05e]. Warrior} [Mar04b, PC04]. Was [Kam14a, Ras04a, BLA21]. Watch [Ano03d]. Watchdogs [NN18e]. Waterfall [LN04, MSNL04]. Watson [NN12b]. Watt [Mac03a]. Wattenberg [Ano10]. Way [Ano20a, MK14, BLA21, Kam21a, Mat24, Bel08]. Wayne [Bro03]. Ways [FPWM09, Lim17a, Mat17a]. Weasel [Sta07a]. Web [Bal09, BLA21, Dwy08, Paq06, Ram08, Sta07u, Wal07, Eng05, Ano03i, Ano12b, AHMT03, Bos05, Bur03, CL11, Ebe05, GBM15, Ing07, JWC08, Ker14, Kil08a, KR03, Kum05, Mat13, NN07f, Pas04, Sav15, SSS06, Ses04, Sou08, Sta04b, Sta05i, Sta06e, TR03, Vin03, YRZB16, Zel07, vD15]. Weber [Sch04b]. WebFountain [Ano03i]. WebRTC [BLA21]. Website [Hin08, HP08]. WEBVACs [Wan15]. Week [Mat17e]. Wellings [Wol08]. Wells [Par06, Kir04]. Wenke [Hen08]. Went [Kam21c]. Were [GRA^{+}14, Leh03]. Werner [Kil20, O’H06v]. Wesley [Bel08, Cha05a, Edw05, Hac06, Hin08, Hor03, Jov06, Kar05, Llo06, Mar05a, Men04, Par07, Pos04, Spi07, Sta06l, Sta06k, Tom06]. Westgate [Lec07]. Wheel [Kam13a]. Where [Ano04i, Ano20a, BLK15, Dec04, NN08b, Fra20c]. whether [CB23]. which [For07]. While [NH20, RCF^{+}21]. whining [Pat05]. White [Ram08]. Whither [NN09e]. Whitney [Can09]. whiz [Bel06]. Who [ATM22, EMC19, Wad14, NN16a]. Whole [NG22]. Whole-program [NG22]. Wi [Puz04, VGM04, Ano03f, Ano04i, Coo23]. WI-FI [Ano03f, Ano04i, Coo23]. Wi-Foo [Puz04, VGM04]. Widgets [Sta05b]. Wieringa [Ber03c]. Wiggins [SNW^{+}23]. Wiki [Paq06, Ebe05]. Wikibooks [Sta05c]. Wikipedia [LPS24]. Wiley [And04, Bel05b, Ber09, Las06]. Wiley-Interscience [Las06]. Will [BG11, Dur10, Hum17, KMNNM22, Kam11b, Kro17, AhCD22, MR21, DiB04, KB07a]. William [Par07]. wind [Cra20]. Windows [Koh06, Nau08b, Par07, Tom06, JR05, Mac08, MB05, Sta05d, Sta07q, Wol04a]. Windows-to-Linux [Wol04a]. WinFS [KTS^{+}04]. Wing [SNW^{+}23]. winter [VTA23]. Wire [Sta07p]. wired [Cre07a].

Wireless [All03b, Ber09, Cha05b, Kuc06, Lab08, Puz04, Sta07a, Vir05, VGM04, Zen05, Zhu04]. Without [Chr09, KB04e, NN03c, Gus05]. WLAN [Rit03]. Woes [Ras03c]. Wolfe [SMM^{+}04]. Won’t [Ano12b, Hal05]. Woodstock [Wol03a]. Words [SM05a]. Work [Gra19, Hum17, Lim18b, Lim20b]. Workers [Ano03d, BY22, Mat23a]. Workflow [De 06]. Working [Ber05, Fea05, Rus23]. Workload [Mub18]. Works [McM15, Lim24]. World [Ano04h, BCJ^{+}18, Don03, Fra20e, Gro03a, GBG^{+}04, HD22, KP23, Leh03, NN23b, Sch17a, Sit18, Sta05g, Kil04, CB08, Hal07, Har05a, Jen02, Kel22b, Sta07v, Ull21]. Worlds [Wal08]. Worm [Bla14]. Worst [NN19d]. Would [Coa03c, KB04e, Mat16d, Kam21b]. WPF [Nau08b, Mac08]. Write [Hel19c, See04]. Writing [HOSC16, NN18f, Pat04, Spi03]. Wrong [CJWG11, Kam10, Kam21c, Lim16b].

X [Bha07a, Fox08, Jaf06, Kil04, Sta05j, Sta06m, Tom05]. Xen [Fox08, Chi08]. XML [Cha05b, Bur03, Hel17b, SM05b, TR03, WG08, Zhu04]. XP} [Mer04, SR03]. XSS [NN05i].

Y2K [Sta07q]. Year [Sta07q, Cra20, Lim24]. Years [Gra08a]. Yoo [Pet04]. York [Bha07b, Hen08, Whi08]. Yoshiteru [Gro04p]. Yourdan [Wie03]. You’re [Ano04l, Kam10, CB23]. Yours [Gro04l]. YouTube [LPS24].

zealotry [NN12d]. Zealots [ERF^{+}03]. Zeldman [Wal07]. Zeller [Ber07]. Zero [BM22, Kel24, LVDM22]. Zetlin [Goo07]. Zhuk [Cha05b]. Zones [SMM^{+}04]. Zoo [HBO10a, HBO10b].

References

Alagappan:2022:CCK

[AA22] Ram Alagappan and Peter Alvaro. Crash consistency: Keeping data safe in the presence of crashes is a fundamental problem. ACM Queue: Tomorrow’s Computing Today, 20 (4):107–115, July 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                                    tronic). URL https://dl.acm.
                                                    org/doi/10.1145/3561654.

Arnbak:2014:SCH

[AAVV14] Axel Arnbak, Hadi Asghari, Michel Van Eeten, and Nico Van Eijk. Security collapse in the HTTPS market. ACM Queue: Tomorrow’s Computing Today, 12(8):30, August 2014. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Angiuli:2015:HIY

[ABW15] Olivia Angiuli, Joe Blitzstein, and Jim Waldo. How to de-identify your data. ACM Queue: Tomorrow’s Computing Today, 13(8):20, September/October 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                                    URL http://queue.acm.org/
                                                    detail.cfm?id=2838930.

Ahuja:2004:VWI

[AE04] Sudhir R. Ahuja and Robert Ensor. VoIP: What is it good for? ACM Queue: Tomorrow’s Computing Today, 2(6):

48–55, September 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Abts:2012:GTT

[AF12] Dennis Abts and Bob Felderman. A guided tour through data-center networking. ACM Queue: Tomorrow’s Computing Today, 10(5):10, May 2012. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Alvaro:2017:RPT

[AGB17] Peter Alvaro, Sumit Galwani, and Peter Bailis. Research for practice: Tracing and debugging distributed systems; programming by examples. ACM Queue: Tomorrow’s Computing Today, 15(1):60, January 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3074451.

Agnew:2022:FRF

[AHC22] James Agnew, Pat Helland, and Adam Cole. FHIR: Reducing friction in the exchange of healthcare data: a discussion with James Agnew, Pat Helland, and Adam Cole. ACM Queue: Tomorrow’s Computing Today, 20(2):67–88, March 2022. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3534861.

Auer:2022:MPL

[AhCD22] Raphael Auer, Rainer BÃ¶hme, Jeremy Clark, and Didem Demirag. Mapping the privacy landscape for central bank digital currencies: Now is the time to shape what future payment flows will reveal about you. ACM Queue: Tomorrow’s Computing Today, 20 (4):16–38, July 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3561796.

Arsanjani:2003:WSP

[AHMT03] Ali Arsanjani, Brent Hailpern, Joanne Martin, and Peri Tarr. Web services: Promises and compromises. ACM Queue: Tomorrow’s Computing Today, 1 (1):48–58, March 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Arthorne:2004:OEF

[AL04] John Arthorne and Chris Laffra. Official Eclipse 3.0 FAQs. Eclipse series. Addison-Wes-ley, Reading, MA, USA, 2004. ISBN 0-321-26838-5 (paperback). xxxiv + 386 pp. LCCN QA76.76.D47 A776 2004.

Alappatt:2017:NAI

[Ala17] Antony Alappatt. Network applications are interactive. ACM Queue: Tomorrow’s Computing Today, 15(4):60, July 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3145628.

Allman:2003:CES

[All03a] Eric Allman. Curmudgeon:

The economics of spam. ACM Queue: Tomorrow’s Computing Today, 1(9):80, 78–79, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Allman:2003:CWN

[All03b] Eric Allman. Curmudgeon:

Wireless networking considered flaky. ACM Queue: Tomorrow’s Computing Today, 1(7): 96, 95, October 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Allman:2003:CCD

[All03c] Eric Allman. Interview: a conversation with Chris DiBona. ACM Queue: Tomorrow’s Computing Today, 1(5):10, 12–19, July/August 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Allman:2003:ICP

[All03d] Eric Allman. Interview: a conversation with Peter Ford. ACM Queue: Tomorrow’s Computing Today, 1(8):18, 20–27, November 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Allman:2003:SSS

[All03e] Eric Allman. Spam, spam, spam, spam, spam, the FTC, and spam. ACM Queue: Tomorrow’s Computing Today, 1 (6):62–69, September/October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Allman:2004:IJG

[All04] Eric Allman. Interview: a conversation with James Gosling. ACM Queue: Tomorrow’s Computing Today, 2(5):24–33, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Allman:2006:CC

[All06a] Eric Allman. Complying with compliance. ACM Queue: Tomorrow’s Computing Today, 4 (7):18–21, September 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Allman:2006:MAW

[All06b] Eric Allman. E-mail authentication: what, why, how? ACM Queue: Tomorrow’s Computing Today, 4(9):30–34, November 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Allman:2008:TJG

[All08] Eric Allman. A tribute to Jim Gray. ACM Queue: Tomorrow’s Computing Today, 6(3):

56–57, May/June 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Allman:2012:MTD

[All12] Eric Allman. Managing technical debt. ACM Queue: Tomorrow’s Computing Today, 10 (3):10, March 2012. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Alonso:2018:FDC

[Alo18] Gustavo Alonso. FPGAs in data centers. ACM Queue: Tomorrow’s Computing Today, 16 (2):60, March 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3231573.

Allman:2003:EVO

[AM03] Eric Allman and Marshall Kirk McKusick. From the Editors: Viewing open source with an open mind. ACM Queue: Tomorrow’s Computing Today, 1 (5):6–7, July/August 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Allman:2004:EOS

[AM04] Eric Allman and Marshall Kirk McKusick. From the Editors: Open source revisited. ACM Queue: Tomorrow’s Computing Today, 2(3):8–9, May 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anderson:2003:GPM

[And03] Dave Anderson. You don’t know jack about disks. ACM Queue: Tomorrow’s Computing Today, 1(4):20–30, June 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anderson:2004:BRR

[And04] Dennis Anderson. Book review: RFID Handbook: Fundamentals and Applications in Contactless Smart Cards and Identification by Klaus Finkenzeller, John Wiley & Sons, 2003, $125, ISBN 0-470-84402-7. ACM Queue: Tomorrow’s Computing Today, 2(3):74, May 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2003:Ca

[Ano03a] Anonymous. Calendar. ACM Queue: Tomorrow’s Computing Today, 1(7):92, October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2003:Cb

[Ano03b] Anonymous. Calendar. ACM Queue: Tomorrow’s Computing Today, 1(8):74, November 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2003:Cc

[Ano03c] Anonymous. Calendar. ACM Queue: Tomorrow’s Computing Today, 1(9):77, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2003:LDH

[Ano03d] Anonymous. Latest developer headlines: Watch out for Bluetooth sniffers; the Samba Team announces Samba 3.0.0 RC1; Aussie government to see Microsoft source code; U.S. tech workers training replacements; ‘Apache Geronimo’ looks to J2EE implementation. ACM Queue: Tomorrow’s Computing Today, 1(??):??, ???? 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2003:NCT

[Ano03e] Anonymous. News 2.0: China’s tech blast-off; thou shalt not print; dust dreams. ACM Queue: Tomorrow’s Computing Today, 1(9):8, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2003:NGG

[Ano03f] Anonymous. News 2.0: Gates to get ATM free?; RFIDs reverse roles; Moore’s Law of Cyber-Crime: Say it isn’t so; WI-FI + VoIP = my new phone? ACM Queue: Tomorrow’s Computing Today, 1(8):8, November 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2003:NOO

[Ano03g] Anonymous. News 2.0: Outfoxing outsourcing; Australia pushes for less spam; there’s a bug in my lunchbox. ACM Queue: Tomorrow’s Computing Today, 1(6):8, September/October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2003:NSC

[Ano03h] Anonymous. News 2.0: Sugar cubes rock; can a personal server really save your day?; hot chips run cool. ACM Queue: Tomorrow’s Computing Today, 1 (7):6, October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2003:NUP

[Ano03i] Anonymous. News 2.0: Uncrackable passwords; WebFountain drinks down the Web; embracing open source in India. ACM Queue: Tomorrow’s Computing Today, 1(5): 8, July/August 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:Ca

[Ano04a] Anonymous. Calendar. ACM Queue: Tomorrow’s Computing Today, 1(10):90, February 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:Cb

[Ano04b] Anonymous. Calendar. ACM Queue: Tomorrow’s Computing Today, 2(1):94, March 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:Cd

[Ano04c] Anonymous. Calendar. ACM Queue: Tomorrow’s Computing Today, 2(2):82, April 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:Ce

[Ano04d] Anonymous. Calendar. ACM Queue: Tomorrow’s Computing Today, 2(3):76, May 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:Cf

[Ano04e] Anonymous. Calendar. ACM Queue: Tomorrow’s Computing Today, 2(4):78, June 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:Cg

[Ano04f] Anonymous. Calendar. ACM Queue: Tomorrow’s Computing Today, 2(5):78, July 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:Ch

[Ano04g] Anonymous. Calendar. ACM Queue: Tomorrow’s Computing Today, 2(6):78, September 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:NBB

[Ano04h] Anonymous. News 2.0: Big book to better the world; A cardboard computer?; obje— that’s French for interoperability. ACM Queue: Tomorrow’s Computing Today, 2(3):11, May 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:NBF

[Ano04i] Anonymous. News 2.0: Boba Fett on the Microsoft payroll?; Wi-Fi goes WAPI; photons and molecules: a new partnership; SETI where are you, Mersenne prime? ACM Queue: Tomorrow’s Computing Today, 1 (10):8, February 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:NCF

[Ano04j] Anonymous. News 2.0: Communications from a field near you; light-speed disk drives; handdelivered curried spam?; mind control for the masses? ACM Queue: Tomorrow’s Computing Today, 2(4):12, June 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:Na

[Ano04k] Anonymous. News 2.0: Leave it to the Irish; the Tammy Faye scam; throttling back network attacks. ACM Queue: Tomorrow’s Computing Today, 2 (5):12, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:NOS

[Ano04l] Anonymous. News 2.0: Office suites, meet NBOR; security comes first, eventually; three strikes and you’re out; LOTR SneakerNet iPod. ACM Queue: Tomorrow’s Computing Today, 2(2):9, April 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:NSS

[Ano04m] Anonymous. News 2.0: Smoke signals, cellphone style; file sharing: Crime and punishment unplugged; musican composition by Any Body. ACM Queue: Tomorrow’s Computing Today, 2 (6):11, September 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:NFH

[Ano04n] Anonymous. News 2.0: The farmer and the hacker; milk and algorithms make for strong bones; keep your distance. ACM Queue: Tomorrow’s Computing Today, 2(1):8, March 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:WYHa

[Ano04o] Anonymous. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 2(5):22, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2004:WYHb

[Ano04p] Anonymous. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 2 (6):20, September 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2009:C

[Ano09] Anonymous. Calendar. ACM Queue: Tomorrow’s Computing Today, 7(2):76, February 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2010:CJH

[Ano10] Anonymous. A conversation with Jeff Heer, Martin Wattenberg, and Fernanda Viégas. ACM Queue: Tomorrow’s Computing Today, 8(3):1–8, March 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2011:JGT

[Ano11] Anonymous. Jim Gray — A talk with THE SQL Guru and architect. ACM Queue: Tomorrow’s Computing Today, 9 (4):40, April 2011. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2012:NSI

[Ano12a] Anonymous. Node summit: The importance of cross platform. ACM Queue: Tomorrow’s Computing Today, 10(12):??, December 2012. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2012:WWS

[Ano12b] Anonymous. The Web won’t be safe or secure until we break it. ACM Queue: Tomorrow’s Computing Today, 10(11): ??, November 2012. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2013:NLP

[Ano13] Anonymous. Node at LinkedIn:

The pursuit of thinner, lighter, faster. ACM Queue: Tomorrow’s Computing Today, 11(12): ??, December 2013. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2567673. A

discussion with Kiran Prasad, Kelly Norton, and Terry Coatta.

Anonymous:2014:RKS

[Ano14a]  Anonymous.  Release the

kraken: a story of Node.js in the enterprise (PayPal). ACM Queue: Tomorrow’s Computing Today, 12(2):80, February 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2014:ROT

[Ano14b] Anonymous. Riding the n(ode) train (Groupon). ACM Queue: Tomorrow’s Computing Today, 12(2):90, February 2014. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2014:NJD

[Ano14c] Anonymous. Node.js in the digital media universe. ACM Queue: Tomorrow’s Computing Today, 12(2):110, February 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2020:ATS

[Ano20a] Anonymous. Always-on timeseries database: Keeping up where there’s no way to catch up: a discussion with Theo Schlossnagle, Justin Sheehy, and Chris McCubbin. ACM Queue: Tomorrow’s Computing Today, 18(6):115–135, December 2020. URL https:

//dl.acm.org/doi/10.1145/
3442632.3442634.

Anonymous:2020:DS

[Ano20b] Anonymous. Demystifying Stablecoins. ACM Queue: Tomorrow’s Computing Today, 18 (1):??, January 2020. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3387945.3388781.

Anonymous:2021:PAC

[Ano21]  Anonymous.  Patent absur-

dity: a case when ignorance is the best policy. ACM Queue: Tomorrow’s Computing Today, 19(4):19–22, July 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487019.3489047.

Alvaro:2023:ATD

[AR23] Peter Alvaro and Manuel Rigger. Automatically testing database systems: DBMS testing with test oracles, transaction history, and fuzzing. ACM Queue: Tomorrow’s Computing Today, 21(6):128–135, November 2023. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3639449.

Agarwal:2007:PF

[ARB07] Naveen Agarwal, Scott Renfro, and Arturo Bejar. Phishing forbidden. ACM Queue: Tomorrow’s Computing Today, 5 (5):28–32, July 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Arnold:2005:PPT

[Arn05] Ken Arnold. Programmers are people, too. ACM Queue: Tomorrow’s Computing Today, 3 (5):54–59, June 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ambler:2006:RDE

[AS06a] Scott W. Ambler and Pramod J. Sadalage. Refactoring databases: evolutionary database design. The Addison-Wesley signature series. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-29353-3 (hardback). xxvii + 350 pp. LCCN QA76.9.D26 A52 2006. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip063/2005031959.html.

Anthias:2006:NNR

[AS06b] Taf Anthias and Krishna Sankar. The network’s new role. ACM Queue: Tomorrow’s Computing Today, 4(4):38–46, May 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Asleson:2006:FA

[AS06c] Ryan Asleson and Nathaniel T. Schutta. Foundations of Ajax. Apress, Berkeley, CA, USA, 2006. ISBN 1-59059-582-3. xxi + 273 pp. LCCN ????

                                        URL http://www.apress.com/
                                        book/bookDisplay.html?bID=
                                        10042.

Ascher:2004:RYF

[Asc04] David Ascher. Is OS right for you? (A fictional case study). ACM Queue: Tomorrow’s Computing Today, 2(3): 32–38, May 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Alvaro:2017:AGF

[AT17] Peter Alvaro and Severine Tymon. Abstracting the geniuses away from failure testing. ACM Queue: Tomorrow’s Computing Today, 15(5):10, September/October 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3155114.

Adl-Tabatabai:2006:UC

[ATKS07] Ali-Reza Adl-Tabatabai, Christos Kozyrakis, and Bratin Saha. Unlocking concurrency. ACM Queue: Tomorrow’s Computing Today, 4(10):24–33, December/January 2006–2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Aldrich:2022:TPH

[ATM22] Guenever Aldrich, Danny Tsang, and Jason McKenney. Threepart harmony for program managers who just don’t get it, yet: Open-source software, open standards, and agile software development. ACM Queue: Tomorrow’s Computing Today, 20(6):58–79, 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3576027.

Bacon:2007:RGC

[Bac07] David F. Bacon. Realtime garbage collection. ACM Queue: Tomorrow’s Computing Today, 5(1):40–49, February 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ballance:2003:BRJ

[Bal03] R. Ballance. Book review: Java Precisely, by Peter Seetoft, MIT Press, 2002, $14.95, ISBN 0-262-69276-7. ACM Queue: Tomorrow’s Computing Today, 1(6): 70, September/October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Balaraman:2007:RBP

[Bal07a] S. Balaraman. Review of ‘Beginning POJOs: From Novice to Professional’ by Brian Sam-Bodden, Apress, 2006, $39.99, ISBN 1-59059-596-3. ACM Queue: Tomorrow’s Computing Today, 5(6):45, September 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [SB06].

Balaraman:2007:RPJ

[Bal07b]  S. Balaraman.  Review of

‘Pro JSF and Ajax: Building Rich Internet Components’ by Jonas Jacobi and John Fallows, Apress, 2006, $49.99, ISBN 1-59059-580-7. ACM Queue: Tomorrow’s Computing Today, 5 (3):44, April 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [JF06].

Balaraman:2009:BRR

[Bal09] S. Balaraman. Book review:

Review of Pro Web 2.0 Application Development with GWT by Jeff Dwyer, Apress, 2008, $44.99, ISBN 978-1-59059-985-3. ACM Queue: Tomorrow’s Computing Today, 7(1):40, January 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Dwy08].

Barton:2003:SRL

[Bar03] Jim Barton. From server room to living room. ACM Queue: Tomorrow’s Computing Today, 1(5):20–28, 31–32, July/August 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Barroso:2005:PP

[Bar05]  Luiz André Barroso.  The

price of performance. ACM Queue: Tomorrow’s Computing Today, 3(7):48–53, September 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Barton:2006:TL

[Bar06] Jim Barton. TiVo-lution. ACM Queue: Tomorrow’s Computing Today, 4(3):28–35, April 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bates:2004:CPF

[Bat04a] Rodney Bates. Curmudgeon:

Programming in Franglais.

ACM Queue: Tomorrow’s Computing Today, 2(8):80, 78, November 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bates:2004:OBO

[Bat04b] Rodney Bates. Opinion: Buffer overrun madness. ACM Queue: Tomorrow’s Computing Today, 2(3):12, 14, May 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bates:2004:OSC

[Bat04c]  Rodney Bates.  Opinion:

Schizoid classes. ACM Queue: Tomorrow’s Computing Today, 2(6):12–15, September 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bates:2005:SH

[Bat05] Rodney Bates. Curmudgeon: Syntactic heroin. ACM Queue: Tomorrow’s Computing Today, 3(5):64, 62–63, June 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bartram:2003:DPC

[BB03] Lyn Bartram and Michael Blackstock. Designing portable collaborative networks. ACM Queue: Tomorrow’s Computing Today, 1(3):40–49, May 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Blum:2018:ADP

[BB18] Raymond Blum and Betsy Beyer. Achieving digital permanence. ACM Queue: Tomorrow’s Computing Today, 16(6): 50, November 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3311889.

Barr:2006:AGB

[BC06] Jeff Barr and Luis Felipe Cabrera. AI gets a brain. ACM Queue: Tomorrow’s Computing Today, 4(4):24–29, May 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bennett:2018:HLP

[BCJ^{+}18] Rich Bennett, Craig Callahan, Stacy Jones, Matt Levine, Merrill Miller, and Andy Ozment. How to live in a post-Meltdown and -Spectre world. ACM Queue: Tomorrow’s Computing Today, 16(4):30, July 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3281471.

Beam:2003:CN

[BD03] Michael Beam and James Duncan Davidson. Cocoa in a nutshell. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00462-1. xvii + 545 pp. LCCN QA76.64 .B398 2003.

Beevers:2006:HPT

[Bee06]  Philip Beevers.  A high-

performance team. ACM Queue: Tomorrow’s Computing Today, 4(1):38–43, February 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bejtlich:2006:EDS

[Bej06] Richard Bejtlich. Extrusion detection: security monitoring for internal intrusions. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-34996-2 (paperback). xxviii + 385 pp. LCCN TK5105.59 .B43 2006.

Bell:2003:SSK

[Bel03] Gordon Bell. Sink or swim: Know when it’s time to bail. ACM Queue: Tomorrow’s Computing Today, 1(9):60–67, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bell:2004:DUF

[Bel04a] Alex E. Bell. Death by UML fever. ACM Queue: Tomorrow’s Computing Today, 2(1): 72–80, March 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bell:2004:TPS

[Bel04b] Gordon Bell. A time and a place for standards. ACM Queue: Tomorrow’s Computing Today, 2 (6):66–74, September 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bell:2005:UFD

[Bel05a] Alex E. Bell. UML fever: diagnosis and recovery. ACM Queue: Tomorrow’s Computing Today, 3(2):48–56, March 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bellin:2005:RPM

[Bel05b] David Bellin. Review of PHP5 and MySQL Bible by Tim Converse and Joyce Park, Clark Morgan Wiley, 2004, $49.99, ISBN 0-7645-5746-7. ACM Queue: Tomorrow’s Computing Today, 3(2):58, March 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bell:2006:CSD

[Bel06] Alex E. Bell. Curmudgeon:

Software development amidst the whiz of silver bullets. . .. ACM Queue: Tomorrow’s Computing Today, 4(5):64, 62–63, June 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bell:2007:CDS

[Bel07a] Alex Bell. Curmudgeon: DOA with SOA. ACM Queue: Tomorrow’s Computing Today, 5 (1):56–ff, February 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bell:2007:GCA

[Bel07b] Alex Bell. Opinion: Ground control to Architect Tom . . ..

ACM Queue: Tomorrow’s Computing Today, 5(6):8–10, September 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bellin:2008:RRW

[Bel08] David Bellin. Review of The Rails Way by Obie Fernandez, Addison-Wesley Professional, 2007, $49.99, ISBN: 978-0-321-44561-2. ACM Queue: Tomorrow’s Computing Today, 6(7): 41, November/December 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bell:2013:SI

[Bel13] Alex E. Bell. The software inferno. ACM Queue: Tomorrow’s Computing Today, 11 (11):10, November 2013. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Berger:2003:OSP

[Ber03a] Robert J. Berger. Open spectrum: a path to ubiquitous connectivity. ACM Queue: Tomorrow’s Computing Today, 1 (3):60–68, May 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Berzal:2003:BRA

[Ber03b] Fernando Berzal. Book review: Agile Software Development, Principles, Patterns, and Practices by Robert C. Martin, Prentice Hall, 2002, $55.00, ISBN 0-13-597444-5. ACM Queue: Tomorrow’s Computing Today, 1(9):76, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Berzal:2003:BRD

[Ber03c] Fernando Berzal. Book review: Design Methods for Reactive Systems, by R. J. Wieringa, Morgan Kaufmann, 2002, $64.95, ISBN 1-55860-755-2. ACM Queue: Tomorrow’s Computing Today, 1(7): 90–91, October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Berzal:2005:RWE

[Ber05]  Fernando Berzal.  Review

of “Working Effectively with Legacy Code”, by Michael Feathers, Prentice Hall PTR, 2004, $44.99 ISBN 0-13-117705-2. ACM Queue: Tomorrow’s Computing Today, 3(9): 58, November 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Berg:2006:HAD

[Ber06]  Clifford J. Berg.  High-

assurance design: architecting secure and reliable enterprise applications. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-37577-7. xxvii + 660 pp. LCCN QA76.9.S88 B46 2006. URL http:

//www.loc.gov/catdir/toc/
ecip0515/2005018459.html.

Berzal:2007:RWP

[Ber07] Fernando Berzal. Review of ‘Why Programs Fail: a Guide to Systematic Debugging’ by Andreas Zeller, Morgan Kaufmann, 2005, $54.95, ISBN 1-55860-866-4. ACM Queue: Tomorrow’s Computing Today, 5 (1):50, February 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Zel06].

Berni:2009:RWA

[Ber09] Alessandro Berni. Review of Wireless Ad Hoc and Sensor Networks (1st ed) by Houda Labiod, John Wiley & Sons, Inc., 2008, $165.00, ISBN 1-84821-003-5. ACM Queue: Tomorrow’s Computing Today, 7 (1):40, January 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Lab08].

Bailis:2016:RPD

[BFAP16] Peter Bailis, Camille Fournier, Joy Arulraj, and Andy Pavlo. Research for practice: Distributed consensus and implications of NVM on database management systems. ACM Queue: Tomorrow’s Computing Today, 14(3):40, May 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=2967618.

Bird:2022:TFC

[BFZ^{+}22] Christian Bird, Denae Ford, Thomas Zimmermann, Nicole Forsgren, Eirini Kalliamvakou, Travis Lowdermilk, and Idan Gazit. Taking flight with copilot: Early insights and opportunities of AI-powered pairprogramming tools. ACM Queue: Tomorrow’s Computing Today, 20(6):35–57, 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3582083.

Beeler:2006:RP

[BG06] George W. Beeler, Jr. and Dana Gardner. A requirements primer. ACM Queue: Tomorrow’s Computing Today, 4(7): 22–26, September 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Berriman:2011:HWA

[BG11] G. Bruce Berriman and Steven L.

Groom. How will astronomy archives survive the data tsunami? ACM Queue: Tomorrow’s Computing Today, 9 (10):21, October 2011. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Burns:2016:BOK

[BGO^{+}16] Brendan Burns, Brian Grant, David Oppenheimer, Eric Brewer, and John Wilkes. Borg, Omega, and Kubernetes. ACM Queue: Tomorrow’s Computing Today, 14(1):10, January 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=2898444.

Bhattacharya:2007:RESb

[Bha07a] Shantanu Bhattacharya. Review of ‘Enterprise Service-Oriented Architectures: Concepts, Challenges, Recommendations’ by James McGovern, Oliver Sims, Ashish Jain, and Mark Little, Springer-Verlag, 2006, $59.95, ISBN 1-4020-3704-X. ACM Queue: Tomorrow’s Computing Today, 5(5):46, July 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [MSJL06].

Bhattacharya:2007:RESa

[Bha07b] Shantanu Bhattacharya. Review of ‘Essential Software Architecture’ by Ian Gorton, Springer-Verlag New York Inc., Secaucus, NJ, 2006, $59.95, ISBN 3-540-28713-2. ACM Queue: Tomorrow’s Computing Today, 5 (2):56, March 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Gor06].

Birsan:2005:PIE

[Bir05] Dorian Birsan. On plug-ins and extensible architectures. ACM Queue: Tomorrow’s Computing Today, 3(2):40–46, March 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bailis:2014:NR

[BK14] Peter Bailis and Kyle Kingsbury. The network is reliable.

ACM Queue: Tomorrow’s Computing Today, 12(7):20, July 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bonawitz:2021:FLP

[BKMR21] Kallista Bonawitz, Peter Kairouz, Brendan McMahan, and Daniel Ramage. Federated learning and privacy: Building privacypreserving systems for machine learning and data science on decentralized data. ACM Queue: Tomorrow’s Computing Today, 19(5):87–114, September 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3494834.3500240.

Bryan:2007:DS

[BL07] David A. Bryan and Bruce B.

Lowekamp. Decentralizing SIP. ACM Queue: Tomorrow’s Computing Today, 5(2): 34–41, March 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Black:2004:CTP

[Bla04] Rex Black. Critical testing processes: plan, prepare, perform, perfect. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-201-74868-1. xxxvi + 571 pp. LCCN QA76.76.T48 B552 2003.

Bland:2014:FMT

[Bla14] Mike Bland. Finding more than one worm in the Apple. ACM Queue: Tomorrow’s Computing Today, 12(5):10, May 2014. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Blum:2021:WRC

[BLA21] Niklas Blum, Serge Lachapelle, and Harald Alvestrand. WebRTC — realtime communication for the Open Web Platform: What was once a way to bring audio and video to the web has expanded into more use cases we could ever imagine. ACM Queue: Tomorrow’s Computing Today, 19(1):77–93, February 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3454122.3457587.

Binder:2014:MBT

[BLK15] Robert V. Binder, Bruno Legeard, and Anne Kramer. Modelbased testing: Where does it stand? ACM Queue: Tomorrow’s Computing Today, 13(1): 40, December/January 2014– 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2723708.

Blow:2004:GDH

[Blo04] Jonathan Blow. Game development: Harder than you think. ACM Queue: Tomorrow’s Computing Today, 1(10): 28–37, February 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Beckman:2013:STD

[BM13] Brian Beckman and Erik Meijer. The story of the teapot in DHTML. ACM Queue: Tomorrow’s Computing Today, 11 (2):1–10, February 2013. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bush:2022:ZOH

[BM22] Matthew Bush and Atefeh Mashatan. From zero to one hundred: Demystifying zero trust and its implications on enterprise people, process, and technology. ACM Queue: Tomorrow’s Computing Today, 20 (4):80–106, July 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3561799.

Burgess:2005:LQT

[BNB^{+}05] Philip Burgess, Miriam Amos Nihart, Manfred Berndtgen, Dan Griffin, and Gabe da Silveira. Letters: QA ! = testing; more common than you think; Kay-OK. ACM Queue: Tomorrow’s Computing Today, 3(3):10, April 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Brown:2003:EDA

[BNNC03] David J. Brown, George V.

Neville-Neal, and Terry Coatta. From the Editors: The developer’s art today: Aikido or sumo? ACM Queue: Tomorrow’s Computing Today, 1(6):

6–7, September/October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Boekhoudt:2003:BBT

[Boe03] Caspar Boekhoudt. The big bang theory of IDEs. ACM Queue: Tomorrow’s Computing Today, 1(7):74–82, October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Boer:2009:OCP

[Boe09] Benjamin Boer. The Obama campaign: a programmer’s perspective. ACM Queue: Tomorrow’s Computing Today, 7 (1):36–39, January 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Booch:2004:FR

[Boo04] Grady Booch. The fever is real. ACM Queue: Tomorrow’s Computing Today, 2(1): 81, March 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Borkar:2003:GGC

[Bor03] Shekhar Borkar. Getting gigascale chips: Challenges and opportunities in continuing Moore’s Law. ACM Queue: Tomorrow’s Computing Today, 1 (7):26–33, October 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Borriello:2006:IA

[Bor06] Gaetano Borriello. The invisible assistant. ACM Queue: Tomorrow’s Computing Today, 4(6): 44–49, July/August 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bosworth:2005:LW

[Bos05] Adam Bosworth. Learning from the Web. ACM Queue: Tomorrow’s Computing Today, 3 (8):26–32, October 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bostick:2006:BTS

[Bos06] John Bostick. Box their SOXes off. ACM Queue: Tomorrow’s Computing Today, 4(7): 28–29, September 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bourne:2003:EFR

[Bou03] Steve Bourne. From the Editors: Finding the right questions. ACM Queue: Tomorrow’s Computing Today, 1 (1):4, March 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bourne:2004:ICB

[Bou04] Steve Bourne. Interview: a conversation with Bruce Lindsay. ACM Queue: Tomorrow’s Computing Today, 2(8):

22–33, November 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Bourne:2008:CRS

[Bou08] Steve Bourne. CTO roundtable STORAGE. ACM Queue: Tomorrow’s Computing Today, 6 (6):38–45, October 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Boyd:2008:DPC

[Boy08] Chas. Boyd. Data-parallel computing. ACM Queue: Tomorrow’s Computing Today, 6 (2):30–39, March 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Beeler:2009:OAO

[BP09] Charles Beeler and Craig Partridge. Opinion: All-optical computing and all-optical networks are dead. ACM Queue: Tomorrow’s Computing Today, 7(3):10–11, April 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Baquero:2016:WLC

[BP16] Carlos Baquero and Nuno Preguiça. Why logical clocks are easy. ACM Queue: Tomorrow’s Computing Today, 14 (1):60, January 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2917756.

Berlin:2006:PS

[BR06] Daniel J. Berlin and Garrett Rooney. Practical subversion. The experts voice in open source. Apress, Berkeley, CA, USA, second edition, 2006. ISBN 1-59059-753-2. xix + 304 pp. LCCN ????

Brown:2010:TEE

[BR10] David J. Brown and Charles Reams. Toward energy-efficient computing. ACM Queue: Tomorrow’s Computing Today, 8 (2):30, February 2010. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Brandman:2005:PE

[Bra05] George Brandman. Patching the enterprise. ACM Queue: Tomorrow’s Computing Today, 3 (2):32–39, March 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Brown:2003:ICW

[Bro03] David Brown. Interview: a conversation with Wayne Rossing. ACM Queue: Tomorrow’s Computing Today, 1(6):12, 14–20, September/October 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Broderson:2004:ICT

[Bro04a] Robert Broderson. Interview: a conversation with Teresa Meng. ACM Queue: Tomorrow’s Computing Today, 2(1): 14, 16–21, March 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Brown:2004:OCH

[Bro04b] Aaron B. Brown. Oops! coping with human error in IT systems. ACM Queue: Tomorrow’s Computing Today, 2(8): 34–41, November 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Brown:2004:CSH

[Bro04c] David J. Brown. Curmudgeon: Search hurts. ACM Queue: Tomorrow’s Computing Today, 2 (2):84, 83, April 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Brown:2005:EUS

[Bro05] David J. Brown. From the editors: An update on software updates. ACM Queue: Tomorrow’s Computing Today, 3 (2):10–11, March 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Brown:2010:ICS

[Bro10] David Brown. Interview: a conversation with Steve Furber.

ACM Queue: Tomorrow’s Computing Today, 8(2):1–8, February 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Brook:2015:EPL

[Bro15]  Andrew Brook.  Evolution

and practice: Low-latency distributed applications in finance. ACM Queue: Tomorrow’s Computing Today, 13(4):40, April 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2770868.

Brady:2004:LSP

[BS04] Mike Brady and Vladimir Sanchez. Letters: Sender pays?; errata: [Alexander Nareyek, AI in computer games, ACM queue 1(10) 58–65 (2004)]. ACM Queue: Tomorrow’s Computing Today, 2(3):10, May 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Nar04].

Bailey:2016:FDS

[BS16] Josh Bailey and Stephen Stuart. Faucet: Deploying SDN in the enterprise. ACM Queue: Tomorrow’s Computing Today, 14(5):30, September 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3015763.

Bailis:2016:IRP

[BSP16] Peter Bailis, Justine Sherry, and Simon Peter. Introducing research for practice. ACM Queue: Tomorrow’s Computing Today, 14(2):70, March 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=2949831.

Boehm:2003:BAD

[BT03] Barry W. Boehm and Richard Turner. Balancing agility and discipline: a guide for the perplexed. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-321-18612-5. xxvii + 266 pp. LCCN QA76.76.D47 B635 2003.

Barrows:2006:SCI

[BT06] Ryan Barrows and Jim Traverso. Search considered integral. ACM Queue: Tomorrow’s Computing Today, 4(4):30–36, May 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Baudry:2022:LLS

[BTM22] Benoit Baudry, Tim Toady, and Martin Monperrus. Long live software Easter eggs!: They are as old as software. ACM Queue: Tomorrow’s Computing Today, 20(2):31–42, March 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3534857.

Bueso:2014:STP

[Bue14] Davidlohr Bueso. Scalability techniques for practical synchronization primitives. ACM Queue: Tomorrow’s Computing Today, 12(11):40, November 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2698990.

Burner:2003:DRT

[Bur03] Mike Burner. The deliberate revolution: Transforming integration with XML Web services. ACM Queue: Tomorrow’s Computing Today, 1(1): 28–37, March 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Burns:2007:CRT

[BW07] Alan Burns and Andrew J.

Wellings. Concurrent and real-time programming in Ada 2005. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-86697-9 (hardcover). xiv + 461 pp. LCCN QA76.73.A35 B854 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0803/2007281389-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0803/
2007281389-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0803/2007281389-t.html.

Beschastnikh:2016:DDS

[BWB^{+}16] Ivan Beschastnikh, Patty Wang, Yuriy Brun, Michael D., and Ernst. Debugging distributed systems. ACM Queue: Tomorrow’s Computing Today, 14 (2):50, March 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=2940294.

Butler:2022:WMT

[BY22] Jenna Butler and Catherine Yeh. Walk a mile in their shoes: The Covid pandemic through the lens of four tech workers. ACM Queue: Tomorrow’s Computing Today, 20 (2):43–66, March 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3534860.

Carata:2014:PP

[CAB^{+}14] Lucian Carata, Sherif Akoush, Nikilesh Balakrishnan, Thomas Bytheway, Ripduman Sohan, Margo Seltzer, and Andy Hopper. A primer on provenance. ACM Queue: Tomorrow’s Computing Today, 12(3):10, March 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cannon:2005:PWD

[Can05] J. C. Cannon. Privacy: what developers and IT professionals should know. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-321-22409-4 (paperback). xxxi + 347 pp. LCCN QA76.9.A25 C3645 2005.

Canny:2006:FHC

[Can06a]  John Canny.  The future

of human-computer interaction. ACM Queue: Tomorrow’s Computing Today, 4(6): 24–32, July/August 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cantrill:2006:HPS

[Can06b] Bryan Cantrill. Hidden in plain sight. ACM Queue: Tomorrow’s Computing Today, 4(1): 26–36, February 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cantrill:2008:PFI

[Can08] Bryan Cantrill. A pioneer’s flash of insight. ACM Queue: Tomorrow’s Computing Today, 6(4): 16–17, July/August 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cantrill:2009:ICA

[Can09] Bryan Cantrill. Interview: a conversation with Arthur Whitney. ACM Queue: Tomorrow’s Computing Today, 7(2): 12–13, February 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cantrill:2010:CPC

[Can10] Stephen V. Cantrill. Computers in patient care: The promise and the challenge. ACM Queue:

Tomorrow’s Computing Today, 8(8):20, August 2010. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Carpenter:2006:BFM

[Car07] Brian Carpenter. Better, faster, more secure. ACM Queue: Tomorrow’s Computing Today, 4 (10):42–48, December/January 2006–2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Carnes:2008:OS

[Car08] Donna Carnes. Ode to a sailor. ACM Queue: Tomorrow’s Computing Today, 6(3):58–62, May/ June 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cannon:2006:CD

[CB06] J. C. Cannon and Marilee Byers. Compliance deconstructed. ACM Queue: Tomorrow’s Computing Today, 4(7): 30–37, September 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Crosby:2006:VR

[CB07] Simon Crosby and David Brown. The virtualization reality. ACM Queue: Tomorrow’s Computing Today, 4(10):34–41, December/January 2006–2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cantrill:2008:RWC

[CB08] Bryan Cantrill and Jeff Bonwick. Real-world concurrency. ACM Queue: Tomorrow’s Computing Today, 6(5): 16–25, September 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Collier-Brown:2023:YDK

[CB23]  David Collier-Brown.  You

don’t know jack about application performance: Knowing whether you’re doomed to fail is important when starting a project. ACM Queue: Tomorrow’s Computing Today, 21 (2):54–65, March 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3595862.

Callow:2007:BGS

[CBB07] Mark Callow, Paul Beardow, and David Brittain. Big games, small screens. ACM Queue: Tomorrow’s Computing Today, 5 (7):40–50, November 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cascaval:2008:STM

[CBM^{+}08] Calin Cascaval, Colin Blundell, Maged Michael, Harold W. Cain, Peng Wu, Stefanie Chiras, and Siddhartha Chatterjee. Software transactional memory: why is it only a research toy? ACM Queue: Tomorrow’s Computing Today, 6 (5):46–58, September 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cafarella:2004:BNO

[CC04] Mike Cafarella and Doug Cutting. Building Nutch: Open source search. ACM Queue: Tomorrow’s Computing Today, 2(2):54–61, April 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Chan:2020:DIG

[CC20] Charisma Chan and Beth Cooper. Debugging incidents in Google’s distributed systems. ACM Queue: Tomorrow’s Computing Today, 18(2):??, March

                                        2020. URL https://dl.acm.
                                        org/doi/abs/10.1145/3400899.
                                        3404974.

Cardwell:2016:BCB

[CCG^{+}16] Neal Cardwell, Yuchung Cheng, C. Stephen Gunn, Soheil Hassas Yeganeh, and Van Jacobson. BBR: Congestionbased congestion control. ACM Queue: Tomorrow’s Computing Today, 14(5):50, September 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=3022184.

Coatta:2014:AQT

[CDH14] Terry Coatta, Michael Donat, and Jafar Husain. Automated QA testing at EA: Driven by events. ACM Queue: Tomorrow’s Computing Today, 12 (5):20, May 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cheng:2003:BCI

[CdSH^{+}03] Li-Te Cheng, Cleidson R. B.

de Souza, Susanne Hupfer, John Patterson, and Steven Ross. Building collaboration into IDEs. ACM Queue: Tomorrow’s Computing Today, 1 (9):40–50, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cerf:2014:APP

[Cer14] Vinton G. Cerf. ACM and the professional programmer. ACM Queue: Tomorrow’s Computing Today, 12(7):10, July 2014. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Feng:2003:MCE

[cF03] Wu chun Feng. Making a case for efficient supercomputing. ACM Queue: Tomorrow’s Computing Today, 1(7): 54–64, October 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Casari:2023:BRB

[CFL23] Amanda Casari, Julia Ferraioli, and Juniper Lovato. Beyond the repository: Best practices for open source ecosystems researchers. ACM Queue: Tomorrow’s Computing Today, 21 (2):14–34, March 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3595879.

Coatta:2010:UDA

[CG10] Terry Coatta and Julian Gosper. UX design and agile: a natural fit? ACM Queue: Tomorrow’s Computing Today, 8 (11):50, November 2010. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Crankshaw:2018:PSS

[CG18] Dan Crankshaw and Joseph Gonzalez. Prediction-serving systems. ACM Queue: Tomorrow’s Computing Today, 16 (1):70, January 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3210557.

Czerwonka:2018:CIC

[CGB^{+}18] Jacek Czerwonka, Michaela Greiler, Christian Bird, Lucas Panjer, and Terry Coatta. CodeFlow: Improving the code review process at Microsoft. ACM Queue: Tomorrow’s Computing Today, 16(5):20, September 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3292420.

Chand:2005:RPW

[Cha05a] Don Chand. Review of Privacy: What Developers and IT Professionals Should Know by J. C. Cannon, Addison-Wesley Professional, 2004, $49.99. ISBN 0-321-22409-4. ACM Queue: Tomorrow’s Computing Today, 3(5):60, June 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Chang:2005:RIR

[Cha05b]  R. S. Chang.  Review of

Integration-Ready Architecture and Design: Software Engineering with XML, Java, .NET, Wireless, Speech, and Knowledge Technologies by Jeff Zhuk, Cambridge University Press, 2004, $75.00, ISBN 0-521-52583-7. ACM Queue: Tomorrow’s Computing Today, 3(2): 58, March 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Champy:2006:PP

[Cha06] James Champy. People and process. ACM Queue: Tomorrow’s Computing Today, 4 (2):34–38, March 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Chakrabarti:2007:GCS

[Cha07] Anirban Chakrabarti. Grid computing security. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-44492-0 (hardcover). xiv + 331 pp. LCCN QA76.9.C58 C53 2007.

Chesley:2003:BRS

[Che03]  D. Chesley.  Book review:

802.11 Security, by Bruce Potter and Bob Fleck, O’Reilly, 2002, $34.95, ISBN 0-596-00290-4. ACM Queue: Tomorrow’s Computing Today, 1(6): 70, September/October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Chessin:2010:IEF

[Che10] Steve Chessin. Injecting errors for fun and profit. ACM Queue: Tomorrow’s Computing Today, 8(8):10, August 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Chisnall:2008:DGX

[Chi08] David Chisnall. The Definitive Guide to the Xen Hypervisor. Prentice Hall open source software development series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2008. ISBN 0-13-234971-X. xx + 286 pp. LCCN QA76.9.V5 C427 2008.

Chisa:2014:EPM

[Chi14a]  Ellen Chisa.  Evolution of

the product manager. ACM Queue: Tomorrow’s Computing Today, 12(9):40, September 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=2683579.

Chisnall:2014:TNS

[Chi14b] David Chisnall. There’s no such thing as a general-purpose processor. ACM Queue: Tomorrow’s Computing Today, 12(10): 20, October 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2687011.

Chisnall:2018:CLL

[Chi18] David Chisnall. C is not a lowlevel language. ACM Queue: Tomorrow’s Computing Today, 16(2):10, March 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3212479.

Chisnall:2023:HDI

[Chi23] David Chisnall. How to design an ISA: The popularity of RISC-V has led many to try designing instruction sets. ACM Queue: Tomorrow’s Computing Today, 21(6):27–46, November 2023. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3639445.

Christy:2009:CTR

[Chr09] Peter Christy. Commentary: a trip without a roadmap. ACM Queue: Tomorrow’s Computing Today, 7(1):2:1–2:??, January 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cerf:2011:BWW

[CJWG11] Vint Cerf, Van Jacobson, Nick Weaver, and Jim Gettys. BufferBloat: What’s wrong with the Internet? ACM Queue: Tomorrow’s Computing Today, 9 (12):10, December 2011. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Coates:2004:LBO

[CKG^{+}04] Jeffrey R. Coates, Vadim Kavalerov, Brian Goetz, Marcus Ranum, Ryan McLoughlin, and Robert Angelino. Letters: Buffer overruns and everyone; another look at security problems. ACM Queue: Tomorrow’s Computing Today, 2 (6):10, September 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Charland:2011:MAD

[CL11] Andre Charland and Brian LeRoux. Mobile application development: Web vs. native. ACM Queue: Tomorrow’s Computing Today, 9(4):20, April 2011. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Chen:2021:IML

[CLK^{+}21] Valerie Chen, Jeffrey Li, Joon Sik Kim, Gregory Plumb, and Ameet Talwalkar. Interpretable machine learning: Moving from mythos to diagnostics. QUEUE, 19(6):28– 56, November 2021. URL

https://dl.acm.org/doi/10.
1145/3511299.

Coleman:2003:LOH

[CLNM03] Jim Coleman, R. Bob Lee, Mel Ray Notage, and Bill Mc-Farland. Letters: OpenOffice in home office; the “Smart disk” is already here; 802.11 in action. ACM Queue: Tomorrow’s Computing Today, 1(7):5, October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Coar:2003:SNS

[Coa03a] Ken Coar. The sun never sits on distributed development. ACM Queue: Tomorrow’s Computing Today, 1(9):32–39, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Coata:2003:ADA

[Coa03b] Terry Coata. Another day, another bug. ACM Queue: Tomorrow’s Computing Today, 1(6):58–61, September/October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Coata:2003:EWY

[Coa03c] Terry Coata. From the Editors:

Would you like some data with that? ACM Queue: Tomorrow’s Computing Today, 1(3):4, May 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Coates:2003:OBS

[Coa03d] Josh Coates. Opinion: Big storage: Make or buy? ACM Queue: Tomorrow’s Computing Today, 1(4):6–7, June 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Coates:2004:CBO

[Coa04a] Josh Coates. Curmudgeon: a bigot by any other name. . .. ACM Queue: Tomorrow’s Computing Today, 2(3):80, 78, May 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Coates:2004:CWB

[Coa04b]  Josh Coates.  Curmudgeon:

When bad people happen to good games. ACM Queue: Tomorrow’s Computing Today, 1 (10):92, 91, February 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Coatta:2005:CTT

[Coa05] Terry Coatta. Curmudgeon: Traipsing through the QA tools desert. ACM Queue: Tomorrow’s Computing Today, 3(1): 64, February 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Coatta:2006:QFP

[Coa06]  Terry Coatta.  Q focus:

Purpose-built systems: The (not so) hidden computer. ACM Queue: Tomorrow’s Computing Today, 4(3):22–26, April 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Coatta:2008:ICE

[Coa08] Terry Coatta. Interview: a conversation with Erik Meijer and Jose Blakeley. ACM Queue: Tomorrow’s Computing Today, 6 (3):9–15, May/June 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Coatta:2013:QPK

[Coa13] Terry Coatta. Queue portrait:

Kate Matsudaira. ACM Queue: Tomorrow’s Computing Today, 11(5):10, May 2013. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cobbs:2022:PPW

[Cob22] Archie L. Cobbs. Persistence programming: Are we doing this right? ACM Queue: Tomorrow’s Computing Today, 20 (1):24–39, January 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3526210.

Coffman:2004:YFP

[Cof04] James E. Coffman. Not your father’s PBX? ACM Queue: Tomorrow’s Computing Today, 2 (6):40–47, September 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cohn:2006:AEP

[Coh06] Mike Cohn. Agile estimating and planning. Robert C. Martin series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2006. ISBN 0-13-147941-5 (paperback). xxx + 330 pp. LCCN QA76.76.D47 C6427 2006. URL http:

//www.loc.gov/catdir/toc/
ecip0517/2005023257.html.

Colyer:2019:BUS

[Col19a]  Adrian Colyer.  Back un-

der a SQL umbrella. ACM Queue: Tomorrow’s Computing Today, 17(5):10, September 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3371598.

Colyer:2019:GDD

[Col19b] Adrian Colyer. GAN dissection and datacenter RPCs. ACM Queue: Tomorrow’s Computing Today, 17(2):10, March 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3329783.

Colyer:2019:PML

[Col19c] Adrian Colyer. Putting machine learning into production systems. ACM Queue: Tomorrow’s Computing Today, 17 (4):60, July 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3365847.

Colyer:2019:SN

[Col19d] Adrian Colyer. SageDB and NetAccel. ACM Queue: Tomorrow’s Computing Today, 17 (1):10, January 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3317289.

Colyer:2019:TPO

[Col19e]  Adrian Colyer.  Time pro-

tection in operating systems and speaker legitimacy detection. ACM Queue: Tomorrow’s Computing Today, 17 (3):10, May 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3344779.

Convery:2004:NSA

[Con04] Sean Convery. Network security architectures. Networking technology series. Cisco Press, Indianapolis, IN, USA, 2004. ISBN 1-58705-115-X. xxxix + 739 pp. LCCN TK5105.59 .C5793 2004. Duplicate ISBN with [TCR03].

Cook:2019:ALB

[Coo19] Richard I. Cook. Above the line, below the line. ACM Queue: Tomorrow’s Computing Today, 17(6):??, November/December

2019. URL https://dl.acm.
org/doi/abs/10.1145/3380774.
3380777.

Cooper:2023:DOU

[Coo23] Geoffrey H. Cooper. Device onboarding using FDO and the untrusted installer model: FDO’s untrusted model is contrasted with Wi-Fi Easy Connect to illustrate the advantages of each mechanism. ACM Queue: Tomorrow’s Computing Today, 21 (5):29–53, September 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3631327.

Corchuelo:2007:REV

[Cor07] Rafael Corchuelo. Review of ‘Expert VB 2005 Business Objects’ (second edition) by Rockford Lhotka, Apress, 2006, $59.99, ISBN 1-59059-631-5. ACM Queue: Tomorrow’s Computing Today, 5(4):56, May– June 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Lho06].

Cornwell:2012:ASS

[Cor12] Michael Cornwell. Anatomy of a solid-state drive. ACM Queue: Tomorrow’s Computing Today, 10(10):??, October 2012. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cormode:2017:DS

[Cor17] Graham Cormode. Data sketching. ACM Queue: Tomorrow’s Computing Today, 15(2): 60, March 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3104030.

Cox:2019:SSD

[Cox19] Russ Cox. Surviving software dependencies. ACM Queue: Tomorrow’s Computing Today, 17 (2):80, March 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3344149.

Cavage:2014:BAC

[CP14] Mark Cavage and David Pacheco.

Bringing arbitrary compute to authoritative data. ACM Queue: Tomorrow’s Computing Today, 12(6):40, June 2014. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Choudhury:2018:DCS

[CP18] Diptanu Gon Choudhury and Timothy Perrett. Designing cluster schedulers for Internetscale services. ACM Queue: Tomorrow’s Computing Today, 16 (1):30, January 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3199609.

Cody:2004:LBO

[CPG^{+}04] Robert Cody, Lawrence C. Paulson, Todd Greer, Rodney Bates, Daniel Myer, Gary Hewitt, Jay Michaelson, John C. Nash, and Josh Coates. Letters: Buffer overrun madness; open source:

All grown up?; when is free really Free?; it’s not cool to be a bigot. ACM Queue: Tomorrow’s Computing Today, 2 (5):10, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Converse:2004:PMB

[CPM04] Tim Converse, Joyce Park, and Clark Morgan. PHP5 and MySQL bible. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2004. ISBN 0-7645-5746-

7. xxxvii + 1042 pp. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
bios/wiley047/2004103176.
html;   http://www.loc.gov/
catdir/description/wiley041/
2004103176.html;   http://
www.loc.gov/catdir/toc/wiley041/
2004103176.html.

Crawshaw:2020:EVN

[Cra20] David Crawshaw. Everything VPN is new again: The 24-year-old security model has found a second wind. ACM Queue: Tomorrow’s Computing Today, 18 (5):54–66, October 2020. URL

https://dl.acm.org/doi/10.
1145/3434571.3439745.

Creeger:2005:CMC

[Cre05]  Mache Creeger.  Curmud-

geon: Multicore CPUs for the masses. ACM Queue: Tomorrow’s Computing Today, 3(7): 64, 63, September 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2006:CER

[Cre06a] Mache Creeger. Curmudgeon:

Evolution or revolution? ACM Queue: Tomorrow’s Computing Today, 4(3):56–ff, April 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2006:GBH

[Cre06b] Mache Creeger. geek@home:

Better health care through technology. ACM Queue: Tomorrow’s Computing Today, 4(9): 13–15, November 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2006:GRH

[Cre06c] Mache Creeger. geek@home:

Rationalizing a home terabyte server. ACM Queue: Tomorrow’s Computing Today, 4(7): 14–17, September 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2007:GEW

[Cre07a] Mache Creeger. geek@home:

Embracing wired networks. ACM Queue: Tomorrow’s Computing Today, 5(4):12–15, May– June 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2007:RCH

[Cre07b]  Mache Creeger.  Repurpos-

ing consumer hardware. ACM Queue: Tomorrow’s Computing Today, 5(2):12–13, March 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2008:CRS

[Cre08] Mache Creeger. CTO roundtable storage: part two. ACM Queue: Tomorrow’s Computing Today, 6(7):34–40, November/December 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2009:CCO

[Cre09a] Mache Creeger. Cloud computing: An overview: Part two. ACM Queue: Tomorrow’s Computing Today, 7(5): 2:1–2:??, June 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2009:CRV

[Cre09b] Mache Creeger. CTO roundtable: Virtualization: Part one. ACM Queue: Tomorrow’s Computing Today, 7(1):26–35, January 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2010:CEC

[Cre10a] Mache Creeger. A conversation with Ed Catmull. ACM Queue: Tomorrow’s Computing Today, 8(11):30, November 2010. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2010:CRM

[Cre10b] Mache Creeger. CTO roundtable:

Malware defense. ACM Queue: Tomorrow’s Computing Today, 8(2):40, February 2010. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2010:MEA

[Cre10c] Mache Creeger. Moving to the edge: An ACM CTO roundtable on network virtualization. ACM Queue: Tomorrow’s Computing Today, 8(7):20, July 2010. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2010:MEC

[Cre10d] Mache Creeger. Moving to the edge: CTO roundtable overview. ACM Queue: Tomorrow’s Computing Today, 8 (7):40, July 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2010:TBIa

[Cre10e] Mache Creeger. The theft of business innovation: An ACM– BCS roundtable on threats to global competitiveness. ACM Queue: Tomorrow’s Computing Today, 8(11):10, November 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2010:TBIb

[Cre10f] Mache Creeger. The theft of business innovation: Overview. ACM Queue: Tomorrow’s Computing Today, 8(11):20, November 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2011:ACR

[Cre11a] Mache Creeger. ACM CTO roundtable on mobile devices in the enterprise. ACM Queue: Tomorrow’s Computing Today, 9(8):10, August 2011. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2011:MDE

[Cre11b] Mache Creeger. Mobile devices in the enterprise: CTO roundtable overview. ACM Queue: Tomorrow’s Computing Today, 9(8):20, August 2011. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Creeger:2022:RFH

[Cre22] Mache Creeger. The rise of fully homomorphic encryption: Often called the Holy Grail of cryptography, commercial FHE is near. ACM Queue: Tomorrow’s Computing Today, 20 (4):39–60, July 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3561800.

Crowley:2006:SP

[Cro06] James L. Crowley. Social perception. ACM Queue: Tomorrow’s Computing Today, 4(6): 34–43, July/August 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Christensen:2006:TMI

[CSL^{+}06] Jim Christensen, Jeremy Sussman, Stephen Levy, William E. Bennett, Tracee Vetting Wolf, and Wendy A. Kellogg. Too much information. ACM Queue: Tomorrow’s Computing Today, 4(6):50–57, July/August 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Clem:2021:SAG

[CT21] Timothy Clem and Patrick Thomson. Static analysis at GitHub: an experience report. ACM Queue: Tomorrow’s Computing Today, 19(4):42–67, July 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487019.3487022.

Currid:2004:TOR

[Cur04] Andy Currid. TCP offload to the rescue. ACM Queue: Tomorrow’s Computing Today, 2 (3):58–65, May 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cole:2010:PSK

[CW10] Clem Cole and Russell Williams. Photoshop scalability: Keeping it simple. ACM Queue: Tomorrow’s Computing Today, 8 (9):20, September 2010. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cymru:2006:CE

[Cym06] Team Cymru. Cybercrime: an epidemic. ACM Queue: Tomorrow’s Computing Today, 4 (9):24–35, November 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Dadzie:2005:USP

[Dad05] Joseph Dadzie. Understanding software patching. ACM Queue: Tomorrow’s Computing Today, 3(2):24–30, March 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Date:2016:SYU

[Dat16] Sachin Date. Should you upload or ship big data to the cloud? ACM Queue: Tomorrow’s Computing Today, 14 (2):30, March 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=2933408.

Davidovic:2018:CAS

[DB18] Štěpán Davidovič and Betsy Beyer. Canary analysis service. ACM Queue: Tomorrow’s Computing Today, 16(1): 10, January 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3194655.

DeJong:2006:QFW

[De 06] Peter De Jong. Q focus: Workflow systems: Going with the flow. ACM Queue: Tomorrow’s Computing Today, 4(2): 24–32, March 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Decrem:2004:DLW

[Dec04] Bart Decrem. Desktop Linux:

Where art thou? ACM Queue: Tomorrow’s Computing Today, 2(3):48–56, May 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Degenbaev:2016:ITG

[DEE^{+}16] Ulan Degenbaev, Jochen Eisinger, Manfred Ernst, Ross McIlroy, and Hannes Payer. Idle-time garbage-collection scheduling. ACM Queue: Tomorrow’s Computing Today, 14(3):70, May 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2977741.

Dougherty:2009:UBI

[DFZ^{+}09] Matthew T. Dougherty, Michael J.

Folk, Erez Zadok, Herbert J. Bernstein, Frances C. Bernstein, Kevin W. Eliceiri, Werner Benger, and Christoph Best. Unifying biological image formats with HDF5. ACM Queue: Tomorrow’s Computing Today, 7(9):20, October 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Davidovic:2015:RCA

[DG15] Štěpán Davidovič and Kavita Guliani. Reliable Cron across the planet. ACM Queue: Tomorrow’s Computing Today, 13 (3):30, March 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2745840.

Dhanuskodi:2023:CFC

[DGK^{+}23] Gobikrishna Dhanuskodi, Sudeshna Guha, Vidhya Krishnan, Aruna Manjunatha, Michael O’Connor, Rob Nertney, and Phil Rogers. Creating the first confidential GPUs: The team at NVIDIA brings confidentiality and integrity to user code and data for accelerated computing. ACM Queue: Tomorrow’s Computing Today, 21(4):68–93, July 2023. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3623393.3623391.

Diakopoulos:2015:AAD

[Dia15] Nicholas Diakopoulos. Accountability in algorithmic decisionmaking. ACM Queue: Tomorrow’s Computing Today, 13(9):50, November/December 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2886105.

DiBona:2004:ICW

[DiB04] Chris DiBona. Interview: a conversation with Will Harvey. ACM Queue: Tomorrow’s Computing Today, 1(10): 20–27, February 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ditzel:2003:ICD

[Dit03] David Ditzel. Interview: a conversation with Dan Dobberpuhl. ACM Queue: Tomorrow’s Computing Today, 1(7):14–24, October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Dally:2004:SPP

[DKK^{+}04] William J. Dally, Ujval J. Kapasi, Brucek Khailany, Jung Ho Ahn, and Abhishek Das. Stream processors: Programmability and efficiency. ACM Queue: Tomorrow’s Computing Today, 2 (1):52–62, March 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Danowitz:2012:CDR

[DKM^{+}12] Andrew Danowitz, Kyle Kelley, James Mao, John P. Stevenson, and Mark Horowitz. CPU DB: Recording microprocessor history. ACM Queue: Tomorrow’s Computing Today, 10 (4):18, April 2012. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Diffie:2009:CSP

[DL09] Whitfield Diffie and Susan Landau. Communications surveillance: Privacy and security at risk. ACM Queue: Tomorrow’s Computing Today, 7(8): 10, September 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Delignat-Lavaud:2023:WST

[DLdFV^{+}23] Antoine Delignat-Lavaud, CÃ\begin{matrix}c \\ ◯ \end{matrix}dric Fournet, Kapil Vaswani, Sylvan Clebsch, Maik Riechert, Manuel Costa, and Mark Russinovich. Why should I trust your code?: Confidential computing enables users to authenticate code running in TEEs, but users also need evidence this code is trustworthy. ACM Queue: Tomorrow’s Computing Today, 21(4):94–122, July 2023. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3623460.

Degenbaev:2019:GCJ

[DLP19] Ulan Degenbaev, Michael Lippautz, and Hannes Payer. Garbage collection as a joint venture. ACM Queue: Tomorrow’s Computing Today, 17(1): 60, January 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3325132.

Dowling:2015:UFO

[DM15] Phelim Dowling and Kevin Mc-Grath. Using free and open source tools to manage software quality. ACM Queue: Tomorrow’s Computing Today, 13 (4):20, April 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2767182.

Dillahunt:2017:RPT

[DMBB17] Tawanna Dillahunt, Stefanie Mueller, Patrick Baudisch, and Peter Bailis. Research for practice: Technology for underserved communities; personal fabrication. ACM Queue: Tomorrow’s Computing Today, 15 (2):70, March 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3105830.

Donat:2003:DAW

[Don03] Michael Donat. Debugging in an asynchronous world. ACM Queue: Tomorrow’s Computing Today, 1(6):22–30, September/October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Donat:2005:OAT

[Don05] Michael Donat. Orchestrating an automated test lab. ACM Queue: Tomorrow’s Computing Today, 3(1):46–53, February 2005. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic).

Drepper:2008:CV

[Dre08a] Ulrich Drepper. The cost of virtualization. ACM Queue: Tomorrow’s Computing Today, 6 (1):28–35, January 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Drepper:2008:PPT

[Dre08b] Ulrich Drepper. Parallel programming with transactional memory. ACM Queue: Tomorrow’s Computing Today, 6(5): 38–45, September 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Daries:2014:PAB

[DRW^{+}14] Jon P. Daries, Justin Reich, Jim Waldo, Elise M. Young, Jonathan Whittinghill, Daniel Thomas Seaton, Andrew Dean Ho, and Isaac Chuang. Privacy, anonymity, and big data in the social sciences. ACM Queue: Tomorrow’s Computing Today, 12 (7):30, July 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Deokuliar:2023:ITD

[DSBS23] Harsh Deokuliar, Raghvinder S. Sangwan, Youakim Badr, and Satish M. Srinivasan. Improving testing of deep-learning systems: a combination of differential and mutation testing results in better test data. ACM Queue: Tomorrow’s Computing Today, 21 (5):54–65, September 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3631340.

Durkee:2010:WCC

[Dur10] Dave Durkee. Why cloud computing will never be free. ACM Queue: Tomorrow’s Computing Today, 8(4):1–10, April 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

DeLine:2010:SDC

[DVR10] Robert DeLine, Gina Venolia, and Kael Rowan. Software development with code maps. ACM Queue: Tomorrow’s Computing Today, 8 (7):10, July 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Dwivedi:2006:SSP

[Dwi06] Himanshu Dwivedi. Securing storage: a practical guide to SAN and NAS security. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-34995-4 (hardcover). xxxii + 526 pp. LCCN TK5105.86 .D95 2006. URL http:

//www.loc.gov/catdir/toc/
ecip0516/2005021323.html.

Dwyer:2008:PWA

[Dwy08] Jeff Dwyer. Pro Web 2.0 Application Development with GWT.

Apress, Berkeley, CA, USA, 2008. ISBN 1-59059-985-3. xxvii + 450 pp. LCCN ????

Ebersbach:2005:WWC

[Ebe05] Anja Ebersbach. Wiki: Web collaboration. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-25995-3 (hardcover). ???? pp. LCCN ????

Edwards:2005:RSS

[Edw05] J. S. Edwards. Review of “Slamming Spam: A Guide for System Administrators” by Robert Haskins and Dale Nielsen, Addison-Wesley Professional, 2004, $44.99 ISBN 0-13-146716-6. ACM Queue: Tomorrow’s Computing Today, 3 (8):60, October 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Eilert:2003:LM

[EEMS03] John Eilert, Maria Eisenhaendler, Dorothea Matthaeu, and Ingol Salm, editors. Linux on the mainframe. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-101415-3 (paperback). xxviii + 430 pp. LCCN QA76.76.O63 L54553 2003.

Elberson:2003:LSD

[EGBA03] Matt Elberson, Kurth Guntheroth, Rodney M. Bates, and Donald E. Amby. Letters: Software development tools: Cautionary tales; coding for readability. ACM Queue: Tomorrow’s Computing Today, 1(9):6, December/January 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Eini:2011:PIL

[Ein11] Oren Eini. The pain of implementing LINQ providers. ACM Queue: Tomorrow’s Computing Today, 9(7):10, July 2011. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Erbad:2012:SSB

[EK12] Aiman Erbad and Charles Krasic. Sender-side buffers and the case for multimedia adaptation. ACM Queue: Tomorrow’s Computing Today, 10(10):??, October 2012. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Eure:2005:LRR

[EKG^{+}05] Tim Eure, Ben Kuehlhorn, Vijay Gill, James Synge, Edward Weir, and The Editors. Letters: RFID revisited; Queueing up priority queueing; guru code overload. ACM Queue: Tomorrow’s Computing Today, 3 (1):10, February 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ektare:2005:RDS

[Ekt05] Arun Ektare. Review of Designing Software for the Mobile Context: a Practitioner’s Guide (Computer Communications and Networks) by Roman Longoria (Ed.), Springer-Verlag, 2004, $69.95, ISBN 1-85233-785-0. ACM Queue: Tomorrow’s Computing Today, 3 (1):61, February 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Elerath:2007:HDD

[Ele07] Jon Elerath. Hard disk drives: the good, the bad and the ugly! ACM Queue: Tomorrow’s Computing Today, 5(6): 28–37, September 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Evans:2019:ACH

[EMC19] David Evans, Richard McDonald, and Terry Coatta. Access controls and health care records: Who owns the data? ACM Queue: Tomorrow’s Computing Today, 17(2):50, March 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3339247.

Enguix:2005:REG

[Eng05] Carlos F. Enguix. Review of Explorer’s Guide to the Semantic Web by Thomas B. Passin, Manning Publications, 2004, $39.95, ISBN 1-932394-20-6. ACM Queue: Tomorrow’s Computing Today, 3(3): 61, April 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Eberbach:2003:LSF

[ERF^{+}03] Andrew Eberbach, Jef Raskin, Mike Frisch, Mark Scott, and Harry Rudin. Letters: Slaves of fashion; on heretics and zealots; the FTC and spam. ACM Queue: Tomorrow’s Computing Today, 1(8):6, November 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Eriksen:2016:FS

[Eri16] Marius Eriksen. Functional at scale. ACM Queue: Tomorrow’s Computing Today, 14 (4):60, July 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3001119.

Ernst:2003:PIA

[Ern03] Rolf Ernst. Putting it all together. ACM Queue: Tomorrow’s Computing Today, 1(2): 50–55, April 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Esparrachiari:2018:TCM

[ERR18] Silvia Esparrachiari, Tanya Reilly, and Ashleigh Rentz. Tracking and controlling microservice dependencies. ACM Queue: Tomorrow’s Computing Today, 16(4):10, July 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3277541.

Eaton:2022:DLP

[ESTT22] Brian Eaton, Jeff Stewart, Jon Tedesco, and N. Cihan Tas. Distributed latency profiling through critical path tracing: CPT can provide actionable and precise latency analysis. ACM Queue: Tomorrow’s Computing Today, 20(1):40–79, January 2022. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3526967.

Fata:2018:CCG

[FAHB18] Matt Fata, Philippe-Joseph Arida, Patrick Hahn, and Betsy Beyer. Corp to cloud: Google’s virtual desktops. ACM Queue: Tomorrow’s Computing Today, 16(3):60, May 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3264508.

Feathers:2005:WEL

[Fea05] Michael C. Feathers. Working effectively with legacy code. Robert C. Martin series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-117705-2. xxi + 434 pp. LCCN QA76.76.A65 F43 2005.

Feldman:2003:ICM

[Fel03] Stu Feldman. Interview: a conversation with Mario Mazzola.

ACM Queue: Tomorrow’s Computing Today, 1(3):12, 14–16, May 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Feldman:2004:CAK

[Fel04a] Stuart Feldman. Interview: a conversation with Alan Kay. ACM Queue: Tomorrow’s Computing Today, 2(9):20–30, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Feldman:2004:ICB

[Fel04b] Stuart Feldman. Interview:

a conversation with Brewster Kahle. ACM Queue: Tomorrow’s Computing Today, 2(4): 24, 26–30, 32–33, June 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Feldman:2005:EDC

[Fel05a] Stuart Feldman. From the editors: Distributed computing in the modern enterprise. ACM Queue: Tomorrow’s Computing Today, 3(6):8, July 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Feldman:2005:QAM

[Fel05b] Stuart Feldman. Quality assurance: much more than testing. ACM Queue: Tomorrow’s Computing Today, 3(1): 26–29, February 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ferris:2003:ACO

[Fer03] Paul Ferris. The age of corporate open source enlightenment. ACM Queue: Tomorrow’s Computing Today, 1(5): 34–44, July/August 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Franciscus:2005:SR

[FG05] George Franciscus and Danilo Gurovich. Struts recipes. Manning, Greenwich, CT, USA, 2005. ISBN 1-932394-24-9 (paperback). xxix + 489 pp. LCCN TK5105.8885.S76 F73 2005.

Fatahalian:2008:GCL

[FH08] Kayvon Fatahalian and Mike Houston. GPUs: a closer look. ACM Queue: Tomorrow’s Computing Today, 6(2): 18–28, March 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Finkenzeller:2003:RHF

[Fin03] Klaus Finkenzeller. RFID handbook: fundamentals and applications in contactless smart cards and identification. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, second edition, 2003. ISBN 0-470-84402-7. xviii + 427 pp. LCCN TS160 .F5513 2003. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;

 http://www.loc.gov/catdir/
description/wiley039/2002192439.
html;   http://www.loc.gov/
catdir/toc/wiley031/2002192439.
html.

Forsgren:2017:DM

[FK17] Nicole Forsgren and Mik Kersten. DevOps metrics. ACM Queue: Tomorrow’s Computing Today, 15(6):30, November 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3182626.

Fitzmaurice:2003:SD

[FKB^{+}03] George W. Fitzmaurice, Azam Khan, William Buxton, Gordon Kurtenback, and Ravin Balakrishnan. Sentient data access via a diverse society of devices. ACM Queue: Tomorrow’s Computing Today, 1(8): 52–62, November 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Forsgren:2023:DAS

[FKN^{+}23] Nicole Forsgren, Eirini Kalliamvakou, Abi Noda, Michaela Greiler, Brian Houck, and Margaret-Anne Storey. DevEx in action: a study of its tangible impacts. ACM Queue: Tomorrow’s Computing Today, 21(6): 47–77, November 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3639443.

Florescu:2005:MSS

[Flo05] Daniela Florescu. Managing semi-structured data. ACM Queue: Tomorrow’s Computing Today, 3(8):18–24, October 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Fall:2005:YDK

[FM05] Kevin Fall and Steve McCanne. You don’t know jack about network performance. ACM Queue: Tomorrow’s Computing Today, 3(4):54–59, May 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Falcao:2022:PPC

[FoDF22] Gabriel Falcao and JoÃ£o Dinis Ferreira. To PiM or not to PiM: The case for inmemory inferencing of quantized CNNs at the edge. ACM Queue: Tomorrow’s Computing Today, 20(6):9–34, 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3580503.

Ford:2007:OVC

[For07]  Richard Ford.  Open vs.

closed: which source is more secure? ACM Queue: Tomorrow’s Computing Today, 5 (1):32–38, February 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Foxwell:2008:RDG

[Fox08] Harry J. Foxwell. Review of ‘The Definitive Guide to the Xen Hypervisor’ (Prentice Hall Open Source Software Development Series), David Chisnall, Prentice Hall PTR, 2007, $49.99, ISBN 0-13-234971-X. ACM Queue: Tomorrow’s Computing Today, 6(3):69, May/ June 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Foster:2009:HDM

[FPWM09] Ian Foster, Savas Parastatidis, Paul Watson, and Mark McKeown. How do I model state? Let me count the ways. ACM Queue: Tomorrow’s Computing Today, 7(2):54–55, February 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Frazelle:2018:SMA

[Fra18] Jessie Frazelle. Security for the modern age. ACM Queue: Tomorrow’s Computing Today, 16(5):70, September 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3301253.

Frazelle:2019:OSF

[Fra19a] Jessie Frazelle. Open-source firmware. ACM Queue: Tomorrow’s Computing Today, 17 (3):20, May 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3349301.

Frazelle:2019:OBM

[Fra19b]  Jessie Frazelle.  Opening

up the baseboard management controller. ACM Queue: Tomorrow’s Computing Today, 17(5):??, September 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371595.3378404.

Frazelle:2019:SBP

[Fra19c] Jessie Frazelle. Securing the boot process. ACM Queue: Tomorrow’s Computing Today, 17(6):??, November/December

                                        2019. URL https://dl.acm.
                                        org/doi/abs/10.1145/3380774.
                                        3382016.

Frazelle:2020:BDC

[Fra20a] Jessie Frazelle. Battery Day: a closer look at the technology that makes portable electronics possible. ACM Queue: Tomorrow’s Computing Today, 18 (5):5–25, October 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434571.3439415.

Frazelle:2020:CML

[Fra20b] Jessie Frazelle. Chipping away at Moore’s Law. ACM Queue: Tomorrow’s Computing Today, 18(1):??, January 2020. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/abs/10.1145/3387945.
                                        3388515.

Frazelle:2020:LPG

[Fra20c]  Jessie Frazelle.  Let’s play

global thermonuclear energy: It’s important to know where your power comes from. ACM Queue: Tomorrow’s Computing Today, 18(6):5–20, December 2020. URL https:

//dl.acm.org/doi/10.1145/
3442632.3447956.

Frazelle:2020:LDB

[Fra20d] Jessie Frazelle. The life of a data byte: Be kind and rewind. ACM Queue: Tomorrow’s Computing Today, 18 (3):5–28, June 2020. URL

https://dl.acm.org/doi/10.
1145/3411757.3419941.

Frazelle:2020:WAM

[Fra20e] Jessie Frazelle. Out-of-this-world additive manufacturing: From thingamabobs to rockets, 3D printing takes many forms. ACM Queue: Tomorrow’s Computing Today, 18 (4):5–20, August 2020. URL

https://dl.acm.org/doi/10.
1145/3424302.3430113.

Frazelle:2020:PP

[Fra20f]  Jessie Frazelle.  Power to

the people. ACM Queue: Tomorrow’s Computing Today, 18(2):??, March 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3400899.3402527.

Frazelle:2021:CMC

[Fra21a] Jessie Frazelle. Chip measuring contest: The benefits of purpose-built chips. ACM Queue: Tomorrow’s Computing Today, 19(5):5–21, September 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3494834.3501254.

Frazelle:2021:NEM

[Fra21b] Jessie Frazelle. A new era for mechanical CAD: Time to move forward from decades-old design. ACM Queue: Tomorrow’s Computing Today, 19(2):5–16, April 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3466132.3469844.

Freedman:2003:OSE

[Fre03] Avi Freedman. Opinion: Securing the edge. ACM Queue: Tomorrow’s Computing Today, 1 (1):6, 8–9, March 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Feamster:2013:RS

[FRZ13] Nick Feamster, Jennifer Rexford, and Ellen Zegura. The road to SDN. ACM Queue: Tomorrow’s Computing Today, 11(12):??, December 2013. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2560327.

Frantz:2004:DPP

[FS04] Gene Frantz and Ray Simar. DSP: Of processors and processing. ACM Queue: Tomorrow’s Computing Today, 2 (1):22–30, March 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Forsgren:2021:SDP

[FSM^{+}21] Nicole Forsgren, Margaret-Anne Storey, Chandra Maddila, Thomas Zimmermann, Brian Houck, and Jenna Butler. The SPACE of developer productivity: There’s more to it than you think. ACM Queue: Tomorrow’s Computing Today, 19(1):20–48, February 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3454122.3454124.

Fedorova:2009:MPE

[FSSP09] Alexandra Fedorova, Juan Carlos Saez, Daniel Shelepov, and Manuel Prieto. Maximizing power efficiency with asymmetric multicore systems. ACM Queue: Tomorrow’s Computing Today, 7(10):30, November 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Foster:2005:DED

[FT05] Ian Foster and Steven Tuecke.

Describing the elephant: the different faces of IT as service. ACM Queue: Tomorrow’s Computing Today, 3(6):26–29, July 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Fischl:2007:MSM

[FT07] Jason Fischl and Hannes Tschofenig. Making SIP make cents. ACM Queue: Tomorrow’s Computing Today, 5(2): 42–49, March 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gehani:2021:DBP

[GAI^{+}21] Ashish Gehani, Raza Ahmad, Hassan Irshad, Jianqiao Zhu, and Jignesh Patel. Digging into Big Provenance (with SPADE): a user interface for querying provenance. ACM Queue: Tomorrow’s Computing Today, 19(3):77–106, June 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3475965.3476885.

Garfinkel:2018:UDR

[GAM18] Simson Garfinkel, John M. Abowd, and Christian Martindale. Understanding database reconstruction attacks on public data. ACM Queue: Tomorrow’s Computing Today, 16(5): 50, September 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3295691.

Garfinkel:2007:DME

[Gar07a] Simson L. Garfinkel. Document & media exploitation. ACM Queue: Tomorrow’s Computing Today, 5(7):22–30, November 2007. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic).

Garrett:2007:P

[Gar07b] Matthew Garrett. Powering down. ACM Queue: Tomorrow’s Computing Today, 5(7): 16–21, November 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gatlin:2004:TTD

[Gat04] Kang Su Gatlin. Trials and tribulations of debugging concurrency. ACM Queue: Tomorrow’s Computing Today, 2 (7):66–73, October 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grude:2004:LSC

[GBG^{+}04] Ulrich Grude, Rodney Bates, Annelise Gimpel, Stan Kelly-Bootle, Henri Socha, Bob Supnik, and Ian Phillips. Letters: Schizoid classes; don’t forget to vote!; something about simulators; the bigots of the world. ACM Queue: Tomorrow’s Computing Today, 2(8):12, November 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). Socha and Supnik discuss the origins of little-endian addressing in DEC PDP-11 and Intel x86.

Guha:2015:SOE

[GBM15] R. V. Guha, Dan Brickley, and Steve MacBeth. Schema.org: Evolution of structured data on the Web. ACM Queue: Tomorrow’s Computing Today, 13(9):10, November/December 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2857276.

Gray:2005:CA

[GC05] Jim Gray and Mark Compton. A call to arms. ACM Queue: Tomorrow’s Computing Today, 3(3):30–38, April 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gasevic:2006:MDA

[GDD06] Dragan Gašević, Dragan Djurić, and Vladan Devedzic. Model driven architecture and ontology development. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-32180-2 (hardback). xvii + 311 pp. LCCN QA76.76.D47 G37 2006; QA76.76.D47 G38 2006.

Guevara:2020:DPP

[GDWC20] Miguel Guevara, Damien Desfontaines, Jim Waldo, and Terry Coatta. Differential privacy: The pursuit of protections by default: a discussion with Miguel Guevara, Damien Desfontaines, Jim Waldo, and Terry Coatta. ACM Queue: Tomorrow’s Computing Today, 18 (5):93–112, October 2020. URL

https://dl.acm.org/doi/10.
1145/3434571.3439229.

Geer:2006:PK

[Gee06] Daniel E. Geer. Playing for keeps. ACM Queue: Tomorrow’s Computing Today, 4(9): 42–48, November 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Geer:2007:ES

[Gee07] Daniel E. Geer. The evolution of security. ACM Queue: Tomorrow’s Computing Today, 5(3):30–35, April 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gerstel:2005:SSD

[Ger05] Tom Gerstel. Streams and standards: delivering mobile video. ACM Queue: Tomorrow’s Computing Today, 3(4): 48–53, May 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gettys:2012:HBD

[Get12] Jim Gettys. Home bufferbloat demonstration videos. ACM Queue: Tomorrow’s Computing Today, 10(2):10, February 2012. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gray:2008:FDO

[GF08] Jim Gray and Bob Fitzgerald. Flash disk opportunity for server applications. ACM Queue: Tomorrow’s Computing Today, 6(4):18–23, July/August 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Green:2015:NLT

[GHM15] Spence Green, Jeffrey Heer, and Christopher D. Manning. Natural language translation at the intersection of AI and HCI. ACM Queue: Tomorrow’s Computing Today, 13(6):30, June 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2798086.

Ghosh:2011:DU

[Gho11] Debasish Ghosh. DSL for the uninitiated. ACM Queue: Tomorrow’s Computing Today, 9 (6):10, June 2011. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gibian:2007:RPS

[Gib07] Marc S. Gibian. Review of ‘Practical Subversion’, Second Edition, by Daniel Berlin and Garrett Rooney, Apress, 2006, $34.99, ISBN 1-59059-753-2. ACM Queue: Tomorrow’s Computing Today, 5(7):51, November 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [BR06].

Gill:2004:LPQ

[Gil04] Vijay Gill. Lack of priority queuing considered harmful. ACM Queue: Tomorrow’s Computing Today, 2(8):

64–69, November 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gill:2014:DSL

[Gil14] Andy Gill. Domain-specific languages and code synthesis using Haskell. ACM Queue: Tomorrow’s Computing Today, 12 (4):30, April 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gittleman:2006:RFA

[Git06] Arthur Gittleman. Review of “Foundations of Ajax” by Ryan Asleson and Nathaniel Schutta, APress, 2005, $39.99, ISBN 1-59059-582-3. ACM Queue: Tomorrow’s Computing Today, 4 (1):62, February 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [AS06c].

Gazis:2022:MWK

[GK22] Alexandros Gazis and Eleftheria Katsiri. Middleware 101: What to know now and for the future. ACM Queue: Tomorrow’s Computing Today, 20 (1):10–23, January 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3526211.

Green:2011:CGF

[GL11] Robert Green and Henry Ledgard. Coding guidelines: Finding the art in the science.

ACM Queue: Tomorrow’s Computing Today, 9(11):10, November 2011. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gaud:2015:CMM

[GLF^{+}15] Fabien Gaud, Baptiste Lepers, Justin Funston, Mohammad Dashti, Alexandra Fedorova, Vivien Quéma, Renaud Lachaize, and Mark Roth. Challenges of memory management on modern NUMA system. ACM Queue: Tomorrow’s Computing Today, 13(8):70, September/October 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2852078.

Godard:2003:DLE

[God03] Ivan Godard. Division of labor in embedded systems. ACM Queue: Tomorrow’s Computing Today, 1(2):32–41, April 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Godbolt:2019:OCC

[God19] Matt Godbolt. Optimizations in C++ compilers. ACM Queue: Tomorrow’s Computing Today, 17(5):20, September 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3372264.

Goff:2004:NDC

[Gof04] Max K. Goff. Network distributed computing: fitscapes and fallacies. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-100152-3 (paperback). xxii + 280 pp. LCCN TK5105.5 .G58 2004.

Goldner:2003:EI

[Gol03] Jeffrey S. Goldner. The emergence of iSCSI. ACM Queue: Tomorrow’s Computing Today, 1(4):44–53, June 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Goldberg:2014:WIT

[Gol14] Sharon Goldberg. Why is it taking so long to secure Internet routing? ACM Queue: Tomorrow’s Computing Today, 12(8):20, August 2014. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Goodman:2007:RGG

[Goo07]  Brian D. Goodman.  Re-

view of ‘The Geek Gap: Why Business and Technology Professionals Don’t Understand Each Other and Why They Need Each Other to Survive’ by Bill Pfleging and Minda Zetlin, Prometheus Books, 2006, $25.00, ISBN 1-59102-415-3. ACM Queue: Tomorrow’s Computing Today, 5(2): 56, March 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [PZ06].

Gorton:2006:ESA

[Gor06] Ian Gorton. Essential software architecture. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-28713-2 (hardback). xviii + 283 pp. LCCN QA76.754 .G67 2006.

                                        URL http://www.loc.gov/
                                        catdir/enhancements/fy0625/
                                        2006921741-d.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy0814/2006921741-b.html;
                                         http://www.loc.gov/catdir/
                                        toc/fy0613/2006921741.html.

Gorelik:2007:OSA

[Gor07]  Vlad Gorelik.  One step

ahead. ACM Queue: Tomorrow’s Computing Today, 5(1): 24–31, February 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Guo:2024:TAU

[GPPV24] Jinnan Guo, Peter Pietzuch, Andrew Paverd, and Kapil Vaswani. Trustworthy AI using Confidential Federated Learning: Federated learning and confidential computing are not competing technologies. ACM Queue: Tomorrow’s Computing Today, 22(2), March 2024. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3665220.

Gunther:2015:HSS

[GPT15] Neil Gunther, Paul Puglia, and Kristofer Tomasette. Hadoop superlinear scalability. ACM Queue: Tomorrow’s Computing Today, 13(5):20, May 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2789974.

Gray:2005:ICT

[Gra05] Jim Gray. Interview: a conversation with Tim Bray. ACM Queue: Tomorrow’s Computing Today, 3(1):20–25, February 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Graefe:2008:FMR

[Gra08a]  Goetz Graefe.  The five-

minute rule 20 years later: and how flash memory changes the rules. ACM Queue: Tomorrow’s Computing Today, 6(4): 40–52, July/August 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gray:2008:DCE

[Gra08b] Jim Gray. Distributed computing economics. ACM Queue: Tomorrow’s Computing Today, 6(3):63–68, May/June 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Golab:2014:ECW

[GRA^{+}14] Wojciech Golab, Muntasir R. Rahman, Alvin AuYoung, Kimberly Keeton, and Xiaozhou (Steve) Li. Eventually consistent: Not what you were expecting? ACM Queue: Tomorrow’s Computing Today, 12(1):30, January 2014. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grayson:2019:CWH

[Gra19] Marisa R. Grayson. Cognitive work of hypothesis exploration during anomaly response. ACM Queue: Tomorrow’s Computing Today, 17(6): ??, November/December 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3380774.3380778.

Gregg:2010:VSL

[Gre10] Brendan Gregg. Visualizing system latency. ACM Queue: Tomorrow’s Computing Today, 8 (5):30, May 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gregg:2016:FG

[Gre16] Brendan Gregg. The flame graph. ACM Queue: Tomorrow’s Computing Today, 14 (2):10, March 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=2927301.

Grossman:2003:ENW

[Gro03a] Edward Grossman. From the Editors: New world order. ACM Queue: Tomorrow’s Computing Today, 1(9):5, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grover:2003:MSP

[Gro03b] Andrew Grover. Modern system power management. ACM Queue: Tomorrow’s Computing Today, 1(7):66–72, October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:Ca

[Gro04a] Edward Grossman. Calendar.

ACM Queue: Tomorrow’s Computing Today, 2(7):76, October 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:Cb

[Gro04b] Edward Grossman. Calendar.

ACM Queue: Tomorrow’s Computing Today, 2(8):72, November 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:Cd

[Gro04c] Edward Grossman. Calendar.

ACM Queue: Tomorrow’s Computing Today, 2(9):77, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:Cc

[Gro04d] Edward Grossman. Careers. ACM Queue: Tomorrow’s Computing Today, 2(8):74–77, November 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:ECA

[Gro04e] Edward Grossman. From the Editors: Calling all cars [overstrike] readers. ACM Queue: Tomorrow’s Computing Today, 2(6):8, September 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:EFG

[Gro04f] Edward Grossman. From the Editors: Fund and games and software development. ACM Queue: Tomorrow’s Computing Today, 1(10):5, February 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:ERI

[Gro04g] Edward Grossman. From the Editors: RFID isn’t science fiction. ACM Queue: Tomorrow’s Computing Today, 2(7):10, October 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:ESE

[Gro04h] Edward Grossman. From the Editors: Search — an enterprising affair. ACM Queue: Tomorrow’s Computing Today, 2(2):6, April 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:BPL

[Gro04i] Edward Grossman. From the Editors: The big programming languages issue. ACM Queue: Tomorrow’s Computing Today, 2(9):10, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:EGC

[Gro04j] Edward Grossman. From the Editors: The guru code. ACM Queue: Tomorrow’s Computing Today, 2(8):10, November 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:ENS

[Gro04k] Edward Grossman. From the Editors: The new screen of death. ACM Queue: Tomorrow’s Computing Today, 2(4):8, June 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:EVY

[Gro04l] Edward Grossman. From the Editors: Virtually yours. ACM Queue: Tomorrow’s Computing Today, 2(5):8, July 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:L

[Gro04m] Edward Grossman. Letters. ACM Queue: Tomorrow’s Computing Today, 2(7):12, October 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:N

[Gro04n] Edward Grossman. News 2.0. ACM Queue: Tomorrow’s Computing Today, 2(7):13, October 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:RDR

[Gro04o] Edward Grossman. Review of Dynamic Reconfiguration: Architectures and Algorithms by Ramachandran Vaidyanathan and Jerry L. Trahan. Plenum Publishing, 2003, ISBN 0-306-48189-8. ACM Queue: Tomorrow’s Computing Today, 2 (7):74, October 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:RIB

[Gro04p] Edward Grossman. Review of Immunity-based systems by Yoshiteru Ishida. Springer-Verlag, 2004, ISBN 3-540-00896-9. ACM Queue: Tomorrow’s Computing Today, 2(7):74, October 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:WYHa

[Gro04q] Edward Grossman. What’s on your hard drive? ACM Queue:

Tomorrow’s Computing Today, 2(7):14, October 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:WYHb

[Gro04r] Edward Grossman. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 2(8):16, November 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2004:WYHc

[Gro04s] Edward Grossman. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 2(9):14, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2005:Ca

[Gro05a] Edward Grossman. Calendar.

ACM Queue: Tomorrow’s Computing Today, 3(1):62, February 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2005:Cb

[Gro05b] Edward Grossman. Calendar.

ACM Queue: Tomorrow’s Computing Today, 3(2):60, March 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2005:Cc

[Gro05c] Edward Grossman. Calendar.

ACM Queue: Tomorrow’s Computing Today, 3(3):62, April 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2005:Cd

[Gro05d]  Edward Grossman.  Calen-

dar. ACM Queue: Tomorrow’s Computing Today, 3(4):61, May 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2005:EAL

[Gro05e] Edward Grossman. From the Editors: I am an abstraction layer. ACM Queue: Tomorrow’s Computing Today, 3(3):8, April 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2005:EPQ

[Gro05f] Edward Grossman. From the Editors: Puttin’ the Queue in QA. ACM Queue: Tomorrow’s Computing Today, 3(1): 8, February 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2005:WYHa

[Gro05g] Edward Grossman. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 3(1):11, February 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2005:WYHb

[Gro05h] Edward Grossman. What’s on your hard drive? ACM Queue:

Tomorrow’s Computing Today, 3(2):18, March 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2005:WYHc

[Gro05i] Edward Grossman. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 3(3):14, April 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Grossman:2005:WYHd

[Gro05j] Edward Grossman. What’s on your hard drive. ACM Queue: Tomorrow’s Computing Today, 3(4):15, May 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Garfinkel:2023:SYT

[GS23] Simson Garfinkel and Jon Stewart. Sharpening your tools:

Updating bulk extractor for

the 2020s. ACM Queue: Tomorrow’s Computing Today, 21 (1):30–56, January 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3587827.

Goodson:2007:SSC

[GSI07] Garth Goodson, Sai Susharla, and Rahul Iyer. Standardizing storage clusters. ACM Queue: Tomorrow’s Computing Today, 5(6):20–27, September 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ghosh:2003:AEC

[GT03] Ashish Ghosh and Shigeyoshi Tsutsui, editors. Advances in evolutionary computing: theory and applications. Natural computing series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-43330-9. xvi + 1006 pp. LCCN QA76.618 .A215 2003. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/fy036/2002029164.html.

Garcia-Tobin:2024:ESA

[GTK24] Charles Garcia-Tobin and Mark Knight. Elevating security with Arm CCA: Attestation and verification are integral to adopting confidential computing. ACM Queue: Tomorrow’s Computing Today, 22(2):39–56, 2024. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664291.

Gardiner:2021:QST

[GTNNM21] Michael Gardiner, Alexander Truskovsky, George Neville-Neil, and Atefeh Mashatan. Quantum-safe trust for vehicles: The race is already on. ACM Queue: Tomorrow’s Computing Today, 19(2):93–115, April 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3466132.3467860.

Gumz:2004:BRC

[Gum04a] Joy Gumz. Book review: Computer and Intrusion Forensics edited by George Mohay, Alison Anderson, Byron Collie, Olivier de Vel, and Rod McKemmish, Artech House, 2003, $79.00, ISBN 1-580-53369-8. ACM Queue: Tomorrow’s Computing Today, 2(2):81, April 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gumz:2004:BRH

[Gum04b] Joy Gumz. Book review: Hacking Exposed: Network Security Secrets and Solutions, 4th ed., by Stuart McClure, Joel Scambray, and George Kurtz, McGraw-Hill, 2003, $49.99, ISBN 0-07-222742-7. ACM Queue: Tomorrow’s Computing Today, 1(10):88, February 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Guo:2017:MJF

[Guo17]  Robert Guo.  MongoDB’s

JavaScript fuzzer. ACM Queue: Tomorrow’s Computing Today, 15(1):30, January 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3059007.

Gustafsson:2005:TP

[Gus05] Andreas Gustafsson. Threads without the pain. ACM Queue: Tomorrow’s Computing Today, 3(9):42–47, November 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gunteroth:2005:LEP

[GWE^{+}05] Kurt Gunteroth, Greg Wilson, Michael Erdmann, Dmitriy Vasilev, Ed Ravin, and Vivek Rao. Letters: Extensible programming, extended edition; A lesson in self-healing; Fortran lives. ACM Queue: Tomorrow’s Computing Today, 3(2): 12–14, March 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Hacken:2006:RHA

[Hac06] George Hacken. Review of “High-Assurance Design: Architecting Secure and Reliable Enterprise Applications by Clifford Berg,” Addison-Wesley Professional, 2005, $54.99, ISBN 0-321-37577-7. ACM Queue: Tomorrow’s Computing Today, 4 (4):50, May 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Ber06].

Halevy:2005:WYD

[Hal05] Alon Halevy. Why your data won’t mix. ACM Queue: Tomorrow’s Computing Today, 3 (8):50–58, October 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Hallinan:2007:ELP

[Hal07] Christopher Hallinan. Embedded Linux primer: a practical, real-world approach. Prentice Hall open source software development series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2007. ISBN 0-13-167984-8 (paperback). xxix + 537 pp. LCCN QA76.76.O63 H34462 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0612/2006012886.html.

Hamilton:2005:ICP

[Ham05] James Hamilton. Interview: a conversation with Pat Selinger. ACM Queue: Tomorrow’s Computing Today, 3(3):18–28, April 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2014:NBN

[Ham14] Eran Hammer. Notes from the battlefield (Node.js at Walmart). ACM Queue: Tomorrow’s Computing Today, 12(2): 70, February 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Hanrahan:2009:CDSa

[Han09a] Pat Hanrahan. A conversation with David Shaw. ACM Queue: Tomorrow’s Computing Today, 7(8):20–21, September 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Hanrahan:2009:CDSb

[Han09b] Pat Hanrahan. A conversation with David Shaw — audio clip. ACM Queue: Tomorrow’s Computing Today, 7(8):??, September 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Harr:2003:ICJ

[Har03a] Randy Harr. Interview: a conversation with Jim Ready. ACM Queue: Tomorrow’s Computing Today, 1(2):6, 8, 10–13, April 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Harr:2003:SSE

[Har03b] Randy Harr. Storage — n sides to every story. ACM Queue: Tomorrow’s Computing Today, 1(4):4, June 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Harr:2005:EFS

[Har05a] Randy Harr. From the editors: On feeling secure in an unsafe world. ACM Queue: Tomorrow’s Computing Today, 3(5):8, June 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Harris:2005:GHE

[Har05b] Shon Harris, editor. Gray hat hacking: the ethical hacker’s handbook. All-in-one. Osborne/ McGraw-Hill, Berkeley, CA, USA, 2005. ISBN 0-07-225709-1 (paperback). xx + 434 pp. LCCN QA76.9.A25 G743 2005.

Harris:2015:DBE

[Har15] Rich Harris. Dismantling the barriers to entry. ACM Queue: Tomorrow’s Computing Today, 13(5):30, May 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2790378.

Hartmann:2016:SE

[Har16] Heinrich Hartmann. Statistics for engineers. ACM Queue: Tomorrow’s Computing Today, 14 (1):20, January 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2903468.

Hays:2004:DBF

[Hay04] W. Patrick Hays. DSPs: Back to the future. ACM Queue: Tomorrow’s Computing Today, 2 (1):42–51, March 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Heer:2010:TTVa

[HBO10a] Jeffrey Heer, Michael Bostock, and Vadim Ogievetsky. A tour through the visualization zoo. ACM Queue: Tomorrow’s Computing Today, 8(4):40, April 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Heer:2010:TTVb

[HBO10b] Jeffrey Heer, Michael Bostock, and Vadim Ogievetsky. A tour through the visualization zoo. ACM Queue: Tomorrow’s Computing Today, 8(5):20, May 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Hughes:2022:CIT

[HD22] James P. Hughes and Whitfield Diffie. The challenges of IoT, TLS, and random number generators in the real world: Bad random numbers are still with us and are proliferating in modern systems. ACM Queue: Tomorrow’s Computing Today, 20(3):18–40, May 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546933.

Helland:2015:ICE

[Hel15]  Pat Helland.  Immutability

changes everything. ACM Queue: Tomorrow’s Computing Today, 13(9):40, November/December 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2884038.

Helland:2016:LBD

[Hel16a] Pat Helland. Life beyond distributed transactions. ACM Queue: Tomorrow’s Computing Today, 14(5):70, September 2016. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3025012.

Helland:2016:PB

[Hel16b] Pat Helland. The power of babble. ACM Queue: Tomorrow’s Computing Today, 14 (4):70, July 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3003188.

Helland:2016:SSS

[Hel16c] Pat Helland. The singular success of SQL. ACM Queue: Tomorrow’s Computing Today, 14 (3):80, May 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2983199.

Helland:2016:SDS

[Hel16d] Pat Helland. Standing on distributed shoulders of giants. ACM Queue: Tomorrow’s Computing Today, 14(2):80, March 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2953944.

Helland:2017:SEF

[Hel17a] Pat Helland. Side effects, front and center! ACM Queue: Tomorrow’s Computing Today, 15 (2):50, March 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3099561.

Helland:2017:XJL

[Hel17b] Pat Helland. XML and JSON are like cardboard. ACM Queue: Tomorrow’s Computing Today, 15(4):50, July 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3143320.

Helland:2018:CE

[Hel18a] Pat Helland. Consistently eventual. ACM Queue: Tomorrow’s Computing Today, 16(2): 40, March 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3226077.

Helland:2018:ION

[Hel18b] Pat Helland. Identity by any other name. ACM Queue: Tomorrow’s Computing Today, 16 (6):70, November 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3314115.

Helland:2018:MYS

[Hel18c] Pat Helland. Mind your state for your state of mind. ACM Queue: Tomorrow’s Computing Today, 16(3):10, May 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=3236388.

Helland:2019:ESL

[Hel19a] Pat Helland. Extract, shoehorn, and load. ACM Queue: Tomorrow’s Computing Today, 17 (2):60, March 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3339880.

Helland:2019:STD

[Hel19b] Pat Helland. Space time discontinuum. ACM Queue: Tomorrow’s Computing Today, 17(5): 30, September 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3372732.

Helland:2019:WAV

[Hel19c] Pat Helland. Write amplification versus read perspiration. ACM Queue: Tomorrow’s Computing Today, 17 (4):40, July 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3364509.

Helland:2020:BAL

[Hel20a] Pat Helland. Baleen Analytics: Large-scale filtering of data provides serendipitous surprises. ACM Queue: Tomorrow’s Computing Today, 18(6): 21–30, December 2020. URL

https://dl.acm.org/doi/10.
1145/3442632.3446917.

Helland:2020:DOV

[Hel20b] Pat Helland. Data on the outside vs. data on the inside: Data kept outside SQL has different characteristics from data kept inside. ACM Queue: Tomorrow’s Computing Today, 18 (3):43–60, June 2020. URL

https://dl.acm.org/doi/10.
1145/3411757.3415014.

Helland:2021:AMP

[Hel21a] Pat Helland. ACID: My personal “C“ change: How could I miss such a simple thing? ACM Queue: Tomorrow’s Computing Today, 19(2):17–20, April 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3466132.3469647.

Helland:2021:DGS

[Hel21b] Pat Helland. Don’t get stuck in the “Con” game: Consistency, convergence, and confluence are not the same! eventual consistency and eventual convergence aren’t the same as confluence, either. ACM Queue: Tomorrow’s Computing Today, 19(3):16–35, June 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3475965.3480470.

Helland:2021:FFF

[Hel21c]  Pat Helland.  Fail-fast is

failing… fast!: Changes in compute environments are placing pressure on tried-and-true distributed-systems solutions. ACM Queue: Tomorrow’s Computing Today, 19(1):5–15, February 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3454122.3458812.

Helland:2022:ACW

[Hel22a] Pat Helland. Autonomous computing: We frequently compute across autonomous boundaries but the implications of the patterns to ensure independence are rarely discussed. ACM Queue: Tomorrow’s Computing Today, 20(1):80–104, January 2022. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3526208.

Helland:2022:IPL

[Hel22b] Pat Helland. I’m probably less deterministic than I used to be: Embracing randomness is necessary in cloud environments. ACM Queue: Tomorrow’s Computing Today, 20 (3):5–13, May 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546935.

Henning:2004:MMM

[Hen04] Michi Henning. Massively multiplayer middleware. ACM Queue: Tomorrow’s Computing Today, 1(10):38–45, February 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Henning:2006:RFC

[Hen06] Michi Henning. The rise and fall of CORBA. ACM Queue: Tomorrow’s Computing Today, 4(5):28–34, June 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Henning:2007:ADM

[Hen07] Michi Henning. API design matters. ACM Queue: Tomorrow’s Computing Today, 5(4): 24–36, May–June 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Henderson:2008:RBD

[Hen08] David B. Henderson. Review of Botnet Detection: Countering the Largest Security Threat (Advances in Information Security) Wenke Lee, Cliff Wang, David Dagon, Springer-Verlag New York Inc., 2007, $99.00, ISBN: 0-387-68766-1. ACM Queue: Tomorrow’s Computing Today, 6(4):53, July/August 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Hidgson:2004:BRL

[Hid04] J. P. E. Hidgson. Book review: Linux on the Mainframe by John Eilert, Maria Eisenhaendler, Dorothea Matthaeu, and Ingol Salm, Prentice Hall Professional Technical Reference, 2003, $49.99, ISBN 0-13-101415-3. ACM Queue: Tomorrow’s Computing Today, 2 (1):92, March 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Hildebrand:2003:NIA

[Hil03] Joe Hildebrand. Nine IM accounts and counting. ACM Queue: Tomorrow’s Computing Today, 1(8):44–50, November 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Hill:2008:REI

[Hil08] David G. Hill. Review of Eating the IT Elephant: Moving from Greenfield Development to Brownfield (1st ed.), IBM Press, 2008, $29.99, ISBN 0-13-713012-0. ACM Queue: Tomorrow’s Computing Today, 6(5):59, September 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Hinze:2008:RRB

[Hin08] Annika Hinze. Review of RailsSpace: Building a Social Networking Website with Ruby on Rails, (Addison-Wesley Professional Ruby Series) Michael Hartl, Aurelius Prochazka, Addison-Wesley Professional, 2007, $44.99, ISBN 0-321-48079-1. ACM Queue: Tomorrow’s Computing Today, 6(4): 53, July/August 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Hirvensalo:2004:QC

[Hir04] Mika Hirvensalo. Quantum computing. Natural computing series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-40704-9. xiii + 214 pp. LCCN QA76.889 .H57 2004.

Hopkins:2008:EIE

[HJ08] Richard Hopkins and Kevin Jenkins. Eating the IT elephant: moving from greenfield development to brownfield. IBM Press/Pearson plc, Upper Saddle River, NJ, USA, 2008. ISBN 0-13-713012-0 (paperback). xxx + 222 pp. LCCN T58.5 .H69 2008.

Haber:2010:CSA

[HKM10] Eben M. Haber, Eser Kandogan, and Paul Maglio. Collaboration in system administration. ACM Queue: Tomorrow’s Computing Today, 8(12):10, December 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Hoglund:2004:ESH

[HM04] Greg Hoglund and Gary Mc-Graw. Exploiting software: how to break code. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-201-78695-8 (paperback). xxxv + 471 pp. LCCN QA76.9.A25 H635 2004. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        toc/ecip0411/2003025556.html.

Haskins:2005:SSG

[HN05] Robert Haskins and Dale (Dale P.) Nielsen. Slamming spam: a guide for system administrators. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-13-146716-6. xxiv + 396 pp. LCCN TK5105.743 .H37 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
http://www.slammingspam.
com.

Hoffman:2005:MYS

[Hof05] Bill Hoffman. Monitoring, at your service. ACM Queue: Tomorrow’s Computing Today, 3(10):34–43, December 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Holzner:2004:E

[Hol04]  Steven Holzner.  Eclipse.

O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00641-1. xiv + 317 pp. LCCN QA76.73.J38 H63 2004.

Horvath:2003:BRL

[Hor03] Joan Horvath. Book review:

Lean Software Development: An Agile Toolkit, by Mary Poppendieck and Tom Poppendieck, Addison-Wesley, 2003, $39.99, ISBN 0-321-15078-3. ACM Queue: Tomorrow’s Computing Today, 1(7):90–91, October 2003. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic).

Hunt:2016:RFF

[HOSC16] Pete Hunt, Paul O’Shannessy, Dave Smith, and Terry Coatta. React: Facebook’s functional turn on writing JavaScript. ACM Queue: Tomorrow’s Computing Today, 14(4):40, July 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2994373.

Hartl:2008:RBS

[HP08] Michael Hartl and Aurelius Prochazka. RailsSpace: building a social networking Website with Ruby on Rails. Addison-Wesley professional Ruby series. Addison-Wesley, Reading, MA, USA, 2008. ISBN 0-321-48079-1 (paperback). xxi + 537 pp. LCCN TK5105.888 .H374 2008. URL http:

//www.loc.gov/catdir/toc/
ecip0715/2007011979.html.

Hallem:2003:USD

[HPE03] Seth Hallem, David Park, and Dawson Engler. Uprooting software defects at the source. ACM Queue: Tomorrow’s Computing Today, 1(8):64–71, November 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Hensley:2013:PSD

[HSN13] Zachary Hensley, Jibonananda Sanyal, and Joshua New. Provenance in sensor data management. ACM Queue: Tomorrow’s Computing Today, 11(12): ??, December 2013. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2574836.

Hubbard:2004:OSC

[Hub04] Jordan Hubbard. Open source to the core. ACM Queue: Tomorrow’s Computing Today, 2 (3):24–31, May 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Humble:2017:CDS

[Hum17] Jez Humble. Continuous delivery sounds great, but will it work here? ACM Queue: Tomorrow’s Computing Today, 15 (6):70, November 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3190610.

Helland:2017:TBF

[HWH17] Pat Helland, Simon Weaver, and Ed Harris. Too big NOT to fail. ACM Queue: Tomorrow’s Computing Today, 15(1): 70, January 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3077383.

Hyde:2009:DF

[Hyd09] Julian Hyde. Data in flight.

ACM Queue: Tomorrow’s Computing Today, 7(11):20, December 2009. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic).

Ierusalimschy:2011:PLT

[IdFC11] Roberto Ierusalimschy, Luiz Henrique de Figueiredo, and Waldemar Celes. Passing a language through the eye of a needle. ACM Queue: Tomorrow’s Computing Today, 9(5):20, May 2011. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

itojunHagino:2005:INP

[iiH05] Jun ichiro itojun Hagino. IPv6 network programming. Elsevier/ Digital Press, Amsterdam, The Netherlands; Boston, MA, USA, 2005. ISBN 1-55558-318-0 (paperback). xi + 361 pp. LCCN TK5105.585 .H36 2005. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        description/els051/2004557600.
                                        html.

Imrey:2003:BRC

[Imr03] Lee Imrey. Book review: The Code Book: The Science of Secrecy from Ancient Egypt to Quantum Cryptography by Simon Singh, Delacorte Press, 2000, $15.00, ISBN 0-385-49532-3. ACM Queue: Tomorrow’s Computing Today, 1(9):76, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ingleshwar:2007:UTW

[Ing07] Vikram V. Ingleshwar. Usability testing for the Web. ACM Queue: Tomorrow’s Computing Today, 5(5):34–37, July 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ishida:2004:IBS

[Ish04] Yoshiteru Ishida. Immunitybased systems: a design perspective. Advanced information processing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-00896-9 (hardcover). xii + 177 pp. LCCN QA76.875 .Y67 2004; QA76.875 .I84 2004.

Jacobs:2005:ESS

[Jac05] Dean Jacobs. Enterprise software as service. ACM Queue: Tomorrow’s Computing Today, 3(6):36–42, July 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Jacobs:2009:PBD

[Jac09] Adam Jacobs. The pathologies of big data. ACM Queue: Tomorrow’s Computing Today, 7 (6):1:1–1:10, July 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Jaffe:2006:RPP

[Jaf06] Elliot Jaffe. Review of “Pro Perl” by Peter Wainwright, Apress, 2005, $59.99, ISBN 1-59059-438-X. ACM Queue: Tomorrow’s Computing Today, 4 (2):59, March 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

James:2005:PE

[Jam05] Lance James. Phishing exposed. Syngress Publishing, Inc., Rockland, MA, USA, 2005. ISBN 1-59749-030-X. LCCN HV6773 .J36 2005eb.

                                        URL http://link.library.
                                        utoronto.ca/eir/EIRdetail.
                                        cfm?Resources__ID=232523&
                                        T=resource.

Jania:2003:BMM

[Jan03] Frank Jania. Broadcast messaging: Messaging to the masses. ACM Queue: Tomorrow’s Computing Today, 1(8): 38–43, November 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Jacobson:2023:UCE

[JC23] Ivar Jacobson and Alistair Cockburn. Use cases are essential: Use cases provide a proven method to capture and explain the requirements of a system in a concise and easily understood format. ACM Queue: Tomorrow’s Computing Today, 21(5): 66–86, September 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3631182.

Jenson:2002:SSI

[Jen02] Scott Jenson. The simplicity shift: innovative design tactics in a corporate world. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-52749-X (paperback). viii + 174 pp. LCCN TS171 .J46 2002. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/cam041/2003272889.
html;   http://www.loc.gov/
catdir/samples/cam041/2003272889.
html;   http://www.loc.gov/
catdir/toc/cam041/2003272889.
html.

Jenkins:2005:REN

[Jen05a] G. K. Jenkins. Review of Expert .NET Programming by Simon Robinson, Apress, 2004, $59.99, ISBN 1-59059-222-0. ACM Queue: Tomorrow’s Computing Today, 3(4):60, May 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Jenkins:2005:RLE

[Jen05b] G. K. Jenkins. Review of “The Linux Enterprise Cluster”, by Karl Kopper, No Starch Press, 2005, $49.95 ISBN 1-59327-036-4. ACM Queue: Tomorrow’s Computing Today, 3 (9):58, November 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Jacobi:2006:PJA

[JF06] Jonas Jacobi and John R. Fallows. Pro JSF and Ajax: building rich Internet components. The Expert’s voice in Java. Apress, Berkeley, CA, USA, 2006. ISBN 1-59059-580-7. xxiii + 435 pp. LCCN ????

Jacobson:2014:MLS

[JNSM14] Ivar Jacobson, Pan-Wei Ng, Ian Spence, and Paul E. McMahon. Major-league SEMAT: Why should an executive care? ACM Queue: Tomorrow’s Computing Today, 12(2):20, February 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Johnson:2006:JT

[Joh06] Stephen Johnson. Java in a teacup. ACM Queue: Tomorrow’s Computing Today, 4 (3):36–41, April 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Jovanovic:2006:RRD

[Jov06] Vladan Jovanovic. Review of Refactoring Databases: Evolutionary Database Design by Scott Ambler and Pramodkumar Sadalage, Addison-Wesley Professional, 2006, $49.99, ISBN 0-321-29353-3. ACM Queue: Tomorrow’s Computing Today, 4(7):44, September 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Jover:2020:SAS

[Jov20] Roger Piqueras Jover. Security analysis of SMS as a second factor of authentication: The challenges of multifactor authentication based on SMS, including cellular security deficiencies, SS7 exploits, and SIM swapping. ACM Queue: Tomorrow’s Computing Today, 18 (4):37–60, August 2020. URL

https://dl.acm.org/doi/10.
1145/3424302.3425909.

Joy:2004:RES

[Joy04] M. S. Joy. Review of Eclipse by Steve Holzner. O’Reilly & Associates, 2004, $44.95, ISBN 0-596-00641-1. ACM Queue: Tomorrow’s Computing Today, 2 (9):76, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Johansson:2005:PYW

[JR05] Jesper M. Johansson and Steve Riley. Protect your Windows network: from perimeter to data. Addison-Wesley Microsoft technology series. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-321-33643-7 (paperback). xxvi + 578 pp. LCCN QA76.9.A25 J62 2005.

Jacobson:2014:NSE

[JS14] Ivar Jacobson and Ed Seidewitz. A new software engineering. ACM Queue: Tomorrow’s Computing Today, 12(10): 30, October 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2693160.

Jacobson:2018:TMP

[JS18] Ivar Jacobson and Roly Stimson. Tear down the method prisons! Set free the practices! ACM Queue: Tomorrow’s Computing Today, 16(5): 80, September 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3301760.

Jacobson:2016:UC

[JSK16] Ivar Jacobson, Ian Spence, and Brian Kerr. Use-Case 2.0. ACM Queue: Tomorrow’s Computing Today, 14(1):50, January 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2912151.

Jacobson:2017:TSM

[JSN17] Ivar Jacobson, Ian Spence, and Pan-Wei Ng. Is there a single method for the Internet of Things? ACM Queue: Tomorrow’s Computing Today, 15 (3):20, May 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3123501.

Jacobson:2016:ISA

[JSS16] Ivar Jacobson, Ian Spence, and Ed Seidewitz. Industrial scale agile — from craft to engineering. ACM Queue: Tomorrow’s Computing Today, 14(5): 10, September 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3012428.

Johnson:2024:CCG

[JVG^{+}24] Matthew A. Johnson, Stavros Volos, Ken Gordon, Sean T. Allen, Christoph M. Wintersteiger, Sylvan Clebsch, John Starks, and Manuel Costa. Confidential Container Groups: Implementing confidential computing on Azure container instances. ACM Queue: Tomorrow’s Computing Today, 22(2), March 2024. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664293.

Johnston-Watt:2006:UNM

[JW06] Duncan Johnston-Watt. Under new management. ACM Queue: Tomorrow’s Computing Today, 4(2):50–58, March 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Johnson:2008:EAS

[JWC08] Dave Johnson, Alexei White, and Andre Charland. Enterprise AJAX: strategies for building high performance Web applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2008. ISBN 0-13-224206-0 (paperback). xix + 468 pp. LCCN TK5105.8885.A52 J64 2008. URL http:

//www.loc.gov/catdir/toc/
ecip0715/2007015974.html.

Jerraya:2003:ESS

[JYWV03] Ahmed A. (Ahmed Amine)

Jerraya, Sungjoo Yoo, Norbert When, and Diederik Verkest, editors. Embedded software for SOC. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-7528-6 (hardcover). xix + 530 pp. LCCN QA76.76.D47 E46 2003.

Kleppmann:2022:CRP

[KA22] Martin Kleppmann and Peter Alvaro. Convergence: Research for practice reboot. ACM Queue: Tomorrow’s Computing Today, 20(3):88–95, May 2022. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3546931.

Kaffes:2023:SBS

[KA23] Kostis Kaffes and Peter Alvaro. OS scheduling: Better scheduling policies for modern computing systems. ACM Queue: Tomorrow’s Computing Today, 21 (2):88–95, March 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3595837.

Kamp:2010:YDI

[Kam10] Poul-Henning Kamp. You’re doing it wrong. ACM Queue: Tomorrow’s Computing Today, 8(6):20, June 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kamp:2011:BCT

[Kam11a] Poul-Henning Kamp. B.Y.O.C. (1,342 times and counting).

ACM Queue: Tomorrow’s Computing Today, 9(2):20, February 2011. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kamp:2011:OSW

[Kam11b] Poul-Henning Kamp. The onesecond war (what time will you die?). ACM Queue: Tomorrow’s Computing Today, 9 (4):10, April 2011. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kamp:2011:SIP

[Kam11c] Poul-Henning Kamp. The software industry IS the problem: The time has come for software liability laws. ACM Queue: Tomorrow’s Computing Today, 9 (9):10, September 2011. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kamp:2013:CWS

[Kam13a] Poul-Henning Kamp. Center wheel for success. ACM Queue: Tomorrow’s Computing Today, 11(12):10, December 2013. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kamp:2013:MES

[Kam13b] Poul-Henning Kamp. More encryption is not the solution. ACM Queue: Tomorrow’s Computing Today, 11(7):10, July 2013. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kamp:2014:BQS

[Kam14a] Poul-Henning Kamp. The bikeshed: Quality software costs money — Heartbleed was free. ACM Queue: Tomorrow’s Computing Today, 12(6):10, June 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kamp:2014:PPO

[Kam14b] Poul-Henning Kamp. Please put OpenSSL out of its misery. ACM Queue: Tomorrow’s Computing Today, 12(3): 20, March 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kamp:2014:HIP

[Kam15a] Poul-Henning Kamp. HTTP/ 2.0 — the IETF is phoning it in. ACM Queue: Tomorrow’s Computing Today, 13(1): 10, December/January 2014– 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=2716278.

Kamp:2015:HIP

[Kam15b] Poul-Henning Kamp. HTTP/2.0 — the IETF is phoning it in. ACM Queue: Tomorrow’s Computing Today, 13(2):10, February 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=2716278.

Kamp:2016:MEM

[Kam16] Poul-Henning Kamp. More encryption means less privacy. ACM Queue: Tomorrow’s Computing Today, 14(1):30, January 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2904894.

Kamp:2021:SIS

[Kam21a] Poul-Henning Kamp. The software industry IS STILL the problem: The time is (also) way overdue for IT professional liability. ACM Queue: Tomorrow’s Computing Today, 19(4):23–28, July 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487019.3489045.

Kamp:2021:STC

[Kam21b] Poul-Henning Kamp. Surveillance too cheap to meter: Stopping big brother would require an expensive overhaul of the entire system. QUEUE, 19(6): 22–27, November 2021. URL

https://dl.acm.org/doi/10.
1145/3511661.

Kamp:2021:WWW

[Kam21c] Poul-Henning Kamp. What went wrong?: Why we need an IT accident investigation board. ACM Queue: Tomorrow’s Computing Today, 19(3):40–44, June 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3475965.3475967.

Kamp:2022:COO

[Kam22a] Poul-Henning Kamp. CSRB’s Opus One: Comments on the Cyber Safety Review Board Log4j Event Report. ACM Queue: Tomorrow’s Computing Today, 20(4):5–10, July 2022. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3561797.

Kamp:2022:LAS

[Kam22b] Poul-Henning Kamp. Linear address spaces: Unsafe at any speed. ACM Queue: Tomorrow’s Computing Today, 20 (2):5–10, March 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3534854.

Kamp:2023:DTI

[Kam23] Poul-Henning Kamp. Don’t “Think of the Internet!”: No human right is absolute. ACM Queue: Tomorrow’s Computing Today, 21(3):17–22, May 2023. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3606023.

Kamp:2024:FOS

[Kam24] Poul-Henning Kamp. Free and open source software — and other market failures: Open source is not a goal as much as a means to an end. ACM Queue: Tomorrow’s Computing Today, 22(1):10–16, January 2024. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3649298.

Kaplan:2023:HVI

[Kap23] David Kaplan. Hardware VM isolation in the cloud: Enabling confidential computing with AMD SEV-SNP technology. ACM Queue: Tomorrow’s Computing Today, 21 (4):49–67, July 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3623392.

Karels:2003:COS

[Kar03]  Michael J. Karels.  Com-

mercializing open source software. ACM Queue: Tomorrow’s Computing Today, 1(5): 46–55, July/August 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Karl:2005:ROE

[Kar05] Neil Karl. Review of “Official Eclipse 3.0 FAQs” by John Arthorne and Chris Laffra, Addison-Wesley Professional, 2004, $34.99 ISBN 0-321-26838-5. ACM Queue: Tomorrow’s Computing Today, 3(8):60, October 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kaujalgi:2003:BRB

[Kau03] V. B. Kaujalgi. Book review: The Bug by Ellen Ullman, Nan A. Talese/Doubleday, 2003, $23.95, ISBN 0-385-50860-3. ACM Queue: Tomorrow’s Computing Today, 1(8): 72, November 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2004:CDD

[KB04a] Stan Kelly-Bootle. Curmudgeon: Damnéd digits. ACM Queue: Tomorrow’s Computing Today, 2(1):96, 95, March 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2004:CM

[KB04b] Stan Kelly-Bootle. Curmudgeon: From this moment on. ACM Queue: Tomorrow’s Computing Today, 2(4):80, 79, June 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2004:CLF

[KB04c] Stan Kelly-Bootle. Curmudgeon: Linguae francae. ACM Queue: Tomorrow’s Computing Today, 2(9):80, 78–79, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2004:CVE

[KB04d] Stan Kelly-Bootle. Curmudgeon: Vote early, vote often. ACM Queue: Tomorrow’s Computing Today, 2(6):

80–79, September 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2004:CNS

[KB04e] Stan Kelly-Bootle. Curmudgeon: Without a NULL that string would never end. ACM Queue: Tomorrow’s Computing Today, 2(5):80, 79, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2005:CAS

[KB05a] Stan Kelly-Bootle. Curmudgeon: Anything su doku, I can do better. ACM Queue: Tomorrow’s Computing Today, 3(10): 56, 54, December 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2005:FUU

[KB05b] Stan Kelly-Bootle. Curmudgeon: File under “unknowable!”. ACM Queue: Tomorrow’s Computing Today, 3(3): 64, 63, April 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2005:CCG

[KB05c]  Stanley Kelly-Bootle.  Cur-

mudgeon: Call that gibberish. ACM Queue: Tomorrow’s Computing Today, 3(6):64, 62–63, July 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2006:CHS

[KB06a] Stan Kelly-Bootle. Curmudgeon: But, having said that, . . .. ACM Queue: Tomorrow’s Computing Today, 4(2): 64, 63, March 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2006:CLP

[KB06b] Stan Kelly-Bootle. Curmudgeon: Like a podcast in the sea: mean ol’ LoTech blues. ACM Queue: Tomorrow’s Computing Today, 4(6):64, 62–63, July/August 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2006:CCF

[KB06c] Stan Kelly-Bootle. Curmudgeon: The calculus formally known as pi. ACM Queue: Tomorrow’s Computing Today, 4 (4):56, 54–55, May 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2006:CYC

[KB06d] Stan Kelly-Bootle. Curmudgeon: You can look it up: or maybe not. ACM Queue: Tomorrow’s Computing Today, 4 (8):56, 54–55, October 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2006:CWR

[KB07a] Stan Kelly-Bootle. Curmudgeon: Will the real bots stand up? ACM Queue: Tomorrow’s Computing Today, 4(10): 56, 52–55, December/January 2006–2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2007:CA

[KB07b] Stan Kelly-Bootle. Curmudgeon: Alloneword. ACM Queue: Tomorrow’s Computing Today, 5(4):64, 61–63, May–June 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2007:CBB

[KB07c] Stan Kelly-Bootle. Curmudgeon: As big as a barn? ACM Queue: Tomorrow’s Computing Today, 5(2):64, 62–63, March 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2007:COC

[KB07d] Stan Kelly-Bootle. Curmudgeon: Ode or code?: programmers be used! ACM Queue: Tomorrow’s Computing Today, 5(3):52, 48–51, April 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2007:CSS

[KB07e] Stan Kelly-Bootle. Curmudgeon: Some swans are black. ACM Queue: Tomorrow’s Computing Today, 5(5):52, 49–51, July 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2007:CCD

[KB07f] Stan Kelly-Bootle. Curmudgeon: The code delusion. ACM Queue: Tomorrow’s Computing Today, 5(6):52, 49–51, September 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2007:CUI

[KB07g] Stan Kelly-Bootle. Curmudgeon: Use it or lose it. ACM Queue: Tomorrow’s Computing Today, 5(7):56, 53–55, November 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2008:CAR

[KB08a] Stan Kelly-Bootle. Curmudgeon: Affine romance. ACM Queue: Tomorrow’s Computing Today, 6(5):61–63, September 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2008:CAT

[KB08b] Stan Kelly-Bootle. Curmudgeon: All things being equal? ACM Queue: Tomorrow’s Computing Today, 6(1):56, 54– 55, January 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2008:CSS

[KB08c] Stan Kelly-Bootle. Curmudgeon: Solomon’s sword beats Occam’s razor. ACM Queue: Tomorrow’s Computing Today, 6(2):72, 69–71, March 2008.

CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2008:CFR

[KB08d] Stan Kelly-Bootle. Curmudgeon: The fabrication of reality. ACM Queue: Tomorrow’s Computing Today, 6(4): 56–58, July/August 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2008:CTL

[KB08e] Stan Kelly-Bootle. Curmudgeon: There’s a lot of it about. ACM Queue: Tomorrow’s Computing Today, 6(3):72–ff, May/ June 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2009:COP

[KB09a] Stan Kelly-Bootle. Curmudgeon: One peut-Être, two peut-Être, three peut-Être, more. ACM Queue: Tomorrow’s Computing Today, 7(2): 77–80, February 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2009:SRR

[KB09b] Stan Kelly-Bootle. Curmudgeon: Some rules and restrictions may apply. ACM Queue: Tomorrow’s Computing Today, 7(11):1–3, December 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2009:CFN

[KB09c] Stan Kelly-Bootle. Curmudgeon: The flaws of nature. ACM Queue: Tomorrow’s Computing Today, 7(1):42–45, January 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2010:CFU

[KB10] Stan Kelly-Bootle. Curmudgeon: Facing an uncertain past. ACM Queue: Tomorrow’s Computing Today, 8(9):40, September 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly-Bootle:2011:TBA

[KB11] Stan Kelly-Bootle. Two books alike in dignity. ACM Queue: Tomorrow’s Computing Today, 9(2):10, February 2011. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kleppmann:2019:OEP

[KBS19] Martin Kleppmann, Alastair R. Beresford, and Boerge Svingen. Online event processing. ACM Queue: Tomorrow’s Computing Today, 17(1): 40, January 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3321612.

Kellogg:2003:E

[Kel03] Wendy A. Kellogg. From the Editors. ACM Queue: Tomorrow’s Computing Today, 1 (8):5, November 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly:2019:PMP

[Kel19] Terence Kelly. Persistent memory programming on conventional hardware. ACM Queue: Tomorrow’s Computing Today, 17(4):10, July 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3358957.

Kelly:2020:DC

[Kel20a] Terence Kelly. Decentralized computing. ACM Queue: Tomorrow’s Computing Today, 18 (5):41–53, October 2020. URL

https://dl.acm.org/doi/10.
1145/3434571.3436964.

Kelly:2020:EGS

[Kel20b] Terence Kelly. Efficient graph search. ACM Queue: Tomorrow’s Computing Today, 18(4): 25–36, August 2020. URL

https://dl.acm.org/doi/10.
1145/3424302.3424304.

Kelly:2020:PMP

[Kel20c] Terence Kelly. Is persistent memory persistent? ACM Queue: Tomorrow’s Computing Today, 18(2):??, March 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3400899.3400902.

Kelly:2020:OAL

[Kel20d] Terence Kelly. Offline algorithms in low-frequency trading: Clearing combinatorial auctions. ACM Queue: Tomorrow’s Computing Today, 18(6): 37–51, December 2020. URL

https://dl.acm.org/doi/10.
1145/3442632.3448307.

Kelly:2021:CON

[Kel21a]  Terence Kelly.  Crashproof-

ing the original NoSQL key– value store: an upgrade for the gdbm database. ACM Queue: Tomorrow’s Computing Today, 19(4):5–18, July 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487019.3487353.

Kelly:2021:SML

[Kel21b] Terence Kelly. Steampunk machine learning: Victorian contrivances for modern data science. QUEUE, 19(6):5–17, November 2021. URL https:

//dl.acm.org/doi/10.1145/
3511543.

Kelly:2022:LE

[Kel22a] Terence Kelly. Literate executables. ACM Queue: Tomorrow’s Computing Today, 20(5): 20–32, September 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3570938.

Kelly:2022:PMA

[Kel22b] Terence Kelly. Persistent memory allocation: Leverage to move a world of software. ACM Queue: Tomorrow’s Computing Today, 20(2):16–30, March 2022. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3534855.

Kelly:2023:PJI

[Kel23a] Terence Kelly. Programmer job interviews: The hidden agenda. ACM Queue: Tomorrow’s Computing Today, 21(6): 16–26, November 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3639444.

Kelly:2023:PSC

[Kel23b] Terence Kelly. Protecting secrets from computers. ACM Queue: Tomorrow’s Computing Today, 21(4):23–43, July 2023. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3623614.

Kelly:2024:ZTB

[Kel24]  Terence Kelly.  Zero toler-

ance for bias. ACM Queue: Tomorrow’s Computing Today, 22(2), March 2024. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664645.

Kern:2014:STW

[Ker14] Christoph Kern. Securing the tangled Web. ACM Queue: Tomorrow’s Computing Today, 12(7):40, July 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kern:2024:DES

[Ker24] Christoph Kern. Developer ecosystems for software safety: Continuous assurance at scale. ACM Queue: Tomorrow’s Computing Today, 22(1):73–99, January 2024. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3648601.

Kfir:2019:DCL

[KF19] Shaul Kfir and Camille Fournier. DAML: The contract language of distributed ledgers. ACM Queue: Tomorrow’s Computing Today, 17(3):70, May 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3357728.

Kelly:2021:SAD

[KGM21] Terence Kelly, Weiwei Gu, and Vladimir Maksimovski. Schrödinger’s code: Undefined behavior in theory and practice. ACM Queue: Tomorrow’s Computing Today, 19(2):28–44, April 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3466132.3468263.

Kobayashi-Hillary:2005:PI

[KH05] Mark Kobayashi-Hillary. A passage to India. ACM Queue: Tomorrow’s Computing Today, 3 (1):54–60, February 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Khan:2005:RSR

[Kha05] Sajjad Khan. Review of “Struts Recipes by George Franciscus and Danilo Gurovich”, Manning Publications, 2004, $44.95 ISBN 1-932394-24-9. ACM Queue: Tomorrow’s Computing Today, 3(6):60, July 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kilov:2004:BRS

[Kil04] H. I. Kilov. Book review:

The Simplicity Shift: Innovative Design Tactics in a Corporate World by Scott Jenson, Cambridge University Press, 2002, $25, ISBN 0-521-52749-X. ACM Queue: Tomorrow’s Computing Today, 2(1):92, March 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Killalea:2008:BSW

[Kil08a] Tom Killalea. Building scalable Web services. ACM Queue: Tomorrow’s Computing Today, 6 (6):10–13, October 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Killalea:2008:MV

[Kil08b] Tom Killalea. Meet the virts.

ACM Queue: Tomorrow’s Computing Today, 6(1):14–18, January 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Killalea:2016:HDM

[Kil16] Tom Killalea. The hidden dividends of microservices. ACM Queue: Tomorrow’s Computing Today, 14(3):10, May 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=2956643.

Killalea:2019:VSE

[Kil19] Tom Killalea. Velocity in software engineering. ACM Queue: Tomorrow’s Computing Today, 17(3):40, May 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3352692.

Killalea:2020:SCW

[Kil20] Tom Killalea. A second conversation with Werner Vogels: The Amazon CTO sits with Tom Killalea to discuss designing for evolution at scale. ACM Queue: Tomorrow’s Computing Today, 18(5):67–92, October 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3434571.3434573.

Kirsch:2004:ICM

[Kir04] Steve Kirsch. Interview: a conversation with Matt Wells. ACM Queue: Tomorrow’s Computing Today, 2(2):18, 20–24, April 2004. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic).

Kitson:2005:MMM

[Kit05] Fred Kitson. Mobile media:

making it a reality. ACM Queue: Tomorrow’s Computing Today, 3(4):38–47, May 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kleiman:2003:DNH

[Kle03] Steve Kleiman. DAFS: a new high-performance networked file system. ACM Queue: Tomorrow’s Computing Today, 1 (4):54–62, June 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kleen:2014:SEL

[Kle14] Andi Kleen. Scaling existing lock-based applications with lock elision. ACM Queue: Tomorrow’s Computing Today, 12 (1):20, January 2014. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kaushik:2023:RHS

[KLL23] Divyansh Kaushik, Zachary C.

Lipton, and Alex John London. Resolving the human-subjects status of machine learning’s crowdworkers: What ethical framework should govern the interaction of ML researchers and crowdworkers? ACM Queue: Tomorrow’s Computing Today, 21(6):101–127, November 2023.

CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3639452.

Kleinberg:2009:MCT

[KM09] Samantha Kleinberg and Bud Mishra. Metamorphosis: the coming transformation of translational systems biology. ACM Queue: Tomorrow’s Computing Today, 7(9):40, October 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kaczorowski:2022:OSC

[KMNNM22] Maya Kaczorowski, Falcon Momot, George V. Neville-Neil, and Chris McCubbin. OSS supply-chain security: What will it take? ACM Queue: Tomorrow’s Computing Today, 20(5):86–102, September 2022. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3570923.

Kumar:2013:HMI

[KNR13] Arun Kumar, Feng Niu, and Christopher Ré. Hazy: Making it easier to build and maintain big-data analytics. ACM Queue: Tomorrow’s Computing Today, 11(1):1–17, January 2013. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kohlhepp:2004:RND

[Koh04]  Bayard Kohlhepp.  Review

of Network Distributed Computing: Fitscapes and Fallacies by Max K. Goff. Prentice Hall Professional Technical Reference, 2003, ISBN 0-13-100152-3. ACM Queue: Tomorrow’s Computing Today, 2 (8):70, November 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kohlhepp:2006:RWL

[Koh06]  Bayard Kohlhepp.  Review

of “Windows and Linux Integration: Hands-on Solutions for a Mixed Environment by Jermy Moskowitz and Thomas Boutell,” Sybex Inc., 2005, $59.99, ISBN 0-7821-4428-4. ACM Queue: Tomorrow’s Computing Today, 4(6): 58, July/August 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [MB05].

Kolawa:2004:ODG

[Kol04] Adam Kolawa. Outsourcing:

devising a game plan. ACM Queue: Tomorrow’s Computing Today, 2(8):56–62, November 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kopper:2005:LEC

[Kop05] Karl Kopper. The Linux Enterprise Cluster: build a highly available cluster with commodity hardware and free software. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-036-4. xxiv + 430 pp. LCCN QA76.58 .K67 2005.

Kotsovinos:2010:VBC

[Kot10] Evangelos Kotsovinos. Virtualization: Blessing or curse? ACM Queue: Tomorrow’s Computing Today, 8(11):40, November 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kelly:2023:CNC

[KP23] Terence Kelly and Yekai Pan. Catch-23: The new C standard sets the world on fire. ACM Queue: Tomorrow’s Computing Today, 21(1):12–30, January 2023. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3588242.

Kleijnen:2003:OWS

[KR03] Stans Kleijnen and Srikanth Raju. An open Web services architecture. ACM Queue: Tomorrow’s Computing Today, 1 (1):38–46, March 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kriens:2008:HOC

[Kri08]  Peter Kriens.  How OSGi

changed my life. ACM Queue: Tomorrow’s Computing Today, 6(1):44–51, January 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kromhout:2017:CWF

[Kro17] Bridget Kromhout. Containers will not fix your broken culture (and other hard truths). ACM Queue: Tomorrow’s Computing Today, 15(6): 50, November 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3185224.

Kruchten:2007:VAM

[Kru07] Philippe Kruchten. Voyage in the agile memeplex. ACM Queue: Tomorrow’s Computing Today, 5(5):38–44, July 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Keith:2008:ECO

[KS08] Michael Keith and Randy Stafford. Exposing the cache ORM. ACM Queue: Tomorrow’s Computing Today, 6(3): 38–47, May/June 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kerpelman:2004:LKV

[KTS^{+}04] Charles Kerpelman, The Editors, Martin Smith, Harry Rudin, Yaacov Fenster, Marcus Ranum, Scott Howser, and Alexander Wolfe. Letters: Kode Vicious wants you; deliman’s da man; in want of nothing; the root of the problem, redux; Longhorn and WinFS. ACM Queue: Tomorrow’s Computing Today, 2(9):12, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kuc:2006:ROS

[Kuc06]  Bernard Kuc.  Review of

“Open Source Wireless Java Tools Suite by Robert Virkus,” Apress, 2005, $39.99, ISBN 1-59059-503-3. ACM Queue: Tomorrow’s Computing Today, 4 (4):48, May 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Vir05].

Kumar:2005:WSI

[Kum05] Pankaj Kumar. Web services and IT management. ACM Queue: Tomorrow’s Computing Today, 3(6):44–49, July 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kuzmanovic:2019:NNU

[Kuz19] Aleksander Kuzmanovic. Net neutrality: Unexpected solution to blockchain scaling. ACM Queue: Tomorrow’s Computing Today, 17(1):20, January 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=3319534.

Kamp:2004:BSS

[KW04] Poul-Henning Kamp and Robert Watson. Building systems to be shared, securely. ACM Queue: Tomorrow’s Computing Today, 2(5):42–51, July 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kwon:2017:POC

[KW17] Albert Kwon and James Wilcox.

Private online communication; highlights in systems verification. ACM Queue: Tomorrow’s Computing Today, 15 (4):70, July 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3149411.

Labiod:2008:WAH

[Lab08] Houda Labiod, editor. Wireless ad hoc and sensor networks. The expert’s voice in Web development. ISTE, London, UK, 2008. ISBN 1-84821-003-5. xii + 317 pp. LCCN ????

Lam:2018:URC

[Lam18] Alpha Lam. Using remote cache service for Bazel. ACM Queue: Tomorrow’s Computing Today, 16(4):60, July 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3287302.

Laplante:2003:OSD

[Lap03] Phillip A. Laplante. Opinion: Stand and deliver: Why I hate stand-up meetings. ACM Queue: Tomorrow’s Computing Today, 1(7):7–9, October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Laplante:2004:BBD

[Lap04a] Phillip Laplante. The burning bag of dung: and other environmental antipatterns. ACM Queue: Tomorrow’s Computing Today, 2(7):80–??, October 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Laplante:2004:OFD

[Lap04b] Phillip A. Laplante. Opinion: First, do no harm: a Hippocratic oath for software developers? ACM Queue: Tomorrow’s Computing Today, 2(4):14–18, June 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Laplante:2005:MMF

[Lap05] Phillip Laplante. Curmudgeon: Mal managerium: a field guide. ACM Queue: Tomorrow’s Computing Today, 3(4):64, 62–63, May 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Laplante:2006:CII

[Lap06a] Phillip Laplante. Curmudgeon: It isn’t your father’s realtime anymore. ACM Queue: Tomorrow’s Computing Today, 4 (1):64, 63, February 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Laplante:2006:CJS

[Lap06b] Phillip A. Laplante. Curmudgeon: The joy of spam. ACM Queue: Tomorrow’s Computing Today, 4(9):56, 54–55, November 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Larson:2008:ECP

[Lar08] Jim Larson. Erlang for concurrent programming. ACM Queue: Tomorrow’s Computing Today, 6(5):18–23, September 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Laszlo:2006:RCO

[Las06]  Michael Laszlo.  Review

of “Component-Oriented Programming by Andy Ju An Wang and Kai Qian,” Wiley-Interscience, 2005, $89.95, ISBN 0-471-64446-3. ACM Queue: Tomorrow’s Computing Today, 4(5):59, June 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [WQ05].

Laurie:2004:NF

[Lau04] Ben Laurie. Network forensics. ACM Queue: Tomorrow’s Computing Today, 2(4):50–56, June 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Laurie:2014:CT

[Lau14] Ben Laurie. Certificate transparency. ACM Queue: Tomorrow’s Computing Today, 12 (8):10, August 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Lauinger:2018:TSD

[LCW18] Tobias Lauinger, Abdelberi Chaabane, and Christo Wilson. Thou shalt not depend on me. ACM Queue: Tomorrow’s Computing Today, 16(1):50, January 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=3205288.

Lear:2011:SAS

[Lea11] Christina Lear. System administration soft skills. ACM Queue: Tomorrow’s Computing Today, 9(1):10, January 2011. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Lecarme:2007:RPD

[Lec07] O. Lecarme. Review of ‘Pro Drupal Development’, by John VanDyk and Matt Westgate, Apress 2007, $44.99, ISBN 1-59059-755-9. ACM Queue: Tomorrow’s Computing Today, 5 (7):51, November 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [VW07].

Lehey:2003:CSF

[Leh03] Greg Lehey. Closed source fights back: SCO vs. the world — what were they thinking? ACM Queue: Tomorrow’s Computing Today, 1(5):56–62, July/August 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Leighton:2008:IPI

[Lei08] Tom Leighton. Improving performance on the Internet. ACM Queue: Tomorrow’s Computing Today, 6(6):20–29, October 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Leventhal:2008:FST

[Lev08] Adam Leventhal. Flash storage today. ACM Queue: Tomorrow’s Computing Today, 6(4): 24–30, July/August 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Leventhal:2009:TPR

[Lev09] Adam Leventhal. Triple-parity RAID and beyond. ACM Queue: Tomorrow’s Computing Today, 7(11):30, December 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Levine:2020:CHS

[Lev20] Edlyn V. Levine. The die is cast: Hardware security is not assured. ACM Queue: Tomorrow’s Computing Today, 18(4): 95–109, August 2020. URL

https://dl.acm.org/doi/10.
1145/3424302.3431245.

Levine:2023:CCA

[Lev23] Edlyn V. Levine. Cargo cult AI: Is the ability to think scientifically the defining essence of intelligence? ACM Queue: Tomorrow’s Computing Today, 21(2):66–87, March 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3595860.

Lloyd:2014:DSE

[LFKA14] Wyatt Lloyd, Michael J. Freedman, Michael Kaminsky, and David G. Andersen. Don’t settle for eventual consistency. ACM Queue: Tomorrow’s Computing Today, 12(3):30, March 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Leqi:2021:WCB

[LHML21] Liu Leqi, Dylan Hadfield-Menell, and Zachary C. Lipton. When curation becomes creation: Algorithms, microcontent, and the vanishing distinction between platforms and creators. ACM Queue: Tomorrow’s Computing Today, 19(3):11–15, June 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3475965.3477229.

Lhotka:2006:EVB

[Lho06] Rockford Lhotka. Expert VB 2005 business objects. Apress, Berkeley, CA, USA, second edition, 2006. ISBN 6610724849 (electronic book), 1-59059-631-5 (paperback), 1-4302-0165-7 (electronic book). xxiv + 670 pp. LCCN QA76.73 .B3 INTERNET; QA76.73.B3 L54 2006eb. URL http:

//ezproxy.lib.ucalgary.ca:
2048/login?url=http://www.
myilibrary.com?id=72484;
http://ezproxy.lib.ucalgary.
ca:2048/login?url=http://
www.springerlink.com/openurl.
asp?genre=book&isbn=978-1-
59059-631-9.

Limoncelli:2010:PSV

[Lim10] Thomas A. Limoncelli. A plea to software vendors from sysadmins — 10 do’s and don’ts. ACM Queue: Tomorrow’s Computing Today, 8(12):30, December 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Limoncelli:2015:ASL

[Lim15a] Thomas A. Limoncelli. Automation should be like Iron Man, not Ultron. ACM Queue: Tomorrow’s Computing Today, 13(8):50, September/October 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2841313.

Limoncelli:2015:HSD

[Lim15b] Thomas A. Limoncelli. How sysadmins devalue themselves. ACM Queue: Tomorrow’s Computing Today, 13(9):70, November/December 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2891413.

Limoncelli:2016:OLS

[Lim16a] Thomas A. Limoncelli. 10 optimizations on linear search. ACM Queue: Tomorrow’s Computing Today, 14(4):10, July 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=2984631.

Limoncelli:2016:YLB

[Lim16b] Thomas A. Limoncelli. Are you load balancing wrong? ACM Queue: Tomorrow’s Computing Today, 14(6):10, November 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=3028689.

Limoncelli:2016:SBP

[Lim16c] Thomas A. Limoncelli. The small batches principle. ACM Queue: Tomorrow’s Computing Today, 14(2):60, March 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=2945077.

Limoncelli:2017:FWM

[Lim17a] Thomas A. Limoncelli. Four ways to make CS&IT curricula more immersive. ACM Queue: Tomorrow’s Computing Today, 15(3):50, May 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3130216.

Limoncelli:2017:OEA

[Lim17b] Thomas A. Limoncelli. Operational excellence in April Fools’ pranks. ACM Queue: Tomorrow’s Computing Today, 15(5):60, September/October 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3171529.

Limoncelli:2018:GPM

[Lim18a] Thomas A. Limoncelli. GitOps: a path to more selfservice IT. ACM Queue: Tomorrow’s Computing Today, 16 (3):20, May 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3237207.

Limoncelli:2018:MWB

[Lim18b] Thomas A. Limoncelli. Manual work is a bug. ACM Queue: Tomorrow’s Computing Today, 16 (1):20, January 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3197520.

Limoncelli:2018:SNE

[Lim18c] Thomas A. Limoncelli. SQL is no excuse to avoid DevOps. ACM Queue: Tomorrow’s Computing Today, 16(5): 60, September 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3300018.

Limoncelli:2019:API

[Lim19a] Thomas A. Limoncelli. API practices if you hate your customers. ACM Queue: Tomorrow’s Computing Today, 17(5): 50, September 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3375635.

Limoncelli:2019:DDC

[Lim19b] Thomas A. Limoncelli. Demo data as code. ACM Queue: Tomorrow’s Computing Today, 17(3):50, May 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3355565.

Limoncelli:2019:TTT

[Lim19c] Thomas A. Limoncelli. Tom’s top ten things executives should know about software. ACM Queue: Tomorrow’s Computing Today, 17(1):70, January 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=3325792.

Limoncelli:2020:CUN

[Lim20a] Thomas A. Limoncelli. Communicate using the numbers 1, 2, 3, and more. ACM Queue: Tomorrow’s Computing Today, 18 (1):??, January 2020. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3387945.3387947.

Limoncelli:2020:FNR

[Lim20b] Thomas A. Limoncelli. Five nonobvious remote work techniques: Emulating the efficiency of in-person conversations. ACM Queue: Tomorrow’s Computing Today, 18 (3):29–38, June 2020. URL

https://dl.acm.org/doi/10.
1145/3411757.3417752.

Limoncelli:2020:TSB

[Lim20c] Thomas A. Limoncelli. The time I stole $10,000 from Bell Labs: Or why DevOps encourages us to celebrate outages. ACM Queue: Tomorrow’s Computing Today, 18(5): 26–35, October 2020. URL

https://dl.acm.org/doi/10.
1145/3434571.3434773.

Limoncelli:2022:SYO

[Lim22] Thomas A. Limoncelli. Split your overwhelmed teams: Two teams of five is not the same as one team of ten. ACM Queue: Tomorrow’s Computing Today, 20(5):5–13, September 2022. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3570920.

Limoncelli:2023:IEE

[Lim23a] Thomas A. Limoncelli. Improvement on end-to-end encryption may lead to silent revolution: Researchers are on a brink of what could be the next big improvement in communication privacy. ACM Queue: Tomorrow’s Computing Today, 21 (2):10–13, March 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3590144.

Limoncelli:2023:KWY

[Lim23b] Thomas A. Limoncelli. Knowing what you neeed to know: Personal, team, and organizational effectiveness can be improved with a little preparation. ACM Queue: Tomorrow’s Computing Today, 21(4):5–17, July 2023. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3623425.

Limoncelli:2023:WDT

[Lim23c] Thomas A. Limoncelli. What do trains, horses, and home Internet installation have in common?: Avoid changes midprocess. ACM Queue: Tomorrow’s Computing Today, 21(6): 5–11, November 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3639447.

Limoncelli:2024:MTT

[Lim24] Thomas A. Limoncelli. Make two trips: Larry David’s New Year’s resolution works for IT too. ACM Queue: Tomorrow’s Computing Today, 22(2), March 2024. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3664275.

Lipton:2018:MMI

[Lip18] Zachary C. Lipton. The mythos of model interpretability. ACM Queue: Tomorrow’s Computing Today, 16(3):30, May 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3241340.

Liu:2014:JNU

[Liu14] Alex Liu. JavaScript and the Netflix user interface. ACM Queue: Tomorrow’s Computing Today, 12(9):20, September 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Llanos:2007:RGC

[Lla07] Diego R. Llanos. Review of ‘Grid Computing Security’ by Anirban Chakrabarti, Springer, 2007, $69.96, ISBN 3-540-44492-0. ACM Queue: Tomorrow’s Computing Today, 5(6): 45, September 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Cha07].

Lloret:2006:RSS

[Llo06] Jose Lloret. Review of “Securing Storage: A Practical Guide to SAN and NAS Security by Himanshu Dwivedi,” Addison-Wesley Professional, 2005, $49.99, ISBN 0-321-34995-4. ACM Queue: Tomorrow’s Computing Today, 4(4):50, May 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Dwi06].

Laplante:2004:ODW

[LN04] Phillip A. Laplante and Colin J. Neill. Opinion: The demise of the waterfall model is imminent. ACM Queue: Tomorrow’s Computing Today, 1(10): 10–15, February 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Lutz:2014:USE

[LNV14] Michael J. Lutz, J. Fernando Naveda, and James R. Vallino. Undergraduate software engineering: Addressing the needs of professional software development. ACM Queue: Tomorrow’s Computing Today, 12 (6):30, June 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Lodding:2004:HGB

[Lod04] Kenneth N. Lodding. The hitchhiker’s guide to biomorphic software. ACM Queue: Tomorrow’s Computing Today, 2(4):66–75, June 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Loh:2010:IHP

[Loh10] Eugene Loh. The ideal HPC programming language. ACM Queue: Tomorrow’s Computing Today, 8(6):30, June 2010. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Longoria:2004:DSM

[Lon04] Roman Longoria. Designing software for the mobile context: a practitioner’s guide. Computer communications and networks, 1617-7975. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 1-85233-785-0. xvii + 152 pp. LCCN QA76.59 L65 2004.

Long:2014:MID

[Lon15] Dave Long. META II: Digital vellum in the digital scriptorium. ACM Queue: Tomorrow’s Computing Today, 13(1): 50, December/January 2014– 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2724586.

Loveless:2017:CMI

[Lov17] Jacob Loveless. Cache me if you can. ACM Queue: Tomorrow’s Computing Today, 15 (4):30, July 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3136953.

Luo:2024:PME

[LPS24] Queenie Luo, Michael J. Puett, and Michael D. Smith. A “Perspectival” mirror of the elephant: Investigating language bias on Google, Chat-GPT, YouTube, and Wikipedia. ACM Queue: Tomorrow’s Computing Today, 22(1):23–47, January 2024. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3649303.

Lipton:2019:TTM

[LS19] Zachary C. Lipton and Jacob Steinhardt. Troubling trends in machine learning scholarship. ACM Queue: Tomorrow’s Computing Today, 17(1): 80, January 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3328534.

Leung:2017:TIC

[LSB17] Andrew Leung, Andrew Spyker, and Tim Bozarth. Titus: Introducing containers to the Netflix cloud. ACM Queue: Tomorrow’s Computing Today, 15(5):30, September/October 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3158370.

Lucke:2005:BCL

[Luc05]  Robert W. Lucke.  Build-

ing clustered Linux systems. Hewlett–Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-144853-6. xxxviii + 606 pp. LCCN QA76.76.O63 L838 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/ecip0418/2004014016.html.

Lyman:2003:SHM

[LV03] Peter Lyman and Hal R. Varian. Sidebar: How much storage is enough? ACM Queue: Tomorrow’s Computing Today, 1(4):28–29, June 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Loftus:2022:AZT

[LVDM22] Michael Loftus, Andrew Vezina, Rick Doten, and Atefeh Mashatan. The arrival of zero trust: What does it mean? ACM Queue: Tomorrow’s Computing Today, 20 (4):61–79, July 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3561826.

Lee:2007:BDC

[LWD07] Wenke Lee, Cliff Wang, and David Dagon, editors. Botnet detection: countering the largest security threat, volume 36 of Advances in information security. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-68766-1 (hardcover), 0-387-68768-8 (ebook). xii + 167 pp. LCCN TK5105.59 .B68 2008; QA76.9.A25 B64 2008. URL

http://www.loc.gov/catdir/
enhancements/fy0825/2007936179-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0825/
2007936179-t.html.

Lyon:2004:EDY

[Lyo04] Richard F. Lyon. From the Editors: DSP 4 you. ACM Queue: Tomorrow’s Computing Today, 2(1):6, March 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Martin:2009:CCH

[M^{+}09] Robert C. Martin et al., editors. Clean code: a handbook of agile software craftsmanship. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2009. ISBN 0-13-235088-2 (paperback). xxix + 431 pp. LCCN QA76.76.D47 C583 2009.

Mencer:2020:HSF

[MAB^{+}20] Oskar Mencer, Dennis Allison, Elad Blatt, Mark Cummings, Michael J. Flynn, Jerry Harris, Carl Hewitt, Quinn Jacobson, Maysam Lavasani, Mohsen Moazami, Hal Murray, Masoud Nikravesh, Andreas Nowatzyk, Mark Shand, and Shahram Shirazi. The history, status, and future of FPGAs: Hitting a nerve with field-programmable gate arrays. ACM Queue: Tomorrow’s Computing Today, 18 (3):71–82, June 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3411757.3411759.

MacFaden:2003:ECM

[Mac03a] Mike MacFaden. From the Editors: CPUs with 2,000 MIPS per watt, anyone? ACM Queue: Tomorrow’s Computing Today, 1(7):4, October 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Mohay:2003:CIF

[MAC^{+}03b] George M. Mohay, Alison Anderson, Byron Collie, Olivier de Vel, and Rod McKemmish, editors. Computer and intrusion forensics. Artech House computer security series. Artech House Inc., Norwood, MA, USA, 2003. ISBN 1-58053-369-8. xxi + 395 pp. LCCN QA76.9.A25 C628 2003.

Macri:2004:SP

[Mac04] Dean Macri. The scalability problem. ACM Queue: Tomorrow’s Computing Today, 1 (10):66–73, February 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

MacDonald:2008:PWC

[Mac08] Matthew MacDonald. Pro WPF in C# 2008: Windows Presentation Foundation with .NET 3.5. Apress, Berkeley, CA, USA, second edition, 2008. ISBN 1-59059-955-1. xxix + 1040 pp. LCCN ????

Maddox:2015:TDS

[Mad15] Philip Maddox. Testing a distributed system. ACM Queue: Tomorrow’s Computing Today, 13(7):10, July 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2800697.

Maguire:2019:MHC

[Mag19] Laura M. D. Maguire. Managing the hidden costs of coordination. ACM Queue: Tomorrow’s Computing Today, 17(6): ??, November/December 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3380774.3380779.

Martin:2003:ASD

[Mar03] Robert C. Martin. Agile software development: principles, patterns, and practices. Alan Apt series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-597444-

5. xxii + 529 pp. LCCN QA76.76.D47 M362 2003. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        toc/fy035/2002070056.html.

Margolies:2004:ACL

[Mar04a]  D. Margolies.  The ANSI

Common Lisp reference book. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-220-4. 1000 (est.) pp. LCCN ????

Marien:2004:BR

[Mar04b] Andre Mariën. Book review: Security Warrior by Cyrus Peikari and Anton Chuvakin, O’Reilly, 2004, $44.95, ISBN 0-596-00545-

8. ACM Queue: Tomorrow’s Computing Today, 2(6): 76, September 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Marien:2005:RES

[Mar05a] A. Mariën. Review of Exploiting Software: How to Break Code by Greg Hoglund and Gary McGraw, Addison-Wesley, 2004, $49.99, ISBN 0-201-78695-8. ACM Queue: Tomorrow’s Computing Today, 3(3): 60, April 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Marien:2005:RIN

[Mar05b] A. Mariën. Review of IPv6 Network Programming by Junichiro itojun Hagino, Digital Press, 2004, $49.95, ISBN 1-55558-318-0. ACM Queue: Tomorrow’s Computing Today, 3 (5):60, June 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Mark:2008:FGA

[Mar08] William Mark. Future graphics architectures. ACM Queue: Tomorrow’s Computing Today, 6(2):54–64, March 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Mashey:2004:LLL

[Mas04] John R. Mashey. Languages, levels, libraries, and longevity. ACM Queue: Tomorrow’s Computing Today, 2(9):32–38, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Mashey:2006:LRB

[Mas06] John R. Mashey. The long road to 64 bits. ACM Queue: Tomorrow’s Computing Today, 4 (8):24–35, October 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Matsudaira:2013:MMW

[Mat13] Kate Matsudaira. Making the mobile Web faster. ACM Queue: Tomorrow’s Computing Today, 11(1):1–9, January 2013. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Matsudaira:2015:LSD

[Mat15a] Kate Matsudaira. Lean software development — building and shipping two versions. ACM Queue: Tomorrow’s Computing Today, 13(8):40, September/October 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=2841311.

Matsudaira:2015:PAR

[Mat15b] Kate Matsudaira. The paradox of autonomy and recognition. ACM Queue: Tomorrow’s Computing Today, 13(9):80, November/December 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=2893471.

Matsudaira:2015:SMD

[Mat15c] Kate Matsudaira. The science of managing data science. ACM Queue: Tomorrow’s Computing Today, 13 (4):30, April 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2767971.

Matsudaira:2016:BSA

[Mat16a] Kate Matsudaira. Bad software architecture is a people problem. ACM Queue: Tomorrow’s Computing Today, 14 (3):60, May 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2974011.

Matsudaira:2016:DA

[Mat16b] Kate Matsudaira. Delegation as art. ACM Queue: Tomorrow’s Computing Today, 14(1): 70, January 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2926696.

Matsudaira:2016:FS

[Mat16c] Kate Matsudaira. Fresh starts.

ACM Queue: Tomorrow’s Computing Today, 14(4):50, July 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2996549.

Matsudaira:2016:NTD

[Mat16d] Kate Matsudaira. Nine things I didn’t know I would learn being an engineer manager. ACM Queue: Tomorrow’s Computing Today, 14(2):40, March 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2935693.

Matsudaira:2016:RC

[Mat16e] Kate Matsudaira. Resolving conflict. ACM Queue: Tomorrow’s Computing Today, 14(5): 40, September 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3017847.

Matsudaira:2017:WBI

[Mat17a] Kate Matsudaira. 10 ways to be a better interviewer. ACM Queue: Tomorrow’s Computing Today, 15(3):30, May 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3125635.

Matsudaira:2017:BD

[Mat17b] Kate Matsudaira. Breadth and depth. ACM Queue: Tomorrow’s Computing Today, 15 (4):40, July 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3138817.

Matsudaira:2017:CTL

[Mat17c] Kate Matsudaira. Conversations with technology leaders: Erik Meijer. ACM Queue: Tomorrow’s Computing Today, 15 (2):30, March 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3092954.

Matsudaira:2017:DAL

[Mat17d] Kate Matsudaira. Does anybody listen to you? ACM Queue: Tomorrow’s Computing Today, 15(1):20, January 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3056769.

Matsudaira:2017:HYW

[Mat17e] Kate Matsudaira. How is your week going so far? ACM Queue: Tomorrow’s Computing Today, 15(6):40, November 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3184136.

Matsudaira:2017:VT

[Mat17f] Kate Matsudaira. Views from the top. ACM Queue: Tomorrow’s Computing Today, 15(5):20, September/October 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3156692.

Matsudaira:2018:DPM

[Mat18a] Kate Matsudaira. Design patterns for managing up. ACM Queue: Tomorrow’s Computing Today, 16(6):20, November 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=3308563.

Matsudaira:2018:HCG

[Mat18b] Kate Matsudaira. How to come up with great ideas. ACM Queue: Tomorrow’s Computing Today, 16(1):40, January 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=3203214.

Matsudaira:2018:HGT

[Mat18c] Kate Matsudaira. How to get things done when you don’t feel like it. ACM Queue: Tomorrow’s Computing Today, 16 (4):20, July 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3280677.

Matsudaira:2018:IGF

[Mat18d] Kate Matsudaira. The importance of a great finish. ACM Queue: Tomorrow’s Computing Today, 16(5):30, September 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=3293457.

Matsudaira:2018:SFC

[Mat18e] Kate Matsudaira. The secret formula for choosing the right next role. ACM Queue: Tomorrow’s Computing Today, 16 (3):40, May 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3241970.

Matsudaira:2019:EM

[Mat19a] Kate Matsudaira. The evolution of management. ACM Queue: Tomorrow’s Computing Today, 17(3):30, May 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3350548.

Matsudaira:2019:HCG

[Mat19b] Kate Matsudaira. How to create a great team culture (and why it matters). ACM Queue: Tomorrow’s Computing Today, 17 (1):50, January 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3323993.

Matsudaira:2019:OA

[Mat19c] Kate Matsudaira. Overly attached. ACM Queue: Tomorrow’s Computing Today, 17 (2):30, March 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3335594.

Mattioli:2021:FCC

[Mat21] Michael Mattioli. FPGAs in client compute hardware: Despite certain challenges, FP-GAs provide security and performance benefits over ASICs. QUEUE, 19(6):66–88, November 2021. URL https:

//dl.acm.org/doi/10.1145/
3512327.

Matsudaira:2023:MHT

[Mat23a] Kate Matsudaira. Managing hybrid teams: The combination of on-site and remote workers takes extra effort from team leaders. ACM Queue: Tomorrow’s Computing Today, 21 (3):5–16, May 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606015.

Matsudaira:2023:SMG

[Mat23b] Kate Matsudaira. Software managers’ guide to operational excellence: The secret to being a great engineering leader? Setting up the right checks and balances. ACM Queue: Tomorrow’s Computing Today, 21 (5):5–12, September 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3631176.

Matsudaira:2024:GYP

[Mat24] Kate Matsudaira. Give your project a name: It goes a long way toward creating a cohesive team with strong morale. ACM Queue: Tomorrow’s Computing Today, 22(1):5–9, January 2024. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3649296.

Maurer:2005:Ca

[Mau05a] Jim Maurer. Calendar. ACM Queue: Tomorrow’s Computing Today, 3(5):61, June 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maurer:2005:Cb

[Mau05b] Jim Maurer. Calendar. ACM Queue: Tomorrow’s Computing Today, 3(6):61, July 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maurer:2005:ICD

[Mau05c] Jim Maurer. Interview: a conversation with David Anderson. ACM Queue: Tomorrow’s Computing Today, 3(6):18–25, July 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maurer:2005:ICP

[Mau05d] Jim Maurer. Interview: a conversation with Peter Tippett and Steven Hofmeyr. ACM Queue: Tomorrow’s Computing Today, 3(5):22–32, June 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maurer:2005:ICR

[Mau05e] Jim Maurer. Interview: a conversation with Roger Sessions and Terry Coatta. ACM Queue: Tomorrow’s Computing Today, 3(7):16–25, September 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maurer:2005:NOS

[Mau05f] Jim Maurer. News 2.0: Open Source/2; anti-spam activism … or vigilantism?; ride, robot, ride. ACM Queue: Tomorrow’s Computing Today, 3(7): 8, September 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maurer:2005:WYHa

[Mau05g] Jim Maurer. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 3 (5):16, June 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maurer:2005:WYHb

[Mau05h] Jim Maurer. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 3(6):12, July 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maurer:2005:WYHc

[Mau05i] Jim Maurer. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 3 (7):10, September 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maurer:2007:EGF

[Mau07] James Maurer. From the Editors: Going forward. ACM Queue: Tomorrow’s Computing Today, 5(3):8, April 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maurer:2008:ENE

[Mau08a] James Maurer. From the editors: a new era. ACM Queue: Tomorrow’s Computing Today, 6(3):7, May/June 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maurer:2008:EGD

[Mau08b] James Maurer. From the Editors: Going digital. ACM Queue: Tomorrow’s Computing Today, 6(2):7, March 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maurer:2009:ILA

[Mau09] James Maurer. Introducing … acmqueue. ACM Queue: Tomorrow’s Computing Today, 7 (2):5, February 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maurer:2015:FS

[Mau15] Ben Maurer. Fail at scale.

ACM Queue: Tomorrow’s Computing Today, 13(8):30, September/October 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2839461.

Mayor:2003:SCY

[May03] Michael Mayor. Sidebar: I challenge you to respond? ACM Queue: Tomorrow’s Computing Today, 1(6):66–67, September/October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Murmann:2004:DAA

[MB04] Boris Murmann and Bernhard Boser. Digitally assisted analog integrated circuits. ACM Queue: Tomorrow’s Computing Today, 2(1):64–71, March 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Moskowitz:2005:WLI

[MB05] Jeremy Moskowitz and Thomas Boutell. Windows and Linux integration: hands-on solutions for a mixed environment. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2005. ISBN 0-7921-4428-6, 0-7821-4428-4. xix + 539 pp. LCCN ????

Miller:2023:PFM

[MBNN^{+}23] Ethan Miller, Achilles Benetopoulos, George Neville-Neil, Pankaj Mehra, and Daniel Bittman. Pointers in far memory: a rethink of how data and computations should be organized. ACM Queue: Tomorrow’s Computing Today, 21 (3):75–93, May 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606029.

McCallum:2005:IED

[McC05] Andrew McCallum. Information extraction: distilling structured data from unstructured text. ACM Queue: Tomorrow’s Computing Today, 3(9): 48–57, November 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

McCaffrey:2015:VDS

[McC15] Caitie McCaffrey. The verification of a distributed system. ACM Queue: Tomorrow’s Computing Today, 13(9):60, November/December 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2889274.

McDougall:2005:ESS

[McD05] Richard McDougall. Extreme software scaling. ACM Queue: Tomorrow’s Computing Today, 3(7):36–46, September 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

McGraw:2006:SSB

[McG06] Gary McGraw. Software security: building security in. Addison-Wesley software security series. Addison-Wes-ley, Reading, MA, USA, 2006. ISBN 0-321-35670-5 (paperback). xxxvi + 408 pp. LCCN QA76.9.A25 M4286 2006. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip062/2005031598.html.

McKusick:2003:IAB

[McK03] Marshall Kirk McKusick. Interview with Adam Bosworth. ACM Queue: Tomorrow’s Computing Today, 1(1):12, 14– 21, March 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

McKusick:2005:ICT

[McK05] Kirk McKusick. Interview: a conversation with Tim Marsland. ACM Queue: Tomorrow’s Computing Today, 3(4):20–28, May 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

McKusick:2012:DPF

[McK12] Marshall Kirk McKusick. Disks from the perspective of a file system. ACM Queue: Tomorrow’s Computing Today, 10 (9):1–3, September 2012. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

McMullen:2015:IPW

[McM15] Tyler McMullen. It probably works. ACM Queue: Tomorrow’s Computing Today, 13(8):15, September/October 2015. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2855183.

McTear:2004:SDT

[McT04] Michael McTear. Spoken dialogue technology: toward the conversational user interface. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 1-85233-672-2 (softcover). xii + 432 pp. LCCN QA76.9.U83 M45 2004.

Medeiros:2016:DCW

[Med16] Andre Medeiros. Dynamics of change: Why reactivity matters. ACM Queue: Tomorrow’s Computing Today, 14 (3):50, May 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2971330.

Meijer:2014:CEM

[Mei14a] Erik Meijer. The curse of the excluded middle. ACM Queue: Tomorrow’s Computing Today, 12(4):20, April 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Meijer:2014:EJR

[Mei14b]  Erik Meijer.  Evolution of

Javascript revisited. ACM Queue: Tomorrow’s Computing Today, 12(2):50, February 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2014:WLS

[Mei14c] Erik Meijer. Why large scale mobile and e-commerce apps use Node.js. ACM Queue: Tomorrow’s Computing Today, 12 (2):60, February 2014. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Meijer:2017:MMU

[Mei17] Erik Meijer. Making money using math. ACM Queue: Tomorrow’s Computing Today, 15 (1):10, January 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3055303.

Menon:2004:BRB

[Men04] A. K. Menon. Book review: Balancing Agility and Discipline: A Guide for the Perplexed, by Barry Boehm and Richard Turner, Addison-Wesley Professional, 2003, $29.99, ISBN 0-321-18612-5. ACM Queue: Tomorrow’s Computing Today, 2 (5):76, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Menon:2005:ROS

[Men05] A. K. Menon. Review of “Open Systems and Standards for Software Product Development” by P. A. Dargan, Artech House, 2005, $89.00, ISBN 1-58053-749-9. ACM Queue: Tomorrow’s Computing Today, 3(10): 52, December 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Merritt:2004:BRE

[Mer04] Susan M. Merritt. Book review:

Extreme Programming Refactored: The Case Against XP by Matt Stephens and Doug Rosenberg, APress, 2003, $39.99, ISBN 1-590-59096-1. ACM Queue: Tomorrow’s Computing Today, 2(2):81, April 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Millen:2005:SBE

[MFK05] David Millen, Jonathan Feinberg, and Bernard Kerr. Social bookmarking in the enterprise. ACM Queue: Tomorrow’s Computing Today, 3(9): 28–35, November 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Mondal:2022:PPI

[MGL22] Sutapa Mondal, Mangesh S.

Gharote, and Sachin P. Lodha. Privacy of personal information: Going incog in a goldfish bowl. ACM Queue: Tomorrow’s Computing Today, 20 (3):41–87, May 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546934.

Mashatan:2021:CPQ

[MH21] Atefeh Mashatan and Douglas Heintzman. The complex path to quantum resistance: Is your organization prepared? ACM Queue: Tomorrow’s Computing Today, 19(2):65–92, April 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3466132.3466779.

Michaelson:2004:TNS

[Mic04] Jay Michaelson. There’s no such thing as a free (software) lunch. ACM Queue: Tomorrow’s Computing Today, 2(3): 40–47, May 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Miers:2006:BPB

[Mie06] Derek Miers. Best practice (BPM). ACM Queue: Tomorrow’s Computing Today, 4 (2):40–48, March 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Miller:2008:BRP

[Mil08] J. Miller. Book review: Practical API Design: Confessions of a Java Framework Architect (1st edition), Jaroslav Tulach, Apress, 2008, $74.99, ISBN: 978-1-4302-0973-7. ACM Queue: Tomorrow’s Computing Today, 6(6):54, October 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Millsap:2010:TCA

[Mil10] Cary Millsap. Thinking clearly about performance. ACM Queue: Tomorrow’s Computing Today, 8(9):1–10, September 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Mirian:2019:HH

[Mir19] Ariana Mirian. Hack for hire.

ACM Queue: Tomorrow’s Computing Today, 17(4):50, July 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3365458.

Mitchell:2013:LS

[Mit13] Neil Mitchell. Leaking space.

ACM Queue: Tomorrow’s Computing Today, 11(9):1–14, September 2013. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Maglio:2004:EMW

[MK04] Paul P. Maglio and Eser Kandogan. Error messages: what’s the problem? ACM Queue: Tomorrow’s Computing Today, 2 (8):50–55, November 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Meijer:2014:REE

[MK14] Erik Meijer and Vikram Kapoor.

The responsive enterprise: Embracing the hacker way. ACM Queue: Tomorrow’s Computing Today, 12(10):10, October 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=2685692.

Mukherjee:2004:EST

[MM04] Rajat Mukherjee and Jianchang Mao. Enterprise search: Tough stuff. ACM Queue: Tomorrow’s Computing Today, 2(2):36–46, April 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Meijer:2015:SDS

[MMB15] Erik Meijer, Kevin Millikin, and Gilad Bracha. Spicing up Dart with side effects. ACM Queue: Tomorrow’s Computing Today, 13(3):40, March 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=2747873.

McKusick:2004:TSF

[MNN04] Marshall Kirk McKusick and George V. Neville-Neil. Thread scheduling in FreeBSD 5.2. ACM Queue: Tomorrow’s Computing Today, 2(7):58–64, October 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Montagne:2010:TAC

[Mon10]  Kevin Montagne.  Tackling

architectural complexity with modeling. ACM Queue: Tomorrow’s Computing Today, 8 (9):30, September 2010. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Morrison:2004:BR

[Mor04] Jim Morrison. Blaster revisited. ACM Queue: Tomorrow’s Computing Today, 2(4):34–43, June 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Moreno:2006:RAE

[Mor06] A. Moreno. Review of “Agile Estimating and Planning by Mike Cohn,” Prentice Hall PTR, 2005, $44.99, ISBN 0-13-147941-5. ACM Queue: Tomorrow’s Computing Today, 4 (5):59, June 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Coh06].

Morrison:2016:SSM

[Mor16] Adam Morrison. Scaling synchronization in multicore programs. ACM Queue: Tomorrow’s Computing Today, 14 (4):20, July 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2991130.

Mor:2018:EC

[Mor18] Nitesh Mor. Edge computing. ACM Queue: Tomorrow’s Computing Today, 16(6): 60, November 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3313377.

McKusick:2009:CSG

[MQ09] Marshall Kirk McKusick and Sean Quinlan. Case study: GFS: Evolution on fast-forward. ACM Queue: Tomorrow’s Computing Today, 7(7):10, August 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Molino:2021:DML

[MR21] Piero Molino and Christopher Ré. Declarative machine learning systems: The future of machine learning will depend on it being in the hands of the rest of us. ACM Queue: Tomorrow’s Computing Today, 19(3):46–76, June 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3475965.3479315.

McGovern:2006:ESO

[MSJL06] James McGovern, Oliver Sims, Ashish Jain, and Mark Little, editors. Enterprise service oriented architectures: concepts, challenges, recommendations. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-4020-3704-X. xxxv + 408 pp. LCCN T58.6 .E58 2006. URL

http://www.loc.gov/catdir/
enhancements/fy0824/2007427469-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0824/
2007427469-t.html.

McClure:2003:HEN

[MSK03] Stuart McClure, Joel Scambray, and George Kurtz. Hacking exposed: network security secrets and solutions. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, fourth edition, 2003. ISBN 0-07-222742-7. xxiv + 737 pp. LCCN TK5105359 .M48 2003. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
bios/mh051/2004351501.html;
 http://www.loc.gov/catdir/
description/mh051/2004351501.
html;   http://www.loc.gov/
catdir/toc/mh051/2004351501.
html.

Maxson:2004:LWR

[MSNL04] Lynn H. Maxson, Joseph M.

Saur, Colin Neill, and Phillip Laplante. Letters: Waterfall redux. ACM Queue: Tomorrow’s Computing Today, 2(2):8, April 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Mubeen:2018:WFS

[Mub18] Noor Mubeen. Workload frequency scaling law — derivation and verification. ACM Queue: Tomorrow’s Computing Today, 16(2):50, March 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3229201.

Murphy:2004:ASF

[Mur04] Brendan Murphy. Automating software failure reporting. ACM Queue: Tomorrow’s Computing Today, 2(8):42–48, November 2004. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic).

McFarland:2003:FD

[MW03] Bill McFarland and Michael Wong. The family dynamics of 802.11. ACM Queue: Tomorrow’s Computing Today, 1 (3):28–38, May 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Moshayedi:2008:ES

[MW08] Mark Moshayedi and Patrick Wilkison. Enterprise SSDs. ACM Queue: Tomorrow’s Computing Today, 6(4):32–39, July/August 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Nagy:2022:FFD

[NA22] Stefan Nagy and Peter Alvaro. The fun in fuzzing: The debugging technique comes into its own. ACM Queue: Tomorrow’s Computing Today, 20(6):80– 87, 2022. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3580504.

Nadeem:2021:HCA

[Nad21]  Ayman Nadeem.  Human-

centered approach to staticanalysis-driven developer tools: The future depends on good HCI. ACM Queue: Tomorrow’s Computing Today, 19(4):68–95, July 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487019.3487024.

Nagaraj:2004:BRA

[Nag04] S. V. Nagaraj. Book review:

Advances in Evolutionary Computing: Theory and Applications by Ashish Ghosh and Shigeyoshi Tsutsui, Springer-Verlag, 2003, $109, ISBN 3-540-43330-9. ACM Queue: Tomorrow’s Computing Today, 2 (3):74, May 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Nareyek:2004:ACG

[Nar04] Alexander Nareyek. AI in computer games. ACM Queue: Tomorrow’s Computing Today, 1 (10):58–65, February 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See minor erratum [BS04].

Naugler:2008:RPL

[Nau08a] David Naugler. Review of Pro LINQ Object Relational Mapping in C# 2008 by Vijay Mehta, Apress, 2008, $49.99, ISBN: 978-1-59059-965-5. ACM Queue: Tomorrow’s Computing Today, 6(7):41, November/December 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Naugler:2008:RPW

[Nau08b] David Naugler. Review of Pro WPF in C# 2008: Windows Presentation Foundation with .NET 3.5, (2nd ed.),’ Apress, 2008, $54.99, ISBN 1-59059-955-1. ACM Queue: Tomorrow’s Computing Today, 6(5): 59, September 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Nickolls:2008:SPP

[NBGS08] John Nickolls, Ian Buck, Michael Garland, and Kevin Skadron. Scalable parallel programming with CUDA. ACM Queue: Tomorrow’s Computing Today, 6(2):40–53, March 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Narayanan:2017:BAP

[NC17] Arvind Narayanan and Jeremy Clark. Bitcoin’s academic pedigree. ACM Queue: Tomorrow’s Computing Today, 15 (4):20, July 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3136559.

Nilles:2004:LRR

[NFG04] Jack M. Nilles, Joe Fenn, and Jim Gray. Letters: Required reading for telemanagers; state of the carrier industry. ACM Queue: Tomorrow’s Computing Today, 2(1):7, March 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Navas:2022:OVC

[NG22] Jorge A. Navas and Ashish Gehani. OCCAM-v2: Combining static and dynamic analysis for effective and efficient wholeprogram specialization: Leveraging scalable pointer analysis, value analysis, and dynamic analysis. ACM Queue: Tomorrow’s Computing Today, 20(5): 58–85, September 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3570922.

Noy:2019:ISK

[NGJ^{+}19] Natasha Noy, Yuqing Gao, Anshu Jain, Anant Narayanan, Alan Patterson, and Jamie Taylor. Industry-scale knowledge graphs: Lessons and challenges. ACM Queue: Tomorrow’s Computing Today, 17(2):20, March 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3332266.

Nokleberg:2020:BPA

[NH20] Chris Nokleberg and Brad Hawkes. Best practice: Application frameworks: While powerful, frameworks are not for everyone. ACM Queue: Tomorrow’s Computing Today, 18(6): 52–77, December 2020. URL

https://dl.acm.org/doi/10.
1145/3442632.3447806.

Nielsen:2003:CII

[Nie03] Jakob Nielsen. Curmudgeon:

IM, not IP (information pollution). ACM Queue: Tomorrow’s Computing Today, 1(8): 76, 75, November 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Narayanan:2020:DPP

[NMCK20] Arvind Narayanan, Arunesh Mathur, Marshini Chetty, and Mihir Kshirsagar. Dark patterns: Past, present, and future. ACM Queue: Tomorrow’s Computing Today, 18(2):??, March

                                        2020. URL https://dl.acm.
                                        org/doi/abs/10.1145/3400899.
                                        3400901.

Narayanan:2016:RPC

[NMHB16] Arvind Narayanan, Andrew Miller, Song Han, and Peter Bailis. Research for practice: Cryptocurrencies, blockchains, and smart contracts; hardware for deep learning. ACM Queue: Tomorrow’s Computing Today, 14(6):60, November 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=3043967.

Neville-Neal:2003:CSE

[NN03a] George V. Neville-Neal. Code spelunking: Exploring cavernous code bases. ACM Queue: Tomorrow’s Computing Today, 1(6):42–48, September/October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2003:ETA

[NN03b] George Neville-Neil. From the Editors: The truth about embedded systems. ACM Queue:

Tomorrow’s Computing Today, 1(2):4, April 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2003:PN

[NN03c] George V. Neville-Neil. Programming without a net. ACM Queue: Tomorrow’s Computing Today, 1(2):16–23, April 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2004:ICM

[NN04a] George Neville-Neil. Interview:

a conversation with Mike Deliman. ACM Queue: Tomorrow’s Computing Today, 2(7): 32–39, October 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2004:KV

[NN04b] George Neville-Neil. Kode Vicious. ACM Queue: Tomorrow’s Computing Today, 2(7): 16–18, October 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2004:KVS

[NN04c] George Neville-Neil. Kode Vicious strikes again. ACM Queue: Tomorrow’s Computing Today, 2(8):18–21, November 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2004:KVR

[NN04d] George Neville-Neil. Kode Vicious the return. ACM Queue: Tomorrow’s Computing Today, 2(9):16–18, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2005:KVB

[NN05a] George Neville-Neil. Kode Vicious battles on. ACM Queue: Tomorrow’s Computing Today, 3(3):15–17, April 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2005:KVG

[NN05b] George Neville-Neil. Kode vicious gets dirty. ACM Queue: Tomorrow’s Computing Today, 3(5):18–20, June 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2005:KVKa

[NN05c] George Neville-Neil. Kode Vicious: Kode Vicious cycles on. ACM Queue: Tomorrow’s Computing Today, 3(6):14–16, July 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2005:KVKc

[NN05d] George Neville-Neil. Kode Vicious: Kode Vicious unscripted. ACM Queue: Tomorrow’s Computing Today, 3(8):14–16, October 2005. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2005:KVKb

[NN05e] George Neville-Neil. Kode Vicious: KV the konqueror. ACM Queue: Tomorrow’s Computing Today, 3(7):12–14, September 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2005:KVRb

[NN05f] George Neville-Neil. Kode Vicious reloaded. ACM Queue: Tomorrow’s Computing Today, 3(2):20, 22, March 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2005:KVD

[NN05g] George Neville-Neil. Kode Vicious: The doctor is in. ACM Queue: Tomorrow’s Computing Today, 3(9):14–16, November 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2005:KVU

[NN05h] George Neville-Neil. Kode Vicious unleashed. ACM Queue: Tomorrow’s Computing Today, 3(1):12–13, February 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2005:KVVb

[NN05i] George Neville-Neil. Kode Vicious: Vicious XSS. ACM Queue: Tomorrow’s Computing Today, 3(10):12–15, December 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2005:KVVa

[NN05j] George Neville-Neil. Kode Vicious vs. Mothra. ACM Queue: Tomorrow’s Computing Today, 3(4):16–18, May 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2006:KVF

[NN06a] George Neville-Neil. Kode vicious: Facing the strain. ACM Queue: Tomorrow’s Computing Today, 4(7):12–13, September 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2006:KVG

[NN06b]  George Neville-Neil.  Kode

Vicious: Gettin’ your kode on. ACM Queue: Tomorrow’s Computing Today, 4(1): 12–14, February 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2006:KVH

[NN06c]  George Neville-Neil.  Kode

Vicious: Human-KV interaction. ACM Queue: Tomorrow’s Computing Today, 4(2): 10–12, March 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2006:KVK

[NN06d] George Neville-Neil. Kode Vicious: Kode Vicious bugs out. ACM Queue: Tomorrow’s Computing Today, 4(3):10–12, April 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2006:KVL

[NN06e] George Neville-Neil. Kode Vicious: Logging on with KV. ACM Queue: Tomorrow’s Computing Today, 4(5):9–10, June 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2006:KVPa

[NN06f] George Neville-Neil. Kode Vicious: Phishing for solutions. ACM Queue: Tomorrow’s Computing Today, 4(4):12–13, May 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2006:KVPb

[NN06g] George Neville-Neil. Kode Vicious: Pointless PKI. ACM Queue: Tomorrow’s Computing Today, 4(6):11–12, July/August 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2006:KVS

[NN06h] George Neville-Neil. Kode vicious: Saddle up, aspiring code jockeys. ACM Queue: Tomorrow’s Computing Today, 4 (8):10–11, October 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2006:KVU

[NN06i] George Neville-Neil. Kode vicious: Understanding the problem. ACM Queue: Tomorrow’s Computing Today, 4(9): 10–12, November 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2006:KVP

[NN07a] George Neville-Neil. Kode vicious: Peerless P2P. ACM Queue: Tomorrow’s Computing Today, 4(10):12–13, December/January 2006–2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2007:KVL

[NN07b] George Neville-Neil. Kode vicious: a license to kode. ACM Queue: Tomorrow’s Computing Today, 5(1):12–14, February 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2007:KVAb

[NN07c] George Neville-Neil. Kode vicious: Advice to a newbie. ACM Queue: Tomorrow’s Computing Today, 5(3):12–14, April 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2007:KVAa

[NN07d] George Neville-Neil. Kode vicious: APIs with an appetite. ACM Queue: Tomorrow’s Computing Today, 5(2):8–10, March 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2007:KVN

[NN07e] George Neville-Neil. Kode vicious: The next big thing. ACM Queue: Tomorrow’s Computing Today, 5(6):11–12, September 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2007:BSW

[NN07f] George V. Neville-Neil. Building secure Web applications. ACM Queue: Tomorrow’s Computing Today, 5(5):22–26, July 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2007:KVG

[NN07g] George V. Neville-Neil. Kode vicious: Gettin’ your head straight. ACM Queue: Tomorrow’s Computing Today, 5 (5):9–10, July 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2007:KVK

[NN07h] George V. Neville-Neil. Kode vicious: KV the loudmouth. ACM Queue: Tomorrow’s Computing Today, 5(4):8–10, May– June 2007. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2007:KVT

[NN07i] George V. Neville-Neil. Kode vicious: Take a freaking measurement! ACM Queue: Tomorrow’s Computing Today, 5 (7):7–8, November 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2008:DD

[NN08a] George Neville-Neil. Debugging devices. ACM Queue: Tomorrow’s Computing Today, 6(7): 6–9, November/December 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2008:KVB

[NN08b] George Neville-Neil. Kode vicious: Beautiful code exists, if you know where to look. ACM Queue: Tomorrow’s Computing Today, 6(5):4–7, September 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2008:KVG

[NN08c] George Neville-Neil. Kode vicious: Get real about realtime. ACM Queue: Tomorrow’s Computing Today, 6(6):6–9, October 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2008:KVL

[NN08d] George Neville-Neil. Kode vicious: Latency and livelocks. ACM Queue: Tomorrow’s Computing Today, 6(2):8–10, March 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2008:KVP

[NN08e] George Neville-Neil. Kode vicious: Poisonous programmers. ACM Queue: Tomorrow’s Computing Today, 6(1):5–6, January 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2008:KVS

[NN08f] George Neville-Neil. Kode vicious: Sizing your system. ACM Queue: Tomorrow’s Computing Today, 6(4):6–7, July/August 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2008:KVV

[NN08g] George Neville-Neil. Kode vicious: The virtue of paranoia. ACM Queue: Tomorrow’s Computing Today, 6(3):7–8, May/ June 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2008:CSR

[NN08h] George V. Neville-Neil. Code spelunking redux. ACM Queue: Tomorrow’s Computing Today, 6(7):26–33, November/December 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2009:MEM

[NN09a] George Neville-Neil. Kode vicious: Merge early, merge often. ACM Queue: Tomorrow’s Computing Today, 7(9):60–62, October 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2009:SA

[NN09b] George Neville-Neil. Kode vicious: Standards advice. ACM Queue: Tomorrow’s Computing Today, 7(11):40–43, December 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2009:DTS

[NN09c] George V. Neville-Neil. Kode vicious: Don’t be typecast as a software developer. ACM Queue: Tomorrow’s Computing Today, 7(2):6–9, February 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2009:KVP

[NN09d] George V. Neville-Neil. Kode vicious: Pride and prejudice: (the Vasa). ACM Queue: Tomorrow’s Computing Today, 7 (1):6–7, January 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2009:WS

[NN09e] George V. Neville-Neil. Whither sockets? ACM Queue: Tomorrow’s Computing Today, 7 (4):34–35, May 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2010:KVP

[NN10a] George Neville-Neil. Kode vicious: a paucity of ports. ACM Queue: Tomorrow’s Computing Today, 8(8):30, August 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2010:KVA

[NN10b] George Neville-Neil. Kode vicious: Avoiding obsolescence. ACM Queue: Tomorrow’s Computing Today, 8(4):50, April 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2010:KVB

[NN10c] George Neville-Neil. Kode vicious: Broken builds. ACM Queue: Tomorrow’s Computing Today, 8(3):20, March 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2010:KVC

[NN10d]  George Neville-Neil.  Kode

vicious: Collecting counters. ACM Queue: Tomorrow’s Computing Today, 8(6):10, June 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2010:CI

[NN10e]  George Neville-Neil.  Kode

vicious: Commitment issues. ACM Queue: Tomorrow’s Computing Today, 8(2):20– 22, February 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2010:BSL

[NN10f] George V. Neville-Neil. Bound by the speed of light. ACM Queue: Tomorrow’s Computing Today, 8(12):20, December 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2010:LL

[NN10g] George V. Neville-Neil. Lessons from the letter. ACM Queue: Tomorrow’s Computing Today, 8(7):30, July 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2011:CR

[NN11a] George Neville-Neil. Code rototilling. ACM Queue: Tomorrow’s Computing Today, 9 (12):20, December 2011. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2011:HIS

[NN11b] George Neville-Neil. How to improve security? ACM Queue: Tomorrow’s Computing Today, 9(8):30, August 2011. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2011:PA

[NN11c] George Neville-Neil. Porting with Autotools. ACM Queue: Tomorrow’s Computing Today, 9(3):10, March 2011. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2011:SS

[NN11d] George Neville-Neil. Storage strife. ACM Queue: Tomorrow’s Computing Today, 9 (5):10, May 2011. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2011:WAD

[NN11e] George Neville-Neil. Wanton acts of debuggery. ACM Queue: Tomorrow’s Computing Today, 9(10):30, October 2011. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2012:NPC

[NN12a] George Neville-Neil. A nice piece of code. ACM Queue: Tomorrow’s Computing Today, 10(6):10, June 2012. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2012:QPR

[NN12b] George Neville-Neil. Queue portrait: Robert Watson. ACM Queue: Tomorrow’s Computing Today, 10(10):??, October 2012. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2012:CMC

[NN12c] George V. Neville-Neil. Can more code mean fewer bugs? ACM Queue: Tomorrow’s Computing Today, 10(8):10, August 2012. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2012:NPB

[NN12d] George V. Neville-Neil. The network protocol battle: a tale of hubris and zealotry. ACM Queue: Tomorrow’s Computing Today, 10(1):10, January 2012. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2013:KVL

[NN13a] George Neville-Neil. Kode vicious: A lesson in resource management. ACM Queue: Tomorrow’s Computing Today, 11(8): 10:10–10:12, August 2013. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2013:KVB

[NN13b] George Neville-Neil. Kode vicious: Bugs and bragging rights. ACM Queue: Tomorrow’s Computing Today, 11(10):1–3, October 2013. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2013:KVC

[NN13c] George Neville-Neil. Kode vicious: Cherry-picking and the Scientific Method. ACM Queue:

Tomorrow’s Computing Today, 11(4):10, April 2013. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2013:KVD

[NN13d] George Neville-Neil. Kode vicious: Divided by division. ACM Queue: Tomorrow’s Computing Today, 11(1):1–3, January 2013. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2013:KVN

[NN13e] George Neville-Neil. Kode vicious: The naming of hosts is a difficult matter; also, the perils of premature rebooting. ACM Queue: Tomorrow’s Computing Today, 11(6):10, June 2013. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2013:KVF

[NN13f] George Neville-Neil. Kode vicious: This is the foo field. ACM Queue: Tomorrow’s Computing Today, 11(12):??, December 2013. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2566971.

Neville-Neil:2014:KVF

[NN14a] George Neville-Neil. Kode vicious: Forked over. ACM Queue: Tomorrow’s Computing Today, 12(4):10, April 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2014:KVO

[NN14b] George Neville-Neil. Kode vicious: Outsourcing responsibility. ACM Queue: Tomorrow’s Computing Today, 12 (6):20, June 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2014:KVL

[NN14c] George Neville-Neil. Kode vicious: The logic of logging. ACM Queue: Tomorrow’s Computing Today, 12(2):30, February 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2014:PS

[NN14d] George Neville-Neil. Port squatting. ACM Queue: Tomorrow’s Computing Today, 12(9): 10, September 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neville-Neil:2014:TBF

[NN14e] George Neville-Neil. Too big to fail. ACM Queue: Tomorrow’s Computing Today, 12(11): 10, November 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=2693195.

Neville-Neil:2015:CH

[NN15a]  George Neville-Neil.  Code

hoarding. ACM Queue: Tomorrow’s Computing Today, 13(9):90, November/December 2015. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2897034.

Neville-Neil:2015:KVH

[NN15b] George Neville-Neil. Kode vicious: Hickory dickory doc: On null encryption and automated documentation. ACM Queue: Tomorrow’s Computing Today, 13(6):10, June 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2791303.

Neville-Neil:2015:KVL

[NN15c] George Neville-Neil. Kode vicious: Lazarus code. ACM Queue: Tomorrow’s Computing Today, 13(5):10, May 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2773214.

Neville-Neil:2015:KVP

[NN15d] George Neville-Neil. Kode vicious: Pickled patches. ACM Queue: Tomorrow’s Computing Today, 13(8):90, September/October 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2856150.

Neville-Neil:2015:KVR

[NN15e] George Neville-Neil. Kode vicious: Raw networking. ACM Queue: Tomorrow’s Computing Today, 13(2):20, February 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=2732268.

Neville-Neil:2015:TI

[NN15f]  George Neville-Neil.  Time

is an illusion. ACM Queue: Tomorrow’s Computing Today, 13(9):30, November/December 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=2878574.

Neville-Neil:2016:CPW

[NN16a] George Neville-Neil. The chess player who couldn’t pass the salt. ACM Queue: Tomorrow’s Computing Today, 14(6): 20, November 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3029797.

Neville-Neil:2016:CM

[NN16b] George Neville-Neil. Chilling the messenger. ACM Queue: Tomorrow’s Computing Today, 14(3):20, May 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=2962575.

Neville-Neil:2016:CC

[NN16c]  George Neville-Neil.  Cloud

calipers. ACM Queue: Tomorrow’s Computing Today, 14 (4):30, July 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2993454.

Neville-Neil:2016:GL

[NN16d] George Neville-Neil. GNL is not Linux. ACM Queue: Tomorrow’s Computing Today, 14 (1):40, January 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2909572.

Neville-Neil:2016:UTS

[NN16e] George Neville-Neil. The unholy Trinity of software development. ACM Queue: Tomorrow’s Computing Today, 14(5): 20, September 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3014148.

Neville-Neil:2016:WYT

[NN16f] George Neville-Neil. What are you trying to pull? ACM Queue: Tomorrow’s Computing Today, 14(2):20, March 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2931077.

Neville-Neil:2017:CHC

[NN17a] George Neville-Neil. Cold, hard cache. ACM Queue: Tomorrow’s Computing Today, 15 (4):10, July 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3134436.

Neville-Neil:2017:FEH

[NN17b] George Neville-Neil. Forced exception-handling. ACM Queue: Tomorrow’s Computing Today, 15(1):40, January 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=3064643.

Neville-Neil:2017:IIT

[NN17c] George Neville-Neil. IoT: The Internet of Terror. ACM Queue: Tomorrow’s Computing Today, 15(3):10, May 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3121440.

Neville-Neil:2017:OE

[NN17d] George Neville-Neil. The observer effect. ACM Queue: Tomorrow’s Computing Today, 15 (2):10, March 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3084695.

Neville-Neil:2017:PK

[NN17e] George Neville-Neil. Popping kernels. ACM Queue: Tomorrow’s Computing Today, 15(6): 20, November 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3180444.

Neville-Neil:2017:RAS

[NN17f] George Neville-Neil. Reducing the attack surface. ACM Queue: Tomorrow’s Computing Today, 15(5):40, September/October 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3161600.

Neville-Neil:2018:CG

[NN18a] George Neville-Neil. A chance gardener. ACM Queue: Tomorrow’s Computing Today, 16 (4):50, July 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3286730.

Neville-Neil:2018:ESL

[NN18b] George Neville-Neil. Every silver lining has a cloud. ACM Queue: Tomorrow’s Computing Today, 16(2):20, March 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3215876.

Neville-Neil:2018:KYA

[NN18c] George Neville-Neil. Know your algorithms. ACM Queue: Tomorrow’s Computing Today, 16 (6):40, November 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3310152.

Neville-Neil:2018:OCK

[NN18d]  George Neville-Neil.  The

obscene coupling known as spaghetti code. ACM Queue: Tomorrow’s Computing Today, 16(3):70, May 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3265943.

Neville-Neil:2018:WVS

[NN18e] George Neville-Neil. Watchdogs vs. snowflakes. ACM Queue: Tomorrow’s Computing Today, 16(1):60, January 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=3207170.

Neville-Neil:2018:WTP

[NN18f] George Neville-Neil. Writing a test plan. ACM Queue: Tomorrow’s Computing Today, 16(5): 40, September 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3294770.

Neville-Neil:2019:KA

[NN19a] George Neville-Neil. Koding academies. ACM Queue: Tomorrow’s Computing Today, 17 (4):20, July 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3361872.

Neville-Neil:2019:MM

[NN19b] George Neville-Neil. MUST and MUST NOT. ACM Queue: Tomorrow’s Computing Today, 17 (2):70, March 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3341158.

Neville-Neil:2019:WCG

[NN19c] George Neville-Neil. What is a CSO good for? ACM Queue: Tomorrow’s Computing Today, 17(3):60, May 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3357152.

Neville-Neil:2019:WIA

[NN19d]  George Neville-Neil.  The

worst idea of all time. ACM Queue: Tomorrow’s Computing Today, 17(1):30, January 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3320268.

Neville-Neil:2019:MT

[NN19e] George V. Neville-Neil. Master of tickets. ACM Queue: Tomorrow’s Computing Today, 17(6):??, November 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3380774.3383461.

Neville-Neil:2019:NCS

[NN19f] George V. Neville-Neil. Numbers are for computers, strings are for humans. ACM Queue:

Tomorrow’s Computing Today, 17(5):??, September 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3371595.3379349.

Neville-Neil:2020:BHC

[NN20a] George Neville-Neil. Broken hearts and coffee mugs. ACM Queue: Tomorrow’s Computing Today, 18(2):??, March 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3400899.3407138.

Neville-Neil:2020:SVI

[NN20b] George Neville-Neil. Sanity vs. invisible markings: Tabs vs. spaces. ACM Queue: Tomorrow’s Computing Today, 18 (3):39–42, June 2020. URL

https://dl.acm.org/doi/10.
1145/3411757.3412568.

Neville-Neil:2020:KFK

[NN20c] George V. Neville-Neil. Kabin fever: KV’s guidelines for KFH (koding from home). ACM Queue: Tomorrow’s Computing Today, 18(5):35–40, October 2020. URL https:

//dl.acm.org/doi/10.1145/
3434571.3437846.

Neville-Neil:2020:NPG

[NN20d] George V. Neville-Neil. The non-psychopath’s guide to managing an open-source project: Respect your staff, learn from others, and know when to let go. ACM Queue: Tomorrow’s Computing Today, 18(6): 30–36, December 2020. URL

https://dl.acm.org/doi/10.
1145/3442632.3447645.

Neville-Neil:2020:RKD

[NN20e] George V. Neville-Neil. Removing kode: Dead functions and dead features. ACM Queue: Tomorrow’s Computing Today, 18 (4):21–25, August 2020. URL

https://dl.acm.org/doi/10.
1145/3424302.3426887.

Neville-Neil:2021:AVR

[NN21a] George V. Neville-Neil. Aversion to versions: Resolving code-dependency issues. ACM Queue: Tomorrow’s Computing Today, 19(1):16–19, February 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3454122.3458680.

Neville-Neil:2021:DCU

[NN21b] George V. Neville-Neil. Divide and conquer: The use and limits of bisection. ACM Queue: Tomorrow’s Computing Today, 19(3):37–39, June 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3475965.3477581.

Neville-Neil:2021:GMP

[NN21c] George V. Neville-Neil. Getting off the mad path: Debuggers and assertions. QUEUE, 19(6): 18–21, November 2021. URL

https://dl.acm.org/doi/10.
1145/3511662.

Neville-Neil:2021:UWL

[NN21d] George V. Neville-Neil. I unplugged what?: The lessons here are broader than just a simple “Don’t do that.”. ACM Queue: Tomorrow’s Computing Today, 19(5):22–25, September 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3498578.

Neville-Neil:2021:PDT

[NN21e] George V. Neville-Neil. In praise of the disassembler: There’s much to be learned from the lower-level details of hardware. ACM Queue: Tomorrow’s Computing Today, 19(2):21–28, April 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3466132.3469121.

Neville-Neil:2022:ASR

[NN22a] George V. Neville-Neil. All sliders to the right: Hardware overkill. ACM Queue: Tomorrow’s Computing Today, 20(6): 4–8, 2022. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3580505.

Neville-Neil:2022:ERT

[NN22b] George V. Neville-Neil. The elephant in the room: It’s time to get the POSIX elephant off our necks. ACM Queue: Tomorrow’s Computing Today, 20 (5):14–19, September 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3570921.

Neville-Neil:2022:FHA

[NN22c] George V. Neville-Neil. The four horsemen of an ailing software project: Don’t let the pale rider catch you with an exception. ACM Queue: Tomorrow’s Computing Today, 20 (4):11–15, July 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3561798.

Neville-Neil:2022:PCD

[NN22d] George V. Neville-Neil. The planning and care of data: Rearranging buckets for no good reason. ACM Queue: Tomorrow’s Computing Today, 20 (1):4–9, January 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3526209.

Neville-Neil:2022:SCJ

[NN22e] George V. Neville-Neil. Securing the company jewels: GitHub and runbook security. ACM Queue: Tomorrow’s Computing Today, 20(3):14–17, May 2022. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546936.

Neville-Neil:2022:WSB

[NN22f] George V. Neville-Neil. When should a black box be transparent?: When is a replacement not a replacement? ACM Queue: Tomorrow’s Computing Today, 20(2):11–15, March 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3534856.

Neville-Neil:2023:DDK

[NN23a] George V. Neville-Neil. Dear diary: On keeping a laboratory notebook. ACM Queue: Tomorrow’s Computing Today, 21(5): 22–28, September 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3631181.

Neville-Neil:2023:HAW

[NN23b] George V. Neville-Neil. Halfway around the world: Learn the language, meet the people, eat the food. ACM Queue: Tomorrow’s Computing Today, 21 (4):18–22, July 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3623426.

Neville-Neil:2023:HTT

[NN23c] George V. Neville-Neil. The human touch: There is no substitute for good, direct, honest training. ACM Queue: Tomorrow’s Computing Today, 21 (2):9–13, March 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3595821.

Neville-Neil:2023:TAS

[NN23d] George V. Neville-Neil. Is there another system?: Computer science is the study of what can be automated. ACM Queue: Tomorrow’s Computing Today, 21 (6):12–15, November 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3639446.

Neville-Neil:2023:PPT

[NN23e] George V. Neville-Neil. The parchment path?: Is there ever a time when learning is not of value — for its own sake? ACM Queue: Tomorrow’s Computing Today, 21 (1):5–11, January 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3587482.

Neville-Neil:2023:SKB

[NN23f] George V. Neville-Neil. Stone knives and bear skins. ACM Queue: Tomorrow’s Computing Today, 21(3):29–35, May 2023. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606027.

Neville-Neil:2024:SDO

[NN24a] George V. Neville-Neil. Software drift: Open source forking. ACM Queue: Tomorrow’s Computing Today, 22(1): 17–22, January 2024. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3649297.

Neville-Neil:2024:SSP

[NN24b] George V. Neville-Neil. Structuring success: The problem with software structure is people don’t really learn it until they really need it. ACM Queue: Tomorrow’s Computing Today, 22(2), March 2024. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664276.

Neville-Neil:2015:VSS

[NNB15] George Neville-Neil and Steve Bourne. Version status: Still finding the right questions. ACM Queue: Tomorrow’s Computing Today, 13(8):10, September/October 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2838346.

Nolann:2006:CSC

[Nol06] Greg A. Nolann. Curmudgeon: Seeking compliance nirvana. ACM Queue: Tomorrow’s Computing Today, 4(7): 72–??, September 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Norwalk:2009:MMA

[Nor09] Jeff Norwalk. Making the move to AJAX. ACM Queue: Tomorrow’s Computing Today, 7 (1):1:1–1:??, January 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Noy:2005:OC

[Noy05] Natalya Noy. Order from chaos.

ACM Queue: Tomorrow’s Computing Today, 3(8):42–49, October 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Nukala:2018:WSD

[NR18] Shylaja Nukala and Vivek Rau.

Why SRE documents matter. ACM Queue: Tomorrow’s Computing Today, 16(4):40, July 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3283589.

Noda:2023:DWA

[NSFG23] Abi Noda, Margaret-Anne Storey, Nicole Forsgren, and Michaela Greiler. DevEx: What actually drives productivity: The developer-centric approach to measuring and improving productivity. ACM Queue: Tomorrow’s Computing Today, 21 (2):35–53, March 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3595878.

Nanavati:2015:NVS

[NSWW15] Mihir Nanavati, Malte Schwarzkopf, Jake Wires, and Andrew Warfield. Non-volatile storage. ACM Queue: Tomorrow’s Computing Today, 13(9):20, November/December 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=2874238.

ODell:2009:NFE

[O’D09] Mike O’Dell. Network frontend processors, yet again. ACM Queue: Tomorrow’s Computing Today, 7(3):28–29, April 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

ODell:2017:DM

[O’D17] Devon H. O’Dell. The debugging mindset. ACM Queue: Tomorrow’s Computing Today, 15 (1):50, January 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3068754.

OHanlon:2005:Ca

[O’H05a] Charlene O’Hanlon. Calendar. ACM Queue: Tomorrow’s Computing Today, 3(8):61, October 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2005:Cb

[O’H05b] Charlene O’Hanlon. Calendar. ACM Queue: Tomorrow’s Computing Today, 3(9):60, November 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2005:EN

[O’H05c] Charlene O’Hanlon. From the editors: In with the new. ACM Queue: Tomorrow’s Computing Today, 3(10):8, December 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2005:ESA

[O’H05d] Charlene O’Hanlon. From the editors: Socially acceptable behavior. ACM Queue: Tomorrow’s Computing Today, 3(9): 8, November 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2005:EUR

[O’H05e] Charlene O’Hanlon. From the Editors: Unstructured, but not really. ACM Queue: Tomorrow’s Computing Today, 3(8): 8, October 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2005:CPS

[O’H05f]  Charlene O’Hanlon.  Inter-

view: a conversation with Phil Smoot. ACM Queue: Tomorrow’s Computing Today, 3(10): 16–24, December 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2005:ICR

[O’H05g]  Charlene O’Hanlon.  Inter-

view: a conversation with Ray Ozzie. ACM Queue: Tomorrow’s Computing Today, 3(9): 18–26, November 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2005:WYHa

[O’H05h] Charlene O’Hanlon. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 3(8):12, October 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2005:WYHb

[O’H05i] Charlene O’Hanlon. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 3(9):12, November 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2005:WYH

[O’H05j] Charlene O’Hanlon. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 3(10):11, December 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Olukotun:2005:FM

[OH05k] Kunle Olukotun and Lance Hammond. The future of microprocessors. ACM Queue: Tomorrow’s Computing Today, 3 (7):26–29, September 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:Ca

[O’H06a] Charlene O’Hanlon. Calendar. ACM Queue: Tomorrow’s Computing Today, 4(2): 60–61, March 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:Cb

[O’H06b] Charlene O’Hanlon. Calendar. ACM Queue: Tomorrow’s Computing Today, 4(4):52–53, May 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:Cc

[O’H06c] Charlene O’Hanlon. Calendar. ACM Queue: Tomorrow’s Computing Today, 4(5):60–61, June 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:Cd

[O’H06d] Charlene O’Hanlon. Calendar. ACM Queue: Tomorrow’s Computing Today, 4(6): 59, July/August 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:Ce

[O’H06e] Charlene O’Hanlon. Calendar.

ACM Queue: Tomorrow’s Computing Today, 4(8):52, October 2006. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:Cf

[O’H06f] Charlene O’Hanlon. Calendar. ACM Queue: Tomorrow’s Computing Today, 4(9):52, November 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:EAB

[O’H06g] Charlene O’Hanlon. From the editors: Able bodies. ACM Queue: Tomorrow’s Computing Today, 4(6):8, July/August 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:EA

[O’H06h] Charlene O’Hanlon. From the editors: Accessorizing. ACM Queue: Tomorrow’s Computing Today, 4(5):6, June 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:EPR

[O’H06i] Charlene O’Hanlon. From the editors: Playing by the rules. ACM Queue: Tomorrow’s Computing Today, 4(7):8, September 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:EQR

[O’H06j] Charlene O’Hanlon. From the editors: Quality really is job #1. ACM Queue: Tomorrow’s Computing Today, 4(1):8, February 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:RVP

[O’H06k] Charlene O’Hanlon. From the editors: Reality vs. perception. ACM Queue: Tomorrow’s Computing Today, 4(8):7, October 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:ECM

[O’H06l] Charlene O’Hanlon. From the editors: The criminal mind. ACM Queue: Tomorrow’s Computing Today, 4(9):7, November 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:EDE

[O’H06m] Charlene O’Hanlon. From the editors: The domino effect. ACM Queue: Tomorrow’s Computing Today, 4(4):8, May 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:EPU

[O’H06n] Charlene O’Hanlon. From the Editors: The private universe. ACM Queue: Tomorrow’s Computing Today, 4(3):7, April 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:ICC

[O’H06o] Charlene O’Hanlon. Interview:

a conversation with Chuck Mc-Manis. ACM Queue: Tomorrow’s Computing Today, 4(3): 13–21, April 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:ICDa

[O’H06p]  Charlene O’Hanlon.  Inter-

view: a conversation with David Brown. ACM Queue: Tomorrow’s Computing Today, 4 (8):14–23, October 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:ICDb

[O’H06q] Charlene O’Hanlon. Interview: a conversation with Douglas W. Jones and Peter G. Neumann. ACM Queue: Tomorrow’s Computing Today, 4(9): 16–23, November 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:ICJa

[O’H06r] Charlene O’Hanlon. Interview: a conversation with Jarod Jenson. ACM Queue: Tomorrow’s Computing Today, 4(1): 16–24, February 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:ICJb

[O’H06s] Charlene O’Hanlon. Interview: a conversation with Jordan Cohen. ACM Queue: Tomorrow’s Computing Today, 4(6):

14–23, July/August 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:ICL

[O’H06t] Charlene O’Hanlon. Interview:

a conversation with Leo Chang of ClickShift. ACM Queue: Tomorrow’s Computing Today, 4(5):12–18, June 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:ICS

[O’H06u] Charlene O’Hanlon. Interview:

a conversation with Steve Ross-Talbot. ACM Queue: Tomorrow’s Computing Today, 4 (2):14–23, March 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:ICW

[O’H06v] Charlene O’Hanlon. Interview:

a conversation with Werner Vogels. ACM Queue: Tomorrow’s Computing Today, 4(4):14–22, May 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:RFWa

[O’H06w] Charlene O’Hanlon. Reader files: What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 4 (2):9, March 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:RFWb

[O’H06x] Charlene O’Hanlon. Reader files: What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 4 (8):9, October 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:RFW

[O’H06y] Charlene O’Hanlon. Reader files: What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 4 (9):9, November 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:WYHa

[O’H06z] Charlene O’Hanlon. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 4(1):11, February 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:WYHb

[O’H06-27] Charlene O’Hanlon. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 4(3):9, April 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:WYHc

[O’H06-28] Charlene O’Hanlon. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 4(4):10–11, May 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:WYHd

[O’H06-29] Charlene O’Hanlon. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 4(5):8, June 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:WYHe

[O’H06-30] Charlene O’Hanlon. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 4(6):10, July/August 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:WYHf

[O’H06-31] Charlene O’Hanlon. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 4(7):10, September 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:C

[O’H07a] Charlene O’Hanlon. Calendar.

ACM Queue: Tomorrow’s Computing Today, 4(10):50, December/January 2006–2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:FEF

[O’H07b]  Charlene O’Hanlon.  From

the Editors: Forward thinking. ACM Queue: Tomorrow’s Computing Today, 4(10):8, December/January 2006–2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:ICJ

[O’H07c]  Charlene O’Hanlon.  Inter-

view: a conversation with John Hennessy and David Patterson. ACM Queue: Tomorrow’s Computing Today, 4(10):14–22, December/January 2006–2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHanlon:2006:WYHg

[O’H07d] Charlene O’Hanlon. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 4(10):11, December/January 2006–2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

OHara:2007:TCE

[O’H07e] John O’Hara. Toward a commodity enterprise middleware. ACM Queue: Tomorrow’s Computing Today, 5(4):48–55, May– June 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Olsen:2006:CC

[Ols06] Greg Olsen. From COM to common. ACM Queue: Tomorrow’s Computing Today, 4(5):20–26, June 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Olson:2003:CSR

[OO03] Judith S. Olson and Gary M.

Olson. Culture surprises in remote software development teams. ACM Queue: Tomorrow’s Computing Today, 1(9): 52–59, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ostell:2005:DD

[Ost05] James Ostell. Databases of discovery. ACM Queue: Tomorrow’s Computing Today, 3 (3):40–48, April 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Overton:2017:IG

[Ove17] Mark A. Overton. The IDAR graph. ACM Queue: Tomorrow’s Computing Today, 15 (2):20, March 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3089807.

Owens:2010:SEC

[Owe10] Dustin Owens. Securing elasticity in the cloud. ACM Queue: Tomorrow’s Computing Today, 8(5):1–7, May 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Pacheco:2011:PDD

[Pac11] David Pacheco. Postmortem debugging in dynamic environments. ACM Queue: Tomorrow’s Computing Today, 9(10): 12, October 2011. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Paquette:2006:RWW

[Paq06] Marc Paquette. Review of “Wiki: Web Collaboration” by Anja Ebersbach, Markus Glaser and Richard Heigl, Springer, 2005, $64.95, ISBN 3-540-25995-3. ACM Queue: Tomorrow’s Computing Today, 4(2): 59, March 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Paramithiotti:2006:RGD

[Par06] Andrea Paramithiotti. Review of “Grid Database Design” by April Wells, Auerbach Publications, 2005, $79.95, ISBN 0-8493-2800-4. ACM Queue: Tomorrow’s Computing Today, 4 (3):42, April 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Paramithiotti:2007:RHG

[Par07]  Andrea Paramithiotti.  Re-

view of ‘Hitchhiker’s Guide to Visual Studio and SQL Server: Best Practice Architectures and Examples,’ 7th edition (Microsoft Windows Server System Series) by William Vaugh and Peter Blackburn, Addison-Wesley Professional, 2006, $59.99, ISBN 0-321-24362-5. ACM Queue: Tomorrow’s Computing Today, 5(5):46, July 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [VB07].

Passin:2004:EGS

[Pas04] Thomas B. Passin. Explorer’s guide to the Semantic Web. Manning, Greenwich, CT, USA, 2004. ISBN 1-932394-20-6. xxii + 281 pp. LCCN TK5105.88815 .P37 2004. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/ipg051/2004301798.html;
http://www.loc.gov/catdir/
description/ipg051/2004301798.
html.

Patterson:2003:CJG

[Pat03] David Patterson. Interview:

a conversation with Jim Gray. ACM Queue: Tomorrow’s Computing Today, 1(4):8, 10–12, 14–17, June 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Patterson:2004:WWY

[Pat04] Anna Patterson. Why writing your own search engine is hard. ACM Queue: Tomorrow’s Computing Today, 2(2):48–53, April 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Patterson:2005:CSW

[Pat05] David Patterson. Curmudgeon:

Stop whining about outsourcing! ACM Queue: Tomorrow’s Computing Today, 3(9): 64, 63, November 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Paasch:2014:MT

[PB14] Christoph Paasch and Olivier Bonaventure. Multipath TCP. ACM Queue: Tomorrow’s Computing Today, 12(2):40, February 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Poor:2003:SHN

[PBA03] Robert Poor, Cliff Bowman, and Charlotte Burgess Auburn. Selfhealing networks. ACM Queue: Tomorrow’s Computing Today, 1(3):52–59, May 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Papadopoulos:2007:BBC

[PBK07] Philip Papadopoulos, Greg Bruno, and Mason Katz. Beyond Beowulf clusters. ACM Queue: Tomorrow’s Computing Today, 5(3):36–43, April 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Peikari:2004:SW

[PC04] Cyrus Peikari and Anton Chuvakin. Security warrior. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00545-8. xvii + 531 pp. LCCN TK5105.59 .P44985 2004.

Pillai:2015:CC

[PCA^{+}15] Thanumalayan Sankaranarayana Pillai, Vijay Chidambaram, Ramnatthan Alagappan, Samer Al-Kiswany, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Crash consistency. ACM Queue: Tomorrow’s Computing Today, 13(7):20, July 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2801719.

Peterson:2004:BRE

[Pet04] Gregory D. Peterson. Book review: Embedded Software for SoC , edited by Ahmed Amine Jerraya, Sungjoo Yoo, Norbert When, and Diederik Verkest, Kluwer Academic Publishers, 2003, $174.00, ISBN 1-4020-7528-6. ACM Queue: Tomorrow’s Computing Today, 2 (4):76, June 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Petschulat:2009:OPD

[Pet09] Stephen Petschulat. Other people’s data. ACM Queue: Tomorrow’s Computing Today, 7 (10):20, November 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Petrov:2018:ABM

[Pet18] Alex Petrov. Algorithms behind modern storage systems.

ACM Queue: Tomorrow’s Computing Today, 16(2):30, March 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=3220266.

Potter:2003:S

[PF03] Bruce Potter and Bob Fleck. 802.11 Security. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00290-4. xiii + 176 pp. LCCN TK5105.78 .P68 2003. US$34.95, CDN$54.95, UK£24.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        80211security/.

Phillips:2010:SB

[Phi10a] George Phillips. Simplicity betrayed. ACM Queue: Tomorrow’s Computing Today, 8 (4):10, April 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Phillips:2010:EDT

[Phi10b] Peter Phillips. Enhanced debugging with traces. ACM Queue: Tomorrow’s Computing Today, 8(3):1–6, March 2010. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Prakash:2005:FSR

[PO05] Vipul Ved Prakash and Adam O’Donnell. Fighting spam with reputation systems. ACM Queue: Tomorrow’s Computing Today, 3(9):36–41, November 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Poole:2006:BMR

[Poo06] Damon Poole. Breaking the major release habit. ACM Queue: Tomorrow’s Computing Today, 4(8):46–51, October 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Popa:2024:CCC

[Pop24] Raluca Ada Popa. Confidential computing or cryptographic computing?: Tradeoffs between cryptography and hardware enclaves. ACM Queue: Tomorrow’s Computing Today, 22(2), March 2024. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664295.

Porcino:2004:GGR

[Por04a] Nick Porcino. Gaming graphics:

The road to revolution. ACM Queue: Tomorrow’s Computing Today, 2(2):62–71, April 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Portele:2004:RSD

[Por04b] T. Portele. Review of Spoken Dialogue Technology by Michael F. McTear. Springer-Verlag, 2004, $59.95, ISBN 1-85233-672-2. ACM Queue: Tomorrow’s Computing Today, 2 (9):76, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Pospiech:2004:BRC

[Pos04]  Frank Pospiech.  Book re-

view: Critical Testing Process: Plan, Prepare, Perform, Perfect, by Rex Black, Addison-Wesley Professional, 2003, $49.99, ISBN 0-201-74868-1. ACM Queue: Tomorrow’s Computing Today, 2(5):76, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Phillips:2003:NSC

[PP03a] Peter Phillips and George Phillips. No source code? no problem! ACM Queue: Tomorrow’s Computing Today, 1 (6):50–57, September/October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Poppendieck:2003:LSD

[PP03b] Mary Poppendieck and Tom Poppendieck, editors. Lean Software Development: an Agile Toolkit. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-321-15078-3. 203 pp. UK£30.99.

Phelps:2004:FGM

[PP04] Andrew M. Phelps and David M. Parks. Fun and games: Multilanguage development. ACM Queue: Tomorrow’s Computing Today, 1(10):46–56, February 2004. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic).

Pritchett:2008:BAA

[Pri08] Dan Pritchett. BASE: An ACID alternative. ACM Queue: Tomorrow’s Computing Today, 6 (3):48–55, May/June 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Provos:2009:CWC

[PRM09] Niels Provos, Moheeb Abu Rajab, and Panayiotis Mavrommatis. Cybercrime 2.0: When the Cloud turns dark. ACM Queue: Tomorrow’s Computing Today, 7(2):46–47, February 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Phillips:2009:PBM

[PS09] James C. Phillips and John E.

Stone. Probing biomolecular machines with graphics processors. ACM Queue: Tomorrow’s Computing Today, 7(9):30, October 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Parri:2011:RCP

[PSBG11] Jonathan Parri, Daniel Shapiro, Miodrag Bolic, and Voicu Groza. Returning control to the programmer: SIMD intrinsics for virtual machines. ACM Queue: Tomorrow’s Computing Today, 9(2):30, February 2011. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Purdy:2006:MPM

[Pur06] Mark Purdy. Modern performance monitoring. ACM Queue: Tomorrow’s Computing Today, 4(1):48–57, February 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Putnam:2007:RMD

[Put07] Jeffrey Putnam. Review of ‘Model Driven Architecture and Ontology Development’ by Dragan Gašević, Dragan Djurić, and Vladan Devedzic, Springer, 2006, $59.99, ISBN 3-540-32180-2. ACM Queue: Tomorrow’s Computing Today, 5(1): 50, February 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [GDD06].

Putnam:2008:RPM

[Put08] Jeffrey Putnam. Review of ‘Practical MythTV: Building a PVR and Media Center PC’, Stewart Smith and Michael Still, Apress, 2007, $29.99, ISBN 1-59059-779-6. ACM Queue: Tomorrow’s Computing Today, 6 (3):69, May/June 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Puzmanova:2004:RWF

[Puz04] Rita Puzmanova. Review of Wi-Foo: The Secrets of Wireless Hacking by Andrew Vladimirov, Konstantin V. Gravrilenko, and Andrei A. Mikhailovsky. Pearson Education, 2004, ISBN 0-321-20217-1. ACM Queue: Tomorrow’s Computing Today, 2 (8):70, November 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Pfleging:2006:GGW

[PZ06] Bill Pfleging and Minda Zetlin.

The geek gap: Why business and technology professionals don’t understand each other and why they need each other to survive. Prometheus Books, Amherst, NY, USA, 2006. ISBN 1-59102-415-3. 251 pp. LCCN HD30.2 .P487 2006. URL http:

//www.loc.gov/catdir/toc/
ecip0610/2006008263.html.

Ragan:2006:KSI

[Rag06] Tracy Ragan. Keeping score in the IT compliance game. ACM Queue: Tomorrow’s Computing Today, 4(7):38–43, September 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ramirez:2008:REA

[Ram08] Jose M. Ramirez. Review of ‘Enterprise Ajax: Strategies for Building High-performance Web Applications’ by David Johnson, Alexei White, and Andre Charland, Prentice Hall PTR, 2008, $39.99, ISBN 0-13-224206-0. ACM Queue: Tomorrow’s Computing Today, 6 (2):66, March 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [JWC08].

Ranum:2004:SRP

[Ran04] Marcus J. Ranum. Security: The root of the problem. ACM Queue: Tomorrow’s Computing Today, 2(4):44–49, June 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Rao:2004:ISB

[Rao04a] Ramana Rao. From IR to search, and beyond. ACM Queue: Tomorrow’s Computing Today, 2(3):66–73, May 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Rao:2004:BRM

[Rao04b] Shrisha Rao. Book review: Molecular Computing edited by Tanya Sienko, Andrew Adamatzky, Nicholas G. Rambidi, and Michael Conrad, MIT Press, 2003, $45.00, ISBN 0-262-19487-2. ACM Queue: Tomorrow’s Computing Today, 2 (4):76, June 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Raskin:2003:OSS

[Ras03a] Jef Raskin. Opinion: Silicon superstitions. ACM Queue: Tomorrow’s Computing Today, 1(9):10–12, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Raskin:2003:OUI

[Ras03b] Jef Raskin. Opinion: User interface designers, slaves of fashion. ACM Queue: Tomorrow’s Computing Today, 1(6):10–11, September/October 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Raskin:2003:WI

[Ras03c] Jef Raskin. The woes of IDEs.

ACM Queue: Tomorrow’s Computing Today, 1(3):8, 10–11, May 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Raskin:2004:OWC

[Ras04a] Jef Raskin. Opinion: For want of a comma, the meaning was lost. ACM Queue: Tomorrow’s Computing Today, 2(5):14–16, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Raskin:2004:OGP

[Ras04b] Jef Raskin. Opinion:the GIGO principle and voting machines. ACM Queue: Tomorrow’s Computing Today, 2(2):10–11, April 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Raskin:2005:CCM

[Ras05] Jef Raskin. Curmudgeon: Comments are more important than code. ACM Queue: Tomorrow’s Computing Today, 3(2):

64, 62–63, March 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Rathbun:2015:PPP

[Rat15] Spencer Rathbun. Parallel processing with promises. ACM Queue: Tomorrow’s Computing Today, 13(3):10, March 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2742696.

Bailis:2017:VPD

[RB17] John Regehr and Peter Bailis. Vigorous public debates in academic computer science. ACM Queue: Tomorrow’s Computing Today, 15(3):70, May 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3132371.

Reis:2009:BSL

[RBP09] Charles Reis, Adam Barth, and Carlos Pizano. Browser security: Lessons from Google Chrome. ACM Queue: Tomorrow’s Computing Today, 7 (5):3:1–3:8, June 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Richards:2014:PPP

[RBSH14] John T. Richards, Jonathan Brezin, Calvin B. Swart, and Christine A. Halverson. Productivity in parallel programming: a decade of progress. ACM Queue: Tomorrow’s Computing Today, 12(9):30, September 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Richards:2004:POS

[RC04] John Richards and Jim Christensen. People in our software. ACM Queue: Tomorrow’s Computing Today, 1(10): 80–86, February 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Russinovich:2021:TCC

[RCF^{+}21] Mark Russinovich, Manuel Costa, Cédric Fournet, David Chisnall, Antoine Delignat-Lavaud, Sylvan Clebsch, Kapil Vaswani, and Vikas Bhatia. Toward confidential cloud computing: Extending hardwareenforced cryptographic protection to data while in use. ACM Queue: Tomorrow’s Computing Today, 19(1):49–76, February 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3454122.3456125.

Rabaey:2003:DIC

[RCN03] Jan M. Rabaey, Anantha P.

Chandrakasan, and Borivoje Nikolić. Digital integrated circuits: a design perspective. Prentice Hall electronics and VLSI series. Pearson Education, Upper Saddle River, NJ, USA, second edition, 2003. ISBN 0-13-090996-3, 0-13-597444-5. xxii + 761 + 4 pp. LCCN TK7874.65 R33 2003.

Reed:2019:BFI

[Ree19] J. Paul Reed. Beyond the fix-it treadmill. ACM Queue: Tomorrow’s Computing Today, 17(6): ??, November/December 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3380774.3380780.

Richardson:2006:UEJ

[Ric06] Chris Richardson. Untangling Enterprise Java. ACM Queue: Tomorrow’s Computing Today, 4(5):36–44, June 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Richardson:2008:ODL

[Ric08] Chris Richardson. ORM in dynamic languages. ACM Queue: Tomorrow’s Computing Today, 6(3):28–37, May/June 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Richardson:2014:DD

[Ric14]  Rick Richardson.  Disam-

biguating databases. ACM Queue: Tomorrow’s Computing Today, 12(11):20, November 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=2696453.

Riedel:2003:SSJ

[Rie03] Erik Riedel. Storage systems:

Not just a bunch of disks anymore. ACM Queue: Tomorrow’s Computing Today, 1(4):32–41, June 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ritter:2003:FW

[Rit03] Michael W. Ritter. The future of WLAN. ACM Queue: Tomorrow’s Computing Today, 1(3):18–27, May 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Rixner:2008:NVB

[Rix08] Scot Rixner. Network virtualization: breaking the performance barrier. ACM Queue: Tomorrow’s Computing Today, 6(1):36–ff, January 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Robinson:2004:ENP

[Rob04]  Simon Robinson.  Expert

.NET Programming. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-222-0 (paperback). 750 pp. LCCN QA76.625 .R635 2004b. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://silk.library.umass.
edu:2048/login?url=http://
www.books24x7.com/marc.asp?
isbn=1590592220.

Rogers:2005:LF

[Rog05] Daniel Rogers. Lessons from the floor. ACM Queue: Tomorrow’s Computing Today, 3(10): 26–32, December 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Rose:2003:OHS

[Ros03]  Marshall T. Rose.  Opin-

ion: On helicopters and submarines. ACM Queue: Tomorrow’s Computing Today, 1(8): 10, 12–13, November 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Rosenblum:2004:RVM

[Ros04] Mendel Rosenblum. The reincarnation of virtual machines. ACM Queue: Tomorrow’s Computing Today, 2(5):34–40, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Rosenthal:2010:KBS

[Ros10] David S. H. Rosenthal. Keeping bits safe: How hard can it be? ACM Queue: Tomorrow’s Computing Today, 8(10):1–10, October 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Rosenthal:2011:DDC

[Ros11] David Rosenthal. Deduplicating devices considered harmful. ACM Queue: Tomorrow’s Computing Today, 9(5):30, May 2011. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Roussos:2007:SVG

[Rou07] Kostadis Roussos. Storage virtualization gets smart. ACM Queue: Tomorrow’s Computing Today, 5(6):38–44, September 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ratner:2018:KBC

[RR18] Alex Ratner and Christopher Ré. Knowledge base construction in the machine-learning era. ACM Queue: Tomorrow’s Computing Today, 16 (3):50, May 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3243045.

Rubel:2006:HE

[Rub06] Dan Rubel. The heart of eclipse.

ACM Queue: Tomorrow’s Computing Today, 4(8):36–44, October 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Russell:2004:ICS

[Rus04] James Russell. Interview: a conversation with Sam Leffler. ACM Queue: Tomorrow’s Computing Today, 2(3):16–22, May 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Russell:2008:BOR

[Rus08] Craig Russell. Bridging the object-relational divide. ACM Queue: Tomorrow’s Computing Today, 6(3):18–28, May/ June 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Russinovich:2023:CCE

[Rus23] Mark Russinovich. Confidential computing: Elevating cloud security and privacy: Working toward a more secure and innovative future. ACM Queue: Tomorrow’s Computing Today, 21(4):44–48, July 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3623461.

Ridoux:2010:PRT

[RV10] Julien Ridoux and Darryl Veitch. Principles of robust timing over the Internet. ACM Queue: Tomorrow’s Computing Today, 8(4):1–14, April 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Roeder:2006:LQM

[RW06] Nicholas Roeder and Martin Wheatman. Letters: Questioning MDD; the economics of outsourcing. ACM Queue: Tomorrow’s Computing Today, 4 (2):7, March 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Rys:2011:SS

[Rys11] Michael Rys. Scalable SQL.

ACM Queue: Tomorrow’s Computing Today, 9(4):30, April 2011. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sampemane:2014:IAC

[Sam14] Geetanjali Sampemane. Internal access controls. ACM Queue: Tomorrow’s Computing Today, 12(11):30, November 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2697395.

Sanders:2004:ICD

[San04] Lucy Sanders. Interview: a conversation with Donald Peterson. ACM Queue: Tomorrow’s Computing Today, 2(6): 22–29, September 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sarma:2004:IR

[Sar04]  Sanjay Sarma.  Integrating

RFID. ACM Queue: Tomorrow’s Computing Today, 2(7): 50–57, October 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sienko:2003:MC

[SARC03] Tanya Sienko, Andrew Adamatzky, Nicholas G. Rambidi, and Michael Conrad, editors. Molecular computing. MIT Press, Cambridge, MA, USA, 2003. ISBN 0-262-19487-2 (hardcover). xvii + 257 pp. LCCN QA76.887 .M65 2003. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/fy043/2002044409.html.

Savage:2015:CW

[Sav15] Taylor Savage. Componentizing the Web. ACM Queue: Tomorrow’s Computing Today, 13(8):60, September/October 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2844732.

Saxe:2010:PES

[Sax10] Eric Saxe. Power-efficient software. ACM Queue: Tomorrow’s Computing Today, 8(1):10, January 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sam-Bodden:2006:BPN

[SB06] Brian Sam-Bodden. Beginning POJOs: from novice to professional. Apress, Berkeley, CA, USA, 2006. ISBN 1-59059-596-3 (paperback). LCCN QA76.73.J38 S36 2006eb; QA76.73 .J38 INTERNET. URL http:

//ezproxy.lib.ucalgary.ca:
2048/login?url=http://library.
books24x7.com/library.asp?
bookid=14653.

Serrano:2012:MPH

[SB12] Manuel Serrano and Gérard Berry. Multitier programming in Hop. ACM Queue: Tomorrow’s Computing Today, 10 (7):10, July 2012. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Steel:2017:HPR

[SBB^{+}17] Edward Steel, Yanik Berube, Jonas Bonér, Ken Britton, and Terry Coatta. Hootsuite: In pursuit of reactive systems. ACM Queue: Tomorrow’s Computing Today, 15(3):60, May 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3131240.

Smart:2023:MCS

[SBS^{+}23] Nigel Smart, Joshua W. Baron, Sanjay Saravanan, Jordan Brandt, and Atefeh Mashatan. Multiparty computation: To secure privacy, do the math: a discussion with Nigel Smart, Joshua W. Baron, Sanjay Saravanan, Jordan Brandt, and Atefeh Mashatan. ACM Queue: Tomorrow’s Computing Today, 21(6):78–100, November 2023. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3639448.

Srinivasan:2021:BAS

[SC21] Ramya Srinivasan and Ajay Chander. Biases in AI systems: a survey for practitioners. ACM Queue: Tomorrow’s Computing Today, 19(2):45–64, April 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3466132.3466134.

Stachour:2009:YDK

[SCB09] Paul Stachour and David Collier-Brown. You don’t know jack about software maintenance. ACM Queue: Tomorrow’s Computing Today, 7(9): 50, October 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Schneier:2004:SA

[Sch04a] Bruce Schneier. Sensible authentication. ACM Queue: Tomorrow’s Computing Today, 1 (10):74–78, February 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Schweitzer:2004:BRL

[Sch04b] Martin Schweitzer. Book review: Linux in a Nutshell, 4th ed., by Ellen Siever, Stephen Figgins, and Aaron Weber, O’Reilly Books, 2003, $39.95, ISBN 0-596-00482-6. ACM Queue: Tomorrow’s Computing Today, 1(10):88, February 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Schneier:2005:AT

[Sch05] Bruce Schneier. Attack trends: 2004 and 2005. ACM Queue:

Tomorrow’s Computing Today, 3(5):52–53, June 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Schrock:2009:DAP

[Sch09] Eric Schrock. Debugging AJAX in production. ACM Queue: Tomorrow’s Computing Today, 7 (1):3:1–3:??, January 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Schlossnagle:2016:TF

[Sch16] Theo Schlossnagle. Time, but faster. ACM Queue: Tomorrow’s Computing Today, 14(6): 30, November 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3036398.

Schlossnagle:2017:MDW

[Sch17a] Theo Schlossnagle. Monitoring in a DevOps world. ACM Queue: Tomorrow’s Computing Today, 15(6):10, November 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3178371.

Schwarzkopf:2017:CSD

[Sch17b] Malte Schwarzkopf. Cluster scheduling for data centers. ACM Queue: Tomorrow’s Computing Today, 15(5):70, September/October 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3173558.

Singh:2020:ECL

[SCQT20] Jatinder Singh, Jennifer Cobbe, Do Le Quoc, and Zahra Tarkhani. Enclaves in the clouds: Legal considerations and broader implications. ACM Queue: Tomorrow’s Computing Today, 18(6):78–114, December 2020. URL https:

//dl.acm.org/doi/10.1145/
3442632.3448126.

Shortridge:2023:OCM

[SD23] Kelly Shortridge and Josiah Dykstra. Opportunity cost and missed chances in optimizing cybersecurity: The loss of potential gain from other alternatives when one alternative is chosen. ACM Queue: Tomorrow’s Computing Today, 21 (1):30–56, January 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3588041.

Sloss:2017:CSA

[SDRB17] Benjamin Treynor Sloss, Mike Dahlin, Vivek Rau, and Betsy Beyer. The calculus of service availability. ACM Queue: Tomorrow’s Computing Today, 15 (2):40, March 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3096459.

Seacord:2016:UR

[Sea16] Robert C. Seacord. Uninitialized reads. ACM Queue: Tomorrow’s Computing Today, 14 (6):50, November 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3041020.

Seeley:2003:CSP

[See03] Donn M. Seeley. Coding smart:

People vs. tools. ACM Queue: Tomorrow’s Computing Today, 1(6):32–40, September/October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Seeley:2004:HWF

[See04] Donn Seeley. How not to write Fortran in any language. ACM Queue: Tomorrow’s Computing Today, 2(9):58–65, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Seltzer:2005:BRD

[Sel05] Margo Seltzer. Beyond relational databases. ACM Queue: Tomorrow’s Computing Today, 3(3):50–58, April 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sestoft:2002:JP

[Ses02] Peter Sestoft. Java precisely.

MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-69276-7 (paperback). vii + 118 pp. LCCN QA76.73.J38 S435 2002.

Sessions:2004:FBO

[Ses04] Roger Sessions. Fuzzy boundaries: objects, components, and Web services. ACM Queue: Tomorrow’s Computing Today, 2(9):40–47, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sherburne:2004:YDK

[SF04] Phil Sherburne and Cary Fitzgerald. You don’t know jack about VoIP. ACM Queue: Tomorrow’s Computing Today, 2 (6):30–38, September 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Siever:2003:LN

[SFW03] Ellen Siever, Stephen Figgins, and Aaron Weber. Linux in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 2003. ISBN 0-596-00482-6. xiv + 928 pp. LCCN QA76.76.O63 S5582 2003.

Schmidt:2004:LAF

[SGN04] Douglas C. Schmidt, Aniruddha Gokhale, and Balachandran Natarajan. Leveraging application frameworks. ACM Queue: Tomorrow’s Computing Today, 2(5):66–75, July 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Shahri:2003:BLB

[Sha03] Homayoun Shahri. Blurring lines between hardware and software. ACM Queue: Tomorrow’s Computing Today, 1 (2):42–49, April 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Shahri:2004:MAD

[Sha04a] Homayoun Shahri. Mapping algorithms to DSP architectures. ACM Queue: Tomorrow’s Computing Today, 2(1): 32–41, March 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Shapiro:2004:SHM

[Sha04b] Michael W. Shapiro. Selfhealing in modern operating systems. ACM Queue: Tomorrow’s Computing Today, 2(9): 66–75, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Shapiro:2009:PBL

[Sha09] Mike Shapiro. Purpose-built languages. ACM Queue: Tomorrow’s Computing Today, 7 (1):18–24, January 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sha:2019:REA

[Sha19] Sanjay Sha. The reliability of enterprise applications. ACM Queue: Tomorrow’s Computing Today, 17(5):40, September 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3374665.

Sheehy:2015:TNN

[She15] Justin Sheehy. There is no now. ACM Queue: Tomorrow’s Computing Today, 13(3): 20, March 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2745385.

Simons:2004:EVS

[Sim04] Barbara Simons. Electronic voting systems: the good, the bad, and the stupid. ACM Queue: Tomorrow’s Computing Today, 2(7):20–26, October 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Singh:1999:CBE

[Sin99] Simon Singh. The code book: the evolution of secrecy from Mary, Queen of Scots, to quantum cryptography. Doubleday, New York, NY, USA, 1999. ISBN 0-385-49531-5, 0-385-49532-3. xiii + 402 pp. LCCN Z103 .S56 1999. US$24.95.

Singh:2016:CLL

[Sin16] Satnam Singh. Cluster-level logging of containers with containers. ACM Queue: Tomorrow’s Computing Today, 14 (3):30, May 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2965647.

Sites:2018:BHW

[Sit18] Richard L. Sites. Benchmarking “Hello, World!”. ACM Queue: Tomorrow’s Computing Today, 16(5):10, September 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3291278.

Sutherland:2020:SEC

[SJK20] Jeff Sutherland, Ivar Jacobson, and Brian Kerr. Scrum essentials cards: Experiences of Scrum teams improving with essence. ACM Queue: Tomorrow’s Computing Today, 18 (3):83–106, June 2020. URL

https://dl.acm.org/doi/10.
1145/3411757.3418775.

Steimann:2005:CC

[SK05] Friedrich Steimann and Thomas Kühne. Coding for the code. ACM Queue: Tomorrow’s Computing Today, 3(10):44– 51, December 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Senske:2004:LSD

[SKC^{+}04] Randy Senske, Azam Khan, Taylor Cowan, Phillip Laplante, and The Editors. Letters: Sentient data: The notion of location; stand-up meetings really suck; errata: The Game of Life.

ACM Queue: Tomorrow’s Computing Today, 1(10):6, February 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sicker:2004:VSA

[SL04] Douglas C. Sicker and Tom Lookabaugh. VoIP security: Not an afterthought. ACM Queue: Tomorrow’s Computing Today, 2(6):56–64, September 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sutter:2005:SCR

[SL05] Herb Sutter and James Larus. Software and the concurrency revolution. ACM Queue: Tomorrow’s Computing Today, 3 (7):54–62, September 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stone:2004:IMI

[SM04] John Stone and Sarah Merrion. Instant messaging or instant headache? ACM Queue: Tomorrow’s Computing Today, 2(2):72–80, April 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sanchez:2005:LSI

[SM05a] Vladimir Sanchez and Heidi Markovitz. Letters: Selinger’s inspiring words; plug-in predecessors. ACM Queue: Tomorrow’s Computing Today, 3 (5):12, June 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sperberg-McQueen:2005:XSS

[SM05b] C. M. Sperberg-McQueen. XML < and semi-structured data>. ACM Queue: Tomorrow’s Computing Today, 3(8):34–41, October 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Smaalders:2006:PAP

[Sma06] Bart Smaalders. Performance anti-patterns. ACM Queue: Tomorrow’s Computing Today, 4 (1):44–50, February 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Smith:2009:FPT

[Smi09] Jonathan M. Smith. Fighting physics: a tough battle. ACM Queue: Tomorrow’s Computing Today, 7(3):20–21, April 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sengupta:2004:LCT

[SMM^{+}04] Subir Sengupta, Chris Maeda, Pamela R. Mason, Jayson Durham, Cynthia Browning, Mike Zom, Chris Vickery, Jon Ogden, Stephen Vinowski, Dave McComb, and Peter Thomas. Letters: Crossing time zones; is DSP right for you?; can business process management deliver?; that Kelly-Bootle!; stricken by UML fever; errata [Alexander Wolfe, Toolkit: Get Your Graphics On: OpenGL Advances with the Times, ACM Queue 2(1), 10, 12–13 (2004)]. ACM Queue: Tomorrow’s Computing Today, 2(4):10, June 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Wol04c].

Stark:2023:HFC

[SMM23] Samuel W. Stark, A. Theodore Markettos, and Simon W. Moore. How flexible is CXL’s memory protection?: Replacing a sledgehammer with a scalpel. ACM Queue: Tomorrow’s Computing Today, 21 (3):54–64, May 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606014.

Sloss:2018:MM

[SNR18] Benjamin Treynor Sloss, Shylaja Nukala, and Vivek Rau. Metrics that matter. ACM Queue: Tomorrow’s Computing Today, 16(6):30, November 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3309571.

Spector:2023:MTJ

[SNW^{+}23] Alfred Spector, Peter Norvig, Chris Wiggins, Jeannette Wing, Ben Fried, and Michael Tingley. More than just algorithms: a discussion with Alfred Spector, Peter Norvig, Chris Wiggins, Jeannette Wing, Ben Fried, and Michael Tingley. ACM Queue: Tomorrow’s Computing Today, 21(1):102–129, January 2023. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3587481.

Sohn:2007:UD

[Soh07] David Sohn. Understanding DRM. ACM Queue: Tomorrow’s Computing Today, 5(7): 32–39, November 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Seltzer:2021:CMS

[SOM21] Margo Seltzer, Mike Olson, and Kirk McCusick. A conversation with Margo Seltzer and Mike Olson: The history of Berkeley DB. ACM Queue: Tomorrow’s Computing Today, 19(5):69–86, September 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3494834.3498582.

Souders:2008:HPW

[Sou08] Steve Souders. High performance Web sites. ACM Queue: Tomorrow’s Computing Today, 6(6):30–37, October 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Shortridge:2021:LAN

[SP21] Kelly Shortridge and Ryan Petrich. Lamboozling attackers: a new generation of deception: Software engineering teams can exploit attackers’ human nature by building deception environments. ACM Queue: Tomorrow’s Computing Today, 19(5):26–59, September 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3494834.3494836.

Sparks:2007:SBB

[Spa07] Robert Sparks. SIP: basics and beyond. ACM Queue: Tomorrow’s Computing Today, 5 (2):22–33, March 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Spinellis:2003:RWC

[Spi03] Diomidis Spinellis. Reading, writing, and code. ACM Queue: Tomorrow’s Computing Today, 1(7):84–89, October 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Spinellis:2006:RSS

[Spi07] D. Spinellis. Review of ‘Sustainable Software Development: An Agile Perspective’ by Kevin Tate, Addison-Wesley Professional, 2005, $39.99, ISBN 0-321-28608-1. ACM Queue: Tomorrow’s Computing Today, 4(10):49, December/January 2006–2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Tat06].

Stephens:2003:EPR

[SR03] Matt Stephens and Doug Rosenberg. Extreme programming refactored: the case against XP. Apress, Berkeley, CA, USA, 2003. ISBN 1-59059-096-1. xxviii + 400 pp. LCCN QA76.76.D47 R67 2003.

Smith:2007:PMH

[SS07] Stewart Smith and Michael Still.

Practical MythTV: building a PVR and Media Center PC. Apress, Berkeley, CA, USA, 2007. ISBN 1-59059-779-6. LCCN ????

Svensson:2014:DET

[SSN14] Bo Joel Svensson, Mary Sheeran, and Ryan Newton. Design exploration through codegenerating DSLs. ACM Queue: Tomorrow’s Computing Today, 12(4):40, April 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Scambray:2006:HEW

[SSS06] Joel Scambray, Mike Shema, and Caleb Sima. Hacking exposed: Web applications. Mc-Graw-Hill, New York, NY, USA, second edition, 2006. ISBN 0-07-226299-0 (paperback). xxix + 520 pp. LCCN TK5105.59 .S32 2006.

Stanik:2004:N

[Sta04a] John Stanik. News 2.0: It isn’t going away … but IT jobs are; 4 GHz goes bust; Trojan-spyware deathmatch; desktop search madness. ACM Queue: Tomorrow’s Computing Today, 2 (9):13, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2004:NNW

[Sta04b] John Stanik. News 2.0: New Web provides battery-powered option; there’s nothing like a map; shall I compare thee to a big squeeze?; Linux support: a black hole? ACM Queue: Tomorrow’s Computing Today, 2(8):14, November 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2005:EMA

[Sta05a] John Stanik. From the Editors: Mobile applications get real. ACM Queue: Tomorrow’s Computing Today, 3(4):8, May 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2005:NPR

[Sta05b] John Stanik. News 2.0: a patent reduction in IP squabbles; widgets, Macintosh style; biometrics goes 3D. ACM Queue: Tomorrow’s Computing Today, 3 (2):16, March 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2005:NAS

[Sta05c] John Stanik. News 2.0: Acoustical spying techniques cooked up in the lab; .ODF delivers crushing blow to .DOC; Wikibooks— coming to a classroom near you? ACM Queue: Tomorrow’s Computing Today, 3(9):10, November 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2005:NGC

[Sta05d] John Stanik. News 2.0: Gridding for a cure; wanted: Windows robots; lost in licenses. ACM Queue: Tomorrow’s Computing Today, 3(3):11, April 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2005:NKS

[Sta05e] John Stanik. News 2.0: Kernel solid, but vulnerable; IT security get physical; A warm reception for reject chips. ACM Queue: Tomorrow’s Computing Today, 3(4):14, May 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2005:NLO

[Sta05f] John Stanik. News 2.0: Losing our edge? the real cost of Linux; say no to crackberries. ACM Queue: Tomorrow’s Computing Today, 3(5):14, June 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2005:NOM

[Sta05g] John Stanik. News 2.0: OSx86 for the masses?; OS? we don’t need no stinkin’ OS; real crime descends upon virtual world. ACM Queue: Tomorrow’s Computing Today, 3(8):10, October 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2005:NRW

[Sta05h]  John Stanik.  News 2.0:

“Ransom-ware” on the loose; adopt IPv6 or die; IBM steps up on Linux training. ACM Queue: Tomorrow’s Computing Today, 3(6):10, July 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2005:NWL

[Sta05i] John Stanik. News 2.0: Web 2.0—looking toward the future or reviving the past?; easy come, easy splog; carpal tunnel is so nineties! ACM Queue: Tomorrow’s Computing Today, 3 (10):10, December 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

State:2005:RNS

[Sta05j] Radu State. Review of Network Security Architectures by Sean Convery, Cisco Press, 2004, $55.00, ISBN 1-58705-115-X. ACM Queue: Tomorrow’s Computing Today, 3(1):61, February 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

State:2005:RGH

[Sta05k] Radu State. Review of “Gray Hat Hacking: The Ethical Hacker’s Handbook by Shon Harris, Allen Harper, Chris Eagle, Jonathan Ness, and Michael Lester”, McGraw-Hill Osborne Media, 2004, $49.99 ISBN 0-07-225709-1. ACM Queue: Tomorrow’s Computing Today, 3 (6):60, July 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stange:2006:RPN

[Sta06a] Melissa C. Stange. Review of “Pro .NET 1.1 Remoting, Reflection, and Threading by David Curran, Fabio Ferracchiati, Syed Gilani, et al.,” Apress, 2005, $59.99, ISBN 1-59059-452-5. ACM Queue: Tomorrow’s Computing Today, 4 (6):58, July/August 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2006:NAS

[Sta06b] John Stanik. News 2.0: Attackers set sights on Macs; sneakernet security; IT’ paradox. ACM Queue: Tomorrow’s Computing Today, 4(3):8, April 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2006:NCT

[Sta06c] John Stanik. News 2.0: Carrying the torch of anti-spam vigilantism; taking constrained programming to the next level; the human intelligence revolution. ACM Queue: Tomorrow’s Computing Today, 4(6): 9, July/August 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2006:NIG

[Sta06d]  John Stanik.  News 2.0:

IPv6 gains government traction; search-as-a-service gets serious; when detecting life, is biometrics all thumbs? ACM Queue: Tomorrow’s Computing Today, 4(1):10, February 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2006:NMW

[Sta06e] John Stanik. News 2.0: Measuring the Web’s mood; software slump on the mend? mobile gaming for hackers. ACM Queue: Tomorrow’s Computing Today, 4(5):7, June 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2006:NMP

[Sta06f] John Stanik. News 2.0: Medical profession slow to embrace ’net… or is it?; accessibility features not just for the disabled; emerging markets embrace 64-bit. ACM Queue: Tomorrow’s Computing Today, 4 (9):8, November 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2006:NOS

[Sta06g] John Stanik. News 2.0: Open source gets Mac attack; oh no, not another consortium!; get that chip out of my drink. ACM Queue: Tomorrow’s Computing Today, 4(7):10, September 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2006:NPP

[Sta06h] John Stanik. News 2.0: Programmers at play; hunch engine hones in; the real risks of RFID. ACM Queue: Tomorrow’s Computing Today, 4(4):9, May 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2006:NSS

[Sta06i]  John Stanik.  News 2.0:

Symantec’s semantics; spyware defined—for now; more over spyware, there’s myware. ACM Queue: Tomorrow’s Computing Today, 4(2):8, March 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2006:NML

[Sta06j] John Stanik. News 2.0: The mobile Linux challenge; Google not into googling; keeping online video legit. ACM Queue: Tomorrow’s Computing Today, 4(8):8, October 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

State:2006:RSS

[Sta06k] Radu State. Review of Software Security: Building Security In by Gary McGraw, Addison-Wesley Professional, 2006, $49.99, ISBN 0-321-35670-5. ACM Queue: Tomorrow’s Computing Today, 4(7): 44, September 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

State:2006:RED

[Sta06l] Radu State. Review of “Extrusion Detection: Security Monitoring for Internal Intrusions” by Richard Bejtlich, Adddison-Wesley Professional, 2005, $49.99, ISBN 0-321-34996-2. ACM Queue: Tomorrow’s Computing Today, 4 (3):42, April 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

State:2006:RPE

[Sta06m] Radu State. Review of “Phishing Exposed” by Lance James and Joe Stewart, Syngress Publishing, 2005, $49.95, ISBN 1-59749-030-X. ACM Queue: Tomorrow’s Computing Today, 4 (9):50, November 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Jam05].

State:2006:RSV

[Sta06n] Radu State. Review of “Switching to VoIP” by Ted Wallingford, O’Reilly & Associates, 2005, $39.95, ISBN 0-596-00868-6. ACM Queue: Tomorrow’s Computing Today, 4(1): 62, February 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Wal05].

Stanik:2006:NFW

[Sta07a] John Stanik. News 2.0: Fox and the weasel; down on the wireless farm; second-life commerce meets first-life IRS. ACM Queue: Tomorrow’s Computing Today, 4(10):10, December/January 2006–2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:Ca

[Sta07b] John Stanik. Calendar. ACM Queue: Tomorrow’s Computing Today, 5(1):52, February 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:Cb

[Sta07c] John Stanik. Calendar. ACM Queue: Tomorrow’s Computing Today, 5(2):58, March 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:Cc

[Sta07d] John Stanik. Calendar. ACM Queue: Tomorrow’s Computing Today, 5(3):46, April 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:Cd

[Sta07e] John Stanik. Calendar. ACM Queue: Tomorrow’s Computing Today, 5(4):58, May–June 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:Ce

[Sta07f] John Stanik. Calendar. ACM Queue: Tomorrow’s Computing Today, 5(5):48, July 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:Cf

[Sta07g] John Stanik. Calendar. ACM Queue: Tomorrow’s Computing Today, 5(6):46, September 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:Cg

[Sta07h] John Stanik. Calendar. ACM Queue: Tomorrow’s Computing Today, 5(7):52, November 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:ICCb

[Sta07i] John Stanik. Interview: a conversation with Cory Doctorow and Hal Stern. ACM Queue: Tomorrow’s Computing Today, 5(3):16–23, April 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:ICCa

[Sta07j] John Stanik. Interview: a conversation with Cullen Jennings and Doug Wadkins. ACM Queue: Tomorrow’s Computing Today, 5(2):14–20, March 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:ICJa

[Sta07k] John Stanik. Interview: a conversation with Jamie Butler. ACM Queue: Tomorrow’s Computing Today, 5(1): 16–23, February 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:ICJc

[Sta07l] John Stanik. Interview: a conversation with Jeff Bonwick and Bill Moore. ACM Queue: Tomorrow’s Computing Today, 5 (6):13–19, September 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:ICJb

[Sta07m] John Stanik. Interview: a conversation with Joel Spolsky. ACM Queue: Tomorrow’s Computing Today, 5(5):12–20, July 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:ICMb

[Sta07n] John Stanik. Interview: a conversation with Mary Lou Jepsen. ACM Queue: Tomorrow’s Computing Today, 5(7): 9–15, November 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:ICMa

[Sta07o] John Stanik. Interview: a conversation with Michael Stonebraker and Margo Seltzer. ACM Queue: Tomorrow’s Computing Today, 5(4):16–23, May– June 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:LKC

[Sta07p] John Stanik. Letters: Kudos to curmudgeon; long live the wire. ACM Queue: Tomorrow’s Computing Today, 5(5):8, July 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:NGD

[Sta07q] John Stanik. News 2.0: Googlebombs away; DST: This year’s Y2K?; Linux through Windows. ACM Queue: Tomorrow’s Computing Today, 5(2):7, March 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:NMQ

[Sta07r] John Stanik. News 2.0: Malware — quantity over quality; the home robotics revolution; roboexotica. ACM Queue: Tomorrow’s Computing Today, 5 (1):8, February 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:NVR

[Sta07s] John Stanik. News 2.0: Virtualization reconsidered; the power of PS3; Anti-P2P software targets universities. ACM Queue: Tomorrow’s Computing Today, 5(3):10, April 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:WYH

[Sta07t] John Stanik. What’s on your hard drive? ACM Queue: Tomorrow’s Computing Today, 5 (1):10, February 2007. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

State:2006:RHE

[Sta07u] Radu State. Review of ‘Hacking Exposed: Web Applications’, 2nd edition by Joel Scambray, Mike Shema, and Caleb Sima, McGraw-Hill Osborne Media, 2006, $49.99, ISBN 0-07-226299-0. ACM Queue: Tomorrow’s Computing Today, 4(10): 49, December/January 2006– 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [SSS06].

State:2007:REL

[Sta07v] Radu State. Review of ‘Embedded Linux Primer: A Practical Real-world Approach’ by Christopher Hallinan, Prentice Hall PTR, 2006, $44.99, ISBN 0-13-167984-8. ACM Queue: Tomorrow’s Computing Today, 5 (3):44, April 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Hal07].

Stanik:2008:Ca

[Sta08a] John Stanik. Calendar. ACM Queue: Tomorrow’s Computing Today, 6(2):68, March 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2008:Cb

[Sta08b] John Stanik. Calendar. ACM Queue: Tomorrow’s Computing Today, 6(4):54, July/August 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2008:Cc

[Sta08c] John Stanik. Calendar. ACM Queue: Tomorrow’s Computing Today, 6(5):60, September 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2008:Cd

[Sta08d] John Stanik. Calendar. ACM Queue: Tomorrow’s Computing Today, 6(6):55, October 2008.

CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2008:Ce

[Sta08e] John Stanik. Calendar. ACM Queue: Tomorrow’s Computing Today, 6(7):42, November/December 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2008:ICJ

[Sta08f] John Stanik. Interview: a conversation with Jason Hoffman. ACM Queue: Tomorrow’s Computing Today, 6(1):7–13, January 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2008:ICK

[Sta08g] John Stanik. Interview: a conversation with Kurt Akeley and Pat Hanrahan. ACM Queue: Tomorrow’s Computing Today, 6(2):11–17, March 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2008:ICSa

[Sta08h]  John Stanik.  Interview:

a conversation with Steve Bourne, Eric Allman, and Bryan Cantrill. ACM Queue: Tomorrow’s Computing Today, 6(4): 8–15, July/August 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2008:ICSb

[Sta08i]  John Stanik.  Interview:

a conversation with Steve Bourne, Eric Allman, and Bryan Cantrill. ACM Queue: Tomorrow’s Computing Today, 6(5): 8–15, September 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2009:C

[Sta09a] John Stanik. Calendar. ACM Queue: Tomorrow’s Computing Today, 7(1):41, January 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2009:ICV

[Sta09b] John Stanik. Interview: a conversation with Van Jacobson. ACM Queue: Tomorrow’s Computing Today, 7(1):8–16, January 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stevenson:2003:BRC

[Ste03] William Stevenson. Book review: Cocoa in a Nutshell by Michael Beam and James Duncan Davidson, O’Reilly & Associates, 2003, $39.95, ISBN 0-596-00462-1. ACM Queue: Tomorrow’s Computing Today, 1 (8):72, November 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Steinmann:2007:UCS

[Ste07a] Martin J. Steinmann. Unified communications with SIP. ACM Queue: Tomorrow’s Computing Today, 5(2):50–55, March 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stephenson:2007:ER

[Ste07b]  Ian Stephenson.  Essential

RenderMan. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2007. ISBN 1-84628-344-2 (paperback). xi + 266 pp. LCCN TR897.7 .S75 2007.

Stenn:2014:SNT

[Ste15] Harlan Stenn. Securing the Network Time Protocol. ACM Queue: Tomorrow’s Computing Today, 13(1):20, December/ January 2014–2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2717320.

Stobie:2005:TDB

[Sto05] Keith Stobie. Too darned big to test. ACM Queue: Tomorrow’s Computing Today, 3(1): 30–37, February 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Strong:2005:EGC

[Str05] Paul Strong. Enterprise Grid computing. ACM Queue: Tomorrow’s Computing Today, 3 (6):50–59, July 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Supnik:2004:SVM

[Sup04] Bob Supnik. Simulators: Virtual machines of the past (and future). ACM Queue: Tomorrow’s Computing Today, 2(5): 52–58, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Suver:2005:CCD

[Suv05] Chris Suver. Curmudgeon: The cost of data. ACM Queue: Tomorrow’s Computing Today, 3(8):64, 62–63, October 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sproull:2014:APC

[SW14] Robert Sproull and Jim Waldo. The API performance contract. ACM Queue: Tomorrow’s Computing Today, 12(1):10, January 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sweeney:2013:DOA

[Swe13] Latanya Sweeney. Discrimination in online ad delivery. ACM Queue: Tomorrow’s Computing Today, 11(3):10, March 2013. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Sompolinsky:2017:BUI

[SZ17] Yonatan Sompolinsky and Aviv Zohar. Bitcoin’s underlying incentives. ACM Queue: Tomorrow’s Computing Today, 15(5):50:29–50:52, September/October 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3168362.

Takeuchi:2006:AIC

[Tak06] Len Takeuchi. ASPs: the integration challenge. ACM Queue: Tomorrow’s Computing Today, 4(5):46–52, June 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Tate:2006:SSD

[Tat06] Kevin Tate. Sustainable software development: an agile perspective. Agile software development series. Addison-Wes-ley, Reading, MA, USA, 2006. ISBN 0-321-28608-1. xxxiii + 226 pp. LCCN QA76.76.D47 T38 2006. URL http:

//www.loc.gov/catdir/toc/
ecip0515/2005019120.html.

Tang:2003:BIM

[TB03] John C. Tang and James Bo Begole. Beyond instant messaging. ACM Queue: Tomorrow’s Computing Today, 1(8): 28–37, November 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Trudel:2003:DSE

[TCR03] Bernard Trudel, Sean Convery, and Russell Rice. Designing secure enterprise networks. Cisco Press, Indianapolis, IN, USA, 2003. ISBN 1-58705-115-X. ???? pp. LCCN TK5105.59 .C5793 2004. Duplicate ISBN with [Con04].

Tiganov:2021:DUS

[TDA21] Daniil Tiganov, Lisa Nguyen Quang Do, and Karim Ali. Designing UIs for static analysis tools: Evaluating tool design guidelines with SWAN. ACM Queue: Tomorrow’s Computing Today, 19(4):97–118, July 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487019.3487026.

Tellman:2023:DFC

[Tel23]  Zachary Tellman.  Design-

ing a framework for conversational interfaces: Combining the latest advances in machine learning with earlier approaches. ACM Queue: Tomorrow’s Computing Today, 21 (1):85–101, January 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3587715.

Thompson:2004:PSI

[TF04] Herbert H. Thompson and Richard Ford. Perfect storm: The insider, naivety, and hostility. ACM Queue: Tomorrow’s Computing Today, 2(4):58–65, June 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Thomas:2005:RBC

[Tho05] Marlin Thomas. Review of Building Clustered Linux Systems by Robert W. Lucke, Prentice Hall PTR, 2004, $49.99, ISBN 0-13-144853-6. ACM Queue: Tomorrow’s Computing Today, 3(4):60, May 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Thomson:2021:SAI

[Tho21] Patrick Thomson. Static analysis: an introduction: The fundamental challenge of software engineering is one of complexity. ACM Queue: Tomorrow’s Computing Today, 19(4):29–41, July 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487019.3487021.

Titus:2005:PNR

[Tit05] Tobin Titus, editor. Pro .NET 1.1 remoting, reflection, and threading. The expert’s voice in .NET. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-452-

5. xxiii + 647 pp. LCCN TK5105.597 .P76 2005. URL

http://www.loc.gov/catdir/
enhancements/fy0663/2006296447-
d.html;   http://www.loc.
gov/catdir/toc/fy0608/2006296447.
html.

Tomek:2005:RIP

[Tom05] Andreas Tomek. Review of “Intrusion Prevention and Active Response: Deploying Network and Host IPS” by Michael Rash, Angela D. Orebaugh, Graham Clark, Becky Pinkard, and Jake Babbin, Syngress Publishing, 2005, $49.99, ISBN 1-932266-47-X. ACM Queue: Tomorrow’s Computing Today, 3(10): 52, December 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Tomek:2006:RPY

[Tom06] Andreas Tomek. Review of “Protect your Windows Network: From Perimeter to Data by Jesper Johansson and Steve Riely,” Addison-Wesley Professional, 2005, $49.99, ISBN 0-321-33643-7. ACM Queue: Tomorrow’s Computing Today, 4 (4):42, May 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [JR05].

Toxen:2007:SDS

[Tox07] Bob Toxen. The seven deadly sins of Linux security. ACM Queue: Tomorrow’s Computing Today, 5(4):38–47, May– June 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Toxen:2014:NSS

[Tox14] Bob Toxen. The NSA and Snowden: Securing the allseeing eye. ACM Queue: Tomorrow’s Computing Today, 12 (3):40, March 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Terry:2003:CXW

[TR03] Douglas B. Terry and Venugopalan Ramasubramanian. Caching XML Web services for mobility. ACM Queue: Tomorrow’s Computing Today, 1 (3):70–78, May 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Tredennick:2003:IRS

[TS03] Nick Tredennick and Brion Shimamoto. The inevitability of reconfigurable systems. ACM Queue: Tomorrow’s Computing Today, 1(7):34–43, October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Tulach:2008:PAD

[Tul08] Jaroslav Tulach. Practical API design: confessions of a Java Framework architect. Apress, Berkeley, CA, USA, 2008. ISBN 1-4302-0973-9. LCCN ????

Turnlund:2003:DDL

[Tur03] Michael Turnlund. Distributed development: Lessons learned. ACM Queue: Tomorrow’s Computing Today, 1(9):26–31, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ullman:2003:BN

[Ull03] Ellen Ullman. The bug: a novel. Nan A. Talese, New York, NY, USA, 2003. ISBN 0-385-50860-3. 355 pp. LCCN PS3621.L45 B8 2003. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
bios/random054/2002073289.
html;   http://www.loc.gov/
catdir/description/random0413/
2002073289.html;   http://
www.loc.gov/catdir/samples/
random045/2002073289.html.

Ullrich:2021:RWS

[Ull21] Torsten Ullrich. Real-world string comparison: How to handle Unicode sequences correctly. ACM Queue: Tomorrow’s Computing Today, 19(3):107–116, June 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3475965.3478522.

Anonymous:2014:ENJ

[Vac14] Manish Vachharajani. Embedding Node.js into a highperformance network datapath. ACM Queue: Tomorrow’s Computing Today, 12(2):100, February 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Vachon:2020:IEP

[Vac20] Phil Vachon. The identity in everyone’s pocket: Keeping users secure through their smartphones. ACM Queue: Tomorrow’s Computing Today, 18 (4):61–94, August 2020. URL

https://dl.acm.org/doi/10.
1145/3424302.3428660.

Vachon:2021:KKD

[Vac21] Phil Vachon. The keys to the kingdom: a deleted private key, a looming deadline, and a last chance to patch a new static root of trust into the bootloader. QUEUE, 19(6): 57–65, November 2021. URL

https://dl.acm.org/doi/10.
1145/3511664.

Vachon:2023:SJA

[Vac23a] Phil Vachon. The security jawbreaker: Access to a system should not imply authority to use it. Enter the principle of complete mediation. ACM Queue: Tomorrow’s Computing Today, 21(5):13–21, September 2023. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3631478.

Vachon:2023:SMS

[Vac23b] Phil Vachon. Security mismatch: Security must be a business enabler, not a hinderer. ACM Queue: Tomorrow’s Computing Today, 21 (3):23–28, May 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606028.

Varajao:2021:SDD

[Var21] João Varajão. Software development in disruptive times: Creating a software solution with fast decision capability, agile project management, and extreme lowcode technology. ACM Queue: Tomorrow’s Computing Today, 19(1):94–103, February 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3454122.3458743.

Vasisht:2017:TNC

[Vas17] Deepak Vasisht. Toward a network of connected things. ACM Queue: Tomorrow’s Computing Today, 15(6):60, November 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3186470.

Vaughn:2007:HGV

[VB07] William R. Vaughn and Peter Blackburn. Hitchhiker’s guide to Visual Studio and SQL Server: best practice architectures and examples. Microsoft Windows server system series. Addison-Wesley, Reading, MA, USA, seventh edition, 2007. ISBN 0-321-24362-5 (paperback). xxiv + 1095 pp. LCCN QA76.76.A65 V39 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0620/2006030355.html.

Viredaz:2003:EMH

[VBH03] Marc A. Viredaz, Lawrence S. Brakmo, and William R. Hamburgen. Energy management on handheld devices. ACM Queue: Tomorrow’s Computing Today, 1(7):44–52, October 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Vahdat:2015:PBG

[VCR15] Amin Vahdat, David Clark, and Jennifer Rexford. A purpose-built global network: Google’s move to SDN. ACM Queue: Tomorrow’s Computing Today, 13(8):100, September/October 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2856460.

vanDeursen:2015:BPO

[vD15] Arie van Deursen. Beyond page objects: Testing Web applications with state objects. ACM Queue: Tomorrow’s Computing Today, 13(6):20, June 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2793039.

Vitharana:2024:CAS

[VD24] Padmal Vitharana and Shahir A.

Daya. Challenges in adopting and sustaining microservicebased software development: Organizational challenges can be more difficult than technical ones. ACM Queue: Tomorrow’s Computing Today, 22 (1):48–72, January 2024. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3649402.

Verner:2004:BPC

[Ver04]  Laury Verner.  BPM: The

promise and the challenge. ACM Queue: Tomorrow’s Computing Today, 2(1):82–91, March 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Vladimirov:2004:WF

[VGM04] Andrew A. Vladimirov, Konstantin V. Gavrilenko, and Andrei A. Mikhailovsky. Wi-Foo: Secrets of wireless hacking. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-20217-1. xxvii + 555 pp. LCCN TK5105.59 .V53 2004.

Videla:2017:MWC

[Vid17] Alvaro Videla. Metaphors we compute by. ACM Queue: Tomorrow’s Computing Today, 15 (3):40, May 2017. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

                                        URL http://queue.acm.org/
                                        detail.cfm?id=3127495.

Videla:2021:MCC

[Vid21]  Alvaro Videla.  Meaning

and context in computer programs: Sharing domain knowledge among programmers using the source code as the medium. ACM Queue: Tomorrow’s Computing Today, 19(5):60–68, September 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3494834.3498852.

Videla:2023:EIT

[Vid23] Alvaro Videla. Echoes of intelligence: Textual interpretation and large language models. ACM Queue: Tomorrow’s Computing Today, 21 (3):36–53, May 2023. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606011.

Viega:2005:SPS

[Vie05] John Viega. Security—problem solved? ACM Queue: Tomorrow’s Computing Today, 3 (5):40–50, June 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Vincent:2003:TSW

[Vin03] Christopher Vincent. Toolkit:

Scripting Web service prototypes. ACM Queue: Tomorrow’s Computing Today, 1(1): 22, 23–27, March 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Virkus:2005:PJP

[Vir05] Robert Virkus. Pro J2ME Polish: open source wireless Java tools suite. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-503-3 (hardcover). ???? pp. LCCN QA76.73.J38 V57 2005. URL

http://www.loc.gov/catdir/
enhancements/fy0663/2005016571-
d.html;   http://www.loc.

                                        gov/catdir/toc/ecip0514/2005016571.
                                        html.

Vixie:2007:DC

[Vix07] Paul Vixie. DNS complexity. ACM Queue: Tomorrow’s Computing Today, 5(3):24–29, April 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Vixie:2009:WD

[Vix09] Paul Vixie. What DNS is not. ACM Queue: Tomorrow’s Computing Today, 7(10):10, November 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Vixie:2014:RLS

[Vix14] Paul Vixie. Rate-limiting state. ACM Queue: Tomorrow’s Computing Today, 12(2):10, February 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Vixie:2014:GSG

[Vix15a] Paul Vixie. Go static or go home. ACM Queue: Tomorrow’s Computing Today, 13(1): 30, December/January 2014– 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

                                        /queue.acm.org/detail.cfm?
                                        id=2721993.

Vixie:2015:GSG

[Vix15b] Paul Vixie. Go static or go home. ACM Queue: Tomorrow’s Computing Today, 13(2): 30, February 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=2721993.

Viega:2004:SHT

[VM04] John Viega and Matt Messier.

Security is harder than you think. ACM Queue: Tomorrow’s Computing Today, 2(5): 60–65, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Vogels:2008:BSC

[Vog08a] Werner Vogels. Beyond server consolidation. ACM Queue: Tomorrow’s Computing Today, 6 (1):20–26, January 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Vogels:2008:EC

[Vog08b] Werner Vogels. Eventually consistent. ACM Queue: Tomorrow’s Computing Today, 6(6): 14–19, October 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Vogels:2009:CRC

[VOT^{+}09] Werner Vogels, Greg Olsen, Lew Tucker, Greg Badros, Geir Ramleth, and Steve Bourne. CTO roundtable: Cloud computing. ACM Queue: Tomorrow’s Computing Today, 7 (5):1:1–1:18, June 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Vaidyanathan:2003:DRA

[VT03] Ramachandran Vaidyanathan and Jerry L. Trahan. Dynamic reconfiguration: architectures and algorithms. Series in computer science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48189-8. xxiii + 512 pp. LCCN QA76.9.A73 V35 2003.

Varajao:2023:LCD

[VTA23] João Varajão, António Trigo, and Miguel Almeida. Low-code development productivity: “Is winter coming” for code-based technologies? ACM Queue: Tomorrow’s Computing Today, 21(5):87–107, September 2023. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3631183.

VanDyk:2007:PDD

[VW07] John K. VanDyk and Matt Westgate. Pro Drupal development. Expert’s voice in open source; Books for professionals by professionals. Apress, Berkeley, CA, USA, 2007. ISBN 1-59059-755-9. xxv + 428 pp. LCCN TK5105.888 .V363 2007.

Vaccaro:2019:EMM

[VW19] Michelle Vaccaro and Jim Waldo. The effects of mixing machine learning and human judgment. ACM Queue: Tomorrow’s Computing Today, 17(4):30, July 2019. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3363293.

Woods:2019:RCR

[WA19] David D. Woods and John Allspaw. Revealing the critical role of human performance in software. ACM Queue: Tomorrow’s Computing Today, 17(6): ??, November/December 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3380774.3380776.

Wadlow:2005:AC

[Wad05] Thomas Wadlow. The answer is 42 of course. ACM Queue: Tomorrow’s Computing Today, 3(5):34–39, June 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wadlow:2014:WMY

[Wad14] Thomas Wadlow. Who must you trust? ACM Queue: Tomorrow’s Computing Today, 12 (5):30, May 2014. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wainwright:2005:PP

[Wai05] Peter Wainwright. Pro Perl.

The expert’s voice in open source. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-438-X (paperback). xxiii + 1037 pp. LCCN QA76.73.P22 W33 2005.

Wallingford:2005:SV

[Wal05] Ted Wallingford. Switching to VoIP. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00868-6. xxi + 477 pp. LCCN TK5105.8865 .W35 2005.

Wallace:2007:RDW

[Wal07] Will Wallace. Review of ‘Designing with Web Standards’ (second edition) by Jeffrey Zeldman, Peachpit Press, 2006, $44.99, ISBN 0-321-38555-1. ACM Queue: Tomorrow’s Computing Today, 5(4):56, May– June 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Zel07].

Waldo:2008:SGV

[Wal08] Jim Waldo. Scaling in games & virtual worlds. ACM Queue: Tomorrow’s Computing Today, 6(7):10–16, November/December 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Waldo:2018:HGB

[Wal18] Jim Waldo. A Hitchhiker’s guide to the blockchain universe. ACM Queue: Tomorrow’s Computing Today, 16(6): 10, November 2018. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3305265.

Want:2004:MR

[Wan04] Roy Want. The magic of RFID. ACM Queue: Tomorrow’s Computing Today, 2(7):40–48, October 2004. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic).

Wang:2015:EW

[Wan15] Daniel C. Wang. From the EDVAC to WEBVACs. ACM Queue: Tomorrow’s Computing Today, 13(4):10, April 2015. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=2756508.

Wang:2024:OAG

[Wan24] Yifei Wang. From open access to guarded trust: Experimenting responsibly in the age of data privacy. ACM Queue: Tomorrow’s Computing Today, 22(1): 100–113, January 2024. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3649305.

Wasserman:2003:ICS

[Was03] Tony Wasserman. Interview: a conversation with Steve Hagan. ACM Queue: Tomorrow’s Computing Today, 1(9):18, 20–25, December/January 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Watson:2013:DAC

[Wat13] Robert N. M. Watson. A decade of OS access-control extensibility. ACM Queue: Tomorrow’s Computing Today, 11 (1):1–22, January 2013. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Walls:2005:SA

[WB05] Craig Walls and Ryan Breidenbach. Spring in action. Manning, Greenwich, CT, USA, 2005. ISBN 1-932394-35-4 (paperback). xxviii + 444 pp. LCCN QA76.73.J38 W342 2005.

                                        URL http://www.loc.gov/
                                        catdir/enhancements/fy0626/
                                        2005276355-b.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy0626/2005276355-d.html.

Weathersby:2003:SFD

[Wea03] John M. Weathersby, Jr. Sidebar: a foot in the door: Can open source find traction in government? ACM Queue: Tomorrow’s Computing Today, 1(5): 52–53, July/August 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wells:2005:GDD

[Wel05] April J. Wells. Grid database design. Auerbach Publications, Boca Raton, FL, USA, 2005. ISBN 0-8493-2800-4. xx + 287 pp. LCCN QA76.9.C58 W45 2005.

Wiedemann:2019:DP

[WFW^{+}19] Anna Wiedemann, Nicole Forsgren, Manuel Wiesche, Heiko Gewald, and Helmut Krcmar. The DevOps phenomenon. ACM Queue: Tomorrow’s Computing Today, 17(2):40, March 2019. CODEN AQCUAE.

ISSN 1542-7730 (print), 1542-7749 (electronic). URL http:/

/queue.acm.org/detail.cfm?
id=3338532.

Wadlow:2006:CCM

[WG06] Thomas Wadlow and Vlad Gorelik. Criminal code: the making of a cybercriminal. ACM Queue: Tomorrow’s Computing Today, 4(9):36–41, November 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wilde:2008:XF

[WG08] Erik Wilde and Robert J.

Glushko. XML fever. ACM Queue: Tomorrow’s Computing Today, 6(6):46–53, October 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wadlow:2009:SB

[WG09] Thomas Wadlow and Vlad Gorelik. Security in the browser. ACM Queue: Tomorrow’s Computing Today, 7 (2):40–41, February 2009. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

White:2005:STS

[Whi05] William W. White. Sifting through the software sandbox: SCM meets QA. ACM Queue: Tomorrow’s Computing Today, 3(1):38–44, February 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Whitston:2008:RER

[Whi08]  Howard Whitston.  Review

of ‘Essential RenderMan’ by Ian Stephenson, Springer-Verlag New York, Inc., 2008, $39.95, ISBN 1-84628-344-2. ACM Queue: Tomorrow’s Computing Today, 6(2):66, March 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Ste07b].

Winfield:2005:LRC

[WHT^{+}05] David H. Winfield, Erik Hoel, The Editors, Georges Brigham, and Stan Kelly-Bootle. Letters: Readers’ comments are important, too; APL = A possible lingua franca? ACM Queue: Tomorrow’s Computing Today, 3(4):10, May 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wieringa:2003:DMR

[Wie03] Roel Wieringa. Design methods for reactive systems: Yourdan, Statemate, and the UML. Morgan Kaufmann Publishers, 2929 Campus Drive, Suite 260, San Mateo, CA 94403, USA, 2003. ISBN 1-55860-755-2. xxv + 457 pp. LCCN QA76.76.D47 W525 2003. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/els031/2002112068.
html;   http://www.loc.gov/
catdir/toc/els031/2002112068.
html.

Wilson:2004:EPC

[Wil04] Gregory V. Wilson. Extensible programming for the 21st Century. ACM Queue: Tomorrow’s Computing Today, 2(9): 48–57, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

White:2008:BSB

[WKGD08] Walker White, Christoph Koch, Johannes Gehrke, and Alan Demers. Better scripts, better games. ACM Queue: Tomorrow’s Computing Today, 6 (7):18–25, November/December 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Whitney:2003:SSH

[WNN03] Telle Whitney and George Neville-Neil. SoC: Software, hardware, nightmare, bliss. ACM Queue: Tomorrow’s Computing Today, 1(2):24–31, April 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wolfe:2003:TEP

[Wol03a]  Alexander Wolfe.  Toolkit:

Eclipse: a platform becomes an open-source Woodstock. ACM Queue: Tomorrow’s Computing Today, 1(8):14–16, November 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wolfe:2003:TGT

[Wol03b] Alexander Wolfe. Toolkit: GNU tools, still relevant? ACM Queue: Tomorrow’s Computing Today, 1(9):14, 16–17, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wolfe:2003:TMC

[Wol03c] Alexander Wolfe. Toolkit: Microsoft’s Compact Framework targets smart devices. ACM Queue: Tomorrow’s Computing Today, 1(7):10–12, October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wolfe:2004:TSD

[Wol04a] Alex Wolfe. Toolkit: Samba does Windows-to-Linux dance. ACM Queue: Tomorrow’s Computing Today, 2(5):18–21, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wolfe:2004:TSS

[Wol04b] Alexander Wolfe. There’s still some life left in Ada. ACM Queue: Tomorrow’s Computing Today, 2(7):28–31, October 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wolfe:2004:TGY

[Wol04c] Alexander Wolfe. Toolkit: Get your graphics on: OpenGL advances with the times. ACM Queue: Tomorrow’s Computing Today, 2(1):10, 12–13, March 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See minor erratum [SMM^{+}04].

Wolfe:2004:GTC

[Wol04d] Alexander Wolfe. Toolkit: Grid tools: Coming to a cluster near you. ACM Queue: Tomorrow’s Computing Today, 2(4):20–23, June 2004.

Wolfe:2004:TIS

[Wol04e] Alexander Wolfe. Toolkit: Intel is stealth source of heavy-duty software tools. ACM Queue: Tomorrow’s Computing Today, 2(2):12, 14, 16–17, April 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wolfe:2004:TJJ

[Wol04f] Alexander Wolfe. Toolkit: Java is jumpin’. ACM Queue: Tomorrow’s Computing Today, 1 (10):16–19, February 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wolfe:2004:TLT

[Wol04g]  Alexander Wolfe.  Toolkit:

Longhorn ties platform apps to core operating system. ACM Queue: Tomorrow’s Computing Today, 2(6):16–19, September 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wolf:2006:RSA

[Wol06] Markus Wolf. Review of “Spring in Action” by Craig Walls and Ryan Breidenbach, Manning Publications, 2005, $44.95, ISBN 1-932394-35-4. ACM Queue: Tomorrow’s Computing Today, 4(9):50, November 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [WB05].

Wolf:2008:BRC

[Wol08] Markus Wolf. Book review: Concurrent and Real-time Programming in Ada (3rd edition), Alan Burns, Andy Wellings, Cambridge University Press, 2007, $75.00, ISBN: 978-0-521-86697-2. ACM Queue: Tomorrow’s Computing Today, 6 (6):54, October 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Woods:2004:SVF

[Woo04] William A. Woods. Searching vs. finding. ACM Queue: Tomorrow’s Computing Today, 2(2):26–35, April 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Woods:2010:CDC

[Woo10] Andy Woods. Cooling the data center. ACM Queue: Tomorrow’s Computing Today, 8 (3):10, March 2010. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wang:2005:COP

[WQ05] Andy Ju An Wang and Kai Qian. Component-oriented programming. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2005. ISBN 0-471-64446-

3. xi + 319 pp. LCCN QA76.76.C66 W36 2005. URL

http://www.loc.gov/catdir/
enhancements/fy0616/2004015681-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0616/
2004015681-d.html;   http:
//www.loc.gov/catdir/toc/
ecip0420/2004015681.html.

Waldo:2009:TAR

[WRS^{+}09] Jim Waldo, Alan Ramos, Weina Scott, William Scott, Doug Lloyd, and Katherine O’Leary. A threat analysis of RFID passports. ACM Queue: Tomorrow’s Computing Today, 7(9): 10, October 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Whittaker:2003:BBD

[WT03] James A. Whittaker and Herbert H. Thompson. Black box debugging. ACM Queue: Tomorrow’s Computing Today, 1(9):68–74, December/January 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Ward:2022:RBS

[WWG22] Peter Ward, Paul Wankadia, and Kavita Guliani. Reinventing backend subsetting at Google: Designing an algorithm with reduced connection churn that could replace deterministic subsetting. ACM Queue: Tomorrow’s Computing Today, 20 (5):33–57, September 2022. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3570937.

Xin:2021:ITC

[XMGM21] Reynold Xin, Wes McKinney, Alan Gates, and Chris McCubbin. It takes a community: The open-source challenge. ACM Queue: Tomorrow’s Computing Today, 19(5):115–136, September 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3494834.3498568.

Yu:2016:PDA

[YD16] Meng-Day (Mandel) Yu and Srinivas Devadas. Pervasive, dynamic authentication of physical items. ACM Queue: Tomorrow’s Computing Today, 14(6): 70, November 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3047967.

Yang:2016:RPW

[YRZB16] Jean Yang, Vijay Janapa Reddi, Yuhao Zhu, and Peter Bailis. Research for practice: Web security and mobile Web computing. ACM Queue: Tomorrow’s Computing Today, 14 (4):80, July 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3005356.

Zhang:2016:RPD

[ZAB16] Irene Zhang, Fadel Adib, and Peter Bailis. Research for practice: Distributed transactions and networks as physical sensors. ACM Queue: Tomorrow’s Computing Today, 14(5): 60, September 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3023318.

Zahran:2016:HCH

[Zah16]  Mohamed Zahran.  Hetero-

geneous computing: Here to stay. ACM Queue: Tomorrow’s Computing Today, 14(6): 40, November 2016. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

URL http://queue.acm.org/
detail.cfm?id=3038873.

Zandstra:2004:POP

[Zan04] Matt Zandstra. PHP 5 objects, patterns, and practice. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-380-4 (paperback). xvii + 437 pp. LCCN QA76.73.P224 Z35 2004. URL

http://www.loc.gov/catdir/
enhancements/fy0662/2005280419-
d.html.

Zandstra:2008:POP

[Zan08] Matt Zandstra. PHP objects, patterns, and practice. Apress, Berkeley, CA, USA, second edition, 2008. ISBN 1-59059-909-

8. xxiii + 487 pp. LCCN QA76.73.P224 Z35 2008.

Zeller:2006:WPF

[Zel06]  Andreas Zeller.  Why pro-

grams fail: a guide to systematic debugging. Elsevier/ Morgan Kaufmann, San Francisco, CA, USA, 2006. ISBN 1-55860-866-4, 3-89864-279-8. xxviii + 448 pp. LCCN QA76.9.D43 Z45 2006. URL

http://www.loc.gov/catdir/
enhancements/fy0625/2005049837-
t.html;   http://www.loc.
gov/catdir/enhancements/fy0742/
2005049837-d.html.

Zeldman:2007:DWS

[Zel07] Jeffrey Zeldman. Designing with Web standards. Voices that matter. New Riders Publishing, Carmel, IN, USA, second edition, 2007. ISBN 0-321-38555-1 (paperback). xx + 410 pp. LCCN TK5105.888

Z46 2007. URL http://www.
loc.gov/catdir/toc/fy0705/
2006283016.html.

ZenilChavez:2004:BRQ

[Zen04] Hector Zenil Chavez. Book review: Quantum Computing (Natural Computing Series), 2nd ed., by Mika Hirvensalo, Springer-Verlag, 2004, $54.95, ISBN 3-540-40704-9. ACM Queue: Tomorrow’s Computing Today, 2(6):76, September 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Zenel:2005:EGW

[Zen05] Bruce Zenel. Enterprise-grade wireless. ACM Queue: Tomorrow’s Computing Today, 3 (4):30–37, May 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Zhuk:2004:IRA

[Zhu04] Jeff Zhuk. Integration-ready architecture and design: software engineering with XML, Java, .NET, wireless, speech, and knowledge technologies. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-52583-7 (paperback). xxx + 609 pp. LCCN QA76.758 .Z48 2004.

URL ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/
;   http://www.loc.gov/catdir/
description/cam041/2003065381.
html;   http://www.loc.gov/
catdir/toc/cam041/2003065381.
html.

Zandstra:2009:BR

[ZM09] Matt Zandstra and Robert Martin. Book reviews: PHP Objects, Patterns, and Practice (second edition), Matt Zandstra, Apress, 2007, $44.99, ISBN 1-59059-909-8, Clean Code: A Handbook of Agile Software Craftsmanship, Robert Martin, Prentice Hall PTR, 2008, $42.99, ISBN 0-13-235088-2. ACM Queue: Tomorrow’s Computing Today, 7(2):75, February 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Zan08] and [M^{+}09].